### v2

# On a novel UCP result and its application to inverse conductive scattering

Abstract.
In this paper, we derive a novel Unique Continuation Principle (UCP) for a system of second-order elliptic PDEs system and apply it to investigate inverse problems in conductive scattering. The UCP relaxes the typical assumptions imposed on the domain or boundary with certain interior transmission conditions. This is motivated by the study of the associated inverse scattering problem and enables us to establish several novel unique identifiability results for the determination of generalized conductive scatterers using a single far-field pattern, significantly extending the results in [15, 21]. A key technical advancement in our work is the combination of Complex Geometric Optics (CGO) techniques from [15, 21] with the Fourier expansion method to microlocally analyze corner singularities and their implications for inverse problems. We believe that the methods developed can have broader applications in other contexts.
Keywords:  Unique Continuation Principle, conductive medium scattering, inverse problem, uniqueness, single far-filed measurement, polygonal-nest and polygonal-cell structures
2020 Mathematics Subject Classification: 35Q60, 78A46 (primary); 35P25, 78A05, 81U40 (secondary).
1. Introduction
1.1. The main results.
This paper addresses a novel type of the unique continuation principle (UCP) for a system of partial differential equations (PDEs) and applies this principle to investigate inverse problems in conductive medium scattering. In this subsection, we outline the main results of our study, deferring detailed proofs to subsequent sections. Initially, we introduce the geometric setup and notations for our study.
Let $(r,\vartheta)$ denote the polar coordinates in $\mathbb{R}^{2}$, where $\mathbf{x}=(x_{1},x_{2})=(r\cos\vartheta,r\sin\vartheta)\in\mathbb{R}^{2}$. For a given point $\mathbf{x}_{0}$, we denote $B_{r_{0}}(\mathbf{x}_{0})$ as the open disk centered at $\mathbf{x}_{0}$ with radius $r_{0}\in\mathbb{R}_{+}$, and we simply denote $B_{r_{0}}$ as $B_{r_{0}}(\mathbf{0})$. Let a sector $\mathcal{S}$ be defined as follows:
$\mathcal{S}=\{\mathbf{x}\in\mathbb{R}^{2}~{}|~{}\vartheta_{m}<\arg(x_{1}+\mathrm{i}x_{2})<\vartheta_{M}\},$ (1.1)
where $-\pi\leq\vartheta_{m}<\vartheta_{M}<\pi$, $\vartheta_{M}-\vartheta_{m}\in(0,\pi)$, and $\mathrm{i}=\sqrt{-1}$. Additionally, the two boundary lines of $\mathcal{S}$ are given by
$\Gamma^{+}=\{\mathbf{x}\in\mathbb{R}^{2}~{}|~{}\mathbf{x}=(r\cos\vartheta_{M},r\sin\vartheta_{M})^{\top},r>0\},$
$\Gamma^{-}=\{\mathbf{x}\in\mathbb{R}^{2}~{}|~{}\mathbf{x}=(r\cos\vartheta_{m},r\sin\vartheta_{m})^{\top},r>0\}.$
Furthermore, define
$\mathcal{S}_{r_{0}}=\mathcal{S}\cap B_{r_{0}},\ \Gamma^{\pm}_{r_{0}}=\Gamma^{\pm}\cap B_{r_{0}},\ \Lambda_{r_{0}}=\mathcal{S}\cap\partial B_{r_{0}}.$ (1.2)
The opening angle of $\mathcal{S}_{r_{0}}$ is $\vartheta_{M}-\vartheta_{m}$. If $\frac{\vartheta_{M}-\vartheta_{m}}{\pi}$ is a rational number, $\mathcal{S}_{r_{0}}$ is termed a rational corner; otherwise, it is an irrational corner.
Theorem 1.1 shows a novel type of UCP result associated with a PDE system described by (1.3). The detailed proof of Theorem 1.1 is given in Section 2.
Theorem 1.1.
Let $D$ and $\Omega$ be two bounded Lipschitz domains in $\mathbb{R}^{2}$ with connected complements $\mathbb{R}^{2}\setminus\overline{D}$ and $\mathbb{R}^{2}\setminus\overline{\Omega}$ respectively, where $\Omega\subset D$. Assume $\gamma_{1}\in\mathbb{R}_{+}$ and $\gamma_{2}\in L^{\infty}(D)\cap H^{2}(\Omega)$ with ${\rm supp}(\gamma_{2}-\gamma_{1})\subset\Omega$ are given. Suppose that $\Omega\cap B_{r_{0}}={\mathcal{S}}_{r_{0}}$, where ${\mathcal{S}}_{r_{0}}$ is defined by (1.2), $r_{0}\in\mathbb{R}_{+}$ is sufficiently small and $\mathbf{0}\in\partial\Omega$. The line segements $\Gamma^{\pm}_{r_{0}}$ are defined in (1.2). Consider $\mathbf{u}=(v,w)^{\top}\in H^{1}(D)\times H^{1}(D)$ solving the PDE system:
$\begin{cases}\mathcal{P}\mathbf{u}=0\quad\mbox{in}\quad D,\\ \mathcal{B}\mathbf{u}=0\quad\mbox{on}\quad\Gamma^{\pm}_{r_{0}},\end{cases}$ (1.3)
where
$\mathcal{P}=\left(\begin{matrix}\Delta+\gamma_{1}&0\\ 0&\Delta+\gamma_{2}\end{matrix}\right),\quad\mathcal{B}=\left(\begin{matrix}1&-1\\ \partial_{\nu}+\eta&-\partial_{\nu}\end{matrix}\right).$ (1.4)
Here the constant $\eta\in\mathbb{C}$ is defined on $\Gamma_{r_{0}}^{\pm}$ and $\partial_{\nu}v$ is the exterior normal derivative of $v$ with $\nu$ being the exterior normal of $\partial\Omega$. If $\mathcal{S}_{r_{0}}$ forms an irrational corner and $\eta\neq 0$, then $\mathbf{u}=\mathbf{0}$ in $D$.
Remark 1.1.
UCP is a fundamental concept in the theory of elliptic partial differential equations (PDEs). For instance, UCP can be applied to investigate the solvability and stability of partial differential equations [3, 48, 47]. In Theorem 1.1, we establish a novel UCP for the PDE system (1.3). Specifically, if a solution $\mathbf{u}=(v,w)^{\top}$ to (1.3) exists and satisfies certain boundary conditions described by (1.3) on two intersecting line segments within $D$, forming an irrational convex corner, then $\mathbf{u}$ must be identically zero in $D$. Furthermore, the boundary conditions of $\mathbf{u}$ on two intersecting line segments describes certain transmission properties of $v$ and $w$ across these segments. This type of transmission boundary condition is particularly relevant for studying the conductive scattering problem, which will be discussed in detail in subsection 1.2. To the best of our knowledge, this type of UCP is particularly intriguing compared to existing literature on UCP for elliptic systems. It has important applications in the study of the inverse conductive medium scattering problem (1.9), which will be further elaborated in the subsequent subsection.
The PDE system described by (1.3) in Theorem 1.1 originates from the invisibility phenomena associated with the time-harmonic conductive medium scattering problem (1.7), which will be further elaborated upon in subsection 1.2. The direct scattering problem (1.7) involves solving for the scattered wave $u^{s}$ given the incident wave $u^{i}$ and the conductive medium scatterer $(\Omega,q,\eta)$, where $\Omega$ is a bounded Lipschitz domain in $\mathbb{R}^{2}$ with a connected complement such that $\Omega\Subset B_{R}\ (R\in\mathbb{R}_{+})$ and $u^{i}$ can be either a plane wave of the form
$u^{i}(\mathbf{x})=e^{{\rm i}k\mathbf{x}\cdot\mathbf{d}}$ (1.5)
or a point source of the form
$u^{i}(\mathbf{x};\mathbf{z}_{0})=H_{0}^{1}(k|\mathbf{x}-\mathbf{z}_{0}|)$ (1.6)
associated with a wave number $k\in\mathbb{R}_{+}$. Here, $\mathbf{d}\in\mathbb{S}^{1}$ is the incident direction of the plane wave, $H_{0}^{1}$ denotes the zeroth-order Hankel function of the first kind, and $\mathbf{z}_{0}$ signifies the location of the point source with $\mathbf{z}_{0}\in\mathbb{R}^{2}\setminus\overline{B}_{R}$ with $\Omega\Subset B_{R}$ and $R\in\mathbb{R}_{+}$. The function $q\in L^{\infty}(\Omega)$ characterizes the refractive index of the medium scatterer $\Omega$ with ${\rm supp}(q)\subset\Omega$, and the constant conductive parameter $\eta$ defined on $\partial\Omega$ signifies the conductive property of $\Omega$. On the other hand, the inverse problem corresponding to (1.7) aims to determine the shape of the underlying conductive medium scatterer $(\Omega,q,\eta)$ and its physical configuration $(q,\eta)$ through the measurement of the scattered wave. The novel UCP derived in Theorem 1.1 has several applications in the conductive medium scattering problem (1.7). Specifically, in Theorem 1.2, utilizing the UCP for (1.3), it can be shown that the conductive medium scatterer in (1.7) containing a convex irrational corner cannot be invisible. The detailed proof of Theorem 1.2 is postponed to Section 2.
Theorem 1.2.
Consider the conductive medium scattering problem modeled by (1.7), associated with a given incident wave $u^{i}$ with the wave number $k$, as defined by (1.5) or (1.6). Let the bounded Lipschitz domain $\Omega\subset\mathbb{R}^{2}$ with a connected complement in (1.7) describe the corresponding conductive medium scatterer $(\Omega,q,\eta)$, where ${\rm supp}(q)\subset\Omega$, $q\in L^{\infty}(\Omega)$ and $\eta$ is a nonzero constant. If $\Omega$ has a convex irrational corner $\mathbf{x}_{0}$ with $q\in L^{\infty}(\Omega)\cap H^{2}(\Omega\cap B_{r_{0}}(\mathbf{x}_{0}))$, where $r_{0}\in\mathbb{R}_{+}$ is sufficiently small, then $\Omega$ cannot be invisible for any wave number $k\in\mathbb{R}_{+}$.
Using the unique continuation property (UCP) for (1.3) as established in Theorem 1.1, we can address the unique identifiability for the inverse problem of the conductive medium scattering problem (1.7). Specifically, in Theorem 1.3, we present both local and global unique determination of a conductive medium scatterer based on a single measurement of the scattered wave, where the single measurement refers to that taken under a fixed incident wave. The unique identifiability by a single measurement is challenging and has a long and colorful history in inverse scattering problems, often referred to as Schiffer problem in the literature; see [19, 18] and the references therein for more details. The detailed statement of Theorem 1.3 is provided in Theorem 2.1, along with the corresponding proof.
Theorem 1.3.
Considering the conductive medium scattering problem (1.7), if the measurements of the two scattered wave corresponding to two conductive medium scatterers under a fixed incident wave are identical, then the difference between these two scatterers cannot contain a convex irrational corner. Furthermore, if two convex irrational polygonal conductive medium scatterers, $\Omega_{1}$ and $\Omega_{2}$, produce the same scattered wave measurements for a fixed incident wave, then $\Omega_{1}$ and $\Omega_{2}$ must be identical.
When the conductive medium scatterer associated with (1.7) has a polygonal-cell or polygonal-nest structure (see Definitions 3.2 and 3.3 for rigorous explanations), we can uniquely identify the polygonal-nest structure based on a single measurement of the scattered wave, given some a-priori knowledge of the scatterer (see the detailed a-prior information in Definition 3.5). This identification leverages the novel UCP for (1.3). The detailed statements of these uniqueness results can be found in Theorems 3.1 and 3.3.
In addition to determining the shape of the conductive medium scatterer using the newly derived UCP in Theorem 1.1, we also establish uniqueness results for identifying the piecewise-linear-polynomial refractive index and constant conductive parameter within the conductive polygonal-cell or polygonal-nest medium scatterer. This is achieved under some a priori information based on a single measurement of the scattered wave. These results are detailed in Theorems 3.2 and 3.3.
1.2. The conductive medium scattering and invisibility
The conductive medium scattering problem has important and practical applications, such as the modelling of an electromagnetic object coated with a thin layer of a highly conductive material and magnetotellurics in geophysics[5]. Indeed, this kind of the conductive medium scattering problem can be derived by TM (transverse magnetic) polarisation from the corresponding Maxwell scattering system; see more details in [11, 15].
Let $\Omega$ be a conductive medium scatterer suited in a homogenous isotropic background medium $\mathbb{R}^{2}$ with the medium configurations $q$ and $\eta$, which is a bounded Lipschitz domain with a connected complement $\mathbb{R}^{2}\setminus\overline{\Omega}$. The medium parameter $q\in L^{\infty}(\Omega)$ characterises the refractive index of the medium $\Omega$ with ${\rm supp}(q-1)\subset\Omega$ and the constant $\eta$ defined on $\partial\Omega$ signifies the conductive property of $\Omega$. In what follows, we write $(\Omega;q,\eta)$ to describe the conductive medium scatterer $\Omega$ with the physical configurations $q$ and $\eta$.
For a given incident wave $u^{i}$ impinging on $\Omega$, it generates a scattered wave $u^{s}$. The total wave $u$ is given by $u=u^{i}+u^{s}$. The incident wave $u^{i}$ can be either a plane wave of the form (1.5) or a point source of the form (1.6). The conductive medium scattering problem in $\mathbb{R}^{2}$ can be modeled as:
$\begin{cases}\Delta u^{-}+k^{2}qu^{-}=0&\text{in }\Omega,\\[5.0pt] \Delta u^{+}+k^{2}u^{+}=0&\text{in }\mathbb{R}^{2}\setminus\overline{\Omega},\\[5.0pt] u^{+}=u^{-},\quad\partial_{\nu}u^{-}=\partial_{\nu}u^{+}+\eta u^{+}&\text{on }\partial\Omega,\\[5.0pt] u^{+}=u^{i}+u^{s}&\text{in }\mathbb{R}^{2}\setminus\overline{\Omega},\\[5.0pt] \lim_{r\to\infty}r^{1/2}\left(\partial_{r}u^{s}-iku^{s}\right)=0,&\quad r=|\mathbf{x}|,\end{cases}$ (1.7)
where $k\in\mathbb{R}_{+}$ is the wave number, and $\nu\in\mathbb{S}^{1}$ is the exterior unit normal vector to $\partial\Omega$. The third equation in (1.7) describes the conductive transmission property of the total wave $u$ across $\partial\Omega$. The final equation refers to the Sommerfeld radiation condition, characterizing the outgoing nature of the scattered wave $u^{s}$. Moreover, $u^{\pm}$ and $\partial_{\nu}^{\pm}u$ denote the limits of $u$ and $\partial_{\nu}u$ from the exterior and interior of $\Omega$, respectively.
The direct problem (1.7) aims to solve the scattered wave $u^{s}$ given the conductive medium scatterer $(\Omega;q,\eta)$ and the incident wave $u^{i}$. For the well-posedness of the direct problem (1.7) it was proved that there exists a unique solution $u\in H^{1}_{loc}(\mathbb{R}^{2})$ to (1.7) (cf. [9, 11]). Moreover, the scattered field $u^{s}$ possesses the following asymptotic behavior
$u^{s}(\mathbf{\hat{x}})=\frac{e^{\mathrm{i}k|\mathbf{x}|}}{|\mathbf{x}|^{\frac{1}{2}}}\left(u^{\infty}({\mathbf{\hat{x}}})+\mathcal{O}(|\mathbf{x}|^{-\frac{1}{2}})\right),\ |\mathbf{x}|\to\infty,$ (1.8)
where $\mathbf{\hat{x}}=\frac{\mathbf{x}}{|\mathbf{x}|}$ and $u^{\infty}$ is the far-field pattern of $u^{s}$ associated with $u^{i}$. Here it can be checked that the far-field pattern $u^{\infty}$, which is a real analytic function defined in $\mathbb{S}^{1}$, corresponds one-to-one to $u^{s}$ and encodes the information of the scatterer $(\Omega;q,\eta)$. From (1.8) we define the forward operator by
$\mathcal{F}((\Omega;q,\eta);u^{i})\rightarrow u^{\infty}(\mathbf{\hat{x}}).$ (1.9)
The inverse problem associated with (1.9) can be formulated as
$u^{\infty}(\mathbf{\hat{x}};u^{i})\to(\Omega;q,\eta),$ (1.10)
which intends to identify $\Omega$ and its physical configurations $q$ and $\eta$. This type of inverse problem has practical applications in non-destructive testing and radar [18], etc. In this paper we will mainly focus on the uniquely determination of $(\Omega,q,\eta)$ by a single far-field measurement, namely that the far-field pattern $u^{\infty}(\mathbf{\hat{x}};u^{i})$ is generated by a fixed incident wave $u^{i}$ and a fixed wave number $k$, otherwise we say that it refers to many far-field measurements.
When $u^{\infty}\equiv 0$, no scattering pattern can be observed outside $\Omega$, and hence the scatterer $(\Omega,q,\eta)$ is invisible/transparent with respect to the exterior observation under the wave interrogation by $u^{i}$. Then by Rellich’s theorem, which indicates that $u^{+}=u^{i}$ in $\mathbb{R}^{2}\setminus\overline{\Omega}$, one has $(w,v)=(u^{-}|_{\Omega},u^{i}|_{\Omega})$ satisfies the following transmission eigenvalue problem:
$\begin{cases}\Delta w+k^{2}(1+V)w=0\hskip 25.6073pt\mbox{in}\ \Omega,\\ \Delta v+k^{2}v=0\hskip 66.86414pt\mbox{in}\ \Omega,\\ w=v,\ \partial_{\nu}w=\partial_{\nu}v+\eta v\hskip 19.91684pt\mbox{on}\ \Gamma,\end{cases}$ (1.11)
where $V=q-1$ and $\Gamma=\partial\Omega$.
On the other hand, given the bounded Lipschitz domain $\Omega$, $V\in L^{\infty}(\Omega)$ and the boundary parameter $\eta\in L^{\infty}(\partial\Omega)$, consider the PDE system (1.11). If there exists a nontrivial pair of solutions $(w,v)$ to (1.11) associated with $k$, then $k$ is referred to as a transmission eigenvalue and $(w,v)$ is the corresponding pair of transmission eigenfunctions of (1.11). Here, we would like to emphasise that throughout the article, $\eta$ can be taken to be identically zero, which reduces to the case with a normal scattering inhomogeneity of refractive index $q$, and all the results still hold equally. Hence, $(\Omega;q,\eta)$ is actually a generalized conductive medium scatterer and the transmission eigenvalue problem (1.11) is generalized transmission eigenvalue problem in our study, but we shall not distinguish it.
In the following, we shall investigate the connection between the PDE system (1.3) in Theorem 1.1 and the conductive medium scattering (1.7) when the invisibility of the conductive medium scatterer $(\Omega;q,\eta)$ occurs. Recall that the incident wave $u^{i}$ is given by (1.5) or (1.6). Since $\Omega$ is bounded and has a connected complement, let $D:=B_{R}$ such that $\Omega\Subset B_{R}$, where $B_{R}$ is a disk centered at the origin with radius $R$. As discussed earlier, when $(\Omega;q,\eta)$ is invisible/transparent, we have $u^{\infty}\equiv 0$, and by Rellich’s theorem, it follows that $u^{s}\equiv 0$ in $D\backslash\overline{\Omega}$, which implies that $u^{+}\equiv u^{i}$ in $D\backslash\overline{\Omega}$. According to (1.7), we can readily derive:
$\begin{cases}\mathcal{P}\mathbf{u}=0\quad\mbox{in}\quad D,\\ \mathcal{B}\mathbf{u}=0\quad\mbox{on}\quad\partial\Omega,\end{cases}$ (1.12)
where $\mathbf{u}=(u^{i},u^{-}\chi_{\Omega}+u^{i}\chi_{D\backslash\overline{\Omega}})\in H^{1}(D)\times H^{1}(D)$, $\mathcal{P}$ and $\mathcal{B}$ are defined in (1.4) with $\gamma_{1}=k^{2}$ and $\gamma_{2}=k^{2}q\chi_{\Omega}+k^{2}\chi_{D\backslash\overline{\Omega}}$. When $\Omega$ has a convex corner, meaning there are two intersecting line segments lying on $\partial\Omega$ where the intersection point belongs to $\partial\Omega$, we can directly obtain (1.3) from (1.12).
1.3. Connections to existing results and main contributions
The UCP is a fundamental theory with various applications in solvability, controllability, optimal stability, and inverse problems for elliptic partial differential equations (cf. [18, 31, 32]). Several forms of UCP exist for second-order elliptic partial differential equations, including the weak UCP, strong UCP, UCP for Cauchy data, and UCP across a hypersurface [26, 48, 38, 4]. There are multiple approaches to establishing UCP for elliptic equations, such as doubling inequalities, Carleman inequalities, and three spheres inequalities. For further discussions on this topic, we refer to [16, 26, 25, 31, 32, 33, 34, 42, 46] and references therein for a historical overview. Recently, the UCP result for the (fractional) Laplacian equation in Euclidian space or a Riemannian manifold can be found in [27, 37]. For elliptic PDE systems, [44] proved that the zero set of least energy solutions for Lane-Emden systems with Neumann boundary conditions has zero measure. The weak UCP for a general second-order elliptic system was established in [30] under certain assumptions. Moreover, [47] derived UCP results for elliptic PDE systems, including strong UCP, weak UCP, and UCP for local Cauchy data.
As discussed in the previous subsection, the invisibility or transparency in inverse scattering problems implies that the scatterer cannot be detected by external measurements, as the probing incident wave remains unperturbed and the corresponding scattered field is identically zero. However, when the scatterer exhibits geometrical singularities, such as corners or points of high curvature, as shown in [8, 12, 41, 24, 6] for time-harmonic acoustic scattering, invisibility or transparency cannot occur for any frequency associated with an incident plane wave. Recently, new methods based on free boundary problems have been introduced to investigate invisibility or transparency in isotropic and anisotropic acoustic medium scattering [13, 45, 14, 35].
The quest to determine the geometrical and physical properties of scatterers through a single far-field measurement boasts a rich and extensive history in the realm of inverse scattering [2, 7, 18, 19, 43, 39]. This pursuit, often referred to as the Schiffer problem [19], aims to uniquely identify the location and shape of scatterers based on a single far-field measurement. When a scatterer is impenetrable, it is termed an obstacle; otherwise, it is considered a medium. In the inverse obstacle scattering, Lax and Phillips [36] demonstrated that the shape of a sound-soft obstacle can be uniquely identified through infinite far-field measurements. However, achieving uniqueness with just a single far-field measurement has thus far required additional a-priori knowledge of the geometry, such as the size or shape of the scatterer(cf. [20, 17, 39]).
In the context of conductive medium scattering described by (1.7), uniqueness results for determining $\Omega$ based on the far-field patterns of all incident plane waves at a fixed frequency, namely infinitely many far-field measurements, have been obtained in [11, 10, 29]. The unique identifiability based on a single far-field measurement for recovering a convex polygon $\Omega$ and the constant boundary parameter $\eta$ has been established in [21] under the generic non-vanishing assumption on the corresponding total wave field, which can be fulfilled under certain physical scenarios. When the refractive index $q$ and the conductive boundary parameter $\eta$ are piecewise constants, characterized by polygonal-nest or polygonal-cell structures, the corresponding unique determination for $q$ and $\eta$ by a single far-field pattern under the generic condition has been derived in [15], whereas for the case $\eta\equiv 0$, the single far-field measurement uniquely identifiability for recovering $q$ appeared in [7].
The main contributions of this paper can be summarized in several key points:
  1. a)
    We derive a novel type of UCP result for an elliptic second order PDEs system (1.3), where the components $v$ and $w$ of the solution $\mathbf{u}$ to (1.3) satisfy a transmission condition across two intersecting line segments within the underlying domain. These two intersecting line segments form an irrational corner.
  2. b)
    We reveal that when the conductive medium scatter $(\Omega,q,\eta)$ has a convex irrational corner $\mathbf{x}_{0}$ and the physical parameter $q$ satisfies $H^{2}$ regularity near $\mathbf{x}_{0}$, then $\Omega$ radiates any wave number. Furthermore, we demonstrate that an admissible complex irrational polygonal scatterer can be uniquely determined using only a single far-field measurement, independent of its medium content.
  3. c)
    We demonstrate that both the shape and medium parameters of a conductive scatterer within the polygonal-nest or polygonal-cell structure can be uniquely determined through a single far-field measurement.
For the main contribution a), in contrast to the existing weak, strong, or Cauchy-type UCPs for elliptic systems in the literature, we consider the UCP for the case where the components of the solution to (1.3) fulfill a transmission boundary condition on two intersecting line segments within the domain, forming an irrational corner. This transmission boundary condition arises from the invisibility or transparency phenomena in conductive medium scattering, which holds both practical and physical significance. We will discuss this aspect in more detail in Section 2. To the best of our knowledge, this represents a novel type of UCP for second-order elliptic PDE systems, playing a crucial role in investigating invisibility and uniqueness results in inverse conductive medium scattering problems.
For the main contribution b), we can substitute the technical condition on the a-priori information in [21], which requires the non-vanishing of the total wave at the corner points, with an irrationality condition related to the interior angles of the conductive polygonal medium $\Omega$. Although ensuring the non-vanishing nature of the total wave at the corner point is feasible in generic physical scenarios, the a-priori information on the irrationality of the interior angles of the underlying scatterer is less restrictive, making the new global unique identifiability more applicable in practice. We establish this uniqueness result by using the novel UCP result described in Theorem 1.1. Furthermore, using the novel UCP established in Theorem 1.1, we can show that if a conductive medium scatterer $\Omega$ has a convex irrational corner on the boundary, where the reflective index $q$ has $H^{2}$ regularity near the corner, then $\Omega$ radiates any incident wave $u^{i}$ defined by (1.5) or (1.6).
For the main contribution c), we can uniquely determine the refractive index in linear polynomial form and the constant conductive boundary parameter of the medium scatterer, whereas in [15] both the refractive index and the conductive parameter are constants. Moreover, we can relax the admissibility conditions in [15], which requires the non-vanishing of the total wave field at the corner points, to requiring the non-vanishing of either the total wave field or the gradient of the total wave field. In fact, the complexity inherent in the conductive polygonal-nest and polygonal-cell medium scatterer necessitates a detailed study of the singular behavior of the solution to a coupled conductive transmission PDE problem near a corner; see Lemma 3.2. Specifically, given that the refractive index is assumed to be a linear polynomial, it becomes crucial to demonstrate the Hölder regularity of the gradient of the total wave as derived from the direct problem (1.7), particularly near corners; see Lemma 3.1. With this regularity established, we propose another admissible condition, which describes the non-vanishing property of the gradient of the total wave field at the corner point. This admissible condition is generic, which can be fulfilled in more general physical scenarios; see more detailed discussions in Remarks 3.1 and 3.2. Under this generic admissible condition, we can then proceed recursively to determine the interior polygonal-nest structure.
The remainder of the paper is organized as follows. In Section 2, we present the proofs of Theorems 1.1-1.3. In Section 3, we present our main results regarding the uniqueness of recovering the conductive polygonal-nest and polygonal-cell medium scatterers, including the scatterers shape, the partition of their interior polygonal-nest and polygonal-cell structure and the material parameters. Section 4 is devoted to the proofs of the main theorems of Section 3.
2. The proofs of Theorems 1.1-1.3
In this section, we aim to provide the proofs of Theorems 1.1-1.3. Specifically, we present a novel UCP result for an elliptic second-order PDE system and establish both local and global uniqueness results for the shape of an admissible complex conductive polygonal medium scatterer $\Omega$, utilizing only a single far-field measurement. To prove Theorem 1.1, we first introduce some auxiliary lemmas. Drawing upon the results in [12, 41], we present the following lemma, which constructs a complex geometrical optics (CGO) solution.
Lemma 2.1.
[12, 41] Let $\widetilde{q}\in H^{1,1+\varepsilon_{0}}(\mathbb{R}^{2})$, where $\varepsilon_{0}\in(0,1)$. Suppose $\mathbf{d}\in\mathbb{S}^{1}$ and $\mathbf{d}^{\perp}\in\mathbb{S}^{1}$ satisfy that $\mathbf{d}\cdot\mathbf{d}^{\perp}=0$. Define $\rho=-\tau(\mathbf{d}+\mathrm{i}\mathbf{d}^{\perp})$, where $\tau\in\mathbb{R}_{+}$. If $\tau$ is sufficiently large, then there exists a complex geometrical optics (CGO) solution $u_{0}$ of the form
$u_{0}(\mathbf{x})=(1+\psi(\mathbf{x}))e^{\rho\cdot\mathbf{x}},$ (2.1)
which satisfies
$(\Delta+k^{2}\widetilde{q})u_{0}=0\mbox{ in }\mathbb{R}^{2}.$ (2.2)
Furthermore, it holds that
$\|\psi(\mathbf{x})\|_{H^{1,8}(\mathbb{R}^{2})}=\mathcal{O}(\tau^{-\frac{2}{3}}).$
Proposition 2.1.
[22] For a given $\zeta\in(0,e)$ and $s>0$, it yields that
$\left|\int_{\zeta}^{\infty}r^{s}e^{-\mu r}\mathrm{d}r\right|\leq\frac{2}{\Re\mu}e^{-\frac{\zeta}{2}\Re\mu},$
$\int_{0}^{\zeta}r^{s}e^{-\mu r}\mathrm{d}r=\frac{\Gamma(s+1)}{\mu^{s+1}}+\mathcal{O}(\frac{2}{\Re\mu}e^{-\frac{\zeta}{2}\Re\mu}),$ (2.3)
as $\Re{\mu}\to\infty$, where $\Gamma(s)$ stands for the Gamma function.
Moreover, let $u_{0}$ be the CGO solution with the parameter $\tau$ defined in (2.1). Suppose the corner ${\mathcal{S}}_{r_{0}}$ is defined as in (1.2). Then, there exists a vector $\mathbf{d}\in\mathbb{S}^{1}$ such that
$0<\varsigma<\mathbf{d}\cdot\mathbf{\hat{x}}\leq 1,\quad\forall\mathbf{x}\in{\mathcal{S}}_{r_{0}}\backslash\{\mathbf{0}\},\quad\mathbf{\hat{x}}=\mathbf{x}/|\mathbf{x}|,$ (2.4)
where $\varsigma$ is a positive constant depending on ${\mathcal{S}}_{r_{0}}$ and $\mathbf{d}$. For sufficiently large $\tau$, one has
$\int_{\mathcal{S}_{r_{0}}}|\mathbf{x}|^{s}|u_{0}|\mathrm{d}\mathbf{x}\lesssim\tau^{-(s+\frac{29}{12})}+\tau^{-(s+2)}+\mathcal{O}(\tau^{-1}e^{-\frac{1}{2}\varsigma r_{0}\tau}),$
$\|u_{0}(\mathbf{x})\|_{H^{1}(\Lambda_{r_{0}})}\lesssim(1+\tau)(1+\tau^{-\frac{2}{3}})e^{-\varsigma r_{0}\tau},$ (2.5)
$\left|\int_{\Gamma_{r_{0}}^{\pm}}|\mathbf{x}|^{s}\psi(\mathbf{x})e^{\rho\cdot\mathbf{x}}\mathrm{d}\sigma\right|\lesssim\tau^{-(s+\frac{7}{6})}+\mathcal{O}(\tau^{-\frac{7}{6}}e^{-\frac{1}{2}\varsigma r_{0}\tau}),$
where $e^{\rho\cdot\mathbf{x}}\ \mbox{and}\ \psi(\mathbf{x})$ are defined in (2.1). Throughout the rest of this paper, $\lesssim$” means that we only perform the leading asymptotic analysis by neglecting a generic positive constant $C$ with respect to $\tau\to\infty$, where $C$ is not a function of $\tau$. This convention is similar for $\gtrsim$”.
Lemma 2.2.
[18, Section 3.4] Suppose that $v$ is a solution to Helmholtz equation $\Delta+\gamma_{1}v=0$ in $\mathcal{S}_{r_{0}}$, where $\gamma_{1}\in\mathbb{R}_{+}$ and $\mathcal{S}_{r_{0}}$ is defined in (1.2). If $v$ is analytic in $\mathcal{S}_{r_{0}}$, then $v$ has the following spherical wave expansion in polar coordinates around the origin:
$v=\sum_{n=0}^{\infty}(a_{n}e^{\mathrm{i}n\vartheta}+b_{n}e^{-\mathrm{i}n\vartheta})J_{n}(\sqrt{\gamma_{1}}r),\ r=|\mathbf{x}|,$ (2.6)
where $\mathbf{x}=(x_{1},x_{2})=r(\cos\vartheta,\sin\vartheta)\in\mathbb{R}^{2}$, $a_{n}$ and $b_{n}$ are constants and
$J_{n}(t)=\sum_{p=0}^{\infty}\frac{(-1)^{p}}{2^{n+2p}p{!}(n+p)!}t^{n+2p}$
is the Bessel functions of order $n$.
We are in the position to give the proof of Theorem 1.1.
The proof of Theorem 1.1.
According to the statement of Theorem 1.1, we know that the solution $(v,w)$ to (1.3) satisfy the following systerm:
$\begin{cases}\Delta w+\gamma_{2}w=0\hskip 54.34486pt\mbox{in}\ \mathcal{S}_{r_{0}},\\ \Delta v+\gamma_{1}v=0\hskip 59.18178pt\mbox{in}\ \mathcal{S}_{r_{0}},\\ w=v,\ \partial_{\nu}w=\partial_{\nu}v+\eta v\hskip 11.38092pt\mbox{on}\ \Gamma_{r_{0}}^{\pm}.\end{cases}$ (2.7)
By the interior regularity of the elliptic equation, we know that $v$ is analytic in $B_{r_{0}}$. When $v$ is analytic, it was proven in [22, Theorem 3.2] that $v(\mathbf{0})=0$. Using (2.6), one has
$a_{0}+b_{0}=0,$
which indicates that
$v$ $=\sum_{n=1}^{\infty}(a_{n}e^{\mathrm{i}n\vartheta}+b_{n}e^{-\mathrm{i}n\vartheta})J_{n}(\sqrt{\gamma_{1}}r)$
$=\frac{k}{2}(a_{1}e^{\mathrm{i}\vartheta}+b_{1}e^{-\mathrm{i}\vartheta})r+\mathfrak{J_{1}}+\mathfrak{J_{2}}.$ (2.8)
Here
$\mathfrak{J_{1}}$ $=\sum_{p=1}^{\infty}(a_{1}e^{\mathrm{i}\vartheta}+b_{1}e^{-\mathrm{i}\vartheta})\frac{(-1)^{p}(\sqrt{\gamma_{1}}r)^{2p+1}}{2^{2p+1}p!(p+1)!},\ \mathfrak{J_{2}}=\sum_{n=2}^{\infty}\sum_{p=0}^{\infty}(a_{n}e^{\mathrm{i}n\vartheta}+b_{n}e^{-\mathrm{i}\vartheta})\frac{(-1)^{p}(\sqrt{\gamma_{1}}r)^{n+2p}}{2^{n+2p}p!(n+p)!}.$
In what follows, we shall prove this theorem by using mathematical induction. Namely, we shall prove that $a_{\ell}=b_{\ell}=0$ for $\ell\in\mathbb{N}$. Since $\gamma_{2}\in H^{2}(\mathcal{S}_{r_{0}})$, let $\widetilde{\gamma_{2}}$ be the Sobolev extension of $\gamma_{2}$ in $\mathbb{R}^{2}$. According to Lemma 2.1, there exists the CGO solution $u_{0}$ given by (2.1), which satisfies
$(\Delta+\widetilde{\gamma_{2}})u_{0}=0\ \mbox{in}\ \mathbb{R}^{2}.$
Hence, it yields that
$(\Delta+\gamma_{2})u_{0}=0\mbox{ in }\mathcal{S}_{r_{0}}.$ (2.9)
Denoting $u=w-v$ and using (2.7), one has
$\begin{cases}\Delta u+\gamma_{2}u=(\gamma_{1}-\gamma_{2})v,\hskip 21.33955pt\mbox{in}\ \mathcal{S}_{r_{0}},\\ u=0,\ \partial_{\nu}u=\eta v,\hskip 51.21504pt\mbox{on}\ \Gamma_{r_{0}}^{\pm}.\end{cases}$ (2.10)
Combing (2.10) with (2.9), utilizing Green’s formula, we derive the following integral equation
$\eta\int_{\Gamma^{\pm}_{r_{0}}}ve^{\rho\cdot\mathbf{x}}\mathrm{d}\sigma$ $=-\eta\int_{\Gamma^{\pm}_{r_{0}}}v\psi(\mathbf{x})e^{\rho\cdot\mathbf{x}}\mathrm{d}\sigma+\int_{\mathcal{S}_{r_{0}}}(\gamma_{1}-\gamma_{2})vu_{0}\mathrm{d}\mathbf{x}$
$+\int_{\Lambda_{r_{0}}}\partial u_{0}u-\partial_{\nu}uu_{0}\mathrm{d}\sigma.$ (2.11)
Substituting (2.8) into (2.11), we have
$\frac{{\sqrt{\gamma_{1}}}\eta}{2}\int_{\Gamma_{r_{0}}^{\pm}}(a_{1}e^{\mathrm{i}\vartheta}+b_{1}e^{-\mathrm{i}\vartheta})re^{\rho\cdot\mathbf{x}}\mathrm{d}\sigma=\sum_{i=1}^{5}\mathcal{G}_{i},$ (2.12)
where
$\mathcal{G}_{1}$ $=-\eta\int_{\Gamma_{r_{0}}^{\pm}}v\psi(\mathbf{x})e^{\rho\cdot\mathbf{x}}\mathrm{d}\sigma,\ \mathcal{G}_{2}=\int_{\mathcal{S}_{r_{0}}}(\gamma_{1}-\gamma_{2})vu_{0}\mathrm{d}\mathbf{x},\ \mathcal{G}_{3}=\int_{\Lambda_{r_{0}}}\partial_{\nu}u_{0}u-\partial_{\nu}uu_{0}\mathrm{d}\sigma,$
$\mathcal{G}_{4}$ $=-\eta\sum_{p=1}^{\infty}\frac{(-1)^{p}(\sqrt{\gamma_{1}})^{2p+1}}{2^{2p+1}p!(p+1)!}\int_{\Gamma_{r_{0}}^{\pm}}(a_{1}e^{\mathrm{i}\vartheta}+b_{1}e^{-\mathrm{i}\vartheta})r^{2p+1}e^{\rho\cdot\mathbf{x}}\mathrm{d}\sigma,$
$\mathcal{G}_{5}$ $=-\eta\sum_{n=2}^{\infty}\sum_{p=0}^{\infty}\frac{(-1)^{p}(\sqrt{\gamma_{1}})^{n+2p}}{2^{n+2p}p!(n+p)!}\int_{\Gamma_{r_{0}}^{\pm}}(a_{n}e^{\mathrm{i}n\vartheta}+b_{n}e^{-\mathrm{i}n\vartheta})r^{n+2p}e^{\rho\cdot\mathbf{x}}\mathrm{d}\sigma.$
For a given sector $\mathcal{S}_{r_{0}}$, let $\mathbf{d}=(\cos\varphi,\sin\varphi)^{\top}$ be the unit vector satisfying (2.4). Once $\mathbf{d}$ is fixed, then there are two choices of $\mathbf{d}^{\perp}$ as follows,
$\mathbf{d}^{\perp}=(-\sin\varphi,\cos\varphi)^{\top},$ (2.13)
and
$\mathbf{d}^{\perp}=(\sin\varphi,-\cos\varphi)^{\top}.$ (2.14)
Taking $\mathbf{d}^{\perp}$ in the form of (2.13), we can directly observe that the left-hand side of (2.12) satisfies:
$\int_{\Gamma_{r_{0}}^{\pm}}$ $(a_{1}e^{\mathrm{i}\vartheta}+b_{1}e^{-\mathrm{i}\vartheta})re^{\rho\cdot\mathbf{x}}\mathrm{d}\sigma=(a_{1}e^{\mathrm{i}\vartheta_{m}}+b_{1}e^{-\mathrm{i}\vartheta_{m}})\int_{0}^{r_{0}}re^{-\tau e^{\mathrm{i}(\vartheta_{m}-\varphi)r}}\mathrm{d}r$
$+(a_{1}e^{\mathrm{i}\vartheta_{M}}+b_{1}e^{-\mathrm{i}\vartheta_{M}})\int_{0}^{r_{0}}re^{-\tau e^{\mathrm{i}(\vartheta_{M}-\varphi)r}}\mathrm{d}r$
$=(a_{1}e^{\mathrm{i}\vartheta_{m}}+b_{1}e^{-\mathrm{i}\vartheta_{m}})\frac{\Gamma(2)}{\tau^{2}e^{2\mathrm{i}(\vartheta_{m}-\varphi)}}+(a_{1}e^{\mathrm{i}\vartheta_{M}}+b_{1}e^{-\mathrm{i}\vartheta_{M}})\frac{\Gamma(2)}{\tau^{2}e^{2\mathrm{i}(\vartheta_{M}-\varphi)}}$
$+\mathcal{O}\left(\frac{1}{\tau\cos(\vartheta_{m}-\varphi)}e^{-\frac{r_{0}}{2}\tau\cos(\vartheta_{m}-\varphi)}+\frac{1}{\tau\cos(\vartheta_{M}-\varphi)}e^{-\frac{r_{0}}{2}\tau\cos(\vartheta_{M}-\varphi)}\right).$ (2.15)
Without loss of generality, we assume that $\cos(\vartheta_{m}-\varphi)\leq\cos(\vartheta_{M}-\varphi)$. Hence
$\frac{1}{\tau\cos(\vartheta_{m}-\varphi)}e^{-\frac{r_{0}}{2}\tau\cos(\vartheta_{m}-\varphi)}+\frac{1}{\tau\cos(\vartheta_{M}-\varphi)}e^{-\frac{r_{0}}{2}\tau\cos(\vartheta_{M}-\varphi)}$
$=\frac{1}{\tau\cos(\vartheta_{m}-\varphi)}e^{-\frac{r_{0}}{2}\lambda\tau}\left(1+\frac{\cos(\vartheta_{m}-\varphi)}{\cos(\vartheta_{M}-\varphi)}e^{-\frac{r_{0}}{2}\tau(\cos(\vartheta_{M}-\varphi)-\cos(\vartheta_{m}-\varphi))}\right),$
where and in what follows, $\lambda=\cos(\vartheta_{m}-\varphi)\in\mathbb{R}_{+}.$ For convenience, we denote the last term in (2.15) as $\mathcal{O}(\frac{1}{\tau}e^{-\frac{r_{0}}{2}\lambda\tau})$ in the rest of the paper .
Similarly, it can be shown directly that the left-hand side of (2.12) with $\mathbf{d}^{\perp}$ defined in (2.14) fulfills the following estimates
$\int_{\Gamma_{r_{0}}^{\pm}}$ $(a_{1}e^{\mathrm{i}\vartheta}+b_{1}e^{-\mathrm{i}\vartheta})re^{\rho\cdot\mathbf{x}}\mathrm{d}\sigma=(a_{1}e^{\mathrm{i}\vartheta_{m}}+b_{1}e^{-\mathrm{i}\vartheta_{m}})\int_{0}^{r_{0}}re^{-\tau e^{\mathrm{i}(\varphi-\vartheta_{m})r}}\mathrm{d}r$
$+(a_{1}e^{\mathrm{i}\vartheta_{M}}+b_{1}e^{-\mathrm{i}\vartheta_{M}})\int_{0}^{r_{0}}re^{-\tau e^{\mathrm{i}(\varphi-\vartheta_{M})r}}\mathrm{d}r$
$=(a_{1}e^{\mathrm{i}\vartheta_{m}}+b_{1}e^{-\mathrm{i}\vartheta_{m}})\frac{\Gamma(2)}{\tau^{2}e^{2\mathrm{i}(\varphi-\vartheta_{m})}}+(a_{1}e^{\mathrm{i}\vartheta_{M}}+b_{1}e^{-\mathrm{i}\vartheta_{M}})\frac{\Gamma(2)}{\tau^{2}e^{2\mathrm{i}(\varphi-\vartheta_{M})}}$
$+\mathcal{O}(\frac{1}{\tau}e^{-\frac{r_{0}}{2}\lambda\tau}).$
Next we estimate each term $\mathcal{G}_{i}$ ($i=1,\dots,5$) on the right-hand side of (2.12) with respect to $\tau$ as $\tau\rightarrow\infty$. Using Proposition 2.1, we have
$|\mathcal{G}_{1}|$ $\leq\left|\int_{\Gamma_{r_{0}}^{\pm}}v\psi e^{\rho\cdot\mathbf{x}}\mathrm{d}\sigma\right|$
$\leq|\eta|\sum_{n=1}^{\infty}\sum_{p=0}^{\infty}\frac{(-1)^{p}(\sqrt{\gamma_{1}})^{n+2p}(|a_{n}|+|b_{n}|)}{2^{n+2p}p!(n+p)!}\int_{\Gamma_{r_{0}}^{\pm}}r^{n+2p}|\psi(\mathbf{x})e^{\rho\cdot\mathbf{x}}|\mathrm{d}\sigma$
$\lesssim|\eta|\sum_{n=1}^{\infty}\sum_{p=0}^{\infty}\frac{(-1)^{p}(\sqrt{\gamma_{1}})^{n+2p}(|a_{n}|+|b_{n}|)}{2^{n+2p}p!(n+p)!}\left(\tau^{-(n+2p+\frac{7}{6})}+\mathcal{O}(\tau^{-\frac{7}{6}}e^{-\frac{1}{2}\varsigma r_{0}\tau})\right)$
$\lesssim\tau^{-\frac{13}{6}}+\mathcal{O}(\tau^{-\frac{7}{6}}e^{-\frac{1}{2}\varsigma r_{0}\tau}),$
and
$|\mathcal{G}_{2}|$ $\leq|\gamma_{1}-\gamma_{2}|\int_{\mathcal{S}_{r_{0}}}|vu_{0}|\mathrm{d}\mathbf{x}$
$\leq|\gamma_{1}-\gamma_{2}|\sum_{n=1}^{\infty}\sum_{p=0}^{\infty}\frac{(-1)^{p}(\sqrt{\gamma_{1}})^{n+2p}(|a_{n}|+|b_{n}|)}{2^{n+2p}p!(n+p)!}\int_{\mathcal{S}_{r_{0}}}r^{n+2p}|u_{0}|\mathrm{d}\mathbf{x}$
$\lesssim\tau^{-\frac{41}{12}}+\tau^{-3}+\mathcal{O}(\tau^{-1}e^{-\frac{1}{2}\varsigma r_{0}\tau}).$
Moreover, there holds that
$|\mathcal{G}_{3}|$ $\leq\|\partial_{\nu}u_{0}\|_{L^{2}(\Lambda_{r_{0}})}\|u\|_{L^{2}(\Lambda_{r_{0}})}+\|u_{0}\|_{H^{\frac{1}{2}}(\Lambda_{r_{0}})}\|\partial_{\nu}u\|_{H^{-\frac{1}{2}}(\Lambda_{r_{0}})}.$
$\leq C(\|\partial_{\nu}u_{0}\|_{L^{2}(\Lambda_{r_{0}})}+\|u_{0}\|_{H^{1}(\Lambda_{r_{0}})})\|u\|_{H^{1}(\Lambda_{r_{0}})}$
$\leq\|u_{0}\|_{H^{1}(\Lambda_{r_{0}})}+\|\nabla u_{0}\|_{L^{2}(\Lambda_{r_{0}})}\lesssim(1+\tau)(1+\tau^{-\frac{2}{3}})e^{-\varsigma r_{0}\tau}.$
As for the last two terms of $\mathcal{G}_{4}\ \mbox{and}\ \mathcal{G}_{5}$, similarly we can deduce that
$|\mathcal{G}_{4}+\mathcal{G}_{5}|$ $\lesssim\tau^{-3}+\mathcal{O}(\tau^{-1}e^{-\frac{1}{2}\varsigma r_{0}\tau}).$ (2.16)
Combining (2.15)--(2.16) with (2.12) and letting $\tau\to\infty$, this shows that
$\begin{cases}(a_{1}e^{\mathrm{i}\vartheta_{m}}+b_{1}e^{-\mathrm{i}\vartheta_{m}})e^{2\mathrm{i}(\varphi-\vartheta_{m})}+(a_{1}e^{\mathrm{i}\vartheta_{M}}+b_{1}e^{-\mathrm{i}\vartheta_{M}})e^{2\mathrm{i}(\varphi-\vartheta_{M})}=0,\\ (a_{1}e^{\mathrm{i}\vartheta_{m}}+b_{1}e^{-\mathrm{i}\vartheta_{m}})e^{2\mathrm{i}(\vartheta_{m}-\varphi)}+(a_{1}e^{\mathrm{i}\vartheta_{M}}+b_{1}e^{-\mathrm{i}\vartheta_{M}})e^{2\mathrm{i}(\vartheta_{M}-\varphi)}=0.\end{cases}$
By multiplying $e^{-2\mathrm{i}\varphi}$ on both two sides of the first identity above and $e^{2\mathrm{i}\varphi}$ on the second, we get that
$A\left(\begin{matrix}a_{1}\\ b_{1}\end{matrix}\right)=\mathbf{0},$
where
$A=\left(\begin{matrix}e^{-\mathrm{i}\vartheta_{m}}+e^{-\mathrm{i}\vartheta_{M}}&e^{-3\mathrm{i}\vartheta_{m}}+e^{-3\mathrm{i}\vartheta_{M}}\\ e^{3\mathrm{i}\vartheta_{m}}+e^{-3\mathrm{i}\vartheta_{M}}&e^{\mathrm{i}\vartheta_{m}}+e^{\mathrm{i}\vartheta_{M}}.\end{matrix}\right)$
Since $\vartheta_{M}-\vartheta_{m}\not=\frac{\pi}{2}$ by noting $\mathcal{S}_{r_{0}}$ is irrational, after some calculations, it shows that
$\det(A)=2(\cos(\vartheta_{M}-\vartheta_{m})-\cos(3(\vartheta_{M}-\vartheta_{m})))\not=0,$
which implies that $a_{1}=b_{1}=0.$
Suppose that
$a_{j}=b_{j}=0,\quad j=1,\ldots,\ell,\quad\ell\geq 1,$ (2.17)
then we shall verify that $a_{\ell+1}=b_{\ell+1}=0$. Under the assumption (2.17), using (2.6), we have
$v$ $=\sum_{n=\ell+1}^{\infty}(a_{n}e^{\mathrm{i}n\vartheta}+b_{n}e^{-\mathrm{i}n\vartheta})J_{n}(\sqrt{\gamma_{1}}r)=\frac{(\sqrt{\gamma_{1}})^{\ell+1}(a_{\ell+1}e^{\mathrm{i}(\ell+1)\vartheta}+b_{\ell+1}e^{-\mathrm{i}(\ell+1)\vartheta})}{2^{\ell+1}(\ell+1)!}r^{\ell+1}$
$+\sum_{p=1}^{\infty}\frac{(-1)^{p}(\sqrt{\gamma_{1}})^{2p+\ell+1}(a_{\ell+1}e^{\mathrm{i}(\ell+1)\vartheta}+b_{\ell+1}e^{-\mathrm{i}(\ell+1)\vartheta})}{2^{n+2p}p!(\ell+1+p)!}r^{2p+\ell+1}$
$+\sum_{n=\ell+2}\sum_{p=0}^{\infty}\frac{(-1)^{p}(\sqrt{\gamma_{1}})^{2p+n}(a_{n}e^{\mathrm{i}n\vartheta}+b_{n}e^{-\mathrm{i}n\vartheta})}{2^{n+2p}p!(n+p)!}r^{2p+n}.$
From this, we can then derive the following integral equation
$\frac{\eta(\sqrt{\gamma_{1}})^{\ell+1}}{2^{\ell+1}(\ell+1)!}\int_{\Gamma_{r_{0}}^{\pm}}(a_{\ell+1}e^{\mathrm{i}(\ell+1)\vartheta}+b_{\ell+1}e^{-\mathrm{i}(\ell+1)\vartheta})r^{\ell+1}e^{\rho\cdot\mathbf{x}}\mathrm{d}\sigma$
$\ =-\eta\int_{\Gamma_{r_{0}}^{\pm}}v\psi(\mathbf{x})e^{\rho\cdot\mathbf{x}}\mathrm{d}\sigma+\int_{\mathcal{S}_{r_{0}}}(\gamma_{1}-\gamma_{2})vu_{0}\mathrm{d}\mathbf{x}+\int_{\Lambda_{r_{0}}}\partial vu_{0}-\partial_{\nu}uu_{0}\mathrm{d}\sigma$
$\ -\sum_{p=1}^{\infty}\frac{\eta(-1)^{p}(\sqrt{\gamma_{1}})^{2p+\ell+1}}{2^{\ell+1+2p}p!(\ell+1+p)!}\int_{\Gamma_{r_{0}}^{\pm}}(a_{\ell+1}e^{\mathrm{i}(\ell+1)\vartheta}+b_{\ell+1}e^{-\mathrm{i}(\ell+1)\vartheta})r^{2p+\ell+1}e^{\rho\cdot\mathbf{x}}\mathrm{d}\sigma$
$\ -\sum_{n=\ell+2}^{\infty}\sum_{p=0}^{\infty}\frac{\eta(-1)^{p}(\sqrt{\gamma_{1}})^{2p+n}}{2^{n+2p}p!(n+p)!}\int_{\Gamma_{r_{0}}^{\pm}}(a_{n}e^{\mathrm{i}n\vartheta}+b_{n}e^{-\mathrm{i}n\vartheta})r^{2p+n}e^{\rho\cdot\mathbf{x}}\mathrm{d}\sigma.$
Then, by an asymptotic analysis similar to that used to prove $a_{1}=b_{1}=0$, we can conclude that
$\begin{cases}(a_{\ell+1}e^{\mathrm{i}(\ell+1)\vartheta_{m}}+b_{\ell+1}e^{-\mathrm{i}(\ell+1)\vartheta_{m}})e^{-\mathrm{i}(\ell+2)(\vartheta_{m}-\varphi)}\\ \quad+(a_{\ell+1}e^{\mathrm{i}(\ell+1)\vartheta_{M}}+b_{\ell+1}e^{-\mathrm{i}(\ell+1)\vartheta_{M}})e^{-\mathrm{i}(\ell+2)(\vartheta_{M}-\varphi)}=0,\\ (a_{\ell+1}e^{\mathrm{i}(\ell+1)\vartheta_{m}}+b_{\ell+1}e^{-\mathrm{i}(\ell+1)\vartheta_{m}})e^{-\mathrm{i}(\ell+2)(\varphi-\vartheta_{m})}\\ \quad+(a_{\ell+1}e^{\mathrm{i}(\ell+1)\vartheta_{M}}+b_{\ell+1}e^{-\mathrm{i}(\ell+1)\vartheta_{M}})e^{-\mathrm{i}(\ell+2)(\varphi-\vartheta_{M})}=0,\end{cases}$
which further induces that
$B\left(\begin{matrix}a_{\ell+1}\\ b_{\ell+1}\end{matrix}\right)=\mathbf{0},$
where the coefficient matrix $B$ is
$B=\left(\begin{matrix}e^{-\mathrm{i}\vartheta_{m}}+e^{-\mathrm{i}\vartheta_{M}}&e^{-(2\ell+3)\mathrm{i}\vartheta_{m}}+e^{-(2\ell+3)\mathrm{i}\vartheta_{M}}\\ e^{(2\ell+3)\mathrm{i}\vartheta_{m}}+e^{(2\ell+3)\mathrm{i}\vartheta_{M}}&e^{\mathrm{i}\vartheta_{m}}+e^{\mathrm{i}\vartheta_{M}}\end{matrix}\right).$
Using $\mathcal{S}_{r_{0}}$ is irrational, we know that $\vartheta_{M}-\vartheta_{m}\not=\frac{\alpha}{\ell+1}\pi,$ where $\alpha=1,2,\dots,\ell$, and $\vartheta_{M}-\vartheta_{m}\not=\frac{\sigma}{\ell+2}\pi,$ where $\sigma=1,2,\dots,\ell+1.$ Then the determinant of $B$ satisfies
$\det(B)=2(\cos(\vartheta_{M}-\vartheta_{m})-\cos((2\ell+3)(\vartheta_{M}-\vartheta_{m})))\not=0,$
which implies that $a_{\ell+1}=b_{\ell+1}=0$. By mathematical induction, we deduce that $v\equiv 0$ in $B_{r_{0}}$. By unique continuation principle for elliptic PDE, we know that $v\equiv 0$ in $D$. Due to the boundary conditions in (2.7), one has $w=\partial_{\nu}w=0$ on $\Gamma_{r_{0}}$, which implies that $w\equiv 0$ in $D$ by using unique continuation principle for elliptic equation.
The proof is complete. ∎
In the following we will prove Theorem 1.2.
The proof of Theorem 1.2.
We prove this theorem using a contradiction argument. Assume that $\Omega$ has an irrational convex corner and is invisible. Since the Laplacian $\Delta$ is invariant under rigid motion, this corner of $\Omega$ can be described by $\mathcal{S}_{r_{0}}$ as given by (1.2). Since $u^{\infty}\equiv 0$, by Rellich’s theorem, one has the PDE system (1.3), where $v=u^{i}$, $w=u^{-}\chi_{\Omega}+u^{i}\chi_{D\backslash\overline{\Omega}}$, $\gamma_{1}=k^{2}$ and $\gamma_{2}=k^{2}q\chi_{\Omega}+k^{2}\chi_{D\backslash\overline{\Omega}}$. Here $D:=B_{R}$ is chosen such that $\Omega\Subset D$. By applying Theorem 1.1, we conclude that $u^{i}\equiv 0$ in $D$, leading to a contradiction by unique continuation principle. ∎
To address the inverse problem (1.10), it is essential to define the admissible class of scatterers. Definition 2.1 introduces the concept of an irrational polygon, which will be used in Definition 2.2.
Definition 2.1.
Suppose $\Omega\subset\mathbb{R}^{2}$ is a polygon. Let $\omega=\lambda\cdot\pi$, where $\lambda\in(0,2)$, be an interior angle of $\Omega$. An angle $\omega$ is considered irrational if $\lambda$ is an irrational number; otherwise, it is considered rational. Furthermore, if $\lambda\in(0,1)$ is an irrational number, then the angle $\omega$ is considered convex irrational. The polygon $\Omega$ is termed an irrational polygon if each of its interior angles is an irrational angle; otherwise, it is considered rational.
Definition 2.2.
We define $\Omega$ as an admissible complex scatterer if it comprises finitely many disjointed scatterers, denoted as follows:
$(\Omega;q,\eta)=\bigcup_{\ell=1}^{N}(\Omega_{\ell};q_{\ell},\eta_{\ell}),\ \Omega_{\ell}\cap\Omega_{m}=\emptyset\ \mbox{for}\ \ell\not=m,$
where $N\in\mathbb{N}$, and each $\Omega_{\ell}$ is a bounded Lipschtiz domain. Here, $q=\sum_{\ell=1}^{N}q_{\ell}\chi_{\Omega_{\ell}}$ with ${\rm supp}(q_{\ell})\subset\Omega_{\ell}$ and $q_{\ell}\in H^{2}(\Omega_{\ell})$. Additionally, $\eta=\sum_{\ell=1}^{N}\eta_{\ell}\chi_{\partial\Omega_{\ell}}$ and $\eta_{\ell}$ is a constant for $\ell=1,\ldots,N$.
Moreover, if each $\Omega_{\ell}$ is an irrational polygon, then $\Omega$ is considered to be an admissible complex irrational polygonal scatterer. Similarly, if each $\Omega_{\ell}$ is a convex irrational polygon, then $\Omega$ is considered to be an admissible complex convex irrational polygonal scatterer.
The following theorem provides a local uniqueness result for determining the shape of a scatterer through a single far-field measurement, offering a more rigorous version of Theorem 1.3. It asserts that the difference set between two admissible complex polygonal scatterers cannot contain a convex irrational corner if the far-field measurements associated with these two scatterers, for a fixed incident wave, are identical. Furthermore, we establish the global unique identification of the shape of admissible complex convex irrational polygonal scatterers using only a single far-field measurement.
Theorem 2.1.
Consider the conductive medium scattering described by (1.7) with the given incident wave $u^{i}$. Let $u_{\ell}^{s}\in H^{1}_{\rm loc}(\mathbb{R}^{2})$ represent the scattered wave associated with $u^{i}$ and an admissible complex polygonal scatterer $(\Omega_{\ell};q_{\ell},\eta_{\ell})$ be defined in Definition 2.2, where
$(\Omega_{\ell};q_{\ell},\eta_{\ell})=\bigcup_{i=1}^{N_{\ell}}(\Omega_{i,\ell};q_{i,\ell},\eta_{i,\ell}),\quad N_{\ell}\in\mathbb{N}{+},\quad\ell=1,2.$
Let $u^{\infty}_{\ell}$ $(\ell=1,2)$, defined in (1.8), be the far-field pattern corresponding to $u_{\ell}^{s}$, respectively.
If
$u^{\infty}_{1}(\mathbf{\hat{x}};u^{i})=u^{\infty}_{2}(\mathbf{\hat{x}};u^{i}),$ (2.18)
for all $\mathbf{\hat{x}}\in\mathbb{S}^{1}$ and a fixed incident wave $u^{i}$, then $\Omega_{1}\Delta\Omega_{2}=(\Omega_{1}\backslash\Omega_{2})\cup(\Omega_{2}\backslash\Omega_{1})$ cannot contain a convex irrational corner.
Moreover, if (2.18) is satisfied and $\Omega_{\ell}$ is an admissible complex convex irrational polygonal scatterer $(\ell=1,2)$, then $\Omega_{1}=\Omega_{2}$. In other words, $N_{1}=N_{2}$ and $\Omega_{i,1}=\Omega_{i,2}$ for $i=1,\ldots,N_{1}$.
Proof.
We prove the first conclusion of this theorem by contradiction. When (2.18) holds, assume that $\Omega_{1}\Delta\Omega_{2}$ contains an irrational corner. Underlying the fact that $\Delta$ is variant under the rigid motion and by the definition of an admissible complex irrational polygonal scatterer, we also assume that there exists an irrational corner $\mathcal{S}_{r_{0}}$ such that $\mathcal{S}_{r_{0}}\Subset\Omega_{1}\setminus\overline{\Omega_{2}}$, where $r_{0}\in\mathbb{R}_{+}$ is sufficiently small. Let $B_{r_{0}}$ be a disk with $r_{0}\in\mathbb{R}_{+}$ be chosen such that $B_{r_{0}}\Subset\Omega_{1}\setminus\overline{\Omega_{2}}$ and $\mathcal{S}_{r_{0}}\subset B_{r_{0}}$.
Let $u_{1}(\mathbf{x})$ and $u_{2}(\mathbf{x})$ be total wave fields associated $(\Omega_{1};q_{1},\eta_{1})$ and $(\Omega_{2};q_{2},\eta_{2})$, respectively. Then using (2.18), by Rellich lemma, we have $u_{1}=u_{2}$ in $\mathbb{R}^{2}\setminus\overline{\Omega_{1}\cup\Omega_{2}}$. By virtue of the trace theorem, we have
$u^{-}_{1}=u^{+}_{1}=u^{+}_{2}\ \mbox{on}\ \Gamma_{r_{0}}^{\pm}.$ (2.19)
Utilizing the conductive boundary conditions of (1.7) and (2.19), since $u_{1}$ and $u_{2}$ satisfy (1.7), we can therefore obtain that
$\begin{cases}\Delta u^{-}_{1}+k^{2}(q_{1}\chi_{\mathcal{S}_{r_{0}}}+\chi_{B_{r_{0}}\backslash\overline{\mathcal{S}_{r_{0}}}})u^{-}_{1}=0,\hskip 45.52458pt\mbox{in}\ B_{r_{0}},\\ \Delta u^{+}_{2}+k^{2}u^{+}_{2}=0,\hskip 139.41832pt\mbox{in}\ B_{r_{0}},\\ u^{-}_{1}=u^{+}_{2},\ \partial_{\nu}u^{-}_{1}=\partial_{\nu}u^{+}_{2}+\eta u^{+}_{2},\hskip 73.97733pt\mbox{on}\ \Gamma_{r_{0}}^{\pm}.\end{cases}$
It is clear that $u_{2}$ is analytic in $\mathcal{S}_{r_{0}}$. By applying Theorem 1.1, we have
$u_{2}\equiv 0\ \mbox{in}\ \mathcal{S}_{r_{0}}.$
Using the unique continuation principle and noting $u_{2}=u^{i}+u_{2}^{s}$, due to the fact that the scattered wave $u_{2}^{s}$ fulfills the Sommerfeld radiation condition, we get the contradiction.
The second conclusion can be directly obtained by using the definition of an admissible complex convex irrational polygonal scatterer. ∎
3. Unique recovery results for the conductive polygonal-nest and polygonal-cell medium scatterers
In this section, we introduce the concepts of conductive polygonal-nest and polygonal-cell medium scatterers. We then formulate the corresponding inverse problems according to (1.10) when the scatterer in the direct scattering problem (1.7) exhibits a polygonal-nest or polygonal-cell structure. To establish the uniqueness result for determining the shape of the scatterer and its polygonal-nest or polygonal-cell structure using a single far-field measurement, we introduce the concept of the admissible class and prove that the gradient of the total wave field has Hölder regularity near the corners. Additionally, we demonstrate that the reflective index, represented by a linear polynomial, and the constant conductive boundary parameter can be uniquely identified through a single far-field measurement once the shape is determined.
Definition 3.1.
Let $\Omega$ be a bounded, simply connected Lipschitz domain with a connected complement $\mathbb{R}^{2}\setminus\overline{\Omega}$. We say that $\Omega$ has a polygonal-nest structure if there exist open, convex, simply connected polygons $\Sigma_{i}$, for $i=1,\dots,N\in\mathbb{N}$, such that
$\Sigma_{N}\Subset\Sigma_{N-1}\Subset\dots\Subset\Sigma_{2}\Subset\Sigma_{1}=\Omega.$
A schematic illustration of a simply connected Lipschitz domain with a polygonal-nest structure is displayed in Figure 1(a).
Definition 3.2.
Let $\Omega$ be a bounded, simply connected Lipschitz domain with a connected complement $\mathbb{R}^{2}\setminus\overline{\Omega}$. We say that $\Omega$ has a polygonal-cell structure if there exist open, convex, simply connected polygons $\Sigma_{i}$, for $i=1,\dots,N\in\mathbb{N}$, satisfying the following conditions:
  • (a)
    $\Sigma_{i}\subset\Omega$ and $\Sigma_{i}\cap\Sigma_{j}=\emptyset$ for $i\neq j$,
  • (b)
    $\cup_{i=1}^{N}\overline{\Sigma_{i}}=\overline{\Omega}$,
  • (c)
    For each $\Sigma_{i}$, there exists at least one vertex $\mathbf{x}_{p,i}$ such that the two edges of $\partial\Sigma_{i}$ associated with $\mathbf{x}_{p,i}$ are denoted by $\Gamma_{i}^{\pm}\subset\partial\Omega$.
A schematic illustration of a simply connected Lipschitz domain with a polygonal-cell structure is displayed in Figure 1(b).
(a) A polygonal-nest structure.
(b) A polygonal-cell structure.
Figure 1. The schematic illustrations of the polygonal-nest and the polygonal-cell structures.
Definition 3.3.
Let $(\Omega;q,\eta)$ be a conductive medium scatterer, if it satisfies the following conditions:
  • (a)
    $\Omega$ has a polygonal-nest structure defined in Definition 3.1;
  • (b)
    each $\Sigma_{i}$ is a medium scatterer such that $\mathcal{D}_{i}:=\Sigma_{i}\setminus\overline{\Sigma_{i+1}}$ with physical configurations $q_{i}$ and $\eta_{i}$ (on $\partial\Sigma_{i}$), which is denoted as $(\mathcal{D}_{i};q_{i},\eta_{i})$ with $\Sigma_{N+1}=\emptyset$, where $q_{i}$ is a polynomial of order $1$ with the form given by
    $q_{i}=q_{i}^{(0)}+q_{i}^{(1)}x_{1}+q_{i}^{(2)}x_{2},\quad q_{i}^{(j)}\in\mathbb{C},\quad(j=0,1,2)$ (3.1)
    in $\mathcal{D}_{i}$, and $\eta_{i}$ is a constant on $\partial\Sigma_{i}$ for $i=1,\dots,N$,
then we say that $(\Omega;q,\eta)$ possesses a polygonal-nest structure.
Throughout of the rest paper, we rewrite a conductive polygonal-nest medium $(\Omega;q,\eta)$ given by Definition 3.3 as
$(\Omega;q,\eta)=\bigcup_{i=1}^{N}(\mathcal{D}_{i};q_{i},\eta_{i})$ (3.2)
and $\Omega=\bigcup_{i=1}^{N}\mathcal{D}_{i},\ q=\sum_{i=1}^{N}q_{i}\chi_{\mathcal{D}_{i}},\eta=\sum_{i=1}^{N}\eta_{i}\chi_{\partial\Sigma_{i}}$.
Considering the direct scattering problem (1.7) associated with a conductive polygonal-nest medium scatterer $(\Omega;q,\eta)$ described by (3.2), let $u\in H^{1}_{loc}(\mathbb{R}^{2})$ be the corresponding total wave field. In view of the polygonal-nest structure of $(\Omega;q,\eta)$, the conductive transmission boundary conditions of the total wave $u$ across $\partial\Sigma_{i+1}$ are given by
$u_{i}|_{\partial\Sigma_{i+1}}=u_{i+1}|_{\partial\Sigma_{i+1}},\ (\partial_{\nu}u_{i}+\eta_{i+1}u_{i})|_{\partial\Sigma_{i+1}}=\partial_{\nu}u_{i+1}|_{\partial\Sigma_{i+1}},$ (3.3)
where $u_{i}=u|_{\mathcal{D}_{i}},i=0,1,\dots,N-1$ and $\Sigma_{0}=\mathbb{R}^{2}\setminus\overline{\Sigma_{1}}$. Indeed, the well-posedness of the direct scattering problem can be obtained by using a similar variational argument in [11], where the detailed proof is omitted. In this paper, our main focus is the corresponding inverse problems associated with a conductive polygonal-nest medium scatterer $(\Omega;q,\eta)$, which aims to determine the polygonal-nest structure of $\Omega$ and the physical configurations $q_{\ell}$ and $\eta_{\ell}$ from the knowledge of the far-field pattern $u^{\infty}(\mathbf{\hat{x}};u^{i})$ with a fixed incident wave $u^{i}$. Here, $u^{\infty}(\mathbf{\hat{x}};u^{i})$ is the far-field pattern of $u$ given by (1.8). The above inverse problems can be formulated by
$u^{\infty}(\mathbf{\hat{x}};u^{i})\rightarrow\bigcup_{\ell=1}^{N}(\mathcal{D}_{\ell};q_{\ell},\eta_{\ell}),\ \forall\mathbf{\hat{x}}\in\mathbb{S}^{1}.$ (3.4)
Similarly, we can define a conductive polygonal-cell medium scatterer $\Omega$ as follows.
Definition 3.4.
Let $(\Omega;q,\eta)$ be a conductive medium scatterer, if the following conditions are fulfilled:
  • (a)
    $\Omega$ has a polygonal-cell structure defined in Definition 3.2;
  • (b)
    each $\Sigma_{i}$ is a medium scatterer with physical configurations $q_{i}$ and $\tilde{\eta}$ and is denoted as $(\Sigma_{i};q_{i},\tilde{\eta})$, $q_{i}$ is a polynomial of order $1$ in the form of (3.1) in $\Sigma_{i}$ and $\tilde{\eta}$ is a constant on $\partial\Sigma_{i}$ for $i=1,\dots,N$,
then $(\Omega;q,\eta)$ is said to have a polygonal-cell structure.
We write a conductive polygonal-cell medium scatterer $(\Omega;q,\eta)$ as
$(\Omega;q,\eta)=\bigcup_{i=1}^{N}(\Sigma_{i};q_{i},\tilde{\eta}),$ (3.5)
where $\Omega=\cup_{i=1}^{N}\Sigma_{i},\ q=\sum_{i=1}^{N}q_{i}\chi_{\Sigma_{i}},\ \eta=\sum_{i=1}^{N}\tilde{\eta}\chi_{\partial\Sigma_{i}}.$ The inverse problems for a conductive polygonal-cell medium scatterer by a single far-field measurement can be described in a similar way as for the case with a conductive polygonal-nest medium scatterer. Namely, one aims to determine the shape of $\Omega$ and physical configurations $q_{i}$ and $\tilde{\eta}$ through the measurement $u^{\infty}(\mathbf{\hat{x}};u^{i})$ associated with a fixed incident wave $u^{i}$, which can be formulated as follows:
$u^{\infty}(\mathbf{\hat{x}};u^{i})\rightarrow\bigcup_{i=1}^{N}(\Sigma_{i};q_{i},\tilde{\eta}),\ \forall\mathbf{\hat{x}}\in\mathbb{S}^{1}.$ (3.6)
Here, $u^{\infty}(\mathbf{\hat{x}};u^{i})$ is the far-field pattern given by (1.8), associated with the total wave $u$, where $u$ is the solution of (1.7) corresponding to the conductive polygonal-cell medium scatterer $\bigcup_{i=1}^{N}(\Sigma_{i};q_{i},\tilde{\eta})$.
In order to introduce the admissible class for studying the inverse problems, we need to study the Hölder-regularity of the total wave field $u$ which is associated with the scattering problem (1.7) near the corner $\mathcal{S}_{r_{0}}$. Then we have the following lemma, which shows that there holds that $u^{-}\in C^{1,\alpha}(\overline{\mathcal{S}_{r_{0}}})$, where $\alpha\in(0,1)$.
Lemma 3.1.
Let the corner $\mathcal{S}_{2r_{0}}=\mathcal{S}\cap B_{2r_{0}}$, where $\mathcal{S}$ is given by (1.1) and $B_{2r_{0}}$ is a disk centered at the origin with the radius $2r_{0}$ ($r_{0}\in\mathbb{R}_{+}$). Denote $\Gamma^{\pm}_{2r_{0}}=\Gamma^{\pm}\cap B_{2r_{0}}$, where $\Gamma^{\pm}$ is the boundary line of $\mathcal{S}$. Suppose that $u\in H^{1}(B_{2r_{0}})$ with $\ u^{+}=u|_{B_{2r_{0}}\setminus\overline{\mathcal{S}_{2r_{0}}}},\ u^{-}=u|_{\mathcal{S}_{2r_{0}}}$ satisfies that
[…]
where $k$ is a positive constant and $q$ is in the form of (3.1). Assume that $u^{-}$ and $u^{+}$ are real analytic in $\mathcal{S}_{2r_{0}}$ and $B_{2r_{0}}\setminus\overline{\mathcal{S}_{2r_{0}}}$, respectively. Then there exist $\alpha\in(0,1)$ such that $u^{-}\in C^{1,\alpha}(\overline{\mathcal{S}_{r_{0}}}).$
Before proving Lemma 3.1, we first introduce a triangle $V_{2r_{0}}$ described below, which will be used in a subsequent argument. Let us pick up the points $\mathbf{y}_{+}$ and $\mathbf{y}_{-}$ on $\Gamma_{2r_{0}}^{+}$ and $\Gamma_{2r_{0}}^{-}$, respectively, where $|\mathbf{y}_{\pm}|\in(r_{0},2r_{0})$. Denote $\Lambda_{2r_{0}}^{{}^{\prime}}=\mathbf{y}_{+}-\mathbf{y}_{-}$. Then the triangle domain $V_{2r_{0}}$ is formed by three vectors $\mathbf{y}_{+}$, $\mathbf{y}_{-}$ and $\Lambda_{2r_{0}}^{{}^{\prime}}$. It is clear that $\overline{\mathcal{S}_{r_{0}}}\subset\overline{V_{2r_{0}}}\subset\overline{\mathcal{S}_{2r_{0}}}$. We can refer to Fig. 2 for a geometric illustration of $V_{2r_{0}}$.
Figure 2. A schematic illustration of $\mathcal{S}_{2r_{0}}$, $V_{2r_{0}}$ and $\mathcal{S}_{r_{0}}$.
The proof of Lemma 3.1.
Denote $\overline{u}$ by the analytic extension of $u^{+}|_{B_{2r_{0}}\setminus\overline{\mathcal{S}}_{2r_{0}}}$ in $B_{2r_{0}}$ since it is real analytic in $B_{2r_{0}}\setminus\overline{\mathcal{S}}_{2r_{0}}$. By virtue of the boundary condition of (3.7), one has $\overline{u}=u^{-}=u^{+}\ \mbox{on}\ \Gamma_{2r_{0}}^{\pm}.$
Let $w=u^{-}-\overline{u},$ then from (3.7) it can be directly calculated that
$\begin{cases}\Delta w=f,\hskip 15.6491pt\mbox{in}\ V_{2r_{0}},\\ w=\varphi,\hskip 22.76228pt\mbox{on}\ \Lambda^{{}^{\prime}}_{2r_{0}},\\ w=0,\hskip 24.18501pt\mbox{on}\ \Gamma_{2r_{0}}^{\pm},\end{cases}$
[…]
$w({\bf x})=w_{\rm reg}({\bf x})+\sum_{0<\lambda_{\ell}<3-\frac{2}{p}}\zeta^{\ell}r^{\lambda_{\ell}}\sin(\lambda_{\ell}\vartheta),\ w_{\rm reg}\in W^{3,p}(V_{2r_{0}}),\ {\bf x}=r(\cos\vartheta,\sin\vartheta),$
where $\zeta$ is a constant depending only on $f$ and $\varphi$, and $\lambda_{\ell}=\frac{\ell\pi}{\vartheta_{M}-\vartheta_{m}},\ \ell\in\mathbb{N}$. If $\vartheta_{M}-\vartheta_{m}\in(0,\pi/2]$, then we know that $w({\bf x})=w_{\rm reg}({\bf x})$. If $\vartheta_{M}-\vartheta_{m}\in(\pi/2,\pi)$, it can be verified that $\frac{2}{3-\frac{\pi}{\vartheta_{M}-\vartheta_{m}}}\in(1,2)$. Hence we can choose $p\in(1,\frac{2}{3-\frac{\pi}{\vartheta_{M}-\vartheta_{m}}})$. This indicates that $\lambda_{\ell}>3-\frac{2}{p}$ for any $\ell\in\mathbb{N}$. Therefore, $w=w_{\rm reg}\in W^{3,p}(V_{2r_{0}})$. By virtue of the Sobolev embedding theorem (cf. [28, Theorem 7.26]), we have that $w=w_{\rm reg}\in C^{1,\alpha}(\overline{V_{2r_{0}}})$ for $\alpha\in(0,1)$.
The proof is complete. ∎
In the following, we will introduce the admissible class for studying the inverse problems (3.4) and (3.6), which will be used to establish recovery uniqueness results in the subsequent theorems.
Definition 3.5.
Suppose $(\Omega;q,\eta)$ represents a conductive polygonal-nest or polygonal-cell medium scatterer as defined in Definitions 3.1 or 3.2. Let $u$ denote the total wave field of (1.7) associated with $(\Omega;q,\eta)$.
  • Polygonal-Nest Admissibility:
    1. (1)
      For a conductive polygonal-nest medium scatterer $(\Omega;q,\eta)$, it is considered admissible if it satisfies one of the following conditions:
      • (I)
        $u(\mathbf{x}_{p})\not=0,$ (3.8)
      • (II)
        $\nabla u(\mathbf{x}_{p})\neq 0,\ u(\mathbf{x}_{p})=0,$ (3.9)
      where $\mathbf{x}_{p}\in{\mathcal{V}}(\Sigma_{i})$ and ${\mathcal{V}}(\Sigma_{i})=\{\mathbf{x}_{p,i}\}_{p=1}^{\ell_{i}}(3\leq\ell_{i}\in\mathbb{Z}_{+})$ is the vertex set of the convex polygon $\Sigma_{i}$. Here, $\{\Sigma_{i}\}_{i=1}^{N}$ is the polygonal-nest structure of $\Omega$ described by Definition 3.1.
    2. (2)
      If the condition (3.9) is satisfied, it is additionally required that $\angle(\Gamma_{p}^{-},\Gamma_{p}^{+})\in(\frac{\pi}{3},\frac{2\pi}{3})\setminus\{\frac{\pi}{2}\}$, where $\Gamma_{p}^{\pm}$ represent two edges of $\Sigma_{i}$, and $\Gamma_{p}^{-}\cap\Gamma_{p}^{+}=\mathbf{x}_{p}$ with $\angle(\Gamma_{p}^{-},\Gamma_{p}^{+})$ being the intersection angle between $\Gamma_{p}^{-}$ and $\Gamma_{p}^{+}$.
  • Polygonal-Cell Admissibility:
    1. (1)
      For a conductive polygonal-cell medium scatterer $(\Omega;q,\eta)$, it is considered admissible if it satisfies condition (3.8) or (3.9), where $\mathbf{x}_{p}\in{\mathcal{V}}(\Omega)$. Here, ${\mathcal{V}}(\Omega)$ represents the vertex set of the polygon $\Omega$.
    2. (2)
      Similarly, if the condition (3.9) is satisfied, it is additionally required that $\angle(\Gamma_{p}^{-},\Gamma_{p}^{+})\in(\frac{\pi}{3},\frac{2\pi}{3})\setminus\{\frac{\pi}{2}\}$, where $\Gamma_{p}^{\pm}$ represent two edges of $\Omega$, and $\Gamma_{p}^{-}\cap\Gamma_{p}^{+}=\mathbf{x}_{p}$ with $\angle(\Gamma_{p}^{-},\Gamma_{p}^{+})$ being the intersection angle between $\Gamma_{p}^{-}$ and $\Gamma_{p}^{+}$.
Remark 3.1.
The non-void assumptions (3.8) and (3.9) in Definition 3.5 represent generic physical conditions. Indeed, the generic assumptions (3.8) and (3.9) can be fulfilled in certain physical scenarios. For example, when
$k\cdot{\rm diam}(\Omega)\ll 1,$ (3.10)
from a physical perspective, the scattered wave $u^{s}$ can be neglected. In fact, as rigorously justified in [11], it holds that
$\|u^{s}\|_{H^{1}(B_{R})}\leq C(\|\eta u^{i}\|_{H^{-\frac{1}{2}}(\partial\Omega)}+k^{2}\|qu^{i}\|_{L^{2}(\Omega)}),\quad\Omega\subset B_{R},$
where $C$ is a positive number and $B_{R}$ is a ball centered at the origin with radius $R$. Hence, if (3.10) is satisfied and the conductive boundary parameter $\eta$ is sufficiently small, the plane incident wave $u^{i}$ dominates the total wave $u=u^{i}+u^{s}$. It is evident that $u^{i}$ fulfills (3.8) (or (3.9)) , which implies that the total wave $u$ generically satisfies (3.8)or (3.9). On the other hand, from a physical standpoint, in conductive medium scattering, if the total wave does not satisfies (3.8) at a corner point, it implies that the scattered wave and the incident wave cancel each other at those points. Similarly, if the admissible condition (3.9) is violated, it indicates that either the scattered wave and the incident wave cancel each other, or the gradient of the scattered wave and the incident wave cancel each other at those points. Such cancellation phenomena, as described above, are rare in typical scattering scenarios. However, we will not delve deeply into this matter in this paper.
Remark 3.2.
In previous works such as [21, 15], the admissible condition (3.8) plays a crucial role in deriving corresponding uniqueness results. However, in this paper, when (3.8) is violated, we propose another admissible condition, denoted as (3.9). There are three distinct cases for (3.9), characterized as follows:
  • a)
    $\partial_{1}u(\mathbf{x}_{p})\neq 0$, $u(\mathbf{x}_{p})=\partial_{2}u(\mathbf{x}_{p})=0$,
  • b)
    $\partial_{2}u(\mathbf{x}_{p})\neq 0$, $u(\mathbf{x}_{p})=\partial_{1}u(\mathbf{x}_{p})=0$,
  • c)
    $\partial_{1}u(\mathbf{x}_{p})\neq 0$, $\partial_{2}u(\mathbf{x}_{p})\neq 0$, $u(\mathbf{x}_{p})=0$.
Since (3.9) encompasses more cases compared to (3.8), it is more likely to be satisfied in practical scenarios. This increased applicability of (3.9) enhances the practical relevance of our uniqueness results.
In Theorem 3.1, a local uniqueness result for determining the shape of an admissible conductive polygonal-nest or polygonal-cell medium scatterer by a single far-field measurement is established. Furthermore, when the admissible scatterer has a polygonal-nest structure, a global shape identifiability by a single far-field measurement is obtained.
Theorem 3.1.
Let $(\Omega_{\ell};q_{\ell},\eta_{\ell}),\ell=1,2$ be two admissible conductive polygonal-nest or polygonal-cell medium scatterer satisfying (1.7), $u^{\infty}_{\ell}(\mathbf{\hat{x}};u^{i})$ be the far-field pattern associated with the scatterer $(\Omega_{\ell};q_{\ell},\eta_{\ell})$ and the incident field $u^{i}$. If
$u^{\infty}_{1}(\mathbf{\hat{x}};u^{i})=u^{\infty}_{2}(\mathbf{\hat{x}};u^{i})$
is fulfilled for all $\mathbf{\hat{x}}\in\mathbb{S}^{1}$ and a fixed incident wave $u^{i}$, then $\Omega_{1}\Delta\Omega_{2}$ cannot have a convex corner. Moreover, if $\Omega_{1}$ and $\Omega_{2}$ are two admissible conductive polygonal-nest medium scatterers, one holds that
$\partial\Omega_{1}=\partial\Omega_{2}.$
Proof.
We only present the proof for the polygonal-nest case, since the polygonal-cell case can be proven in a similar manner. We proceed to prove this theorem by contradiction. Let us assume that there exists a corner in $\Omega_{1}\Delta\Omega_{2}$. According to (3.2), we can express the admissible conductive polygonal-nest medium scatterers as $(\Omega_{\ell};q_{\ell},\eta_{\ell})$ ($\ell=1,2$) as follows
$(\Omega_{\ell};q_{\ell},\eta_{\ell})=\bigcup_{i=1}^{N_{\ell}}(\mathcal{D}_{i,\ell};q_{i,\ell},\eta_{i,\ell}),$
and
$\Omega_{\ell}=\bigcup_{i=1}^{N_{\ell}}\mathcal{D}_{i,\ell},\ q_{j}=\sum_{i=1}^{N_{\ell}}q_{i,\ell}\chi_{\mathcal{D}_{i,\ell}},\ \eta_{\ell}=\sum_{i=1}^{N_{\ell}}\eta_{i,\ell}\chi_{\partial\Sigma_{i,\ell}}.$ (3.11)
Since $-\Delta$ is invariant under rigid motion, without loss of generality, let us assume that $\bf 0$ is the vertex of the corner $\Omega_{1}\cap\mathcal{S}$ such that $\mathbf{0}\in\partial\Omega_{1}$ but $\mathbf{0}\notin\partial\overline{\Omega_{2}}$, where $\mathbf{0}$ represents the origin. Moreover, we assume that the corner $\Omega_{1}\cap\mathcal{S}=\mathcal{S}_{r_{0}}\Subset\Sigma_{1,1}$, where $\Sigma_{1,1}$ is defined in Definition 3.1 and (3.11).
Let $u_{1}(\mathbf{x})$ and $u_{2}(\mathbf{x})$ be the total wave fields associated with $\Omega_{1}$ and $\Omega_{2}$, respectively. Given $u^{\infty}_{1}=u_{2}^{\infty}$ for all $\mathbf{\hat{x}}\in\mathbb{S}^{1}$ and by Rellich lemma, we conclude that $u^{s}_{1}=u^{s}_{2}$ in $\mathbb{R}^{2}\setminus(\overline{\Omega_{1}\cup\Omega_{2}})$. This implies that $u_{1}(\mathbf{x})=u_{2}(\mathbf{x})$ for all $\mathbf{\hat{x}}\in\mathbb{R}^{2}\setminus(\overline{\Omega_{1}\cup\Omega_{2}})$. Then $u_{1}^{-}$ and $u_{2}^{+}$ satisfy the following PDE system,
$\begin{cases}\Delta u_{1}^{-}+k^{2}q_{1}u_{1}^{-}=0,\hskip 73.97733pt\mbox{in}\ \mathcal{S}_{r_{0}},\\ \Delta u_{2}^{+}+k^{2}u_{2}^{+}=0,\hskip 83.93553pt\mbox{in}\ \mathcal{S}_{r_{0}},\\ u_{1}^{-}=u_{2}^{+},\ \partial_{\nu}u_{1}^{-}=\partial_{\nu}u_{2}^{+}+\eta_{1}u_{2}^{+},\hskip 14.22636pt\mbox{on}\ \Gamma_{r_{0}}^{\pm}.\end{cases}$
Let $\bar{u}=u_{1}^{-}-u_{2}^{+}$. Then it yields that
[…]
Since $q_{1}\in H^{2}(\mathcal{S}_{r_{0}})$, let $\widetilde{q}_{1}$ be the Sobolev extension of $q_{1}$ in $\mathbb{R}^{2}$. According to Lemma 2.1, there exists the CGO solution $u_{0}$ given by (2.1), which satisfies (2.2). Hence it yields that
$(\Delta+k^{2}q_{1})u_{0}=0\mbox{ in }\mathcal{S}_{r_{0}}.$
By virtue of the CGO solution defined in (2.1), using Green’s formula, one has
$\eta_{1}\int_{\Gamma_{r_{0}}^{\pm}}u_{2}^{+}u_{0}\mathrm{d}\mathbf{x}=k^{2}\int_{\mathcal{S}_{r_{0}}}(1-q_{1})u_{2}^{+}u_{0}\mathrm{d}\mathbf{x}+\int_{\Lambda_{r_{0}}}\partial_{\nu}u_{0}\bar{u}-\partial_{\nu}\bar{u}u_{0}\mathrm{d}\sigma.$ (3.12)
According to Lemma 3.1, we know that $u_{2}^{+}\in C^{1,\alpha}(\overline{\mathcal{S}_{r_{0}}}),\alpha\in(0,1)$, which implies the following expansion
$u_{2}^{+}=u_{2}^{+}(\mathbf{0})+\partial_{1}u_{2}^{+}(\mathbf{0})x_{1}+\partial_{2}u_{2}^{+}(\mathbf{0})x_{2}+\delta u_{2}^{+},\ |\delta u_{2}^{+}|\leq|\mathbf{x}|^{1+\alpha}\|u_{2}^{+}\|_{C^{1,\alpha}}.$ (3.13)
According to the admissibility conditions in Definition 3.5, we will divide the proof into four cases, as follows.
Case (a): $u_{2}^{+}(\mathbf{0})\not=0$. Combing (3.13) with (3.12), we can deduce that
$\eta_{1}u_{2}^{+}(\mathbf{0})$ $\int_{\Gamma_{r_{0}}^{\pm}}e^{\rho\cdot\mathbf{x}}\mathrm{d}\mathbf{x}=\sum_{i=1}^{4}I_{i},$ (3.14)
where
$I_{1}$ $=-\eta_{1}u_{2}^{+}(\mathbf{0})\int_{\Gamma_{r_{0}}^{\pm}}\psi(\mathbf{x})e^{\rho\cdot\mathbf{x}}\mathrm{d}\sigma,\ I_{2}=k^{2}\int_{\mathcal{S}_{r_{0}}}(1-q_{1})u_{2}^{+}u_{0}\mathrm{d}\mathbf{x}$
$I_{3}$ $=-\eta_{1}\int_{\Gamma_{r_{0}}^{\pm}}(\partial_{1}u_{2}^{+}(\mathbf{0})x_{1}+\partial_{2}u_{2}^{+}(\mathbf{0})x_{2}+\delta u_{2}^{+})u_{0}\mathrm{d}\sigma,$
$I_{4}$ $=\int_{\Lambda_{r_{0}}}\partial_{\nu}u_{0}\bar{u}-\partial_{\nu}\bar{u}u_{0}\mathrm{d}\sigma.$ (3.15)
By (2.5), we have
$|I_{1}|$ $\lesssim\tau^{-\frac{7}{6}}+\mathcal{O}(\tau^{-\frac{7}{6}}e^{-\frac{1}{2}\varsigma r_{0}\tau}),\ |I_{2}|\lesssim\tau^{-2}+\tau^{-\frac{29}{12}}+\mathcal{O}(\tau^{-1}e^{-\frac{1}{2}\varsigma\tau}),$
$|I_{3}|$ $\lesssim\tau^{-2}+\tau^{-(2+\alpha)}+\mathcal{O}(\tau^{-1}e^{-\frac{1}{2}\varsigma r_{0}\tau}),\ |I_{4}|\lesssim(1+\tau)(1+\tau^{-\frac{2}{3}})e^{-\varsigma r_{0}\tau}.$ (3.16)
Let $\mathbf{d}$ be given in the form of (2.14). We can then directly derive that
$|\eta_{1}u_{2}^{+}(\mathbf{0})|$ $\left|\int_{\Gamma_{r_{0}}^{\pm}}e^{\rho\cdot\mathbf{x}}\mathrm{d}\sigma\right|\gtrsim\frac{\Gamma(1)|\eta_{1}u_{2}^{+}(\mathbf{0})||e^{\mathrm{i}\vartheta_{m}}+e^{\mathrm{i}\vartheta_{M}}|}{|\tau e^{\mathrm{i}\varphi}|}-\mathcal{O}(\tau^{-1}e^{-\frac{1}{2}\lambda r_{0}\tau}),$ (3.17)
here and in what follows, the term $\mathcal{O}(\tau^{-1}e^{-\frac{1}{2}\lambda r_{0}\tau})$ is defined in (2.15). Recall that $\mathcal{S}_{r_{0}}$ is defined by (1.2), where the polar angle of $\Gamma_{r_{0}}^{-}$ is $\vartheta_{m}$ and the polar angle of $\Gamma_{r_{0}}^{+}$ is $\vartheta_{M}$. Combining (3.15)- (3.17) with (3.14), then multiplying $\tau$ and letting $\tau\to\infty$, we see that
$|\eta_{1}u_{2}^{+}(\mathbf{0})||e^{\mathrm{i}\vartheta_{m}}+e^{\mathrm{i}\vartheta_{M}}|=0.$
Since
$|e^{\mathrm{i}\vartheta_{m}}+e^{\mathrm{i}\vartheta_{M}}|=\left(2(\cos(\vartheta_{M}-\vartheta_{m})+1)\right)^{1/2}\not=0\quad\mbox{for}\quad\vartheta_{M}-\vartheta_{m}\in(0,\pi),$
it follows that
$|\eta_{1}u_{2}^{+}(\mathbf{0})|=0.$
However, this contradicts to the admissible condition $u_{2}^{+}(\mathbf{0})\neq 0$ due to $\eta_{1}\neq 0$.
Case (b): $\partial_{1}u_{2}^{+}(\mathbf{0})\not=0$, $u_{2}^{+}(\mathbf{0})=\partial_{2}u_{2}^{+}(\mathbf{0})=0$. Similar to (3.14), we have the following integral equality:
$\eta_{1}\partial_{1}u_{2}^{+}(\mathbf{0})$ $\int_{\Gamma_{r_{0}}^{\pm}}e^{\rho\cdot\mathbf{x}}x_{1}\mathrm{d}\sigma=-\eta_{1}\partial_{1}u_{2}^{+}(\mathbf{0})\int_{\Gamma_{r_{0}}^{\pm}}\psi(\mathbf{x})e^{\rho\cdot\mathbf{x}}x_{1}\mathrm{d}\sigma$
$+\int_{\mathcal{S}_{r_{0}}}(1-q_{1})u_{2}^{+}u_{0}\mathrm{d}\mathbf{x}-\eta_{1}\int_{\Gamma_{r_{0}}^{\pm}}\delta u_{2}^{+}u_{0}\mathrm{d}\sigma+I_{4}.$ (3.18)
By direct calculations we know that the right-hand side of (3.18) satisfies
$|RHS|\lesssim\tau^{-\frac{13}{6}}+\tau^{-(2+\alpha)}+(1+\tau)(1+\tau^{-\frac{2}{3}})e^{-\varsigma r_{0}\tau}+\mathcal{O}(\tau^{-1}e^{-\frac{1}{2}\varsigma r_{0}\tau}).$ (3.19)
Furthermore, we have the following inequality about the left-hand side of (3.18)
$|LHS|\gtrsim\frac{\Gamma(2)|\eta_{1}\partial_{1}u_{2}^{+}(\mathbf{0})|}{\tau^{2}}\left|\frac{\cos\vartheta_{M}}{((\mathbf{d}+\mathrm{i}\mathbf{d})\cdot\mathbf{\hat{x}}_{1})^{2}}+\frac{\cos\vartheta_{m}}{((\mathbf{d}+\mathrm{i}\mathbf{d})\cdot\mathbf{\hat{x}}_{2})^{2}}\right|-\mathcal{O}(\tau^{-1}e^{-\lambda r_{0}\tau}),$ (3.20)
where $\mathbf{\hat{x}}_{1}\in\Gamma_{r_{0}}^{+}$ and $\mathbf{\hat{x}}_{2}\in\Gamma_{r_{0}}^{-}$. Let $\mathbf{d}^{\perp}$ be in the form of (2.14), then we have
$\left|\frac{\cos\vartheta_{M}}{((\mathbf{d}+\mathrm{i}\mathbf{d})\cdot\mathbf{\hat{x}}_{1})^{2}}+\frac{\cos\vartheta_{m}}{((\mathbf{d}+\mathrm{i}\mathbf{d})\cdot\mathbf{\hat{x}}_{2})^{2}}\right|=|\cos\vartheta_{M}e^{-2\mathrm{i}\vartheta_{m}}$ $+\cos\vartheta_{m}e^{-2\mathrm{i}\vartheta_{M}}|\not=0$
for $-\pi<\vartheta_{m}<\vartheta_{M}<\pi,\ \vartheta_{M}-\vartheta_{m}\in(0,\pi)$. Substituting (3.19) and (3.20) into (3.18), then multiplying $\tau^{2}$ on both sides and taking $\tau\to\infty$, we get
$|\eta_{1}\partial_{1}u_{2}^{+}(\mathbf{0})|=0,$
which implies that $\partial_{1}u_{2}^{+}(\mathbf{0})=0$ by noting $\eta_{1}\not=0$. Hence, we obtain a contradiction.
Case (c): $\partial_{2}u_{2}^{+}(\mathbf{0})\not=0,$ $u_{2}^{+}(\mathbf{0})=\partial_{1}u_{2}^{+}(\mathbf{0})=0$. Similar to (3.14), it holds that
$\eta_{1}\partial_{2}u_{2}^{+}(\mathbf{0})$ $\int_{\Gamma_{r_{0}}^{\pm}}e^{\rho\cdot\mathbf{x}}x_{2}\mathrm{d}\sigma=-\eta_{1}\partial_{2}u_{2}^{+}(\mathbf{0})\int_{\Gamma_{r_{0}}^{\pm}}\psi(\mathbf{x})e^{\rho\cdot\mathbf{x}}x_{2}\mathrm{d}\sigma$
$+\int_{\mathcal{S}_{r_{0}}}(1-q_{1})u_{2}^{+}u_{0}\mathrm{d}\mathbf{x}-\eta_{1}\int_{\Gamma_{r_{0}}^{\pm}}\delta u_{2}^{+}u_{0}\mathrm{d}\sigma+I_{4}.$ (3.21)
Similar to (3.18), for the right-hand and left-hand sides of (3.21), we have
$|RHS|\lesssim\tau^{-\frac{13}{6}}+\tau^{-(2+\alpha)}+(1+\tau)(1+\tau^{-\frac{2}{3}})e^{-\varsigma r_{0}\tau}+\mathcal{O}(\tau^{-1}e^{-\frac{1}{2}\varsigma r_{0}\tau})$
and
$|LHS|\gtrsim\frac{\Gamma(2)|\eta_{1}\partial_{2}u_{2}^{+}(\mathbf{0})|}{\tau^{2}}\left|\frac{\sin\vartheta_{M}}{((\mathbf{d}+\mathrm{i}\mathbf{d}^{\perp})\cdot\mathbf{\hat{x}}_{1})^{2}}+\frac{\sin\vartheta_{m}}{((\mathbf{d}+\mathrm{i}\mathbf{d}^{\perp})\cdot\mathbf{\hat{x}}_{2})^{2}}\right|-\mathcal{O}(\tau^{-1}e^{-\lambda r_{0}\tau})$ (3.22)
where $\mathbf{\hat{x}}_{1}\in\Gamma_{r_{0}}^{+},\ \mathbf{\hat{x}}_{2}\in\Gamma_{r_{0}}^{-}$. Let $\mathbf{d}^{\perp}$ be defined as in (2.14), then it follows that:
$\left|\frac{\sin\vartheta_{M}}{((\mathbf{d}+\mathrm{i}\mathbf{d}^{\perp})\cdot\mathbf{\hat{x}}_{1})^{2}}+\frac{\sin\vartheta_{m}}{((\mathbf{d}+\mathrm{i}\mathbf{d}^{\perp})\cdot\mathbf{\hat{x}}_{2})^{2}}\right|=|\sin\vartheta_{M}e^{-2\mathrm{i}\vartheta_{m}}+\sin\vartheta_{m}e^{-2\mathrm{i}\vartheta_{M}}|\not=0,$
for $-\pi<\vartheta_{m}<\vartheta_{M}<\pi,\ \vartheta_{M}-\vartheta_{m}\in(0,\pi)$ and $\vartheta_{M}-\vartheta_{m}\neq\frac{\pi}{2}.$ Combining (3.21)-(3.22)and taking $\tau\to\infty$, one has
$|\eta_{1}\partial_{2}u_{2}^{+}(\mathbf{0})|=0,$
Hence, it yields that $\partial_{2}u_{2}^{+}(\mathbf{0})=0$ since $\eta_{1}\neq 0$, which contradicts to $\partial_{2}u_{2}^{+}(\mathbf{0})\not=0$.
Case (d): $\partial_{1}u_{2}^{+}(\mathbf{0})\not=0$,$\partial_{2}u_{2}^{+}(\mathbf{0})\not=0$, $u_{2}^{+}(\mathbf{0})=0$. Similar to (3.14), we have
$\eta_{1}\int_{\Gamma_{r_{0}}^{\pm}}$ $(\partial_{1}u_{2}^{+}(\mathbf{0})x_{1}+\partial_{2}u_{2}^{+}(\mathbf{0})x_{2})e^{\rho\cdot\mathbf{x}}\mathrm{d}\sigma=\int_{\mathcal{S}_{r_{0}}}(1-q_{1})u_{2}^{+}u_{0}\mathrm{d}\sigma-\eta_{1}\int_{\Gamma_{r_{0}}^{\pm}}\delta u_{2}^{+}u_{0}\mathrm{d}\sigma$
$-\eta_{1}\int_{\Gamma_{r_{0}}^{\pm}}(\partial_{1}u_{2}^{+}(\mathbf{0})x_{1}+\partial_{2}u_{2}^{+}(\mathbf{0})x_{2})e^{\rho\cdot\mathbf{x}}\psi(\mathbf{x})\mathrm{d}\sigma+I_{4}.$ (3.23)
By careful calculations, for the right-hand and left-hand sides of (3.23), it is evident that
$|RHS|\lesssim\tau^{-\frac{13}{6}}+\tau^{-(\alpha+2)}+(1+\tau)(1+\tau^{-\frac{2}{3}})e^{-\varsigma r_{0}\tau}+\mathcal{O}(\tau^{-\frac{7}{6}}e^{-\frac{1}{2}\varsigma r_{0}\tau}),$
and
$|LHS|\gtrsim\frac{\Gamma(2)|\eta_{1}||\mathcal{N}|}{\tau^{2}|(\mathbf{d}+\mathrm{i}\mathbf{d}^{\perp})\cdot\mathbf{\hat{x}}_{1}|^{2}|(\mathbf{d}+\mathrm{i}\mathbf{d}^{\perp})\cdot\mathbf{\hat{x}}_{2}|^{2}}-\mathcal{O}(\tau^{-1}e^{-\lambda r_{0}\tau}).$
Similarly, it yields that
$|\mathcal{N}|=0$
as $\tau\to\infty$. When $\mathbf{d}^{\perp}$ is given in the form of (2.13), we have
$\mathcal{N}=$ $(\partial_{1}u_{2}^{+}(\mathbf{0})\cos\vartheta_{M}+\partial_{2}u_{2}^{+}(\mathbf{0})\sin\vartheta_{M})e^{2\mathrm{i}\vartheta_{m}}$
$\quad+(\partial_{1}u_{2}^{+}(\mathbf{0})\cos\vartheta_{m}+\partial_{2}u_{2}^{+}(\mathbf{0})\sin\vartheta_{m})e^{2\mathrm{i}\vartheta_{M}},$ (3.24)
And when $\mathbf{d}^{\perp}$ is given by (2.14), one has
$\mathcal{N}=$ $(\partial_{1}u_{2}^{+}(\mathbf{0})\cos\vartheta_{M}+\partial_{2}u_{2}^{+}(\mathbf{0})\sin\vartheta_{M})e^{-2\mathrm{i}\vartheta_{m}}$
$\quad+(\partial_{1}u_{2}^{+}(\mathbf{0})\cos\vartheta_{m}+\partial_{2}u_{2}^{+}(\mathbf{0})\sin\vartheta_{m})e^{-2\mathrm{i}\vartheta_{M}}.$ (3.25)
Let us consider the real and imaginary parts of $\partial_{1}u_{2}^{+}(\mathbf{0})$ and $\partial_{2}u_{2}^{+}(\mathbf{0})$, seperately. Denote $\partial_{1}u_{2}^{+}(\mathbf{0})=a_{1}+b_{1}\mathrm{i}$ and $\partial_{2}u_{2}^{+}(\mathbf{0})=a_{2}+b_{2}\mathrm{i}$, where $a_{i},b_{i}\in\mathbb{R}\ (i=1,2).$ For (3.24), it follows that
$\Re(\mathcal{N})$ $=a_{1}(\cos\vartheta_{M}\cos 2\vartheta_{m}+\cos\vartheta_{m}\cos 2\vartheta_{M})-b_{1}(\cos\vartheta_{M}\sin 2\vartheta_{m}+\cos\vartheta_{m}\sin 2\vartheta_{M})$
$+a_{2}(\sin\vartheta_{M}\cos 2\vartheta_{m}+\sin\vartheta_{m}\cos 2\vartheta_{M})-b_{2}(\sin\vartheta_{M}\sin 2\vartheta_{m}+\sin\vartheta_{m}\sin 2\vartheta_{M}),$ (3.26)
$\Im(\mathcal{N})$ $=a_{1}(\cos\vartheta_{M}\sin 2\vartheta_{m}+\cos\vartheta_{m}\sin 2\vartheta_{M})+b_{1}(\cos\vartheta_{M}\cos 2\vartheta_{m}+\cos\vartheta_{m}\cos 2\vartheta_{M})$
$+a_{2}(\sin\vartheta_{m}\sin 2\vartheta_{M}+\sin\vartheta_{m}\sin 2\vartheta_{M})+b_{2}(\sin\vartheta_{M}\cos 2\vartheta_{m}+\sin\vartheta_{m}\cos 2\vartheta_{M}).$
Similarly, for (3.25), one has
$\Re(\mathcal{N})$ $=a_{1}(\cos\vartheta_{M}\cos 2\vartheta_{m}+\cos\vartheta_{m}\cos 2\vartheta_{M})+b_{1}(\cos\vartheta_{M}\sin 2\vartheta_{m}+\cos\vartheta_{m}\sin 2\vartheta_{M})$
$+a_{2}(\sin\vartheta_{M}\cos 2\vartheta_{m}+\sin\vartheta_{m}\cos 2\vartheta_{M})+b_{2}(\sin\vartheta_{M}\sin 2\vartheta_{m}+\sin\vartheta_{m}\sin 2\vartheta_{M}),$
$\Im(\mathcal{N})$ $=-a_{1}(\cos\vartheta_{M}\sin 2\vartheta_{m}+\cos\vartheta_{m}\sin 2\vartheta_{M})+b_{1}(\cos\vartheta_{M}\cos 2\vartheta_{m}+\cos\vartheta_{m}\cos 2\vartheta_{M})$
$-a_{2}(\sin\vartheta_{m}\sin 2\vartheta_{M}+\sin\vartheta_{m}\sin 2\vartheta_{M})+b_{2}(\sin\vartheta_{M}\cos 2\vartheta_{m}+\sin\vartheta_{m}\cos 2\vartheta_{M}).$ (3.27)
Combining (3.26)-(3.27) with the fact that $|\mathcal{N}|=0$, we can derive that
$\mathcal{M}_{1}\left(\begin{matrix}a_{1}\\ a_{2}\end{matrix}\right)=0,\ \mathcal{M}_{2}\left(\begin{matrix}b_{1}\\ b_{2}\end{matrix}\right)=0,$ (3.28)
where
$\mathcal{M}_{1}=\left(\begin{matrix}\cos\vartheta_{M}\cos 2\vartheta_{m}+\cos\vartheta_{m}\cos 2\vartheta_{M}&\sin\vartheta_{M}\cos 2\vartheta_{m}+\sin\vartheta_{m}\cos 2\vartheta_{M}\\ \cos\vartheta_{M}\sin 2\vartheta_{m}+\cos\vartheta_{m}\sin 2\vartheta_{M}&\sin\vartheta_{M}\sin 2\vartheta_{m}+\sin\vartheta_{m}\sin 2\vartheta_{M}\end{matrix}\right)$
and
$\mathcal{M}_{2}=\left(\begin{matrix}\cos\vartheta_{M}\sin 2\vartheta_{m}+\cos\vartheta_{m}\sin 2\vartheta_{M}&\sin\vartheta_{M}\sin 2\vartheta_{m}+\sin\vartheta_{m}\sin 2\vartheta_{M}\\ \cos\vartheta_{M}\cos 2\vartheta_{m}+\cos\vartheta_{m}\cos 2\vartheta_{M}&\sin\vartheta_{M}\cos 2\vartheta_{m}+\sin\vartheta_{m}\cos 2\vartheta_{M}\end{matrix}\right).$
According to the admissible condition, we know that $\vartheta_{M}-\vartheta_{m}\in(0,\pi)$ and $\vartheta_{M}-\vartheta_{m}\not=\frac{\pi}{2}$. After tedious calculations, we have
$\det(\mathcal{M}_{1})=-\det(\mathcal{M}_{2})=-2\sin^{2}(\vartheta_{M}-\vartheta_{m})\cos(\vartheta_{M}-\vartheta_{m})\neq 0.$
Therefore, by (3.28), we have $a_{i}=b_{i}=0$ for $i=1,2$, which indicates that $\partial_{1}u_{2}^{+}(\mathbf{0})=\partial_{2}u_{2}^{+}(\mathbf{0})=0$. This leads to a contradiction.
The proof is complete. ∎
Remark 3.3.
From the proof of Theorem 3.1, we can see that when the conductive boundary parameter $\eta$ is a Hölder continuous function on the boundary and $\eta$ does not vanish at the vertex of the scatterer, we can also establish the local unique result for determining the shape of an admissible conductive polygonal-nest or polygonal-cell medium scatterer by a single far-field pattern. The corresponding proof can be modified slightly from the proof of Theorem 3.1. Similarly, a global unique result for identifying the shape of an admissible conductive polygonal-nest medium scatterer by a single far-field pattern can be obtained for a Hölder continuous conductive boundary parameter with the non-vanishing property at the vertex.
Considering an admissible conductive polygonal-cell medium scatterer, the following theorem states the unique identifiability for the refractive index and conductive boundary parameter by a single far-field measurement, where we require a-priori knowledge of the shape and structure of the admissible conductive polygonal-cell medium scatterer.
Theorem 3.2.
Suppose that $(\Omega;q_{\ell},\eta_{\ell}),\ell=1,2$ are two admissible conductive polygon-cell medium scatterers associated with the scattering problem (1.7). Suppose that the refractive index $q_{\ell}$ and conductive boundary parameter $\eta_{\ell}$ have a common polygonal-cell structure $\cup_{i=1}^{N}\Sigma_{i}$ described by Definition 3.2. Let $q_{\ell}$ and $\eta_{\ell}$ be defined as follows:
$q_{\ell}=\sum_{i=1}^{N}q_{i,\ell}\chi_{\Sigma_{i}},\ \eta_{\ell}=\sum_{i=1}^{N}\tilde{\eta}_{\ell}\chi_{\Sigma_{i}}.$
The coefficients of $q_{i,\ell}$ are real, where $q_{i,\ell}$ follows the form specified in (3.1), while $\tilde{\eta}_{\ell}$ is a real constant. Let $u_{\ell}^{\infty}(\mathbf{\hat{x}};u^{i})$ be the far-field pattern associated with the incident wave $u^{i}$ corresponding to $(\Omega;q_{\ell},\eta_{\ell})$. Assume that
$u_{1}^{\infty}(\mathbf{\hat{x}};u^{i})=u_{2}^{\infty}(\mathbf{\hat{x}};u^{i})$
for all $\mathbf{\hat{x}}\in\mathbb{S}^{1}$ and a fixed incident wave $u^{i}$. Then we have $q_{1}=q_{2}$ and $\eta_{1}=\eta_{2}.$
Considering an admissible conductive polygonal-nest medium scatterer, we have shown that the shape of the underlying medium scatterer can be uniquely determined by a single far-field measurement in Theorem 3.1. In the following theorem, we will further prove that the polygonal-nest structure of of the underlying scatterer, the corresponding reflective indexes and the conductive boundary parameter can be uniquely determined by a single far-field pattern.
Theorem 3.3.
Suppose $(\Omega_{\ell};q_{\ell},\eta_{\ell}),\ell=1,2$ are two admissible polygonal-nest medium scatterers associated with the scattering problem (1.7), where $\Omega_{1}=\bigcup_{i=1}^{N_{1}}\mathcal{D}_{i,1}$ and $\Omega_{2}=\bigcup_{i=1}^{N_{2}}\mathcal{D}_{i,2}$ with $\mathcal{D}_{i,\ell}=\Sigma_{i,\ell}\setminus\overline{\Sigma_{i,\ell}},i=1,\dots,N_{\ell}$. For each $\ell$, let $q_{\ell}$ and $\eta_{\ell}$ be defined as follows:
$q_{\ell}=\sum_{i=1}^{N_{\ell}}q_{i,\ell}\chi_{\mathcal{D}_{i,\ell}},\ \eta_{\ell}=\sum_{i=1}^{N_{\ell}}\eta_{i,\ell}\chi_{\partial\Sigma_{i,\ell}}.$
The coefficients of $q_{i,\ell}$ are real, where $q_{i,\ell}$ follows the form specified in (3.1), while $\eta_{i,\ell}$ is a real constant. Let $u_{\ell}^{\infty}(\mathbf{\hat{x}};u^{i})$ be the far-field pattern associated with the incident wave $u^{i}$ corresponding to $(\Omega_{\ell};q_{\ell},\eta_{\ell})$. Assume that
$u_{1}^{\infty}(\mathbf{\hat{x}};u^{i})=u_{2}^{\infty}(\mathbf{\hat{x}};u^{i})$
for all $\mathbf{\hat{x}}\in\mathbb{S}^{1}$ and a fixed incident wave $u^{i}$. Then we have $N_{1}=N_{2}=N$, $\partial\Sigma_{i,1}=\partial\Sigma_{i,2}$, $q_{i,1}=q_{i,2}$ and $\eta_{i,1}=\eta_{i,2}$ for $i=1,\dots,N$,
Remark 3.4.
In Theorems 3.2 and 3.3, it is assumed that the coefficients of the reflective index follow the form specified in (3.1) and are real, while the conductive boundary parameter is a real constant. This assumption is also crucial in Lemma 3.2, as the proof of Theorems 3.2 and 3.3 relies on Lemma 3.2. Remark 3.5 suggests that the conclusion of Lemma 3.2 may hold in a more general scenario where the coefficients of the reflective index and the conductive boundary parameter are complex numbers. Hence, it is speculated that the linear polynomial reflective index with complex coefficients can be uniquely determined through a single far-field measurement.
Consider an admissible conductive polygonal-nest or polygonal-cell medium scatterer $(\Omega;q,\eta)$ as defined in Definition 3.5, where $\Omega\Subset\mathbb{R}^{2}$ possesses a polygonal-nest or polygonal-cell structure. Each vertex $\mathbf{x}_{c}$ of $\Omega$ corresponds to a corner with an associated opening angle $\theta\in(0,\pi)$. Utilizing the translation transformation ${\mathbf{y}}=\mathbf{x}-\mathbf{x}_{c}$, we observe that the underlying corner can be represented by (1.2), where $\Gamma_{r_{0}}^{\pm}$ denote two line segment boundaries of $\mathcal{S}_{r_{0}}$. Recall that $\vartheta_{m}$ and $\vartheta_{M}$ are the polar angles of $\Gamma_{r_{0}}^{-}$ and $\Gamma_{r_{0}}^{+}$, respectively, such that $\theta=\vartheta_{M}-\vartheta_{m}$.
Furthermore, we introduce the rotation transformation $\widetilde{\mathbf{x}}=M(\vartheta_{m})\mathbf{y}$, employing the rotation matrix $M(\vartheta_{m})$ given by:
$M(\vartheta_{m})=\begin{bmatrix}\cos(\vartheta_{m})&\sin(\vartheta_{m})\\ -\sin(\vartheta_{m})&\cos(\vartheta_{m})\end{bmatrix}.$ (3.29)
In this new coordinate system $\mathcal{C}_{N}$, we set $\mathbf{x}_{c}=\mathbf{0}$ and $\vartheta_{m}=0$. Proposition 3.1 demonstrates that if two linear polynomials coincide in the new coordinate system $\mathcal{C}_{N}$, they are identical in the original coordinate system. This result is crucial for proving the unique determination of a linear polynomial refractive index for $(\Omega;q,\eta)$. Similarly, Proposition 3.2 investigates the corresponding admissible conditions in the new coordinate system $\mathcal{C}_{N}$ for the admissible conditions (3.8) and (3.9), originally formulated in the old coordinate system. The proofs of Proposition 3.1 and 3.2 are omitted.
Proposition 3.1.
In a fixed coordinate system $\mathcal{C}$, consider two linear polynomials with the form given by
$\alpha(\mathbf{x})=a_{0}+\mathbf{a}^{\top}\mathbf{x},\ \mbox{and}\ \gamma(\mathbf{x})=b_{0}+\mathbf{b}^{\top}\mathbf{x},$
where $\mathbf{a}=(a_{1},a_{2})^{\top}$, $\mathbf{b}=(b_{1},b_{2})^{\top}$ and $\mathbf{x}=(x_{1},x_{2})^{\top}$. Pick up any fixed point $\mathbf{x}_{c}$ in the coordinate system $\mathcal{C}$ and a fixed rotation matrix $M(\vartheta_{m})$ given by (3.29). Here $\vartheta_{m}\in[-\pi,\pi)$ is a fixed angle in the coordinate system $\mathcal{C}$. Let
$\tilde{\mathbf{x}}=M(\vartheta_{m})(\mathbf{x}-\mathbf{x}_{c}),$
and
$p(\tilde{\mathbf{x}})=\alpha(M(\vartheta_{m})^{\top}\tilde{\mathbf{x}}+\mathbf{x}_{c}),\ g(\tilde{\mathbf{x}})=\gamma(M(\vartheta_{m})^{\top}\tilde{\mathbf{x}}+\mathbf{x}_{c}).$
If $p(\tilde{\mathbf{x}})\equiv g(\tilde{\mathbf{x}}),$ then $\alpha(\mathbf{x})=\gamma(\mathbf{x}).$
Proposition 3.2.
Consider a fixed coordinate system $\mathcal{C}$ and an admissible conductive polygonal-nest or polygonal-cell medium scatterer $(\Omega;q,\eta)$ in $\mathcal{C}$ as defined in Definition 3.5, where $\Omega$ possesses a polygonal-nest or polygonal-cell structure. Each vertex $\mathbf{x}_{c}$ of $\Omega$ corresponds to a corner with an associated opening angle $\theta\in(0,\pi)$. Let $u$ represent the total wave field of the direct scattering problem (1.7) associated with $(\Omega;q,\eta)$. Define $U(\tilde{\mathbf{x}})=u(M(\vartheta_{m})^{\top}\tilde{\mathbf{x}}+{\mathbf{x}_{c}})$, where $\vartheta_{m}\in[-\pi,\pi)$ is a fixed angle in the coordinate system $\mathcal{C}$. Then, we have the following relations:
  • 1.
    If $u(\mathbf{x}_{c})=0$, then $U(\mathbf{0})=0$.
  • 2.
    If $\nabla u(\mathbf{x}_{c})\neq 0$ and $u(\mathbf{x}_{c})=0$, then $\nabla U(\mathbf{0})\neq 0$ and $U(\mathbf{0})=0$.
The proofs of Theorems 3.2 and 3.3 are deferred to Section 4. Prior to that, we require the following lemma, which establishes local uniqueness results for identifying the refractive index and conductive boundary parameter of a coupled PDE system near a corner.
Lemma 3.2.
Recall that $\mathcal{S}_{r_{0}},\Gamma_{r_{0}}^{\pm}$ and $B_{r_{0}}$ are defined in (1.2). Suppose that $u_{1}=u_{1}^{-}\chi_{\mathcal{S}_{r_{0}}}+u_{1}^{+}\chi_{B_{r_{0}}\setminus\overline{\mathcal{S}_{r_{0}}}}\in H^{1}(B_{r_{0}})$ and $u_{2}=u^{-}_{2}\chi_{\mathcal{S}_{r_{0}}}+u^{+}_{2}\chi_{B_{r_{0}}\setminus\overline{\mathcal{S}_{r_{0}}}}\in H^{1}(B_{r_{0}})$ satisfy the following PDE system:
$\begin{cases}(\Delta+k^{2}q_{1})u_{1}^{-}=0,\hskip 78.24507pt\mbox{in}\ \mathcal{S}_{r_{0}},\\ (\Delta+k^{2}q_{2})u_{2}^{-}=0,\hskip 78.24507pt\mbox{in}\ \mathcal{S}_{r_{0}},\\ (\Delta+k^{2}q_{3})u^{+}_{1}=0,\hskip 78.24507pt\mbox{in}\ B_{r_{0}}\setminus\overline{\mathcal{S}_{r_{0}}},\\ (\Delta+k^{2}q_{3})u^{+}_{2}=0,\hskip 78.24507pt\mbox{in}\ B_{r_{0}}\setminus\overline{\mathcal{S}_{r_{0}}},\\ u^{+}_{1}=u^{-}_{1},\ \partial_{\nu}u^{-}_{1}=\partial_{\nu}u^{+}_{1}+\eta_{1}u^{+}_{1},\hskip 12.80365pt\mbox{on}\ \Gamma_{r_{0}}^{\pm},\\ u^{+}_{2}=u^{-}_{2},\ \partial_{\nu}u^{-}_{2}=\partial_{\nu}u^{+}_{2}+\eta_{2}u^{+}_{2},\hskip 12.80365pt\mbox{on}\ \Gamma_{r_{0}}^{\pm},\\ \end{cases}$ (3.30)
where $\eta_{1}$ and $\eta_{2}$ are real constants and $q_{1},\ q_{2},\ q_{3}$ with forms
$q_{j}=q_{j}^{(0)}+q_{j}^{(1)}x_{1}+q_{j}^{(2)}x_{2},\quad j=1,2,3,$
where $q_{j}^{(0)}$, $q_{j}^{(1)}$ and $q_{j}^{(2)}$ are real constants, $j=1,2,3$.
Assume that for $i=1$ or $2$, either
$|u_{i}(\mathbf{0})|\neq 0,$ (3.31)
or
$\nabla u_{i}(\mathbf{0})\neq 0,\ u_{i}(\mathbf{0})=0$ (3.32)
is fulfilled. Additionally, if (3.32) holds, it is further required that $\vartheta_{M}-\vartheta_{m}\in(\frac{\pi}{3},\frac{2\pi}{3})\setminus\{\frac{\pi}{2}\}$, where $\vartheta_{M}-\vartheta_{m}$ is the opening angle of $\mathcal{S}_{r_{0}}$.
If
$u^{+}_{1}=u^{+}_{2}\ \mbox{in}\ B_{r_{0}}\setminus\overline{\mathcal{S}_{r_{0}}},$
then one has $\eta_{1}=\eta_{2},\ q_{1}=q_{2}.$ Moreover, it holds that $u_{1}^{-}=u_{2}^{-}$ in $\mathcal{S}_{r_{0}}$.
Proof.
Since $u^{+}_{1}=u_{2}^{+}$ in $B_{r_{0}}\setminus\overline{\mathcal{S}_{r_{0}}}$, we have that $u^{+}_{1}=u^{+}_{2}\ \mbox{on}\ \Gamma_{r_{0}}^{\pm}$ by the trace theorem, which indicates that $u_{1}^{-}=u_{1}^{+}=u_{2}^{+}=u_{2}^{-}\mbox{on}\ \Gamma_{r_{0}}^{\pm}$ according to the transmission boundary condition. Without of loss of generality, we assume that $u_{2}^{-}$ fulfills (3.31) or (3.32). Denote $\tilde{u}:=u^{-}_{1}-u_{2}^{-}$, then it can be shown directly that
[…]
Similar to the proof of Theorem 3.1, we know that there exists the CGO solution $u_{0}$ given by (2.1), which satisfies (2.2) and
$(\Delta+k^{2}q_{1})u_{0}=0\mbox{ in }\mathcal{S}_{r_{0}}.$
Using Green’s formula, we have the following integral identity
$\int_{\mathcal{S}_{r_{0}}}k^{2}(q_{2}-q_{1})u_{2}^{-}u_{0}\mathrm{d}\mathbf{x}=(\eta_{1}-\eta_{2})\int_{\Gamma_{r_{0}}^{\pm}}u_{2}^{-}u_{0}\mathrm{d}\sigma+\int_{\Lambda_{r_{0}}}u_{0}\partial_{\nu}\tilde{u}-\tilde{u}\partial_{\nu}u_{0}\mathrm{d}\sigma.$ (3.33)
According to Lemma 3.1, we have $u_{2}^{-}\in C^{1,\alpha}(\overline{\mathcal{S}_{r_{0}}})$. Therefore, it follows that
$u_{2}^{-}$ $=u_{2}^{-}(\mathbf{0})+\partial_{1}u_{2}^{-}(\mathbf{0})x_{1}+\partial_{2}u_{2}^{-}(\mathbf{0})x_{2}+\delta u_{2}^{-},\ |\delta u_{2}^{-}|\leq|\mathbf{x}|^{1+\alpha}\|u_{2}^{-}\|_{C^{1,\alpha}}.$ (3.34)
For convenience, we denote $\tilde{\delta}u_{2}^{-}=\partial_{1}u_{2}^{-}(\mathbf{0})x_{1}+\partial_{2}u_{2}^{-}(\mathbf{0})x_{2}+\delta u_{2}^{-}.$
First, we are going to prove $\eta_{1}=\eta_{2}$ and $q_{1}=q_{2}$. According to the admissible conditions (3.8) and (3.9), in the rest of this proof, we will consider two distinct cases.
Case (I): $u_{2}^{-}(\mathbf{0})\not=0.$ Using (3.33) and (3.34), we can directly see that
$(\eta_{2}-\eta_{1})u_{2}^{-}(\mathbf{0})\int_{\Gamma_{r_{0}}^{\pm}}e^{\rho\cdot\mathbf{x}}\mathrm{d}\sigma=\sum_{i=1}^{4}I_{i}$ (3.35)
where
$I_{1}$ $=(\eta_{1}-\eta_{2})u_{2}^{-}(\mathbf{0})\int_{\Gamma_{r_{0}}^{\pm}}\psi(\mathbf{x})e^{\rho\cdot\mathbf{x}}\mathrm{d}\sigma,I_{2}=\int_{\mathcal{S}_{r_{0}}}k^{2}(q_{1}-q_{2})u_{2}^{-}u_{0}\mathrm{d}\mathbf{x},$
$I_{3}$ $=(\eta_{1}-\eta_{2})\int_{\Gamma_{r_{0}}^{\pm}}\tilde{\delta}u_{2}^{-}u_{0}(\mathbf{x})\mathrm{d}\sigma,I_{4}=\int_{\Lambda_{r_{0}}}u_{0}\partial_{\nu}\tilde{u}-\tilde{u}\partial_{\nu}u_{0}\mathrm{d}\sigma.$
By virtue of (3.15) and (3.16), we can derive that the right-hand side of (3.35) satisfies
$|RHS|\lesssim\tau^{-\frac{7}{6}}+\tau^{-(2+\alpha)}+\mathcal{O}(\tau^{-1}e^{-\frac{1}{2}\varsigma r_{0}\tau})+(1+\tau)(1+\tau^{-\frac{2}{3}})e^{-\varsigma r_{0}\tau}.$
By (2.14) and (3.17), one has the left-hand side of (3.35) satisfying
$|LHS|\gtrsim|(\eta_{2}-\eta_{1})u_{2}^{-}(\mathbf{0})|\frac{\Gamma(1)|e^{\mathrm{i}\vartheta_{m}}+e^{\mathrm{i}\vartheta_{M}}|}{|\tau e^{\mathrm{i}\varphi}|}-\mathcal{O}(\tau^{-1}e^{-\frac{1}{2}\lambda r_{0}\tau}).$ (3.36)
Similar to the proof of Case(a) in Theorem 3.1, we have
$|(\eta_{1}-\eta_{2})u_{2}^{-}(\mathbf{0})|=0,\ \mbox{for}\ \vartheta_{M}-\vartheta_{m}\in(0,\pi),$
which indicates that $\eta_{1}=\eta_{2}$ since $u_{2}^{-}(\mathbf{0})\neq 0.$
Next, using $\eta_{1}=\eta_{2}$, the integral identity (3.35) can be rewritten as
$k^{2}(q_{2}^{(0)}-q_{1}^{(0)})u_{2}^{-}(\mathbf{0})\int_{\mathcal{S}_{r_{0}}}e^{\rho\cdot\mathbf{x}}\mathrm{d}\mathbf{x}=\sum_{i=1}^{4}\mathcal{P}_{i}+I_{4},$ (3.37)
where
$\mathcal{P}_{1}$ $=k^{2}(q_{1}^{(0)}-q_{2}^{(0)})u_{2}^{-}(\mathbf{0})\int_{\mathcal{S}_{r_{0}}}\psi(\mathbf{x})e^{\rho\cdot\mathbf{x}}\mathrm{d}\mathbf{x},\ \mathcal{P}_{2}=k^{2}(q_{1}^{(0)}-q_{2}^{(0)})\int_{\mathcal{S}_{r_{0}}}\tilde{\delta}u_{2}^{-}u_{0}\mathrm{d}\mathbf{x},$
$\mathcal{P}_{3}$ $=k^{2}(q_{1}^{(1)}-q_{2}^{(1)})u_{2}^{-}(\mathbf{0})\int_{\mathcal{S}_{r_{0}}}x_{1}u_{0}\mathrm{d}\mathbf{x}+k^{2}(q_{1}^{(2)}-q_{2}^{(2)})u_{2}^{-}(\mathbf{0})\int_{\mathcal{S}_{r_{0}}}x_{2}u_{0}\mathrm{d}\mathbf{x},$
$\mathcal{P}_{4}$ $=k^{2}(q_{1}^{(1)}-q_{2}^{(1)})u_{2}^{-}(\mathbf{0})\int_{\mathcal{S}_{r_{0}}}x_{1}\tilde{\delta}u_{2}^{-}u_{0}\mathrm{d}\mathbf{x}+k^{2}(q_{1}^{(2)}-q_{2}^{(2)})u_{2}^{-}(\mathbf{0})\int_{\mathcal{S}_{r_{0}}}x_{2}\tilde{\delta}u_{2}^{-}u_{0}\mathrm{d}\mathbf{x},$
and $I_{4}$ is defined in (3.35). Then by Proposition 2.1, it is directly shown that
$|\mathcal{P}_{1}|$ $\lesssim\tau^{-29/12}+\mathcal{O}(\tau^{-1}e^{-1/2\varsigma r_{0}\tau}),\ |\mathcal{P}_{2}|\lesssim\tau^{-3}+\tau^{-(\alpha+3)}+\mathcal{O}(\tau^{-1}e^{-1/2\varsigma r_{0}\tau}),$
$|\mathcal{P}_{3}|$ $\lesssim\int_{\mathcal{S}_{r_{0}}}|x||u_{0}|\mathrm{d}\mathbf{x}\lesssim\tau^{-3}+\mathcal{O}(\tau^{-1}e^{-1/2\varsigma r_{0}\tau}),$ (3.38)
$|\mathcal{P}_{4}|$ $\lesssim\int_{\mathcal{S}_{r_{0}}}|x||\tilde{\delta}u_{2}^{-}u_{0}|\mathrm{d}\mathbf{x}\lesssim\tau^{-4}+\tau^{-(\alpha+4)}+O(\tau^{-1}e^{-1/2\varsigma r_{0}\tau}).$
Let $\mathbf{d}^{\perp}$ be in the form of (2.14), we find that the left-hand side of (3.37) satisfies
$LHS=k^{2}(q_{0}^{(2)}-q_{0}^{(1)})u_{2}^{-}(\mathbf{0})\frac{\Gamma(2)(e^{2\mathrm{i}\vartheta_{M}}-e^{2\mathrm{i}\vartheta_{m}})}{\tau^{2}e^{2\mathrm{i}\varphi}}+\mathcal{O}(\tau^{-1}e^{-\frac{1}{2}\lambda r_{0}\tau}).$ (3.39)
Similarly, by combining (3.37)-(3.39), as $\tau\to\infty$, we can obtain
$|k^{2}(q_{0}^{(2)}-q_{0}^{(1)})u_{2}^{-}(\mathbf{0})|=0,$
since $|e^{2\mathrm{i}\vartheta_{M}}-e^{2\mathrm{i}\vartheta_{m}}|\neq 0$ when $\vartheta_{M}-\vartheta_{m}\in(0,\pi)$. Therefore, $q_{0}^{(2)}=q_{0}^{(1)}$ holds due to $u_{2}^{-}(\mathbf{0})\neq 0$.
Using $q_{1}^{(0)}=q_{2}^{(0)}$, from (3.37), we have the following integral identity
$k^{2}(q_{2}^{(1)}-q_{1}^{(1)})u_{2}^{-}(\mathbf{0})\int_{\mathcal{S}_{r_{0}}}x_{1}e^{\rho\cdot\mathbf{x}}\mathrm{d}\mathbf{x}+k^{2}(q_{2}^{(2)}-q_{1}^{(2)})u_{2}^{-}(\mathbf{0})\int_{\mathcal{S}_{r_{0}}}x_{2}e^{\rho\cdot\mathbf{x}}\mathrm{d}\mathbf{x}=\sum_{i=1}^{2}\mathcal{E}_{i}+I_{4},$ (3.40)
where
$\mathcal{E}_{1}$ $=k^{2}(q_{1}^{(1)}-q_{2}^{(1)})u_{2}^{-}(\mathbf{0})\int_{\mathcal{S}_{r_{0}}}x_{1}\psi(\mathbf{x})e^{\rho\cdot\mathbf{x}}\mathrm{d}\mathbf{x}+k^{2}(q_{1}^{(2)}-q_{2}^{(2)})u_{2}^{-}(\mathbf{0})\int_{\mathcal{S}_{r_{0}}}x_{2}\psi(\mathbf{x})e^{\rho\cdot\mathbf{x}}\mathrm{d}\mathbf{x},$
$\mathcal{E}_{2}$ $=k^{2}(q_{1}^{(1)}-q_{2}^{(1)})\int_{\mathcal{S}_{r_{0}}}x_{1}\tilde{\delta}u_{2}^{-}u_{0}\mathrm{d}\mathbf{x}+k^{2}(q_{1}^{(2)}-q_{2}^{(2)})\int_{\mathcal{S}_{r_{0}}}x_{2}\tilde{\delta}u_{2}^{-}u_{0}\mathrm{d}\mathbf{x}.$
Utilizing Proposition 2.1, it can be calculated estimates about $\mathcal{E}_{1}$ and $\mathcal{E}_{2}$ as follows,
$|\mathcal{E}_{1}|$ $\lesssim\int_{\mathcal{S}_{r_{0}}}|\mathbf{x}||\psi(\mathbf{x})e^{\rho\cdot\mathbf{x}}|\mathrm{d}\mathbf{x}\lesssim\tau^{-\frac{41}{12}}+\mathcal{O}(\tau^{-1}e^{-\frac{1}{2}\varsigma r_{0}\tau}),$
$|\mathcal{E}_{2}|$ $\lesssim\int_{\mathcal{S}_{r_{0}}}|\mathbf{x}||\tilde{\delta}u_{2}^{-}||u_{0}|\mathrm{d}\mathbf{x}\lesssim\tau^{-(3+\alpha)}+\tau^{-(\frac{53}{12}+\alpha)}+\mathcal{O}(\tau^{-1}e^{-\frac{1}{2}\varsigma r_{0}\tau}).$ (3.41)
Then we can write the left-hand side of (3.40) as
$\overline{q}_{1}\int_{\vartheta_{m}}^{\vartheta_{M}}\mathrm{d}\vartheta\int_{0}^{r_{0}}r^{2}\cos\vartheta e^{-\tau e^{\mathrm{i}(\vartheta-\varphi)r}}\mathrm{d}r+\overline{q}_{2}\int_{\vartheta_{m}}^{\vartheta_{M}}\mathrm{d}\vartheta\int_{0}^{r_{0}}r^{2}\sin\vartheta e^{-\tau e^{\mathrm{i}(\vartheta-\varphi)r}}\mathrm{d}r,$ (3.42)
when $\mathbf{d}^{\perp}$ is selected in the form described in (2.13), and
$\overline{q}_{1}\int_{\vartheta_{m}}^{\vartheta_{M}}\mathrm{d}\vartheta\int_{0}^{r_{0}}r^{2}\cos\vartheta e^{-\tau e^{\mathrm{i}(\varphi-\vartheta)r}}\mathrm{d}r+\overline{q}_{2}\int_{\vartheta_{m}}^{\vartheta_{M}}\mathrm{d}\vartheta\int_{0}^{r_{0}}r^{2}\sin\vartheta e^{-\tau e^{\mathrm{i}(\varphi-\vartheta)r}}\mathrm{d}r,$ (3.43)
when $\mathbf{d}^{\perp}$ is selected in the form of (2.14), where $\overline{q}_{i}:=k^{2}(q_{2}^{(i)}-q_{1}^{(i)})u_{2}^{-}(\mathbf{0}),i=1,2$.
Utilizing (2.1), the first term of (3.42) can be expressed as:
$\frac{\Gamma(3)\overline{q}_{1}e^{3\mathrm{i}\varphi}}{\tau^{3}}$ $\int_{\vartheta_{m}}^{\vartheta_{M}}\cos\vartheta e^{-3\mathrm{i}\vartheta}\mathrm{d}\vartheta+\mathcal{O}(\tau^{-1}e^{-\frac{1}{2}\varsigma r_{0}\tau})$
$=\frac{\Gamma(3)e^{3\mathrm{i}\varphi}}{\tau^{3}}\cdot\frac{1}{8}\left\{(-\sin\vartheta+3\mathrm{i}\cos\vartheta)e^{-3\mathrm{i}\vartheta}\right\}\bigg{|}_{\vartheta_{m}}^{\vartheta_{M}}+\mathcal{O}(\tau^{-1}e^{-\frac{1}{2}\lambda r_{0}\tau}).$ (3.44)
The second term can be simplified as:
$\frac{\Gamma(3)\overline{q}_{2}e^{3\mathrm{i}\varphi}}{\tau^{3}}$ $\int_{\vartheta_{m}}^{\vartheta_{M}}\sin\vartheta e^{-3\mathrm{i}\vartheta}\mathrm{d}\vartheta+\mathcal{O}(\tau^{-1}e^{-\frac{1}{2}\varsigma r_{0}\tau})$
$=\frac{\Gamma(3)e^{3\mathrm{i}\varphi}}{\tau^{3}}\cdot\frac{1}{8}\left\{(\cos\vartheta+3\mathrm{i}\sin\vartheta)e^{-3\mathrm{i}\vartheta}\right\}\bigg{|}_{\vartheta_{m}}^{\vartheta_{M}}+\mathcal{O}(\tau^{-1}e^{-\frac{1}{2}\lambda r_{0}\tau}).$ (3.45)
Similarly, the first term of (3.43) can be expressed as:
$\frac{\Gamma(3)\overline{q}_{1}e^{-3\mathrm{i}\varphi}}{\tau^{3}}$ $\int_{\vartheta_{m}}^{\vartheta_{M}}\cos\vartheta e^{3\mathrm{i}\vartheta}\mathrm{d}\vartheta+\mathcal{O}(\tau^{-1}e^{-\frac{1}{2}\varsigma r_{0}\tau})$
$=\frac{\Gamma(3)e^{-3\mathrm{i}\varphi}}{\tau^{3}}\cdot\frac{1}{8}\left\{-(\sin\vartheta+3\mathrm{i}\cos\vartheta)e^{3\mathrm{i}\vartheta}\right\}\bigg{|}_{\vartheta_{m}}^{\vartheta_{M}}+\mathcal{O}(\tau^{-1}e^{-\frac{1}{2}\lambda r_{0}\tau}).$ (3.46)
The second term of (3.43) can be written as:
$\frac{\Gamma(3)\overline{q}_{2}e^{-3\mathrm{i}\varphi}}{\tau^{3}}$ $\int_{\vartheta_{m}}^{\vartheta_{M}}\sin\vartheta e^{3\mathrm{i}\vartheta}\mathrm{d}\vartheta+\mathcal{O}(\tau^{-1}e^{-\frac{1}{2}\varsigma r_{0}\tau})$
$=\frac{\Gamma(3)e^{-3\mathrm{i}\varphi}}{\tau^{3}}\cdot\frac{1}{8}\left\{(\cos\vartheta-3\mathrm{i}\sin\vartheta)e^{3\mathrm{i}\vartheta}\right\}\bigg{|}_{\vartheta_{m}}^{\vartheta_{M}}+\mathcal{O}(\tau^{-1}e^{-\frac{1}{2}\lambda r_{0}\tau}).$ (3.47)
Combining (3.41)-(3.47) with (3.40), one has
[…]
$\leq\tau^{-\frac{41}{12}}+\tau^{-(3+\alpha)}+(1+\tau)(1+\tau^{2/3})e^{-\varsigma r_{0}\tau}+\mathcal{O}(\tau^{-1}e^{-\frac{1}{2}\varsigma r_{0}\tau})+\mathcal{O}(\tau^{-1}e^{-\lambda r_{0}\tau})$
and
[…]
$\leq\tau^{-\frac{41}{12}}+\tau^{-(3+\alpha)}+(1+\tau)(1+\tau^{2/3})e^{-\varsigma r_{0}\tau}+\mathcal{O}(\tau^{-1}e^{-\frac{1}{2}\varsigma r_{0}\tau})+\mathcal{O}(\tau^{-1}e^{-\lambda r_{0}\tau}).$
By multiplying both sides of (3.48) by $\frac{8\tau^{3}}{\Gamma(3)e^{3\mathrm{i}\varphi}}$, (3.49) by $\frac{8\tau^{3}e^{3\mathrm{i}\varphi}}{\Gamma(3)}$, and letting $\tau\to\infty$, we obtain
$B\left(\begin{matrix}\overline{q}_{1}\\ \overline{q}_{2}\end{matrix}\right)=\mathbf{0},$
where
$B=\left(\begin{matrix}(-\sin\vartheta+3\mathrm{i}\cos\vartheta)e^{-3\mathrm{i}\vartheta}|_{\vartheta_{m}}^{\vartheta_{M}}&(\cos\vartheta+3\mathrm{i}\sin\vartheta)e^{-3\mathrm{i}\vartheta}|_{\vartheta_{m}}^{\vartheta_{M}}\\ -(\sin\vartheta+3\mathrm{i}\cos\vartheta)e^{3\mathrm{i}\vartheta}|_{\vartheta_{m}}^{\vartheta_{M}}&(\cos\vartheta-3\mathrm{i}\sin\vartheta)e^{3\mathrm{i}\vartheta}|_{\vartheta_{m}}^{\vartheta_{M}}\end{matrix}\right).$
Furthermore, it can be calculated that
$\det(B)$ $=(20\sin(3\beta)\sin\beta+12\cos(3\beta)\cos\beta-12)\mathrm{i}$
=$-4(\cos(2\beta)-1)^{2}\mathrm{i}\not=0,\ \beta=\vartheta_{M}-\vartheta_{m}\in(0,\pi).$
Hence we get that $\overline{q}_{1}=0$ and $\overline{q}_{2}=0$, which implies that $q_{1}^{(1)}=q_{2}^{(1)}$ and $q_{1}^{(2)}=q_{2}^{(2)}$ by using $u_{2}^{-}(\mathbf{0})\not=0.$
Case (II): $|\nabla u_{2}^{-}(\mathbf{0})|\not=0$ and $u(\mathbf{0})=0.$ In this scenario, we utilize the rotation transformation $\tilde{\mathbf{x}}=M(\vartheta_{m})^{\top}\mathbf{x}$, where $M(\vartheta_{m})$ is defined by (3.29). Consequently, in the new coordinate system, we denote the linear segment boundary of the corner $\mathcal{S}_{r_{0}}$ by $\tilde{\Gamma}_{r_{0}}^{\pm}$, and the new corner is represented as $\mathcal{S}_{r_{0}}^{{}^{\prime}}$. The boundary part $\Lambda_{r_{0}}$ of $\mathcal{S}_{r_{0}}^{{}^{\prime}}$ is denoted by $\Lambda_{r_{0}}^{{}^{\prime}}$. The polar angle of $\tilde{\Gamma}_{r_{0}}^{-}$ is 00, and the polar angle of $\tilde{\Gamma}_{r_{0}}^{+}$ is $\tilde{\vartheta}=\vartheta_{M}-\vartheta_{m}$.
Denote
$\tilde{q}_{i}(\tilde{\mathbf{x}})=q_{i}(M(\vartheta_{m})^{\top}\tilde{\mathbf{x}}),\ \tilde{q}_{i}=\tilde{q}_{0}^{(i)}+\tilde{q}_{1}^{(i)}\tilde{x}_{1}+\tilde{q}_{2}^{(i)}\tilde{x}_{2},\ i=1,2,3,$
and
$U_{j}^{\pm}(\tilde{\mathbf{x}})=u_{j}^{\pm}(M(\vartheta_{m})^{\top}\tilde{\mathbf{x}}),\ j=1,2.$
Then we can rewrite (3.30) as follows:
$\begin{cases}(\Delta+k^{2}\tilde{q}_{1})U_{1}^{-}=0,\hskip 91.04872pt\mbox{in}\ \mathcal{S}^{{}^{\prime}}_{r_{0}},\\ (\Delta+k^{2}\tilde{q}_{2})U_{2}^{-}=0,\hskip 91.04872pt\mbox{in}\ \mathcal{S}^{{}^{\prime}}_{r_{0}},\\ (\Delta+k^{2}\tilde{q}_{3})U^{+}_{1}=0,\hskip 91.04872pt\mbox{in}\ B_{r_{0}}\setminus\overline{\mathcal{S}^{{}^{\prime}}_{r_{0}}},\\ (\Delta+k^{2}\tilde{q}_{3})U^{+}_{2}=0,\hskip 91.04872pt\mbox{in}\ B_{r_{0}}\setminus\overline{\mathcal{S}^{{}^{\prime}}_{r_{0}}},\\ U^{+}_{1}=U^{-}_{1},\ \partial_{\nu}U^{-}_{1}=\partial_{\nu}U^{+}_{1}+\eta_{1}U^{+}_{1},\hskip 15.6491pt\mbox{on}\ \tilde{\Gamma}_{r_{0}}^{\pm},\\ U^{+}_{2}=U^{-}_{2},\ \partial_{\nu}U^{-}_{2}=\partial_{\nu}U^{+}_{2}+\eta_{2}U^{+}_{2},\hskip 15.6491pt\mbox{on}\ \tilde{\Gamma}_{r_{0}}^{\pm}.\\ \end{cases}$ (3.50)
Denote $\tilde{U}=U_{1}^{-}-U_{2}^{-},$ then one has
[…]
By Lemma 2.1, we know that there exists a CGO solution $\tilde{u}_{0}$ of the form
$\tilde{u}_{0}(\tilde{\mathbf{x}})=(1+\psi(\tilde{\mathbf{x}}))e^{\rho\cdot\tilde{\mathbf{x}}}$
satisfying that
$(\Delta+k^{2}\tilde{q}_{1})\tilde{u}_{0}(\tilde{\mathbf{x}})=0,\ \mbox{in}\ \mathcal{S}_{r_{0}}^{{}^{\prime}}.$
Recall that
$u_{2}^{-}$ $=u_{2}^{-}(\mathbf{0})+\partial_{1}u_{2}^{-}(\mathbf{0})x_{1}+\partial_{2}u_{2}^{-}(\mathbf{0})x_{2}+\delta u_{2}^{-}$
$=u_{2}^{-}(\mathbf{0})+\mathbf{x}^{\top}\nabla_{\mathbf{x}}u_{2}^{-}(\mathbf{0})+\delta u_{2}^{-}.$
Due to $\mathbf{x}=M(\vartheta_{m})^{\top}\tilde{\mathbf{x}}$ and $\nabla_{\mathbf{x}}u_{2}^{-}(\mathbf{0})=M(\vartheta_{m})^{\top}\nabla_{\tilde{\mathbf{x}}}U_{2}^{-}(\mathbf{0})$, then we have
$U_{2}^{-}(\tilde{\mathbf{x}})=U_{2}^{-}(\mathbf{0})+\partial_{1}U_{2}^{-}(\mathbf{0})\tilde{x}_{1}+\partial_{2}U_{2}^{-}(\mathbf{0})\tilde{x}_{2}+\delta U_{2}^{-},|\delta U_{2}^{-}|\leq C\|U_{2}^{-}\|_{C^{1,\alpha}}.$ (3.52)
According to Proposition 3.2, using the condition $|\nabla u_{2}^{-}(\mathbf{0})|\not=0$ and $u(\mathbf{0})=0$, we have $\nabla U_{2}^{-}(\mathbf{0})\neq 0$ and $U_{2}^{-}(\mathbf{0})=0$. In the following, we further divide the rest proof into the following three distinct cases.
Case (i): $\partial_{1}U_{2}^{-}(\mathbf{0})\not=0,\ \partial_{2}U_{2}^{-}(\mathbf{0})\not=0,\ U_{2}^{-}(\mathbf{0})=0.$ Recall that $U_{2}^{-}=\partial_{1}U_{2}^{-}(\mathbf{0})\tilde{x}_{1}+\partial_{2}U_{2}^{-}(\mathbf{0})\tilde{x}_{2}+\delta U_{2}^{-}$, then we have
$\overline{\eta}$ $\int_{\tilde{\Gamma}_{r_{0}}^{\pm}}(\partial_{1}U_{2}^{-}(\mathbf{0})\tilde{x}_{1}+\partial_{2}U_{2}^{-}(\mathbf{0})\tilde{x}_{2})e^{\rho\cdot\tilde{\mathbf{x}}}\mathrm{d}\sigma=\int_{\mathcal{S}_{r_{0}}^{{}^{\prime}}}k^{2}(q_{1}-q_{2})\tilde{u}_{0}U_{2}^{-}\mathrm{d}\tilde{\mathbf{x}}$
$-\overline{\eta}\int_{\tilde{\Gamma}_{r_{0}}^{\pm}}(\partial_{1}U_{2}^{-}(\mathbf{0})\tilde{x}_{1}+\partial_{2}U_{2}^{-}(\mathbf{0})\tilde{x}_{2})\psi(\tilde{\mathbf{x}})e^{\rho\cdot\tilde{\mathbf{x}}}\mathrm{d}\sigma-\overline{\eta}\int_{\tilde{\Gamma}_{r_{0}}^{\pm}}\delta U_{2}^{-}\tilde{u}_{0}\mathrm{d}\sigma+\mathcal{I},$ (3.53)
where $\overline{\eta}=\eta_{2}-\eta_{1}$. Comparing (3.53) with (3.23), then we see that the right-hand and left-hand sides of (3.53) satisfy
$|RHS|\lesssim\tau^{-(2+\alpha)}+\tau^{-\frac{13}{6}}+\mathcal{O}(\tau^{-1}e^{-\frac{1}{2}\varsigma r_{0}\tau})+(1+\tau)(1+\tau^{-\varsigma r_{0}\tau})$
and
$LHS=\frac{\Gamma(2)(\eta_{2}-\eta_{1})\mathcal{Q}}{\tau^{2}((\mathbf{d}+\mathrm{i}\mathbf{d}^{\perp})\cdot\mathbf{\hat{x}}_{1})^{2}((\mathbf{d}+\mathrm{i}\mathbf{d}^{\perp})\cdot\mathbf{\hat{x}}_{2})^{2}}+\mathcal{O}(\tau^{-1}e^{-\frac{1}{2}\lambda r_{0}\tau}),$
where
$\mathcal{Q}$ $=(\partial_{1}U_{2}^{-}(\mathbf{0})\cos\tilde{\vartheta}+\partial_{2}U_{2}^{-}(\mathbf{0})\sin\tilde{\vartheta})+\partial_{1}U_{2}^{-}(\mathbf{0})e^{-2\mathrm{i}\tilde{\vartheta}}$
$=(\partial_{1}U_{2}^{-}(\mathbf{0})\cos\tilde{\vartheta}+\partial_{2}U_{2}^{-}(\mathbf{0})\sin\tilde{\vartheta}+\partial_{1}U_{2}^{-}(\mathbf{0})\cos 2\tilde{\vartheta})-\mathrm{i}\partial_{1}U_{2}^{-}(\mathbf{0})\sin 2\tilde{\vartheta}.$
Since $\partial_{1}U_{2}^{-}(\mathbf{0})\neq 0$, it follows that $\Im(\mathcal{Q})\neq 0$ when $\tilde{\vartheta}\neq\frac{\pi}{2}$. By employing a similar manner as in Case (b) of the proof of Theorem 3.1, it can be demonstrated that $\eta_{1}=\eta_{2}$ since $\tilde{\vartheta}=\vartheta_{M}-\vartheta_{m}\neq\frac{\pi}{2}$.
Next, we will prove that $\tilde{q}_{1}^{(0)}=\tilde{q}_{2}^{(0)}$. Using $\eta_{1}=\eta_{2}$, the integral identity (3.53) can be rewritten as
$\tilde{\sigma}_{0}$ $\left(\partial_{1}U_{2}^{-}(\mathbf{0})\int_{\mathcal{S}_{r_{0}}^{{}^{\prime}}}\tilde{x}_{1}e^{\rho\cdot\tilde{\mathbf{x}}}\mathrm{d}\tilde{\mathbf{x}}+\partial_{2}U_{2}^{-}(\mathbf{0})\int_{\mathcal{S}_{r_{0}}^{{}^{\prime}}}\tilde{x}_{2}e^{\rho\cdot\tilde{\mathbf{x}}}\mathrm{d}\tilde{\mathbf{x}}\right)$
$=-\tilde{\sigma}_{0}\left(\partial_{1}U_{2}^{-}(\mathbf{0})\int_{\mathcal{S}_{r_{0}}^{{}^{\prime}}}\tilde{x}_{1}\psi(\tilde{\mathbf{x}})e^{\rho\cdot\tilde{\mathbf{x}}}\mathrm{d}\tilde{\mathbf{x}}+\partial_{2}U_{2}^{-}(\mathbf{0})\int_{\mathcal{S}_{r_{0}}^{{}^{\prime}}}\tilde{x}_{2}\psi(\tilde{\mathbf{x}})e^{\rho\cdot\tilde{\mathbf{x}}}\mathrm{d}\tilde{\mathbf{x}}\right)$
$-\tilde{\sigma}_{0}\int_{\mathcal{S}_{r_{0}}^{{}^{\prime}}}\delta U_{2}^{-}\tilde{u}_{0}\mathrm{d}\tilde{\mathbf{x}}+k^{2}(\tilde{q}_{1}^{(1)}-\tilde{q}_{2}^{(1)})\int_{\mathcal{S}_{r_{0}}^{{}^{\prime}}}\tilde{x}_{1}U_{2}^{-}\tilde{u}_{0}\mathrm{d}\tilde{\mathbf{x}}$
$+k^{2}(\tilde{q}_{1}^{(2)}-\tilde{q}_{2}^{(2)})\int_{\mathcal{S}_{r_{0}}^{{}^{\prime}}}\tilde{x}_{2}U_{2}^{-}\tilde{u}_{0}\mathrm{d}\tilde{\mathbf{x}}+\mathcal{I}$ (3.54)
where $\tilde{\sigma}_{0}=k^{2}(\tilde{q}_{2}^{(0)}-\tilde{q}_{1}^{(0)})$. Initially, according to Proposition 2.1, for the right-hand side of (3.54), it follows that
$|RHS|\lesssim\tau^{-(3+\alpha)}+\tau^{-\frac{41}{12}}+\mathcal{O}(\tau^{-1}e^{-\frac{1}{2}\varsigma r_{0}\tau})+(1+\tau)(1+\tau^{-\frac{2}{3}})e^{-\varsigma r_{0}\tau}.$
Concerning the left-hand side of (3.54) and letting $\mathbf{d}^{\perp}$ be in the form of (2.14), it follows that
$LHS$ $=\frac{\Gamma(3)\tilde{\sigma}_{0}}{\tau^{3}e^{3\mathrm{i}\varphi}}\left(\partial_{1}U_{2}^{-}(\mathbf{0})\int_{0}^{\tilde{\vartheta}}e^{3\mathrm{i}\vartheta}\cos\vartheta\mathrm{d}\vartheta+\partial_{2}U_{2}^{-}(\mathbf{0})\int_{0}^{\tilde{\vartheta}}e^{3\mathrm{i}\vartheta}\sin\vartheta\mathrm{d}\vartheta\right)$
$+\mathcal{O}(\tau^{-1}e^{-\frac{1}{2}\lambda r_{0}\tau}).$
Denote
$\Psi:=\partial_{1}U_{2}^{-}(\mathbf{0})\int_{0}^{\tilde{\vartheta}}e^{3\mathrm{i}\vartheta}\cos\vartheta\mathrm{d}\vartheta+\partial_{2}U_{2}^{-}(\mathbf{0})\int_{0}^{\tilde{\vartheta}}e^{3\mathrm{i}\vartheta}\sin\vartheta\mathrm{d}\vartheta.$
Multiplying both sides of (3.54) by $\tau^{3}$ and taking the limit as $\tau$ tends to infinity yields $\tilde{\sigma}_{0}\Psi=0$. Direct calculations show that
$\Re(\Psi)$ $=-\partial_{1}U_{2}^{-}(\mathbf{0})(\sin\tilde{\vartheta}\cos 3\tilde{\vartheta}-3\sin 3\tilde{\vartheta}\cos\tilde{\vartheta})$
$+\partial_{2}U_{2}^{-}(\mathbf{0})(\cos\tilde{\vartheta}\cos 3\tilde{\vartheta}+3\sin\tilde{\vartheta}\sin 3\tilde{\vartheta}-1),$
$\Im(\Psi)$ $=-\partial U_{2}^{-}(\mathbf{0})(\sin\tilde{\vartheta}\sin 3\tilde{\vartheta}+3\cos\tilde{\vartheta}\cos 3\tilde{\vartheta}-3)$
$+\partial_{2}U_{2}^{-}(\mathbf{0})(\cos\tilde{\vartheta}\sin 3\tilde{\vartheta}-3\sin\tilde{\vartheta}\cos 3\tilde{\vartheta}).$
Next, we will show that $\Psi\not=0$. By contradiction, suppose that $\Re(\Psi)=\Im(\Psi)=0$, then we have
$M\left(\begin{matrix}\partial_{1}U_{2}^{-}(\mathbf{0})\\ \partial_{2}U_{2}^{-}(\mathbf{0})\end{matrix}\right)=\mathbf{0},$
where
$M=\left(\begin{matrix}3\sin 3\tilde{\vartheta}\cos\tilde{\vartheta}-\sin\tilde{\vartheta}\cos 3\tilde{\vartheta}&\cos\tilde{\vartheta}\cos 3\tilde{\vartheta}+3\sin\tilde{\vartheta}\sin 3\tilde{\vartheta}-1\\ \sin\tilde{\vartheta}\sin 3\tilde{\vartheta}+3\cos\tilde{\vartheta}\cos 3\tilde{\vartheta}-3&\cos\tilde{\vartheta}\sin 3\tilde{\vartheta}-3\sin\tilde{\vartheta}\cos 3\tilde{\vartheta}\end{matrix}\right).$
It can be deduced that
$\det(M)=16\sin^{4}\tilde{\vartheta}\not=0\ \mbox{for}\ \tilde{\vartheta}\in(0,\pi),$
where we get a contradiction. Hence, the proof of $\tilde{q}_{2}^{(0)}=\tilde{q}_{1}^{(0)}$ is complete.
Substituting $\tilde{q}_{2}^{(0)}=\tilde{q}_{1}^{(0)}$ into (3.54), we obtain
$\tilde{\sigma}_{1}$ $\int_{\mathcal{S}_{r_{0}}^{{}^{\prime}}}(\partial_{1}U_{2}^{-}(\mathbf{0})\tilde{x}_{1}^{2}+\partial_{2}U_{2}^{-}(\mathbf{0})\tilde{x}_{1}\tilde{x}_{2})e^{\rho\cdot\tilde{\mathbf{x}}}\mathrm{d}\tilde{\mathbf{x}}+\tilde{\sigma}_{2}\int_{\mathcal{S}_{r_{0}}^{{}^{\prime}}}(\partial_{1}U_{2}^{-}(\mathbf{0})\tilde{x}_{1}\tilde{x}_{2}+\partial_{2}U_{2}^{-}(\mathbf{0})\tilde{x}_{2}^{2})e^{\rho\cdot\tilde{\mathbf{x}}}\mathrm{d}\tilde{\mathbf{x}}$
$\ =-\tilde{\sigma}_{1}\int_{\mathcal{S}_{r_{0}}^{{}^{\prime}}}(\partial_{1}U_{2}^{-}(\mathbf{0})\tilde{x}_{1}^{2}+\partial_{2}U_{2}^{-}(\mathbf{0})\tilde{x}_{1}\tilde{x}_{2})\psi(\tilde{\mathbf{x}})e^{\rho\cdot\tilde{\mathbf{x}}}\mathrm{d}\tilde{\mathbf{x}}-\tilde{\sigma}_{1}\int_{\mathcal{S}_{r_{0}}^{{}^{\prime}}}\tilde{x}_{1}\delta U_{2}^{-}\tilde{u}_{0}\mathrm{d}\mathbf{x}$
$\ -\tilde{\sigma}_{2}\int_{\mathcal{S}_{r_{0}}^{{}^{\prime}}}(\partial_{1}U_{2}^{-}(\mathbf{0})\tilde{x}_{1}\tilde{x}_{2}+\partial_{2}U_{2}^{-}(\mathbf{0})\tilde{x}_{2}^{2})\psi(\tilde{\mathbf{x}})e^{\rho\cdot\tilde{\mathbf{x}}}\mathrm{d}\tilde{\mathbf{x}}-\tilde{\sigma}_{2}\int_{\mathcal{S}_{r_{0}}^{{}^{\prime}}}\tilde{x}_{2}\delta U_{2}^{-}\tilde{u}_{0}\mathrm{d}\tilde{\mathbf{x}}+\mathcal{I},$
here $\tilde{\sigma}_{i}=k^{2}(\tilde{q}_{2}^{(i)}-\tilde{q}_{1}^{(i)}),i=1,2.$ Following a similar argument as in Case (I), we obtain a linear equation:
$Q\left(\begin{matrix}\tilde{q}_{2}^{(1)}-\tilde{q}_{1}^{(1)}\\ \tilde{q}_{2}^{(2)}-\tilde{q}_{1}^{(2)}\end{matrix}\right)=\mathbf{0},$ (3.55)
where
$Q_{11}$ $=\left(\partial_{1}U_{2}^{-}(\mathbf{0})(-\frac{1}{2\mathrm{i}}+\frac{1}{3}(2\mathrm{i}\cos 2\vartheta-\sin 2\vartheta))+\frac{1}{3}\partial_{2}U_{2}^{-}(\mathbf{0})(2\mathrm{i}\sin 2\vartheta+\cos 2\vartheta)\right)e^{-4\mathrm{i}\vartheta}|_{0}^{\tilde{\vartheta}}$
$Q_{12}$ $=\left(\frac{1}{3}\partial_{1}U_{2}^{-}(\mathbf{0})(2\mathrm{i}\sin 2\vartheta+\cos 2\vartheta)+\partial_{2}U_{2}^{-}(\mathbf{0})(-\frac{1}{2\mathrm{i}}-\frac{1}{3}(2\mathrm{i}\cos 2\vartheta-\sin 2\vartheta))\right)e^{-4\mathrm{i}\vartheta}|_{0}^{\tilde{\vartheta}}$
$Q_{21}$ $=\left(\partial_{1}U_{2}^{-}(\mathbf{0})(\frac{1}{2\mathrm{i}}-\frac{1}{3}(\sin 2\vartheta+2\mathrm{i}\cos 2\vartheta))+\frac{1}{3}\partial_{2}U_{2}^{-}(\mathbf{0})(\cos 2\vartheta-2\mathrm{i}\sin\vartheta)\right)e^{4\mathrm{i}\vartheta}|_{0}^{\tilde{\vartheta}}$
$Q_{22}$ $=\left(\frac{1}{3}\partial_{1}U_{2}^{-}(\mathbf{0})(\cos 2\vartheta-2\mathrm{i}\sin 2\vartheta)+\partial_{2}U_{2}^{-}(\mathbf{0})(\frac{1}{2\mathrm{i}}+\frac{1}{3}(\sin 2\vartheta+2\mathrm{i}\cos 2\vartheta))\right)e^{4\mathrm{i}\vartheta}|_{0}^{\tilde{\vartheta}}.$
After tedious calculations, it is revealed that
$\det(Q)$ $=\frac{\mathrm{i}}{9}(\partial_{1}U_{2}^{-}(\mathbf{0})\ \partial_{2}U_{2}^{-}(\mathbf{0}))\mathcal{A}\left(\begin{matrix}\partial_{1}U_{2}^{-}(\mathbf{0})\\ \partial_{2}U_{2}^{-}(\mathbf{0})\end{matrix}\right),$
where
$\mathcal{A}(:,1)=\left(\begin{matrix}-3\cos 4\tilde{\vartheta}-8\cos 2\tilde{\vartheta}-5\sin 4\tilde{\vartheta}\sin 2\tilde{\vartheta}+11\\ 6\sin 2\tilde{\vartheta}-6\sin 4\tilde{\vartheta}+3\sin 4\tilde{\vartheta}\cos 2\tilde{\vartheta}\end{matrix}\right)$
and
$\mathcal{A}(:,2)=\left(\begin{matrix}6\sin 2\tilde{\vartheta}-6\sin 4\tilde{\vartheta}+3\sin 4\tilde{\vartheta}\cos 2\tilde{\vartheta}\\ 3\cos 4\tilde{\vartheta}-8\cos 2\tilde{\vartheta}+\sin 4\tilde{\vartheta}\sin 2\tilde{\vartheta}+5\end{matrix}\right).$
In the following, we will show that $\mathcal{A}$ is positive definite. In fact, we have the first order leading principle minors of $\mathcal{A}$ fulfills that
$D_{1}(\mathcal{A})$ $=2(5\cos 2\tilde{\vartheta}+7)(\cos 2\tilde{\vartheta}-1)^{2}>0,\ \mbox{for}\ \tilde{\vartheta}\in(0,\pi).$
When $\tilde{\vartheta}\in(\frac{\pi}{3},\frac{2\pi}{3})$, we can show that the second order leading principle minors of $\mathcal{A}$ satisfies that
$D_{2}(\mathcal{A})$ $=8(2\cos 2\tilde{\vartheta}+1)(\cos 2\tilde{\vartheta}-1)^{5}>0.$
Given that the conductive boundary parameter is a real constant and the refractive index is a linear polynomial with real coefficients, we can treat the PDE system (3.30) separately by considering its real and imaginary parts. This allows us to assume that the solutions $u_{1}$ and $u_{2}$ of (3.30) are real-valued by separating the real and imaginary parts of $u_{1}$ and $u_{2}$. Consequently, $(\partial_{1}u_{2}^{-}(\mathbf{0})\ \partial_{2}u_{2}^{-}(\mathbf{0}))^{\top}$ becomes a nonzero real vector. Leveraging the positivity of $\mathcal{A}$, we deduce that $Q$ is nonsingular, which can be used to prove that $\tilde{q}_{1}^{(1)}=\tilde{q}_{2}^{(1)}$ and $\tilde{q}_{1}^{(2)}=\tilde{q}_{2}^{(2)}$.
By Proposition 3.1, we ascertain that $q_{1}=q_{2}$ and $\eta_{1}=\eta_{2}$.
Case (ii): $\partial_{1}U_{2}^{-}(\mathbf{0})\not=0,\ \partial_{2}U_{2}^{-}(\mathbf{0})=U_{2}^{-}(\mathbf{0})=0.$ In view of (3.51), we observe that
$(\eta_{2}-\eta_{1})\partial_{1}U_{2}^{-}(\mathbf{0})\int_{\tilde{\Gamma}_{r_{0}}^{\pm}}\tilde{x}_{1}e^{\rho\cdot{\tilde{\mathbf{x}}}}\mathrm{d}\sigma=\sum_{i=1}^{3}\mathcal{F}_{i}+\mathcal{I},$ (3.56)
where
$\mathcal{F}_{1}$ $=(\eta_{1}-\eta_{2})\partial_{1}U_{2}^{-}(\mathbf{0})\int_{\tilde{\Gamma}_{r_{0}}^{\pm}}\tilde{x}_{1}\psi(\tilde{\mathbf{x}})e^{\rho\cdot\tilde{\mathbf{x}}}\mathrm{d}\sigma,\ \mathcal{F}_{2}=(\eta_{1}-\eta_{2})\int_{\tilde{\Gamma}_{r_{0}}^{\pm}}\delta U_{2}^{-}\tilde{u}_{0}\mathrm{d}\sigma$
$\mathcal{F}_{3}$ $=\int_{\mathcal{S}^{{}^{\prime}}_{r_{0}}}k^{2}(\tilde{q}_{1}-\tilde{q}_{2})U_{2}^{-}\tilde{u}_{0}\mathrm{d}\tilde{\mathbf{x}},\ \mathcal{I}=\int_{\Lambda_{r_{0}}^{{}^{\prime}}}\partial_{\nu}\tilde{u}_{0}\tilde{U}-\partial_{\nu}\tilde{U}\tilde{u}_{0}\mathrm{d}\sigma.$
Similar to the proof of Case(b) in Theorem 3.1, using Proposition 2.1, it yields that the right-hand side of (3.56) satisfies
$|RHS|\lesssim\tau^{-\frac{13}{6}}+\tau^{-(2+\alpha)}+\mathcal{O}(\tau^{-1}e^{-\frac{1}{2}\varsigma r_{0}\tau})+(1+\tau)(1+\tau^{-\frac{2}{3}})e^{-\varsigma r_{0}\tau}$ (3.57)
and the left-hand side of (3.56) fulfills that
$\overline{\eta}\int_{\tilde{\Gamma}_{r_{0}}^{\pm}}\tilde{x}_{1}e^{\rho\cdot\tilde{\mathbf{x}}}\mathrm{d}\sigma=\frac{\Gamma(2)\overline{\eta}}{\tau^{2}}\left(\frac{\cos\tilde{\vartheta}}{((\mathbf{d}+\mathrm{i}\mathbf{d}^{\perp})\cdot\hat{\tilde{\mathbf{x}}}_{1})^{2}}+\frac{1}{((\mathbf{d}+\mathrm{i}\mathbf{d}^{\perp})\cdot\hat{\tilde{\mathbf{x}}}_{2})^{2}}\right)+\mathcal{O}(\tau^{-1}e^{-\frac{1}{2}\lambda r_{0}\tau})$
where $\overline{\eta}=(\eta_{2}-\eta_{1})\partial_{1}U_{2}^{-}(\mathbf{0})$ and $\hat{\tilde{\mathbf{x}}}_{1}\in\tilde{\Gamma}_{r_{0}}^{+},\hat{\tilde{\mathbf{x}}}_{2}\in\tilde{\Gamma}_{r_{0}}^{-}$. Let $\mathbf{d}$ and $\mathbf{d}^{\perp}$ be defined in (2.14), then one has
$\left|\frac{\cos\tilde{\vartheta}}{((\mathbf{d}+\mathrm{i}\mathbf{d}^{\perp})\cdot\hat{\tilde{\mathbf{x}}}_{1})^{2}}+\frac{1}{((\mathbf{d}+\mathrm{i}\mathbf{d}^{\perp})\cdot\hat{\tilde{\mathbf{x}}}_{2})^{2}}\right|\not=0,\ \tilde{\vartheta}\in(0,\pi).$ (3.58)
By virtue of (3.56)-(3.58), one can directly derive:
$|\Gamma(2)\overline{\eta}|=0,$
as $\tau\to\infty$. Consequently, $\eta_{1}=\eta_{2}$ follows from $\partial_{1}U_{2}^{-}(\mathbf{0})\neq 0$.
Next, from (3.56), we have the following identity due to $\eta_{1}=\eta_{2}$:
$k^{2}(\tilde{q}_{2}^{(0)}$ $-\tilde{q}_{1}^{(0)})\partial_{1}U_{2}^{-}(\mathbf{0})\int_{\mathcal{S}_{r_{0}}^{{}^{\prime}}}\tilde{x}_{1}e^{\rho\cdot\tilde{\mathbf{x}}}\mathrm{d}\tilde{\mathbf{x}}=k^{2}(\tilde{q}_{1}^{(0)}-\tilde{q}_{2}^{(0)})\partial_{1}U_{2}^{-}(\mathbf{0})\int_{\mathcal{S}_{r_{0}}^{{}^{\prime}}}\tilde{x}_{1}\psi(\tilde{\mathbf{x}})e^{\rho\cdot\tilde{\mathbf{x}}}\mathrm{d}\tilde{\mathbf{x}}$
$+k^{2}(\tilde{q}_{1}^{(0)}-\tilde{q}_{2}^{(0)})\int_{\mathcal{S}_{r_{0}}^{{}^{\prime}}}\delta U_{2}^{-}\tilde{u}_{0}\mathrm{d}\tilde{\mathbf{x}}+k^{2}(\tilde{q}_{1}^{(1)}-\tilde{q}_{2}^{(1)})\int_{\mathcal{S}_{r_{0}}^{{}^{\prime}}}\tilde{x}_{1}U_{2}^{-}\tilde{u}_{0}\mathrm{d}\tilde{\mathbf{x}}$ (3.59)
$+k^{2}(\tilde{q}_{1}^{(2)}-\tilde{q}_{2}^{(2)})\int_{\mathcal{S}_{r_{0}}^{{}^{\prime}}}\tilde{x}_{2}U_{2}^{-}\tilde{u}_{0}\mathrm{d}\tilde{\mathbf{x}}+\mathcal{I},$
where $\mathcal{I}$ is given by (3.56). By (2.14) and direct computations, we can derive that the term on the left-hand side of (3) fulfils that
$LHS=-\frac{\Gamma(3)((\sin\tilde{\vartheta}+3\mathrm{i}\cos\tilde{\vartheta})e^{3\mathrm{i}\tilde{\vartheta}}-3\mathrm{i})}{8\tau^{3}e^{3\mathrm{i}\varphi}}+\mathcal{O}(\tau^{-1}e^{-\frac{1}{2}\lambda r_{0}\tau}),$ (3.60)
and the right-hand side of (3) satisfies
$|RHS|\lesssim\tau^{-\frac{41}{12}}+\tau^{-(\alpha+3)}+(1+\tau)(1+\tau^{-\frac{2}{3}})e^{-\varsigma r_{0}\tau}+\mathcal{O}(\tau^{-1}e^{-\frac{1}{2}\varsigma r_{0}\tau})$ (3.61)
for a sufficient large $\tau$. Additionally, since $\tilde{\vartheta}\in(0,\pi)$, we can deduce
$(\sin\tilde{\vartheta}+3\mathrm{i}\cos\tilde{\vartheta})e^{3\mathrm{i}\tilde{\vartheta}}-3\mathrm{i}=-2\sin 2\tilde{\vartheta}(\cos 2\tilde{\vartheta}+1)+2(\cos 2\tilde{\vartheta})(\cos 2\tilde{\vartheta}-1)\mathrm{i}\not=0.$ (3.62)
Subsequently, by combining (3) - (3.62), and multiplying both sides by $\tau^{3}$ before letting $\tau\to\infty$, we obtain
$k^{2}|(\tilde{q}_{2}^{(0)}-\tilde{q}_{1}^{(0)})\partial_{1}U_{2}^{-}(\mathbf{0})|=0,$
which indicates that $\tilde{q}_{1}^{(0)}=\tilde{q}_{2}^{(0)}$ since $|\partial_{1}U_{2}^{-}(\mathbf{0})|\not=0$.
Substituting $\tilde{q}_{1}^{(0)}=\tilde{q}_{2}^{(0)}$ into (3), we have
$k^{2}$ $(\tilde{q}_{2}^{(1)}-\tilde{q}_{1}^{(1)})\partial_{1}U_{2}^{-}(\mathbf{0})\int_{\mathcal{S}_{r_{0}}^{{}^{\prime}}}\tilde{x}_{1}^{2}e^{\rho\cdot\tilde{\mathbf{x}}}\mathrm{d}\tilde{\mathbf{x}}+k^{2}(\tilde{q}_{2}^{(2)}-\tilde{q}_{1}^{(2)})\partial_{1}U_{2}^{-}(\mathbf{0})\int_{\mathcal{S}_{r_{0}}^{{}^{\prime}}}\tilde{x}_{1}\tilde{x}_{2}e^{\rho\cdot\tilde{\mathbf{x}}}\mathrm{d}\tilde{\mathbf{x}}$
$=k^{2}(\tilde{q}_{1}^{(1)}-\tilde{q}_{2}^{(1)})\partial_{1}U_{2}^{-}(\mathbf{0})\int_{\mathcal{S}_{r_{0}}^{{}^{\prime}}}\tilde{x}_{1}^{2}\psi(\tilde{\mathbf{x}})e^{\rho\cdot\tilde{\mathbf{x}}}\mathrm{d}\tilde{\mathbf{x}}+k^{2}(\tilde{q}_{1}^{(1)}-\tilde{q}_{2}^{(1)})\int_{\mathcal{S}_{r_{0}}^{{}^{\prime}}}\tilde{x}_{1}\delta U_{2}^{-}\tilde{u}_{0}\mathrm{d}\tilde{\mathbf{x}}$
$+k^{2}(\tilde{q}_{1}^{(2)}-\tilde{q}_{2}^{(2)})\partial_{1}U_{2}^{-}(\mathbf{0})\int_{\mathcal{S}_{r_{0}}^{{}^{\prime}}}\tilde{x}_{1}\tilde{x}_{2}\psi(\tilde{\mathbf{x}})e^{\rho\cdot\mathbf{x}}\mathrm{d}\mathbf{x}$
$+k^{2}(\tilde{q}_{1}^{(2)}-\tilde{q}_{2}^{(2)})\int_{\mathcal{S}_{r_{0}}^{{}^{\prime}}}\tilde{x}_{2}\delta U_{2}^{-}\tilde{u}_{0}\mathrm{d}\tilde{\mathbf{x}}+\mathcal{I}.$ (3.63)
By Proposition 2.1, we know that the right-hand side of (3.63) satisfies
$|RHS|\lesssim\tau^{-\frac{53}{12}}+\tau^{-(4+\alpha)}+\mathcal{O}(\tau^{-1}e^{-\frac{1}{2}\varsigma r_{0}\tau})+(1+\tau)(1+\tau^{\frac{2}{3}})e^{-\varsigma r_{0}\tau},$ (3.64)
when $\tau$ is sufficient large. Moreover, we can deduce that the left-hand side of (3.63) is equal to
$\sigma_{1}$ $\frac{\Gamma(4)e^{4\mathrm{i}\varphi}}{\tau^{4}}\frac{1}{4}\left(-\frac{1}{2\mathrm{i}}+\frac{1}{3}(2\mathrm{i}\cos 2\vartheta-\sin 2\vartheta)\right)e^{-4\mathrm{i}\vartheta}\bigg{|}_{0}^{\tilde{\vartheta}}$
$\ +\sigma_{2}\frac{\Gamma(4)e^{4\mathrm{i}\varphi}}{\tau^{4}}\frac{1}{12}(2\mathrm{i}\sin 2\vartheta+\cos 2\vartheta)e^{-4\mathrm{i}\vartheta}\bigg{|}_{0}^{\tilde{\vartheta}}+\mathcal{O}(\tau^{-1}e^{-\frac{1}{2}\lambda r_{0}\tau}),$ (3.65)
when $\mathbf{d}^{\perp}$ is defined in (2.13) and
$\sigma_{1}$ $\frac{\Gamma(4)e^{-4\mathrm{i}\varphi}}{\tau^{4}}\frac{1}{4}\left(\frac{1}{2\mathrm{i}}-\frac{1}{3}(\sin 2\vartheta+2\mathrm{i}\cos 3\vartheta)\right)e^{4\mathrm{i}\vartheta}\bigg{|}_{0}^{\tilde{\vartheta}}$
$\ +\sigma_{2}\frac{\Gamma(4)e^{-4\mathrm{i}\varphi}}{\tau^{4}}\frac{1}{12}(\cos 2\vartheta-2\mathrm{i}\sin 2\vartheta)e^{4\mathrm{i}\vartheta}\bigg{|}_{0}^{\tilde{\vartheta}}+\mathcal{O}(\tau^{-1}e^{-\frac{1}{2}\lambda r_{0}\tau}),$ (3.66)
when $\mathbf{d}^{\perp}$ is defined in (2.14) , where $\sigma_{i}=k^{2}(\tilde{q}_{2}^{(i)}-\tilde{q}_{1}^{(i)})\partial_{1}U_{2}^{-}(\mathbf{0}),i=1,2.$ Combing (3.63)-(3.66) and following a similar argument for the Case (I), we derive that
$D\left(\begin{matrix}\sigma_{1}\\ \sigma_{2}\end{matrix}\right)=\mathbf{0},$
where
$D=\left(\begin{matrix}\frac{1}{4}\left(-\frac{1}{2\mathrm{i}}+\frac{1}{3}(2\mathrm{i}\cos 2\vartheta-\sin 2\vartheta)\right)e^{-4\mathrm{i}\vartheta}|_{0}^{\tilde{\vartheta}}&\frac{1}{12}(2\mathrm{i}\sin 2\vartheta+\cos 2\vartheta)e^{-4\mathrm{i}\vartheta}|_{0}^{\tilde{\vartheta}}\\ \frac{1}{4}\left(\frac{1}{2\mathrm{i}}-\frac{1}{3}(\sin 2\vartheta+2\mathrm{i}\cos 2\vartheta)\right)e^{4\mathrm{i}\vartheta}|_{0}^{\tilde{\vartheta}}&\frac{1}{12}(\cos 2\vartheta-2\mathrm{i}\sin 2\vartheta)e^{4\mathrm{i}\vartheta}|_{0}^{\tilde{\vartheta}}\end{matrix}\right).$
After careful and tedious calculations, we have
$\det(D)=\frac{i}{72}(5\cos^{3}2\tilde{\vartheta}-3\cos^{2}\tilde{\vartheta}-9\cos 2\tilde{\vartheta}+7)\not=0,\ \mbox{for}\ \tilde{\vartheta}\in(0,\pi).$
Therefore, it yields that $\sigma_{1}=\sigma_{2}$, which implies that $\tilde{q}_{1}^{(1)}=\tilde{q}_{2}^{(1)}$ and $\tilde{q}_{1}^{(2)}=\tilde{q}_{2}^{(2)}$.
The proof of Case (ii) is complete.
Case (iii): $\partial_{2}U_{2}^{-}(\mathbf{0})\not=0,\ \partial_{1}U_{2}^{-}(\mathbf{0})=\ U_{2}^{-}(\mathbf{0})=0.$ Direct calculations yield
$(\eta_{2}-\eta_{1})\partial_{2}U_{2}^{-}$ $(\mathbf{0})\int_{\tilde{\Gamma}_{r_{0}}^{\pm}}\tilde{x}_{2}e^{\rho\cdot\tilde{\mathbf{x}}}\mathrm{d}\sigma=(\eta_{1}-\eta_{2})\partial_{2}U_{2}^{-}(\mathbf{0})\int_{\tilde{\Gamma}_{r_{0}}^{\pm}}x_{2}\psi(\tilde{\mathbf{x}})e^{\rho\cdot\tilde{\mathbf{x}}}\mathrm{d}\sigma$
$+\int_{\mathcal{S}_{r_{0}}^{{}^{\prime}}}k^{2}(\tilde{q}_{1}-\tilde{q}_{2})(\partial_{2}U_{2}^{-}(\mathbf{0})\tilde{x}_{2}+\delta U_{2}^{-})\tilde{u}_{0}\mathrm{d}\tilde{\mathbf{x}}+\mathcal{I}.$ (3.67)
Comparing (3.67) with (3.56), then we observe that the the right-hand side of (3.67) can be estimated by (3.57).
Furthermore, by utilizing Proposition 2.1 again, we can deduce that the left-hand side of (3.67) satisfies
$LHS=(\eta_{2}-\eta_{1})\frac{\Gamma(2)\partial_{2}U_{2}^{-}(\mathbf{0})}{\tau^{2}}\left\{\frac{\sin\tilde{\vartheta}}{((\mathbf{d}+\mathrm{i}\mathbf{d}^{\perp})\cdot\hat{\tilde{\mathbf{x}}})^{2}}\right\}+\mathcal{O}(\tau^{-1}e^{-\frac{1}{2}\lambda r_{0}\tau}),$ (3.68)
where $\hat{\tilde{\mathbf{x}}}\in\tilde{\Gamma}_{r_{0}}^{+}$. Therefore, we can directly derive that $\eta_{1}=\eta_{2}$ since $\partial_{2}U_{2}^{-}(\mathbf{0})\neq 0$ and $\sin\tilde{\vartheta}\neq 0,\tilde{\vartheta}\in(0,\pi)$.
Consequently, we obtain
$k^{2}(\tilde{q}_{2}^{(0)}$ $-\tilde{q}_{1}^{(0)})\partial_{2}U_{2}^{-}(\mathbf{0})\int_{\mathcal{S}_{r_{0}}^{{}^{\prime}}}\tilde{x}_{2}e^{\rho\cdot\tilde{\mathbf{x}}}\mathrm{d}\tilde{\mathbf{x}}=k^{2}(\tilde{q}_{1}^{(0)}-\tilde{q}_{2}^{(0)})\partial_{2}U_{2}^{-}(\mathbf{0})\int_{\mathcal{S}_{r_{0}}^{{}^{\prime}}}\tilde{x}_{2}\psi(\tilde{\mathbf{x}})e^{\rho\cdot\tilde{\mathbf{x}}}\mathrm{d}\tilde{\mathbf{x}}$
$+k^{2}(\tilde{q}_{1}^{(0)}-\tilde{q}_{2}^{(0)})\int_{\mathcal{S}_{r_{0}}^{{}^{\prime}}}\delta U_{2}^{-}\tilde{u}_{0}\mathrm{d}\tilde{\mathbf{x}}+k^{2}(\tilde{q}_{1}^{(1)}-\tilde{q}_{2}^{(1)})\int_{\mathcal{S}_{r_{0}}^{{}^{\prime}}}\tilde{x}_{1}U_{2}^{-}\tilde{u}_{0}\mathrm{d}\tilde{\mathbf{x}}$
$+k^{2}(\tilde{q}_{1}^{(2)}-\tilde{q}_{2}^{(2)})\int_{\mathcal{S}_{r_{0}}^{{}^{\prime}}}{\tilde{x}}_{2}U_{2}^{-}\tilde{u}_{0}\mathrm{d}\tilde{\mathbf{x}}+\mathcal{I}.$ (3.69)
Through meticulous calculation, we ascertain that the left-hand side of (3) equals
$LHS=k^{2}(\tilde{q}_{2}^{(0)}-\tilde{q}_{1}^{(0)})\partial_{2}U_{2}^{-}(\mathbf{0})\frac{\Gamma(3)((\cos\tilde{\vartheta}-3\mathrm{i}\sin\tilde{\vartheta})e^{3\mathrm{i}\tilde{\vartheta}}-1)}{\tau^{3}e^{3\mathrm{i}\varphi}}+\mathcal{O}(\tau^{-1}e^{-\frac{1}{2}\lambda r_{0}\tau})$
and the right-hand side of (3) fulfills that (3.61). Given that
$(\cos\tilde{\vartheta}-3\mathrm{i}\sin\tilde{\vartheta})e^{3\mathrm{i}\tilde{\vartheta}}-1=-2\cos 2\tilde{\vartheta}(\cos 2\tilde{\vartheta}-1)+2\sin 2\tilde{\vartheta}(1-\cos 2\tilde{\vartheta})\mathrm{i}\not=0$
for $\tilde{\vartheta}\in(0,\pi)$ and $\partial_{2}U_{2}^{-}(\mathbf{0})\not=0$, we can derive that
$\tilde{q}_{2}^{(0)}-\tilde{q}_{1}^{(0)}=0.$
The proof that $\tilde{q}_{2}^{(0)}=\tilde{q}_{1}^{(0)}$ is complete.
Now we are going to prove $\tilde{q}_{1}^{(1)}=\tilde{q}_{2}^{(1)}$ and $\tilde{q}_{1}^{(2)}=\tilde{q}_{2}^{(2)}.$ It is easy to get
$\tilde{\sigma}_{1}\int_{\mathcal{S}_{r_{0}}^{{}^{\prime}}}$ $\tilde{x}_{1}\tilde{x}_{2}e^{\rho\cdot\tilde{\mathbf{x}}}\mathrm{d}\tilde{\mathbf{x}}+\tilde{\sigma}_{2}\int_{\mathcal{S}_{r_{0}}^{{}^{\prime}}}\tilde{x}_{2}^{2}e^{\rho\cdot\tilde{\mathbf{x}}}\mathrm{d}\tilde{\mathbf{x}}=-\tilde{\sigma}_{1}\int_{\mathcal{S}_{r_{0}}^{{}^{\prime}}}\tilde{x}_{1}\tilde{x}_{2}\psi(\tilde{\mathbf{x}})e^{\rho\cdot\tilde{\mathbf{x}}}\mathrm{d}\tilde{\mathbf{x}}$
$-\tilde{\sigma}_{2}\int_{\mathcal{S}_{r_{0}}^{{}^{\prime}}}\tilde{x}_{2}^{2}\psi(\tilde{\mathbf{x}})e^{\rho\cdot\tilde{\mathbf{x}}}\mathrm{d}\tilde{\mathbf{x}}+k^{2}(\tilde{q}_{1}^{(1)}-\tilde{q}_{2}^{(1)})\int_{\mathcal{S}_{r_{0}}^{{}^{\prime}}}\tilde{x}_{1}\delta U_{2}^{-}\tilde{u}_{0}\mathrm{d}\tilde{\mathbf{x}}$
$+k^{2}(q_{1}^{(2)}-\tilde{q}_{2}^{(2)})\int_{\mathcal{S}_{r_{0}}^{{}^{\prime}}}\tilde{x}_{2}\delta U_{2}^{-}\tilde{u}_{0}\mathrm{d}\tilde{\mathbf{x}}+\mathcal{I},$ (3.70)
here, $\tilde{\sigma}_{i}=k^{2}(\tilde{q}_{2}^{(i)}-\tilde{q}_{1}^{(i)})\partial_{2}U_{2}^{-}(\mathbf{0}),i=1,2$. It is clear that the right-hand side term of (3.70) satisfies (3.64).
Moreover, the left-hand side term of (3.70) fulfills that
$LHS$ $=\tilde{\sigma}_{1}\frac{\Gamma(4)e^{4\mathrm{i}\varphi}}{\tau^{4}}\frac{1}{12}(2\mathrm{i}\sin 2\vartheta+\cos 2\vartheta)e^{-4\mathrm{i}\vartheta}\bigg{|}_{0}^{\tilde{\vartheta}}$
$+\tilde{\sigma}_{2}\frac{\Gamma(4)e^{4\mathrm{i}\varphi}}{\tau^{4}}\frac{1}{4}\left(-\frac{1}{2\mathrm{i}}-\frac{1}{6}(2\mathrm{i}\cos 2\vartheta-\sin 2\vartheta)\right)e^{-4\mathrm{i}\vartheta}\bigg{|}_{0}^{\tilde{\vartheta}}+\mathcal{O}(\tau^{-1}e^{-\frac{1}{2}\lambda r_{0}\tau})$
when $\mathbf{d}^{\perp}=(-\sin\varphi,\cos\varphi)^{\top}$ and
$LHS$ $=\tilde{\sigma}_{1}\frac{\Gamma(4)e^{-4\mathrm{i}\varphi}}{\tau^{4}}\frac{1}{12}(\cos\vartheta-\cos 2\mathrm{i}\sin\vartheta)e^{4\mathrm{i}\vartheta}\bigg{|}_{0}^{\tilde{\vartheta}}$
$+\tilde{\sigma}_{2}\frac{\Gamma(4)e^{-4\mathrm{i}\varphi}}{\tau^{4}}\frac{1}{4}\left(\frac{1}{2\mathrm{i}}+\frac{1}{6}(2\mathrm{i}\cos 2\vartheta+\sin 2\vartheta)\right)e^{4\mathrm{i}\vartheta}\bigg{|}_{0}^{\tilde{\vartheta}}+\mathcal{O}(\tau^{-1}e^{-\frac{1}{2}\lambda r_{0}\tau})$
when $\mathbf{d}^{\perp}=(\sin\varphi,-\cos\varphi)^{\top}$. For a sufficiently large $\tau$, it holds that
$E\left(\begin{matrix}\tilde{\sigma}_{1}\\ \tilde{\sigma}_{2}\end{matrix}\right)=\mathbf{0},$
where
$E=\left(\begin{matrix}\frac{1}{12}(2\mathrm{i}\sin 2\vartheta+\cos 2\vartheta)e^{-4\mathrm{i}\vartheta}|_{0}^{\tilde{\vartheta}}&\frac{1}{4}\left(-\frac{1}{2\mathrm{i}}-\frac{1}{3}(2\mathrm{i}\cos 2\vartheta-\sin 2\vartheta)\right)e^{-4\mathrm{i}\vartheta}|_{0}^{\tilde{\vartheta}}\\ \frac{1}{12}(\cos 2\vartheta-2\mathrm{i}\sin\vartheta)e^{4\mathrm{i}\vartheta}|_{0}^{\tilde{\vartheta}}&\frac{1}{4}\left(\frac{1}{2\mathrm{i}}+\frac{1}{3}(2\mathrm{i}\cos 2\vartheta+\sin 2\vartheta)\right)e^{4\mathrm{i}\vartheta}|_{0}^{\tilde{\vartheta}}\end{matrix}\right).$
Following meticulous calculations, it yields that
$\det(E)=\frac{-\mathrm{i}}{144}(\cos\tilde{\vartheta}-1)^{3}\neq 0,\ \tilde{\vartheta}\in(0,\pi)$
which indicates that $\tilde{\sigma}_{1}=\tilde{\sigma}_{2}=0$, implying $\tilde{q}_{1}^{(i)}=\tilde{q}_{2}^{(i)}$ for $i=1,2$ due to $\partial_{2}U_{2}^{-}(\mathbf{0})\neq 0$. Thus, the proof for Case (ii) is complete.
Finally, we shall prove $u_{1}^{-}=u_{2}^{-}$ in $\mathcal{S}_{r_{0}}$. By the fact that $\eta_{1}=\eta_{2}$ and $q_{1}=q_{2}$ and the boundary condition, then we have
$\begin{cases}\Delta w+k^{2}qw=0\hskip 11.38092pt\mbox{in}\ \mathcal{S}_{r_{0}},\\ w=\partial_{\nu}w=0\hskip 25.6073pt\mbox{on}\ \Gamma_{r_{0}}^{\pm},\end{cases}$
where $w=u_{2}^{-}-u_{1}^{-}.$ Using Holmgren’s principle (cf.[18]), we have $w=0$ in $\mathcal{S}_{r_{0}}$.
The proof is complete. ∎
Remark 3.5.
In the proof of Lemma 3.2, the assumption that the constant conductive boundary parameter is real and coefficients of the linear polynomial refractive index are real is only necessary for Case (i) of Case (I). However, if the conductive boundary parameter is a complex constant and the refractive index is a linear polynomial with complex coefficients, we speculate that the conclusion of Lemma 3.2 still holds in a generic scenario. In fact, in the final step of proving Case (iii), we demonstrated that the determinant of the coefficient matrix in (3.55) is quadratic, with the matrix $\mathcal{A}$ being positive. To address this scenario, let us consider a complex nonzero vector $(\partial_{1}U_{2}^{-}(\mathbf{0}),\ \partial_{2}U_{2}^{-}(\mathbf{0}))^{\top}$ and define
${\bf a}=\Re\left(\left(\begin{matrix}\partial_{1}U_{2}^{-}(\mathbf{0})\\ \partial_{2}U_{2}^{-}(\mathbf{0})\end{matrix}\right)\right),\quad{\bf b}=\Im\left(\left(\begin{matrix}\partial_{1}U_{2}^{-}(\mathbf{0})\\ \partial_{2}U_{2}^{-}(\mathbf{0})\end{matrix}\right)\right).$
Then $\det(Q)=0$ is equivalent to
${\bf a}^{\top}\mathcal{A}{\bf a}={\bf b}^{\top}\mathcal{A}{\bf b},\quad{\bf a}^{\top}\mathcal{A}{\bf b}=0.$ (3.71)
By defining the $\mathcal{A}$-inner product as $(\bf a,\bf b)_{\mathcal{A}}={\bf a}^{\top}\mathcal{A}{\bf b}$ and the $\mathcal{A}$-norm as $\|\bf a\|_{\mathcal{A}}=\sqrt{{\bf a}^{\top}\mathcal{A}{\bf a}}$, (3.71) can be equivalently expressed as
$\|\bf a\|_{\mathcal{A}}=\|\bf b\|_{\mathcal{A}}$ and $({\bf a},{\bf b})_{\mathcal{A}}=0.$ (3.72)
The entries of $\mathcal{A}$ depend solely on the opening angle of $\mathcal{S}_{r_{0}}$. It is important to note that $\bf a$ and $\bf b$ represent the real and imaginary parts of $(\partial_{1}U_{2}^{-}(\mathbf{0})\ \partial_{2}U_{2}^{-}(\mathbf{0}))^{\top}$, respectively, where $u_{2}^{-}(\mathbf{x})=U_{2}^{-}(M(\vartheta_{m})\mathbf{x})$ is a solution of (3.30). When applying Lemma 3.2 to prove Theorems 3.2 and 3.3, $u_{2}^{-}(\mathbf{x})=U_{2}^{-}(M(\vartheta_{m})\mathbf{x})$ represents the total wave in (1.7). From a physical perspective, (3.72) is rarely encountered, as it necessitates a peculiar algebraic relationship between the gradient of the total wave and $\mathcal{A}$. Consequently, $Q$ is generically nonsingular, enabling the unique determination of the linear polynomial reflective index with complex coefficients through a single far-field measurement, as demonstrated in Theorems 3.2 and 3.3. However, a rigorous proof of this speculation is fraught with significant difficulties and hence we defer it to our future study. Nevertheless, we would like to emphasize that the case with $q$ and $\eta$ being real-valued in the current article are physically unobjectionable.
4. The proofs of Theorems 3.2 and 3.3
Proof of Theorem 3.2.
In accordance with Theorem 3.2, the shape of the conductive medium scatterer $\Omega$ and the corresponding polygonal-cell structures $\Sigma_{i},i=1,\dots,N$, are known in advance. We establish uniqueness results for determining the physical parameters of the refractive index and the conductive boundary parameter through a proof by contradiction. Assume that $\tilde{\eta}_{1}\not=\tilde{\eta}_{2}$ or there is an index $i_{0}$ such that $q_{i_{0},1}\not=q_{i_{0},2}$. From Definition 3.2, we know that there exists a vertex $\mathbf{x}_{p_{0}}\in\Sigma_{i_{0}}$ intersected by two adjacent edges $\Gamma^{\pm}$ of $\Omega$. Since $-\Delta$ is invariant under rigid motion, without loss of generality, we assume that $\mathbf{x}_{p_{0}}$ coincides with the origin. Let $r_{0}\in\mathbb{R}_{+}$ be sufficiently small such that $\mathcal{S}_{r_{0}}=\Omega\cap B_{r_{0}}(\mathbf{0})$ and $\Gamma_{r_{0}}^{\pm}=\partial\Omega\cap B_{r_{0}}(\mathbf{0})$. Denote $W=\mathbb{R}^{2}\setminus\overline{\Omega},$ then we have $\Gamma_{r_{0}}^{\pm}\subset\partial W$, see Figure 3 for a schematic illustration.
Figure 3. Schematic illustration of the polygonal-cell structure.
Recall that $u_{1}$ and $u_{2}$ be the total wave fields associated with the admissible conductive polygonal-cell medium scatterers $(\Omega;q_{\ell},\eta_{\ell}),\ell=1,2$. Since $u_{1}^{\infty}(\mathbf{\hat{x}};u^{i})=u_{2}^{\infty}(\mathbf{\hat{x}};u^{i})$ for all $\mathbf{\hat{x}}\in\mathbb{S}^{1},$ by Rellich lemma and the unique continuation principle, we have
$u_{1}=u_{2}\ \mbox{in}\ B_{r_{0}}\setminus\overline{\mathcal{S}_{r_{0}}}\subset W.$
Utilizing the conductive transmission boundary conditions in (1.7), one can claim that
$\begin{cases}u_{1}^{+}=u_{1}^{-},\ \partial_{\nu}u_{1}^{-}=\partial_{\nu}u_{1}^{+}+\tilde{\eta}_{1}u_{1}^{+},\ \mbox{on}\ \Gamma_{r_{0}}^{\pm},\\ u_{2}^{+}=u_{2}^{-},\ \partial_{\nu}u_{2}^{-}=\partial_{\nu}u_{2}^{+}+\tilde{\eta}_{2}u_{2}^{+},\ \mbox{on}\ \Gamma_{r_{0}}^{\pm}.\end{cases}$
Moreover, it holds that
$\begin{cases}\Delta u_{1}^{-}+k^{2}q_{i_{0},1}u_{1}^{-}=0,\ \hskip 14.22636pt\mbox{in}\ \mathcal{S}_{r_{0}},\\ \Delta u_{1}^{+}+k^{2}u_{1}^{+}=0,\hskip 36.98866pt\mbox{in}\ B_{r_{0}}\setminus\overline{\mathcal{S}_{r_{0}}},\\ \Delta u_{2}^{-}+k^{2}q_{i_{0},2}u_{2}^{-}=0,\ \hskip 14.22636pt\mbox{in}\ \mathcal{S}_{r_{0}},\\ \Delta u_{2}^{+}+k^{2}u_{2}^{+}=0,\hskip 36.98866pt\mbox{in}\ B_{r_{0}}\setminus\overline{\mathcal{S}_{r_{0}}}.\end{cases}$
Recall that $\Omega$ is an admissible conductive polygonal-cell medium scatterer. Then by virtue of Lemma 3.2, we have
$q_{i_{0},1}=q_{i_{0},2}\ \mbox{and}\ \tilde{\eta}_{1}=\tilde{\eta}_{2},$
which is a contradiction.
The proof is completed. ∎
The proof of Theorem 3.3..
We prove this theorem by mathematical induction. By Theorem 3.1, we have $\partial\Omega_{1}=\partial\Omega_{2}$, which indicates that $\partial\Sigma_{1,1}=\partial\Sigma_{1,2}.$ Taking a similar argument as proving Theorem 3.2, it can be shown that $q_{1,1}=q_{1,2}$ and $\eta_{1,1}=\eta_{1,2}$. Suppose that there exists an index $n_{0}\in\mathbb{N}_{+}\setminus\{1\}$ such that
$\partial\Sigma_{i,1}=\partial\Sigma_{i,2},\ q_{i,1}=q_{i,2},\ \eta_{i,1}=\eta_{i,2},\ i=1,\dots,n_{0}-1.$
With the help of Lemma 3.2, we can recursively obtain that
$u_{i,1}=u_{i,2}\ \mbox{in}\ \mathcal{D}_{i}=\Sigma_{i}\setminus\overline{\Sigma_{i+1}},\ i=1,\dots,n_{0}-1$ (4.1)
by making use of $u_{1}^{\infty}(\mathbf{\hat{x}};u^{i})=u_{2}^{\infty}(\mathbf{\hat{x}};u^{i})$, where $u_{i,1}=u_{1}|_{\mathcal{D}_{i}}$ and $u_{i,2}=u_{2}|_{\mathcal{D}_{i}}$, with $u_{1}$ and $u_{2}$ being the total wave fields associated with the admissible conductive polygonal-nest medium scatterers $(\Omega_{\ell};q_{\ell},\eta_{\ell}),\ell=1,2.$ We divide the rest proof into two parts.
Part 1. We first prove that $\partial\Sigma_{n_{0},1}=\partial\Sigma_{n_{0},2}$. According to (4.1), using the similar argument in proving Theorem 3.1, one can prove $\partial\Sigma_{n_{0},1}=\partial\Sigma_{n_{0},2}$ directly by contradiction.
Figure 4. Schematic illustration of the polygonal-nest structure.
Part 2. In the following we prove that
$q_{n_{0},1}=q_{n_{0},2},\ \eta_{n_{0},1}=\eta_{n_{0},2}.$ (4.2)
Let $\mathbf{x}_{p_{0}}$ be a vertex of $\Sigma_{n_{0},1}$. Without loss of generality, we may assume that $\mathbf{x}_{p_{0}}=\mathbf{0}$ since $-\Delta$ is invariant under rigid motion. For sufficiently small $r_{0}>0$, assume that $\mathcal{S}_{r_{0}}\Subset\mathcal{D}_{n_{0},1}=\Sigma_{n_{0},1}\setminus\overline{\Sigma_{n_{0}+1,1}}$; refer to Figure 4 for a schematic illustration. Therefore, it yields that
[…]
It is obvious that $u_{n_{0}-1,1}^{+}=u_{n_{0}-1,2}^{+}$ by virtue of (4.1). In view of (4.3), by Lemma 3.2, one has (4.2).
Finally, we can prove $N_{1}=N_{2}$ through a proof by contradiction. In fact, if $N_{1}\not=N_{2}$, we assume, without loss of generality, that $N_{1}>N_{2}$. Therefore, there exists a corner $\Sigma_{N_{1}+1,1}$ with a vertex $\mathbf{x}_{p_{0}}$ inside $\Sigma_{N_{2},2}$. Using a similar argument in proving Theorem 3.1, we can prove that either the total wave field or the gradient of the total wave field vanishes at the vertex of this corner, which contradicts to the admissible condition of a conductive polygonal-nest medium scatterer in Definition 3.5.
The proof is complete. ∎
Acknowledgements
The work of H. Diao is supported by by National Natural Science Foundation of China (No. 12371422) and the Fundamental Research Funds for the Central Universities, JLU (No. 93Z172023Z01). The work of X. Fei is supported by NSFC/RGC Joint Research Grant No. 12161160314. The work of H. Liu is supported by the NSFC/RGC Joint Research Scheme, N_CityU101/21; ANR/RGC Joint Research Scheme, A_CityU203/19; and the Hong Kong RGC General Research Funds (projects 11311122, 12301420 and 11300821).

### v3

@@
 Abstract.
-In this paper, we derive a novel Unique Continuation Principle (UCP) for a system of second-order elliptic PDEs system and apply it to investigate inverse problems in conductive scattering.
+In this paper, we derive a novel Unique Continuation Principle (UCP) for a system of second-order elliptic PDEs and apply it to investigate inverse problems in conductive scattering.
 The UCP relaxes the typical assumptions imposed on the domain or boundary with certain interior transmission conditions.
@@
 $\mathcal{S}=\{\mathbf{x}\in\mathbb{R}^{2}~{}|~{}\vartheta_{m}<\arg(x_{1}+\mathrm{i}x_{2})<\vartheta_{M}\},$ (1.1)
-where $-\pi\leq\vartheta_{m}<\vartheta_{M}<\pi$, $\vartheta_{M}-\vartheta_{m}\in(0,\pi)$, and $\mathrm{i}=\sqrt{-1}$. Additionally, the two boundary lines of $\mathcal{S}$ are given by
+where $-\pi\leq\vartheta_{m}<\vartheta_{M}<\pi$, $\vartheta_{M}-\vartheta_{m}\in(0,\pi)$ and $\mathrm{i}=\sqrt{-1}$. Additionally, the two boundary lines of $\mathcal{S}$ are given by
 $\Gamma^{+}=\{\mathbf{x}\in\mathbb{R}^{2}~{}|~{}\mathbf{x}=(r\cos\vartheta_{M},r\sin\vartheta_{M})^{\top},r>0\},$
@@
 If $\frac{\vartheta_{M}-\vartheta_{m}}{\pi}$ is a rational number, $\mathcal{S}_{r_{0}}$ is termed a rational corner; otherwise, it is an irrational corner.
-Theorem 1.1 shows a novel type of UCP result associated with a PDE system described by (1.3). The detailed proof of Theorem 1.1 is given in Section 2.
+Theorem 1.1 shows a novel type of UCP result associated with a PDEs system described by (1.3). The detailed proof of Theorem 1.1 is given in Section 2.
 Theorem 1.1.
@@
 The line segements $\Gamma^{\pm}_{r_{0}}$ are defined in (1.2).
-Consider $\mathbf{u}=(v,w)^{\top}\in H^{1}(D)\times H^{1}(D)$ solving the PDE system:
+Consider $\mathbf{u}=(v,w)^{\top}\in H^{1}(D)\times H^{1}(D)$ solving the PDEs system:
 $\begin{cases}\mathcal{P}\mathbf{u}=0\quad\mbox{in}\quad D,\\ \mathcal{B}\mathbf{u}=0\quad\mbox{on}\quad\Gamma^{\pm}_{r_{0}},\end{cases}$ (1.3)
@@
 where $V=q-1$ and $\Gamma=\partial\Omega$.
-On the other hand, given the bounded Lipschitz domain $\Omega$, $V\in L^{\infty}(\Omega)$ and the boundary parameter $\eta\in L^{\infty}(\partial\Omega)$, consider the PDE system (1.11).
+On the other hand, given the bounded Lipschitz domain $\Omega$, $V\in L^{\infty}(\Omega)$ and the boundary parameter $\eta\in L^{\infty}(\partial\Omega)$, consider the PDEs system (1.11).
 If there exists a nontrivial pair of solutions $(w,v)$ to (1.11) associated with $k$, then $k$ is referred to as a transmission eigenvalue and $(w,v)$ is the corresponding pair of transmission eigenfunctions of (1.11).
@@
 We will discuss this aspect in more detail in Section 2.
-To the best of our knowledge, this represents a novel type of UCP for second-order elliptic PDE systems, playing a crucial role in investigating invisibility and uniqueness results in inverse conductive medium scattering problems.
+To the best of our knowledge, this represents a novel type of UCP for second-order elliptic PDEs systems, playing a crucial role in investigating invisibility and uniqueness results in inverse conductive medium scattering problems.
 For the main contribution b), we can substitute the technical condition on the a-priori information in [21], which requires the non-vanishing of the total wave at the corner points, with an irrationality condition related to the interior angles of the conductive polygonal medium $\Omega$.
@@
 Furthermore, using the novel UCP established in Theorem 1.1, we can show that if a conductive medium scatterer $\Omega$ has a convex irrational corner on the boundary, where the reflective index $q$ has $H^{2}$ regularity near the corner, then $\Omega$ radiates any incident wave $u^{i}$ defined by (1.5) or (1.6).
-For the main contribution c), we can uniquely determine the refractive index in linear polynomial form and the constant conductive boundary parameter of the medium scatterer, whereas in [15] both the refractive index and the conductive parameter are constants.
-Moreover, we can relax the admissibility conditions in [15], which requires the non-vanishing of the total wave field at the corner points, to requiring the non-vanishing of either the total wave field or the gradient of the total wave field.
+For the main contribution c), we establish a local uniqueness result for determining the shape of the polygonal-nest or polygonal-cell conductive medium scatterer.
+This is achieved by proposing a novel type of admissible condition: the non-anishing assumption of either the total wave field or its gradient.
+In contrast to the previous admissibility conditions in [15], which only required the non-vanishing of the total wave, the new conditions impose more requirements on the total wave.
+From the perspective of physical intuition, these new admissible conditions are more likely to be violated.
+As a result, the newly established uniqueness result is more applicable to general physical scenarios.
+For more in-depth discussions, please refer to Remarks 3.1 and 3.2.
+Furthermore, we can uniquely determine the refractive index in linear polynomial form and constant conductive boundary parameter of the medium scatterer, whereas in [15] both the refractive index and the conductive parameter are constants.
 In fact, the complexity inherent in the conductive polygonal-nest and polygonal-cell medium scatterer necessitates a detailed study of the singular behavior of the solution to a coupled conductive transmission PDE problem near a corner; see Lemma 3.2.
-Specifically, given that the refractive index is assumed to be a linear polynomial, it becomes crucial to demonstrate the Hölder regularity of the gradient of the total wave as derived from the direct problem (1.7), particularly near corners; see Lemma 3.1.
-With this regularity established, we propose another admissible condition, which describes the non-vanishing property of the gradient of the total wave field at the corner point.
-This admissible condition is generic, which can be fulfilled in more general physical scenarios; see more detailed discussions in Remarks 3.1 and 3.2.
-Under this generic admissible condition, we can then proceed recursively to determine the interior polygonal-nest structure.
+We shall proceed recursively to identify the refractive index and constant conductive boundary parameter by leveraging the so-called complex geometrical optics (CGO) solution and the corresponding total wave field to give a dedicated microlocal analysis near the underlying corner.
 The remainder of the paper is organized as follows.
@@
 In this section, we aim to provide the proofs of Theorems 1.1-1.3.
-Specifically, we present a novel UCP result for an elliptic second-order PDE system and establish both local and global uniqueness results for the shape of an admissible complex conductive polygonal medium scatterer $\Omega$, utilizing only a single far-field measurement.
+Specifically, we present a novel UCP result for an elliptic second-order PDEs system and establish both local and global uniqueness results for the shape of an admissible complex conductive polygonal medium scatterer $\Omega$, utilizing only a single far-field measurement.
 To prove Theorem 1.1, we first introduce some auxiliary lemmas.
@@
 Lemma 2.2.
-[18, Section 3.4] Suppose that $v$ is a solution to Helmholtz equation $\Delta+\gamma_{1}v=0$ in $\mathcal{S}_{r_{0}}$, where $\gamma_{1}\in\mathbb{R}_{+}$ and $\mathcal{S}_{r_{0}}$ is defined in (1.2).
+[18, Section 3.4] Suppose that $v$ is a solution to Helmholtz equation $(\Delta+\gamma_{1})v=0$ in $\mathcal{S}_{r_{0}}$, where $\gamma_{1}\in\mathbb{R}_{+}$ and $\mathcal{S}_{r_{0}}$ is defined in (1.2).
 If $v$ is analytic in $\mathcal{S}_{r_{0}}$, then $v$ has the following spherical wave expansion in polar coordinates around the origin:
@@
 $\lesssim|\eta|\sum_{n=1}^{\infty}\sum_{p=0}^{\infty}\frac{(-1)^{p}(\sqrt{\gamma_{1}})^{n+2p}(|a_{n}|+|b_{n}|)}{2^{n+2p}p!(n+p)!}\left(\tau^{-(n+2p+\frac{7}{6})}+\mathcal{O}(\tau^{-\frac{7}{6}}e^{-\frac{1}{2}\varsigma r_{0}\tau})\right)$
-$\lesssim\tau^{-\frac{13}{6}}+\mathcal{O}(\tau^{-\frac{7}{6}}e^{-\frac{1}{2}\varsigma r_{0}\tau}),$
+$\lesssim\tau^{-\frac{13}{6}}+\mathcal{O}(\tau^{-\frac{7}{6}}e^{-\frac{1}{2}\varsigma r_{0}\tau})$
 and
@@
 where
-$A=\left(\begin{matrix}e^{-\mathrm{i}\vartheta_{m}}+e^{-\mathrm{i}\vartheta_{M}}&e^{-3\mathrm{i}\vartheta_{m}}+e^{-3\mathrm{i}\vartheta_{M}}\\ e^{3\mathrm{i}\vartheta_{m}}+e^{-3\mathrm{i}\vartheta_{M}}&e^{\mathrm{i}\vartheta_{m}}+e^{\mathrm{i}\vartheta_{M}}.\end{matrix}\right)$
+$A=\left(\begin{matrix}e^{-\mathrm{i}\vartheta_{m}}+e^{-\mathrm{i}\vartheta_{M}}&e^{-3\mathrm{i}\vartheta_{m}}+e^{-3\mathrm{i}\vartheta_{M}}\\ e^{3\mathrm{i}\vartheta_{m}}+e^{-3\mathrm{i}\vartheta_{M}}&e^{\mathrm{i}\vartheta_{m}}+e^{\mathrm{i}\vartheta_{M}}\end{matrix}\right).$
 Since $\vartheta_{M}-\vartheta_{m}\not=\frac{\pi}{2}$ by noting $\mathcal{S}_{r_{0}}$ is irrational, after some calculations, it shows that
@@
 Here, $u^{\infty}(\mathbf{\hat{x}};u^{i})$ is the far-field pattern given by (1.8), associated with the total wave $u$, where $u$ is the solution of (1.7) corresponding to the conductive polygonal-cell medium scatterer $\bigcup_{i=1}^{N}(\Sigma_{i};q_{i},\tilde{\eta})$.
-In order to introduce the admissible class for studying the inverse problems, we need to study the Hölder-regularity of the total wave field $u$ which is associated with the scattering problem (1.7) near the corner $\mathcal{S}_{r_{0}}$.
-Then we have the following lemma, which shows that there holds that $u^{-}\in C^{1,\alpha}(\overline{\mathcal{S}_{r_{0}}})$, where $\alpha\in(0,1)$.
+In order to introduce the admissible class for studying the inverse problems of the unique identification of the shape, we need to study the Hölder-regularity of the solution of the coupled PDEs system (3.7) near the corner, where $u_{1}$ and $u_{2}$ are the total wave field corresponding to the two different scatterers $\Omega_{1}$ and $\Omega_{2}$, respectively.
+Indeed, the PDEs system (3.7) is derived from the proceed of proving the unique determination of the shape of the scatterer by contradiction (see (3.12) in the proof of Theorem 3.1).
 Lemma 3.1.
@@
 Denote $\Gamma^{\pm}_{2r_{0}}=\Gamma^{\pm}\cap B_{2r_{0}}$, where $\Gamma^{\pm}$ is the boundary line of $\mathcal{S}$.
-Suppose that $u\in H^{1}(B_{2r_{0}})$ with $\ u^{+}=u|_{B_{2r_{0}}\setminus\overline{\mathcal{S}_{2r_{0}}}},\ u^{-}=u|_{\mathcal{S}_{2r_{0}}}$ satisfies that
+Suppose that $u_{1},u_{2}\in H^{1}(\mathcal{S}_{2r_{0}})$ satisfy
 […]
 where $k$ is a positive constant and $q$ is in the form of (3.1).
-Assume that $u^{-}$ and $u^{+}$ are real analytic in $\mathcal{S}_{2r_{0}}$ and $B_{2r_{0}}\setminus\overline{\mathcal{S}_{2r_{0}}}$, respectively.
-Then there exist $\alpha\in(0,1)$ such that $u^{-}\in C^{1,\alpha}(\overline{\mathcal{S}_{r_{0}}}).$
+If $u_{2}$ is real analytic in $\mathcal{S}_{2r_{0}}$, then $u_{1}\in C^{1,\alpha}(\overline{\mathcal{S}_{r_{0}}})$, where $\alpha\in(0,1)$.
 Before proving Lemma 3.1, we first introduce a triangle $V_{2r_{0}}$ described below, which will be used in a subsequent argument.
@@
 The proof of Lemma 3.1.
-Denote $\overline{u}$ by the analytic extension of $u^{+}|_{B_{2r_{0}}\setminus\overline{\mathcal{S}}_{2r_{0}}}$ in $B_{2r_{0}}$ since it is real analytic in $B_{2r_{0}}\setminus\overline{\mathcal{S}}_{2r_{0}}$.
-By virtue of the boundary condition of (3.7), one has $\overline{u}=u^{-}=u^{+}\ \mbox{on}\ \Gamma_{2r_{0}}^{\pm}.$
-Let $w=u^{-}-\overline{u},$ then from (3.7) it can be directly calculated that
+Let $w=u_{1}-u_{2},$ then from (3.7) it can be directly calculated that
 $\begin{cases}\Delta w=f,\hskip 15.6491pt\mbox{in}\ V_{2r_{0}},\\ w=\varphi,\hskip 22.76228pt\mbox{on}\ \Lambda^{{}^{\prime}}_{2r_{0}},\\ w=0,\hskip 24.18501pt\mbox{on}\ \Gamma_{2r_{0}}^{\pm},\end{cases}$
@@
       • (I)
-        $u(\mathbf{x}_{p})\not=0,$ (3.8)
+        $\lim_{\rho\to 0}\frac{1}{m(B(\mathbf{x}_{p},\rho))}\int_{B(\mathbf{x}_{p},\rho)}|u(\mathbf{x}_{p})|\not=0,$ (3.8)
       • (II)
-        $\nabla u(\mathbf{x}_{p})\neq 0,\ u(\mathbf{x}_{p})=0,$ (3.9)
-      where $\mathbf{x}_{p}\in{\mathcal{V}}(\Sigma_{i})$ and ${\mathcal{V}}(\Sigma_{i})=\{\mathbf{x}_{p,i}\}_{p=1}^{\ell_{i}}(3\leq\ell_{i}\in\mathbb{Z}_{+})$ is the vertex set of the convex polygon $\Sigma_{i}$.
+        if
+        $\lim_{\rho\to 0}\frac{1}{m(B(\mathbf{x}_{p},\rho))}\int_{B(\mathbf{x}_{p},\rho)}|u(\mathbf{x}_{p})|\mathrm{d}\mathbf{x}=0,\ \mbox{then}$
+        $\lim_{\rho\to 0}\frac{1}{m(B(\mathbf{x}_{p},\rho))}\int_{B(\mathbf{x}_{p},\rho)}|\nabla u(\mathbf{x}_{p})|\mathrm{d}\mathbf{x}\neq 0,$ (3.9)
+      where $m(B(\mathbf{x}_{p},\rho))$ is the measure of $B(\mathbf{x}_{p},\rho)$, $\mathbf{x}_{p}\in{\mathcal{V}}(\Sigma_{i})$ and ${\mathcal{V}}(\Sigma_{i})=\{\mathbf{x}_{p,i}\}_{p=1}^{\ell_{i}}(3\leq\ell_{i}\in\mathbb{Z}_{+})$ is the vertex set of the convex polygon $\Sigma_{i}$.
 Here, $\{\Sigma_{i}\}_{i=1}^{N}$ is the polygonal-nest structure of $\Omega$ described by Definition 3.1.
     2. (2)
-      If the condition (3.9) is satisfied, it is additionally required that $\angle(\Gamma_{p}^{-},\Gamma_{p}^{+})\in(\frac{\pi}{3},\frac{2\pi}{3})\setminus\{\frac{\pi}{2}\}$, where $\Gamma_{p}^{\pm}$ represent two edges of $\Sigma_{i}$, and $\Gamma_{p}^{-}\cap\Gamma_{p}^{+}=\mathbf{x}_{p}$ with $\angle(\Gamma_{p}^{-},\Gamma_{p}^{+})$ being the intersection angle between $\Gamma_{p}^{-}$ and $\Gamma_{p}^{+}$.
+      If the condition (3.9) is satisfied, it is additionally required that $\angle(\Gamma_{p}^{-},\Gamma_{p}^{+})\in(0,\pi)\setminus\{\frac{\pi}{2}\}$, where $\Gamma_{p}^{\pm}$ represent two edges of $\Sigma_{i}$, and $\Gamma_{p}^{-}\cap\Gamma_{p}^{+}=\mathbf{x}_{p}$ with $\angle(\Gamma_{p}^{-},\Gamma_{p}^{+})$ being the intersection angle between $\Gamma_{p}^{-}$ and $\Gamma_{p}^{+}$.
   • Polygonal-Cell Admissibility:
@@
     2. (2)
-      Similarly, if the condition (3.9) is satisfied, it is additionally required that $\angle(\Gamma_{p}^{-},\Gamma_{p}^{+})\in(\frac{\pi}{3},\frac{2\pi}{3})\setminus\{\frac{\pi}{2}\}$, where $\Gamma_{p}^{\pm}$ represent two edges of $\Omega$, and $\Gamma_{p}^{-}\cap\Gamma_{p}^{+}=\mathbf{x}_{p}$ with $\angle(\Gamma_{p}^{-},\Gamma_{p}^{+})$ being the intersection angle between $\Gamma_{p}^{-}$ and $\Gamma_{p}^{+}$.
+      Similarly, if the condition (3.9) is satisfied, it is additionally required that $\angle(\Gamma_{p}^{-},\Gamma_{p}^{+})\in(0,\pi)\setminus\{\frac{\pi}{2}\}$, where $\Gamma_{p}^{\pm}$ represent two edges of $\Omega$, and $\Gamma_{p}^{-}\cap\Gamma_{p}^{+}=\mathbf{x}_{p}$ with $\angle(\Gamma_{p}^{-},\Gamma_{p}^{+})$ being the intersection angle between $\Gamma_{p}^{-}$ and $\Gamma_{p}^{+}$.
 Remark 3.1.
@@
 In previous works such as [21, 15], the admissible condition (3.8) plays a crucial role in deriving corresponding uniqueness results.
-However, in this paper, when (3.8) is violated, we propose another admissible condition, denoted as (3.9).
+However, in this paper, when (3.8) is violated, we propose another admissible condition (3.9) to derive the unique determination of the shape of the conductive medium scatterer.
 There are three distinct cases for (3.9), characterized as follows:
   • a)
-    $\partial_{1}u(\mathbf{x}_{p})\neq 0$, $u(\mathbf{x}_{p})=\partial_{2}u(\mathbf{x}_{p})=0$,
+    $\lim_{\rho\to 0}\frac{1}{m(B(\mathbf{x}_{p},\rho))}\int_{B(\mathbf{x}_{p},\rho)}|\partial_{1}u(\mathbf{x}_{p})|\mathrm{d}\mathbf{x}\neq 0$,
+    $\lim_{\rho\to 0}\frac{1}{m(B(\mathbf{x}_{p},\rho))}\int_{B(\mathbf{x}_{p},\rho)}|u(\mathbf{x}_{p})|\mathrm{d}\mathbf{x}=\lim_{\rho\to 0}\frac{1}{m(B(\mathbf{x}_{p},\rho))}\int_{B(\mathbf{x}_{p},\rho)}|\partial_{2}u(\mathbf{x}_{p})|\mathrm{d}\mathbf{x}=0$,
   • b)
-    $\partial_{2}u(\mathbf{x}_{p})\neq 0$, $u(\mathbf{x}_{p})=\partial_{1}u(\mathbf{x}_{p})=0$,
+    $\lim_{\rho\to 0}\frac{1}{m(B(\mathbf{x}_{p},\rho))}\int_{B(\mathbf{x}_{p},\rho)}|\partial_{2}u(\mathbf{x}_{p})|\mathrm{d}\mathbf{x}\neq 0$,
+    $\lim_{\rho\to 0}\frac{1}{m(B(\mathbf{x}_{p},\rho))}\int_{B(\mathbf{x}_{p},\rho)}|u(\mathbf{x}_{p})|\mathrm{d}\mathbf{x}=\lim_{\rho\to 0}\frac{1}{m(B(\mathbf{x}_{p},\rho))}\int_{B(\mathbf{x}_{p},\rho)}|\partial_{1}u(\mathbf{x}_{p})|\mathrm{d}\mathbf{x}=0$,
   • c)
-    $\partial_{1}u(\mathbf{x}_{p})\neq 0$, $\partial_{2}u(\mathbf{x}_{p})\neq 0$, $u(\mathbf{x}_{p})=0$.
+    $\lim_{\rho\to 0}\frac{1}{m(B(\mathbf{x}_{p},\rho))}\int_{B(\mathbf{x}_{p},\rho)}|\partial_{1}u(\mathbf{x}_{p})|\mathrm{d}\mathbf{x}\neq 0$,
+    $\lim_{\rho\to 0}\frac{1}{m(B(\mathbf{x}_{p},\rho))}\int_{B(\mathbf{x}_{p},\rho)}|\partial_{2}u(\mathbf{x}_{p})|\mathrm{d}\mathbf{x}\neq 0$, $\lim_{\rho\to 0}\frac{1}{m(B(\mathbf{x}_{p},\rho))}\int_{B(\mathbf{x}_{p},\rho)}|u(\mathbf{x}_{p})|\mathrm{d}\mathbf{x}=0$.
 Since (3.9) encompasses more cases compared to (3.8), it is more likely to be satisfied in practical scenarios.
@@
 Then $u_{1}^{-}$ and $u_{2}^{+}$ satisfy the following PDE system,
-$\begin{cases}\Delta u_{1}^{-}+k^{2}q_{1}u_{1}^{-}=0,\hskip 73.97733pt\mbox{in}\ \mathcal{S}_{r_{0}},\\ \Delta u_{2}^{+}+k^{2}u_{2}^{+}=0,\hskip 83.93553pt\mbox{in}\ \mathcal{S}_{r_{0}},\\ u_{1}^{-}=u_{2}^{+},\ \partial_{\nu}u_{1}^{-}=\partial_{\nu}u_{2}^{+}+\eta_{1}u_{2}^{+},\hskip 14.22636pt\mbox{on}\ \Gamma_{r_{0}}^{\pm}.\end{cases}$
+$\begin{cases}\Delta u_{1}^{-}+k^{2}q_{1}u_{1}^{-}=0,\hskip 73.97733pt\mbox{in}\ \mathcal{S}_{r_{0}},\\ \Delta u_{2}^{+}+k^{2}u_{2}^{+}=0,\hskip 83.93553pt\mbox{in}\ \mathcal{S}_{r_{0}},\\ u_{1}^{-}=u_{2}^{+},\ \partial_{\nu}u_{1}^{-}=\partial_{\nu}u_{2}^{+}+\eta_{1}u_{2}^{+},\hskip 14.22636pt\mbox{on}\ \Gamma_{r_{0}}^{\pm}.\end{cases}$ (3.12)
 Let $\bar{u}=u_{1}^{-}-u_{2}^{+}$. Then it yields that
@@
 By virtue of the CGO solution defined in (2.1), using Green’s formula, one has
-$\eta_{1}\int_{\Gamma_{r_{0}}^{\pm}}u_{2}^{+}u_{0}\mathrm{d}\mathbf{x}=k^{2}\int_{\mathcal{S}_{r_{0}}}(1-q_{1})u_{2}^{+}u_{0}\mathrm{d}\mathbf{x}+\int_{\Lambda_{r_{0}}}\partial_{\nu}u_{0}\bar{u}-\partial_{\nu}\bar{u}u_{0}\mathrm{d}\sigma.$ (3.12)
-According to Lemma 3.1, we know that $u_{2}^{+}\in C^{1,\alpha}(\overline{\mathcal{S}_{r_{0}}}),\alpha\in(0,1)$, which implies the following expansion
-$u_{2}^{+}=u_{2}^{+}(\mathbf{0})+\partial_{1}u_{2}^{+}(\mathbf{0})x_{1}+\partial_{2}u_{2}^{+}(\mathbf{0})x_{2}+\delta u_{2}^{+},\ |\delta u_{2}^{+}|\leq|\mathbf{x}|^{1+\alpha}\|u_{2}^{+}\|_{C^{1,\alpha}}.$ (3.13)
+$\eta_{1}\int_{\Gamma_{r_{0}}^{\pm}}u_{1}^{-}u_{0}\mathrm{d}\mathbf{x}=k^{2}\int_{\mathcal{S}_{r_{0}}}(1-q_{1})u_{1}^{-}u_{0}\mathrm{d}\mathbf{x}+\int_{\Lambda_{r_{0}}}\partial_{\nu}u_{0}\bar{u}-\partial_{\nu}\bar{u}u_{0}\mathrm{d}\sigma.$ (3.13)
+According to Lemma 3.1, we know that $u_{1}^{-}\in C^{1,\alpha}(\overline{\mathcal{S}_{r_{0}}}),\alpha\in(0,1)$, which implies the following expansion
+$u_{1}^{-}=u_{1}^{-}(\mathbf{0})+\partial_{1}u_{1}^{-}(\mathbf{0})x_{1}+\partial_{2}u_{1}^{-}(\mathbf{0})x_{2}+\delta u_{1}^{-},\ |\delta u_{1}^{-}|\leq|\mathbf{x}|^{1+\alpha}\|u_{2}^{+}\|_{C^{1,\alpha}}.$ (3.14)
 According to the admissibility conditions in Definition 3.5, we will divide the proof into four cases, as follows.
-Case (a): $u_{2}^{+}(\mathbf{0})\not=0$. Combing (3.13) with (3.12), we can deduce that
-$\eta_{1}u_{2}^{+}(\mathbf{0})$ $\int_{\Gamma_{r_{0}}^{\pm}}e^{\rho\cdot\mathbf{x}}\mathrm{d}\mathbf{x}=\sum_{i=1}^{4}I_{i},$ (3.14)
+Case (a): $u_{1}^{-}(\mathbf{0})\not=0$. Combing (3.14) with (3.13), we can deduce that
+$\eta_{1}u_{1}^{-}(\mathbf{0})$ $\int_{\Gamma_{r_{0}}^{\pm}}e^{\rho\cdot\mathbf{x}}\mathrm{d}\mathbf{x}=\sum_{i=1}^{4}I_{i},$ (3.15)
 where
-$I_{1}$ $=-\eta_{1}u_{2}^{+}(\mathbf{0})\int_{\Gamma_{r_{0}}^{\pm}}\psi(\mathbf{x})e^{\rho\cdot\mathbf{x}}\mathrm{d}\sigma,\ I_{2}=k^{2}\int_{\mathcal{S}_{r_{0}}}(1-q_{1})u_{2}^{+}u_{0}\mathrm{d}\mathbf{x}$
-$I_{3}$ $=-\eta_{1}\int_{\Gamma_{r_{0}}^{\pm}}(\partial_{1}u_{2}^{+}(\mathbf{0})x_{1}+\partial_{2}u_{2}^{+}(\mathbf{0})x_{2}+\delta u_{2}^{+})u_{0}\mathrm{d}\sigma,$
+$I_{1}$ $=-\eta_{1}u_{1}^{-}(\mathbf{0})\int_{\Gamma_{r_{0}}^{\pm}}\psi(\mathbf{x})e^{\rho\cdot\mathbf{x}}\mathrm{d}\sigma,\ I_{2}=k^{2}\int_{\mathcal{S}_{r_{0}}}(1-q_{1})u_{1}^{-}u_{0}\mathrm{d}\mathbf{x}$
+$I_{3}$ $=-\eta_{1}\int_{\Gamma_{r_{0}}^{\pm}}(\partial_{1}u_{1}^{-}(\mathbf{0})x_{1}+\partial_{2}u_{1}^{-}(\mathbf{0})x_{2}+\delta u_{1}^{-})u_{0}\mathrm{d}\sigma,$
 $I_{4}$ $=\int_{\Lambda_{r_{0}}}\partial_{\nu}u_{0}\bar{u}-\partial_{\nu}\bar{u}u_{0}\mathrm{d}\sigma.$ (3.16)
@@
 Let $\mathbf{d}$ be given in the form of (2.14). We can then directly derive that
-$|\eta_{1}u_{2}^{+}(\mathbf{0})|$ $\left|\int_{\Gamma_{r_{0}}^{\pm}}e^{\rho\cdot\mathbf{x}}\mathrm{d}\sigma\right|\gtrsim\frac{\Gamma(1)|\eta_{1}u_{2}^{+}(\mathbf{0})||e^{\mathrm{i}\vartheta_{m}}+e^{\mathrm{i}\vartheta_{M}}|}{|\tau e^{\mathrm{i}\varphi}|}-\mathcal{O}(\tau^{-1}e^{-\frac{1}{2}\lambda r_{0}\tau}),$ (3.17)
+$|\eta_{1}u_{1}^{-}(\mathbf{0})|$ $\left|\int_{\Gamma_{r_{0}}^{\pm}}e^{\rho\cdot\mathbf{x}}\mathrm{d}\sigma\right|\gtrsim\frac{\Gamma(1)|\eta_{1}u_{2}^{+}(\mathbf{0})||e^{\mathrm{i}\vartheta_{m}}+e^{\mathrm{i}\vartheta_{M}}|}{|\tau e^{\mathrm{i}\varphi}|}-\mathcal{O}(\tau^{-1}e^{-\frac{1}{2}\lambda r_{0}\tau}),$ (3.18)
 here and in what follows, the term $\mathcal{O}(\tau^{-1}e^{-\frac{1}{2}\lambda r_{0}\tau})$ is defined in (2.15).
@@
 Combining (3.16)- (3.18) with (3.15), then multiplying $\tau$ and letting $\tau\to\infty$, we see that
-$|\eta_{1}u_{2}^{+}(\mathbf{0})||e^{\mathrm{i}\vartheta_{m}}+e^{\mathrm{i}\vartheta_{M}}|=0.$
+$|\eta_{1}u_{1}^{-}(\mathbf{0})||e^{\mathrm{i}\vartheta_{m}}+e^{\mathrm{i}\vartheta_{M}}|=0.$
 Since
@@
 it follows that
-$|\eta_{1}u_{2}^{+}(\mathbf{0})|=0.$
-However, this contradicts to the admissible condition $u_{2}^{+}(\mathbf{0})\neq 0$ due to $\eta_{1}\neq 0$.
-Case (b): $\partial_{1}u_{2}^{+}(\mathbf{0})\not=0$, $u_{2}^{+}(\mathbf{0})=\partial_{2}u_{2}^{+}(\mathbf{0})=0$. Similar to (3.14), we have the following integral equality:
-$\eta_{1}\partial_{1}u_{2}^{+}(\mathbf{0})$ $\int_{\Gamma_{r_{0}}^{\pm}}e^{\rho\cdot\mathbf{x}}x_{1}\mathrm{d}\sigma=-\eta_{1}\partial_{1}u_{2}^{+}(\mathbf{0})\int_{\Gamma_{r_{0}}^{\pm}}\psi(\mathbf{x})e^{\rho\cdot\mathbf{x}}x_{1}\mathrm{d}\sigma$
-$+\int_{\mathcal{S}_{r_{0}}}(1-q_{1})u_{2}^{+}u_{0}\mathrm{d}\mathbf{x}-\eta_{1}\int_{\Gamma_{r_{0}}^{\pm}}\delta u_{2}^{+}u_{0}\mathrm{d}\sigma+I_{4}.$ (3.18)
+$|\eta_{1}u_{1}^{-}(\mathbf{0})|=0.$
+However, this contradicts to the admissible condition $u_{1}^{-}(\mathbf{0})\neq 0$ due to $\eta_{1}\neq 0$.
+Case (b): $\partial_{1}u_{1}^{-}(\mathbf{0})\not=0$, $u_{1}^{-}(\mathbf{0})=\partial_{2}u_{1}^{-}(\mathbf{0})=0$. Similar to (3.15), we have the following integral equality:
+$\eta_{1}\partial_{1}u_{1}^{-}(\mathbf{0})$ $\int_{\Gamma_{r_{0}}^{\pm}}e^{\rho\cdot\mathbf{x}}x_{1}\mathrm{d}\sigma=-\eta_{1}\partial_{1}u_{1}^{-}(\mathbf{0})\int_{\Gamma_{r_{0}}^{\pm}}\psi(\mathbf{x})e^{\rho\cdot\mathbf{x}}x_{1}\mathrm{d}\sigma$
+$+\int_{\mathcal{S}_{r_{0}}}(1-q_{1})u_{1}^{-}u_{0}\mathrm{d}\mathbf{x}-\eta_{1}\int_{\Gamma_{r_{0}}^{\pm}}\delta u_{1}^{-}u_{0}\mathrm{d}\sigma+I_{4}.$ (3.19)
 By direct calculations we know that the right-hand side of (3.19) satisfies
@@
 Furthermore, we have the following inequality about the left-hand side of (3.19)
-$|LHS|\gtrsim\frac{\Gamma(2)|\eta_{1}\partial_{1}u_{2}^{+}(\mathbf{0})|}{\tau^{2}}\left|\frac{\cos\vartheta_{M}}{((\mathbf{d}+\mathrm{i}\mathbf{d})\cdot\mathbf{\hat{x}}_{1})^{2}}+\frac{\cos\vartheta_{m}}{((\mathbf{d}+\mathrm{i}\mathbf{d})\cdot\mathbf{\hat{x}}_{2})^{2}}\right|-\mathcal{O}(\tau^{-1}e^{-\lambda r_{0}\tau}),$ (3.20)
+$|LHS|\gtrsim\frac{\Gamma(2)|\eta_{1}\partial_{1}u_{1}^{-}(\mathbf{0})|}{\tau^{2}}\left|\frac{\cos\vartheta_{M}}{((\mathbf{d}+\mathrm{i}\mathbf{d})\cdot\mathbf{\hat{x}}_{1})^{2}}+\frac{\cos\vartheta_{m}}{((\mathbf{d}+\mathrm{i}\mathbf{d})\cdot\mathbf{\hat{x}}_{2})^{2}}\right|-\mathcal{O}(\tau^{-1}e^{-\lambda r_{0}\tau}),$ (3.21)
 where $\mathbf{\hat{x}}_{1}\in\Gamma_{r_{0}}^{+}$ and $\mathbf{\hat{x}}_{2}\in\Gamma_{r_{0}}^{-}$. Let $\mathbf{d}^{\perp}$ be in the form of (2.14), then we have
@@
 Substituting (3.20) and (3.21) into (3.19), then multiplying $\tau^{2}$ on both sides and taking $\tau\to\infty$, we get
-$|\eta_{1}\partial_{1}u_{2}^{+}(\mathbf{0})|=0,$
-which implies that $\partial_{1}u_{2}^{+}(\mathbf{0})=0$ by noting $\eta_{1}\not=0$. Hence, we obtain a contradiction.
-Case (c): $\partial_{2}u_{2}^{+}(\mathbf{0})\not=0,$ $u_{2}^{+}(\mathbf{0})=\partial_{1}u_{2}^{+}(\mathbf{0})=0$. Similar to (3.14), it holds that
-$\eta_{1}\partial_{2}u_{2}^{+}(\mathbf{0})$ $\int_{\Gamma_{r_{0}}^{\pm}}e^{\rho\cdot\mathbf{x}}x_{2}\mathrm{d}\sigma=-\eta_{1}\partial_{2}u_{2}^{+}(\mathbf{0})\int_{\Gamma_{r_{0}}^{\pm}}\psi(\mathbf{x})e^{\rho\cdot\mathbf{x}}x_{2}\mathrm{d}\sigma$
-$+\int_{\mathcal{S}_{r_{0}}}(1-q_{1})u_{2}^{+}u_{0}\mathrm{d}\mathbf{x}-\eta_{1}\int_{\Gamma_{r_{0}}^{\pm}}\delta u_{2}^{+}u_{0}\mathrm{d}\sigma+I_{4}.$ (3.21)
+$|\eta_{1}\partial_{1}u_{1}^{-}(\mathbf{0})|=0,$
+which implies that $\partial_{1}u_{1}^{-}(\mathbf{0})=0$ by noting $\eta_{1}\not=0$. Hence, we obtain a contradiction.
+Case (c): $\partial_{2}u_{1}^{-}(\mathbf{0})\not=0,$ $u_{1}^{-}(\mathbf{0})=\partial_{1}u_{1}^{-}(\mathbf{0})=0$. Similar to (3.15), it holds that
+$\eta_{1}\partial_{2}u_{1}^{-}(\mathbf{0})$ $\int_{\Gamma_{r_{0}}^{\pm}}e^{\rho\cdot\mathbf{x}}x_{2}\mathrm{d}\sigma=-\eta_{1}\partial_{2}u_{1}^{-}(\mathbf{0})\int_{\Gamma_{r_{0}}^{\pm}}\psi(\mathbf{x})e^{\rho\cdot\mathbf{x}}x_{2}\mathrm{d}\sigma$
+$+\int_{\mathcal{S}_{r_{0}}}(1-q_{1})u_{1}^{-}u_{0}\mathrm{d}\mathbf{x}-\eta_{1}\int_{\Gamma_{r_{0}}^{\pm}}\delta u_{1}^{-}u_{0}\mathrm{d}\sigma+I_{4}.$ (3.22)
 Similar to (3.19), for the right-hand and left-hand sides of (3.22), we have
@@
 and
-$|LHS|\gtrsim\frac{\Gamma(2)|\eta_{1}\partial_{2}u_{2}^{+}(\mathbf{0})|}{\tau^{2}}\left|\frac{\sin\vartheta_{M}}{((\mathbf{d}+\mathrm{i}\mathbf{d}^{\perp})\cdot\mathbf{\hat{x}}_{1})^{2}}+\frac{\sin\vartheta_{m}}{((\mathbf{d}+\mathrm{i}\mathbf{d}^{\perp})\cdot\mathbf{\hat{x}}_{2})^{2}}\right|-\mathcal{O}(\tau^{-1}e^{-\lambda r_{0}\tau})$ (3.22)
+$|LHS|\gtrsim\frac{\Gamma(2)|\eta_{1}\partial_{2}u_{1}^{-}(\mathbf{0})|}{\tau^{2}}\left|\frac{\sin\vartheta_{M}}{((\mathbf{d}+\mathrm{i}\mathbf{d}^{\perp})\cdot\mathbf{\hat{x}}_{1})^{2}}+\frac{\sin\vartheta_{m}}{((\mathbf{d}+\mathrm{i}\mathbf{d}^{\perp})\cdot\mathbf{\hat{x}}_{2})^{2}}\right|-\mathcal{O}(\tau^{-1}e^{-\lambda r_{0}\tau})$ (3.23)
 where $\mathbf{\hat{x}}_{1}\in\Gamma_{r_{0}}^{+},\ \mathbf{\hat{x}}_{2}\in\Gamma_{r_{0}}^{-}$. Let $\mathbf{d}^{\perp}$ be defined as in (2.14), then it follows that:
@@
 for $-\pi<\vartheta_{m}<\vartheta_{M}<\pi,\ \vartheta_{M}-\vartheta_{m}\in(0,\pi)$ and $\vartheta_{M}-\vartheta_{m}\neq\frac{\pi}{2}.$ Combining (3.22)-(3.23)and taking $\tau\to\infty$, one has
-$|\eta_{1}\partial_{2}u_{2}^{+}(\mathbf{0})|=0,$
-Hence, it yields that $\partial_{2}u_{2}^{+}(\mathbf{0})=0$ since $\eta_{1}\neq 0$, which contradicts to $\partial_{2}u_{2}^{+}(\mathbf{0})\not=0$.
-Case (d): $\partial_{1}u_{2}^{+}(\mathbf{0})\not=0$,$\partial_{2}u_{2}^{+}(\mathbf{0})\not=0$, $u_{2}^{+}(\mathbf{0})=0$. Similar to (3.14), we have
-$\eta_{1}\int_{\Gamma_{r_{0}}^{\pm}}$ $(\partial_{1}u_{2}^{+}(\mathbf{0})x_{1}+\partial_{2}u_{2}^{+}(\mathbf{0})x_{2})e^{\rho\cdot\mathbf{x}}\mathrm{d}\sigma=\int_{\mathcal{S}_{r_{0}}}(1-q_{1})u_{2}^{+}u_{0}\mathrm{d}\sigma-\eta_{1}\int_{\Gamma_{r_{0}}^{\pm}}\delta u_{2}^{+}u_{0}\mathrm{d}\sigma$
-$-\eta_{1}\int_{\Gamma_{r_{0}}^{\pm}}(\partial_{1}u_{2}^{+}(\mathbf{0})x_{1}+\partial_{2}u_{2}^{+}(\mathbf{0})x_{2})e^{\rho\cdot\mathbf{x}}\psi(\mathbf{x})\mathrm{d}\sigma+I_{4}.$ (3.23)
+$|\eta_{1}\partial_{2}u_{1}^{-}(\mathbf{0})|=0,$
+Hence, it yields that $\partial_{2}u_{1}^{-}(\mathbf{0})=0$ since $\eta_{1}\neq 0$, which contradicts to $\partial_{2}u_{1}^{-}(\mathbf{0})\not=0$.
+Case (d): $\partial_{1}u_{1}^{-}(\mathbf{0})\not=0$,$\partial_{2}u_{1}^{-}(\mathbf{0})\not=0$, $u_{1}^{-}(\mathbf{0})=0$. Similar to (3.15), we have
+$\eta_{1}\int_{\Gamma_{r_{0}}^{\pm}}$ $(\partial_{1}u_{1}^{-}(\mathbf{0})x_{1}+\partial_{2}u_{1}^{-}(\mathbf{0})x_{2})e^{\rho\cdot\mathbf{x}}\mathrm{d}\sigma=\int_{\mathcal{S}_{r_{0}}}(1-q_{1})u_{1}^{-}u_{0}\mathrm{d}\sigma-\eta_{1}\int_{\Gamma_{r_{0}}^{\pm}}\delta u_{1}^{-}u_{0}\mathrm{d}\sigma$
+$-\eta_{1}\int_{\Gamma_{r_{0}}^{\pm}}(\partial_{1}u_{1}^{-}(\mathbf{0})x_{1}+\partial_{2}u_{1}^{-}(\mathbf{0})x_{2})e^{\rho\cdot\mathbf{x}}\psi(\mathbf{x})\mathrm{d}\sigma+I_{4}.$ (3.24)
 By careful calculations, for the right-hand and left-hand sides of (3.24), it is evident that
@@
 as $\tau\to\infty$. When $\mathbf{d}^{\perp}$ is given in the form of (2.13), we have
-$\mathcal{N}=$ $(\partial_{1}u_{2}^{+}(\mathbf{0})\cos\vartheta_{M}+\partial_{2}u_{2}^{+}(\mathbf{0})\sin\vartheta_{M})e^{2\mathrm{i}\vartheta_{m}}$
-$\quad+(\partial_{1}u_{2}^{+}(\mathbf{0})\cos\vartheta_{m}+\partial_{2}u_{2}^{+}(\mathbf{0})\sin\vartheta_{m})e^{2\mathrm{i}\vartheta_{M}},$ (3.24)
+$\mathcal{N}=$ $(\partial_{1}u_{1}^{-}(\mathbf{0})\cos\vartheta_{M}+\partial_{2}u_{1}^{-}(\mathbf{0})\sin\vartheta_{M})e^{2\mathrm{i}\vartheta_{m}}$
+$\quad+(\partial_{1}u_{1}^{-}(\mathbf{0})\cos\vartheta_{m}+\partial_{2}u_{1}^{-}(\mathbf{0})\sin\vartheta_{m})e^{2\mathrm{i}\vartheta_{M}},$ (3.25)
 And when $\mathbf{d}^{\perp}$ is given by (2.14), one has
-$\mathcal{N}=$ $(\partial_{1}u_{2}^{+}(\mathbf{0})\cos\vartheta_{M}+\partial_{2}u_{2}^{+}(\mathbf{0})\sin\vartheta_{M})e^{-2\mathrm{i}\vartheta_{m}}$
-$\quad+(\partial_{1}u_{2}^{+}(\mathbf{0})\cos\vartheta_{m}+\partial_{2}u_{2}^{+}(\mathbf{0})\sin\vartheta_{m})e^{-2\mathrm{i}\vartheta_{M}}.$ (3.25)
-Let us consider the real and imaginary parts of $\partial_{1}u_{2}^{+}(\mathbf{0})$ and $\partial_{2}u_{2}^{+}(\mathbf{0})$, seperately.
-Denote $\partial_{1}u_{2}^{+}(\mathbf{0})=a_{1}+b_{1}\mathrm{i}$ and $\partial_{2}u_{2}^{+}(\mathbf{0})=a_{2}+b_{2}\mathrm{i}$, where $a_{i},b_{i}\in\mathbb{R}\ (i=1,2).$ For (3.24), it follows that
+$\mathcal{N}=$ $(\partial_{1}u_{1}^{-}(\mathbf{0})\cos\vartheta_{M}+\partial_{2}u_{1}^{-}(\mathbf{0})\sin\vartheta_{M})e^{-2\mathrm{i}\vartheta_{m}}$
+$\quad+(\partial_{1}u_{1}^{-}(\mathbf{0})\cos\vartheta_{m}+\partial_{2}u_{1}^{-}(\mathbf{0})\sin\vartheta_{m})e^{-2\mathrm{i}\vartheta_{M}}.$ (3.26)
+Let us consider the real and imaginary parts of $\partial_{1}u_{1}^{-}(\mathbf{0})$ and $\partial_{2}u_{1}^{-}(\mathbf{0})$, seperately.
+Denote $\partial_{1}u_{1}^{-}(\mathbf{0})=a_{1}+b_{1}\mathrm{i}$ and $\partial_{2}u_{1}^{-}(\mathbf{0})=a_{2}+b_{2}\mathrm{i}$, where $a_{i},b_{i}\in\mathbb{R}\ (i=1,2).$ For (3.25), it follows that
 $\Re(\mathcal{N})$ $=a_{1}(\cos\vartheta_{M}\cos 2\vartheta_{m}+\cos\vartheta_{m}\cos 2\vartheta_{M})-b_{1}(\cos\vartheta_{M}\sin 2\vartheta_{m}+\cos\vartheta_{m}\sin 2\vartheta_{M})$
@@
 $\det(\mathcal{M}_{1})=-\det(\mathcal{M}_{2})=-2\sin^{2}(\vartheta_{M}-\vartheta_{m})\cos(\vartheta_{M}-\vartheta_{m})\neq 0.$
-Therefore, by (3.28), we have $a_{i}=b_{i}=0$ for $i=1,2$, which indicates that $\partial_{1}u_{2}^{+}(\mathbf{0})=\partial_{2}u_{2}^{+}(\mathbf{0})=0$. This leads to a contradiction.
+Therefore, by (3.29), we have $a_{i}=b_{i}=0$ for $i=1,2$, which indicates that $\partial_{1}u_{1}^{-}(\mathbf{0})=\partial_{2}u_{1}^{-}(\mathbf{0})=0$. This leads to a contradiction.
 The proof is complete. ∎
@@
 $q_{\ell}=\sum_{i=1}^{N}q_{i,\ell}\chi_{\Sigma_{i}},\ \eta_{\ell}=\sum_{i=1}^{N}\tilde{\eta}_{\ell}\chi_{\Sigma_{i}}.$
-The coefficients of $q_{i,\ell}$ are real, where $q_{i,\ell}$ follows the form specified in (3.1), while $\tilde{\eta}_{\ell}$ is a real constant.
+The coefficients of $q_{i,\ell}$ $\in\mathbb{C}$, where $q_{i,\ell}$ follows the form specified in (3.1), while $\tilde{\eta}_{\ell}$ $\in\mathbb{C}$.
 Let $u_{\ell}^{\infty}(\mathbf{\hat{x}};u^{i})$ be the far-field pattern associated with the incident wave $u^{i}$ corresponding to $(\Omega;q_{\ell},\eta_{\ell})$.
@@
 $q_{\ell}=\sum_{i=1}^{N_{\ell}}q_{i,\ell}\chi_{\mathcal{D}_{i,\ell}},\ \eta_{\ell}=\sum_{i=1}^{N_{\ell}}\eta_{i,\ell}\chi_{\partial\Sigma_{i,\ell}}.$
-The coefficients of $q_{i,\ell}$ are real, where $q_{i,\ell}$ follows the form specified in (3.1), while $\eta_{i,\ell}$ is a real constant.
+The coefficients of $q_{i,\ell}$ $\in\mathbb{C}$ , where $q_{i,\ell}$ follows the form specified in (3.1), while $\eta_{i,\ell}$ $\in\mathbb{C}$.
 Let $u_{\ell}^{\infty}(\mathbf{\hat{x}};u^{i})$ be the far-field pattern associated with the incident wave $u^{i}$ corresponding to $(\Omega_{\ell};q_{\ell},\eta_{\ell})$.
@@
 for all $\mathbf{\hat{x}}\in\mathbb{S}^{1}$ and a fixed incident wave $u^{i}$.
-Then we have $N_{1}=N_{2}=N$, $\partial\Sigma_{i,1}=\partial\Sigma_{i,2}$, $q_{i,1}=q_{i,2}$ and $\eta_{i,1}=\eta_{i,2}$ for $i=1,\dots,N$,
-Remark 3.4.
-In Theorems 3.2 and 3.3, it is assumed that the coefficients of the reflective index follow the form specified in (3.1) and are real, while the conductive boundary parameter is a real constant.
-This assumption is also crucial in Lemma 3.2, as the proof of Theorems 3.2 and 3.3 relies on Lemma 3.2.
-Remark 3.5 suggests that the conclusion of Lemma 3.2 may hold in a more general scenario where the coefficients of the reflective index and the conductive boundary parameter are complex numbers.
-Hence, it is speculated that the linear polynomial reflective index with complex coefficients can be uniquely determined through a single far-field measurement.
-Consider an admissible conductive polygonal-nest or polygonal-cell medium scatterer $(\Omega;q,\eta)$ as defined in Definition 3.5, where $\Omega\Subset\mathbb{R}^{2}$ possesses a polygonal-nest or polygonal-cell structure.
-Each vertex $\mathbf{x}_{c}$ of $\Omega$ corresponds to a corner with an associated opening angle $\theta\in(0,\pi)$.
-Utilizing the translation transformation ${\mathbf{y}}=\mathbf{x}-\mathbf{x}_{c}$, we observe that the underlying corner can be represented by (1.2), where $\Gamma_{r_{0}}^{\pm}$ denote two line segment boundaries of $\mathcal{S}_{r_{0}}$.
-Recall that $\vartheta_{m}$ and $\vartheta_{M}$ are the polar angles of $\Gamma_{r_{0}}^{-}$ and $\Gamma_{r_{0}}^{+}$, respectively, such that $\theta=\vartheta_{M}-\vartheta_{m}$.
-Furthermore, we introduce the rotation transformation $\widetilde{\mathbf{x}}=M(\vartheta_{m})\mathbf{y}$, employing the rotation matrix $M(\vartheta_{m})$ given by:
-$M(\vartheta_{m})=\begin{bmatrix}\cos(\vartheta_{m})&\sin(\vartheta_{m})\\ -\sin(\vartheta_{m})&\cos(\vartheta_{m})\end{bmatrix}.$ (3.29)
-In this new coordinate system $\mathcal{C}_{N}$, we set $\mathbf{x}_{c}=\mathbf{0}$ and $\vartheta_{m}=0$.
-Proposition 3.1 demonstrates that if two linear polynomials coincide in the new coordinate system $\mathcal{C}_{N}$, they are identical in the original coordinate system.
-This result is crucial for proving the unique determination of a linear polynomial refractive index for $(\Omega;q,\eta)$.
-Similarly, Proposition 3.2 investigates the corresponding admissible conditions in the new coordinate system $\mathcal{C}_{N}$ for the admissible conditions (3.8) and (3.9), originally formulated in the old coordinate system.
-The proofs of Proposition 3.1 and 3.2 are omitted.
-Proposition 3.1.
-In a fixed coordinate system $\mathcal{C}$, consider two linear polynomials with the form given by
-$\alpha(\mathbf{x})=a_{0}+\mathbf{a}^{\top}\mathbf{x},\ \mbox{and}\ \gamma(\mathbf{x})=b_{0}+\mathbf{b}^{\top}\mathbf{x},$
-where $\mathbf{a}=(a_{1},a_{2})^{\top}$, $\mathbf{b}=(b_{1},b_{2})^{\top}$ and $\mathbf{x}=(x_{1},x_{2})^{\top}$.
-Pick up any fixed point $\mathbf{x}_{c}$ in the coordinate system $\mathcal{C}$ and a fixed rotation matrix $M(\vartheta_{m})$ given by (3.29).
-Here $\vartheta_{m}\in[-\pi,\pi)$ is a fixed angle in the coordinate system $\mathcal{C}$.
-Let
-$\tilde{\mathbf{x}}=M(\vartheta_{m})(\mathbf{x}-\mathbf{x}_{c}),$
-and
-$p(\tilde{\mathbf{x}})=\alpha(M(\vartheta_{m})^{\top}\tilde{\mathbf{x}}+\mathbf{x}_{c}),\ g(\tilde{\mathbf{x}})=\gamma(M(\vartheta_{m})^{\top}\tilde{\mathbf{x}}+\mathbf{x}_{c}).$
-If $p(\tilde{\mathbf{x}})\equiv g(\tilde{\mathbf{x}}),$ then $\alpha(\mathbf{x})=\gamma(\mathbf{x}).$
-Proposition 3.2.
-Consider a fixed coordinate system $\mathcal{C}$ and an admissible conductive polygonal-nest or polygonal-cell medium scatterer $(\Omega;q,\eta)$ in $\mathcal{C}$ as defined in Definition 3.5, where $\Omega$ possesses a polygonal-nest or polygonal-cell structure.
-Each vertex $\mathbf{x}_{c}$ of $\Omega$ corresponds to a corner with an associated opening angle $\theta\in(0,\pi)$.
-Let $u$ represent the total wave field of the direct scattering problem (1.7) associated with $(\Omega;q,\eta)$.
-Define $U(\tilde{\mathbf{x}})=u(M(\vartheta_{m})^{\top}\tilde{\mathbf{x}}+{\mathbf{x}_{c}})$, where $\vartheta_{m}\in[-\pi,\pi)$ is a fixed angle in the coordinate system $\mathcal{C}$.
-Then, we have the following relations:
-  • 1.
-    If $u(\mathbf{x}_{c})=0$, then $U(\mathbf{0})=0$.
-  • 2.
-    If $\nabla u(\mathbf{x}_{c})\neq 0$ and $u(\mathbf{x}_{c})=0$, then $\nabla U(\mathbf{0})\neq 0$ and $U(\mathbf{0})=0$.
+Then we have $N_{1}=N_{2}=N$, $\partial\Sigma_{i,1}=\partial\Sigma_{i,2}$, $q_{i,1}=q_{i,2}$ and $\eta_{i,1}=\eta_{i,2}$ for $i=1,\dots,N$.
 The proofs of Theorems 3.2 and 3.3 are deferred to Section 4.
@@
 Recall that $\mathcal{S}_{r_{0}},\Gamma_{r_{0}}^{\pm}$ and $B_{r_{0}}$ are defined in (1.2).
-Suppose that $u_{1}=u_{1}^{-}\chi_{\mathcal{S}_{r_{0}}}+u_{1}^{+}\chi_{B_{r_{0}}\setminus\overline{\mathcal{S}_{r_{0}}}}\in H^{1}(B_{r_{0}})$ and $u_{2}=u^{-}_{2}\chi_{\mathcal{S}_{r_{0}}}+u^{+}_{2}\chi_{B_{r_{0}}\setminus\overline{\mathcal{S}_{r_{0}}}}\in H^{1}(B_{r_{0}})$ satisfy the following PDE system:
+Suppose that $u_{1}=u_{1}^{-}\chi_{\mathcal{S}_{r_{0}}}+u_{1}^{+}\chi_{B_{r_{0}}\setminus\overline{\mathcal{S}_{r_{0}}}}\in H^{1}(B_{r_{0}})$ and $u_{2}=u^{-}_{2}\chi_{\mathcal{S}_{r_{0}}}+u^{+}_{2}\chi_{B_{r_{0}}\setminus\overline{\mathcal{S}_{r_{0}}}}\in H^{1}(B_{r_{0}})$ satisfy the following PDEs system:
 $\begin{cases}(\Delta+k^{2}q_{1})u_{1}^{-}=0,\hskip 78.24507pt\mbox{in}\ \mathcal{S}_{r_{0}},\\ (\Delta+k^{2}q_{2})u_{2}^{-}=0,\hskip 78.24507pt\mbox{in}\ \mathcal{S}_{r_{0}},\\ (\Delta+k^{2}q_{3})u^{+}_{1}=0,\hskip 78.24507pt\mbox{in}\ B_{r_{0}}\setminus\overline{\mathcal{S}_{r_{0}}},\\ (\Delta+k^{2}q_{3})u^{+}_{2}=0,\hskip 78.24507pt\mbox{in}\ B_{r_{0}}\setminus\overline{\mathcal{S}_{r_{0}}},\\ u^{+}_{1}=u^{-}_{1},\ \partial_{\nu}u^{-}_{1}=\partial_{\nu}u^{+}_{1}+\eta_{1}u^{+}_{1},\hskip 12.80365pt\mbox{on}\ \Gamma_{r_{0}}^{\pm},\\ u^{+}_{2}=u^{-}_{2},\ \partial_{\nu}u^{-}_{2}=\partial_{\nu}u^{+}_{2}+\eta_{2}u^{+}_{2},\hskip 12.80365pt\mbox{on}\ \Gamma_{r_{0}}^{\pm},\\ \end{cases}$ (3.30)
-where $\eta_{1}$ and $\eta_{2}$ are real constants and $q_{1},\ q_{2},\ q_{3}$ with forms
+where $\eta_{1}$, $\eta_{2}$ $\in\mathbb{C}$ and $q_{1},\ q_{2},\ q_{3}$ with forms
 $q_{j}=q_{j}^{(0)}+q_{j}^{(1)}x_{1}+q_{j}^{(2)}x_{2},\quad j=1,2,3,$
-where $q_{j}^{(0)}$, $q_{j}^{(1)}$ and $q_{j}^{(2)}$ are real constants, $j=1,2,3$.
-Assume that for $i=1$ or $2$, either
-$|u_{i}(\mathbf{0})|\neq 0,$ (3.31)
-or
-$\nabla u_{i}(\mathbf{0})\neq 0,\ u_{i}(\mathbf{0})=0$ (3.32)
-is fulfilled.
-Additionally, if (3.32) holds, it is further required that $\vartheta_{M}-\vartheta_{m}\in(\frac{\pi}{3},\frac{2\pi}{3})\setminus\{\frac{\pi}{2}\}$, where $\vartheta_{M}-\vartheta_{m}$ is the opening angle of $\mathcal{S}_{r_{0}}$.
+where $q_{j}^{(0)}$, $q_{j}^{(1)}$ and $q_{j}^{(2)}$ $\in\mathbb{C}$, $j=1,2,3$.
+Assume that
+$\lim_{\rho\to 0}\frac{1}{m(B(\mathbf{0},\rho))}\int_{B(\mathbf{x}_{p},\rho)}|u_{i}(\mathbf{0})|\mathrm{d}\mathbf{x}\neq 0,\ i=1\ \mbox{or}\ 2.$ (3.31)
 If
@@
 Since $u^{+}_{1}=u_{2}^{+}$ in $B_{r_{0}}\setminus\overline{\mathcal{S}_{r_{0}}}$, we have that $u^{+}_{1}=u^{+}_{2}\ \mbox{on}\ \Gamma_{r_{0}}^{\pm}$ by the trace theorem, which indicates that $u_{1}^{-}=u_{1}^{+}=u_{2}^{+}=u_{2}^{-}\mbox{on}\ \Gamma_{r_{0}}^{\pm}$ according to the transmission boundary condition.
-Without of loss of generality, we assume that $u_{2}^{-}$ fulfills (3.31) or (3.32).
+Without loss of generality, we assume that $u_{2}^{-}$ fulfills (3.31).
 Denote $\tilde{u}:=u^{-}_{1}-u_{2}^{-}$, then it can be shown directly that
@@
 $\int_{\mathcal{S}_{r_{0}}}k^{2}(q_{2}-q_{1})u_{2}^{-}u_{0}\mathrm{d}\mathbf{x}=(\eta_{1}-\eta_{2})\int_{\Gamma_{r_{0}}^{\pm}}u_{2}^{-}u_{0}\mathrm{d}\sigma+\int_{\Lambda_{r_{0}}}u_{0}\partial_{\nu}\tilde{u}-\tilde{u}\partial_{\nu}u_{0}\mathrm{d}\sigma.$ (3.32)
-According to Lemma 3.1, we have $u_{2}^{-}\in C^{1,\alpha}(\overline{\mathcal{S}_{r_{0}}})$. Therefore, it follows that
-$u_{2}^{-}$ $=u_{2}^{-}(\mathbf{0})+\partial_{1}u_{2}^{-}(\mathbf{0})x_{1}+\partial_{2}u_{2}^{-}(\mathbf{0})x_{2}+\delta u_{2}^{-},\ |\delta u_{2}^{-}|\leq|\mathbf{x}|^{1+\alpha}\|u_{2}^{-}\|_{C^{1,\alpha}}.$ (3.34)
-For convenience, we denote $\tilde{\delta}u_{2}^{-}=\partial_{1}u_{2}^{-}(\mathbf{0})x_{1}+\partial_{2}u_{2}^{-}(\mathbf{0})x_{2}+\delta u_{2}^{-}.$
-First, we are going to prove $\eta_{1}=\eta_{2}$ and $q_{1}=q_{2}$. According to the admissible conditions (3.8) and (3.9), in the rest of this proof, we will consider two distinct cases.
-Case (I): $u_{2}^{-}(\mathbf{0})\not=0.$ Using (3.33) and (3.34), we can directly see that
+First, we demonstrate that the solution $u_{i}^{-}\in H_{1}(\mathcal{S}_{r_{0}}),i=1,2$ of the system (3.30) belong to $C^{\alpha}(\overline{\mathcal{S}_{r_{0}}})$.
+From [39, 43], we know that $u_{i}^{-}$ admits a decomposition into a regular part and a singular part in the following form:
+$u_{i}^{-}=u_{i,reg}^{-}+u_{i,sing}^{-}.$
+The regular part $u_{i,reg}^{-}\in H^{2}(\mathcal{S}_{r_{0}})$ and hence by the Sobolev embedding theorem, it follows that $u_{i,reg}^{-}\in C^{\beta}(\mathcal{S}_{r_{0}}),\beta\in(0,1]$.
+Moreover, $u_{{i,sing}}^{-}\in C^{\alpha}(\overline{\mathcal{S}_{r_{0}}})$, where $\alpha\in(0,1)$ is the Hölder exponent depending on the opening angle of $\mathcal{S}_{r_{0}}$.
+Therefore, $u_{i}^{-}\in C^{\alpha}(\overline{\mathcal{S}_{r_{0}}}).$ Thus, $u_{2}^{-}$ satisfies the following expansion:
+$u_{2}^{-}$ $=u_{2}^{-}(\mathbf{0})+\delta u_{2}^{-},\ |\delta u_{2}^{-}|\leq|\mathbf{x}|^{\alpha}\|u_{2}^{-}\|_{C^{\alpha}}.$ (3.33)
+Now, we are going to prove that $\eta_{1}=\eta_{2}$ and $q_{1}=q_{2}$ in two parts.
+Part 1: $\eta_{1}=\eta_{2}$. Using (3.32) and (3.33), we can directly see that
 $(\eta_{2}-\eta_{1})u_{2}^{-}(\mathbf{0})\int_{\Gamma_{r_{0}}^{\pm}}e^{\rho\cdot\mathbf{x}}\mathrm{d}\sigma=\sum_{i=1}^{4}I_{i}$ (3.34)
@@
 $I_{1}$ $=(\eta_{1}-\eta_{2})u_{2}^{-}(\mathbf{0})\int_{\Gamma_{r_{0}}^{\pm}}\psi(\mathbf{x})e^{\rho\cdot\mathbf{x}}\mathrm{d}\sigma,I_{2}=\int_{\mathcal{S}_{r_{0}}}k^{2}(q_{1}-q_{2})u_{2}^{-}u_{0}\mathrm{d}\mathbf{x},$
-$I_{3}$ $=(\eta_{1}-\eta_{2})\int_{\Gamma_{r_{0}}^{\pm}}\tilde{\delta}u_{2}^{-}u_{0}(\mathbf{x})\mathrm{d}\sigma,I_{4}=\int_{\Lambda_{r_{0}}}u_{0}\partial_{\nu}\tilde{u}-\tilde{u}\partial_{\nu}u_{0}\mathrm{d}\sigma.$
+$I_{3}$ $=(\eta_{1}-\eta_{2})\int_{\Gamma_{r_{0}}^{\pm}}\delta u_{2}^{-}u_{0}(\mathbf{x})\mathrm{d}\sigma,I_{4}=\int_{\Lambda_{r_{0}}}u_{0}\partial_{\nu}\tilde{u}-\tilde{u}\partial_{\nu}u_{0}\mathrm{d}\sigma.$
 By virtue of (3.16) and (3.17), we can derive that the right-hand side of (3.34) satisfies
@@
 which indicates that $\eta_{1}=\eta_{2}$ since $u_{2}^{-}(\mathbf{0})\neq 0.$
-Next, using $\eta_{1}=\eta_{2}$, the integral identity (3.35) can be rewritten as
+Part 2: $q_{1}=q_{2}$. Using $\eta_{1}=\eta_{2}$, the integral identity (3.34) can be rewritten as
 $k^{2}(q_{2}^{(0)}-q_{1}^{(0)})u_{2}^{-}(\mathbf{0})\int_{\mathcal{S}_{r_{0}}}e^{\rho\cdot\mathbf{x}}\mathrm{d}\mathbf{x}=\sum_{i=1}^{4}\mathcal{P}_{i}+I_{4},$ (3.36)
 where
-$\mathcal{P}_{1}$ $=k^{2}(q_{1}^{(0)}-q_{2}^{(0)})u_{2}^{-}(\mathbf{0})\int_{\mathcal{S}_{r_{0}}}\psi(\mathbf{x})e^{\rho\cdot\mathbf{x}}\mathrm{d}\mathbf{x},\ \mathcal{P}_{2}=k^{2}(q_{1}^{(0)}-q_{2}^{(0)})\int_{\mathcal{S}_{r_{0}}}\tilde{\delta}u_{2}^{-}u_{0}\mathrm{d}\mathbf{x},$
+$\mathcal{P}_{1}$ $=k^{2}(q_{1}^{(0)}-q_{2}^{(0)})u_{2}^{-}(\mathbf{0})\int_{\mathcal{S}_{r_{0}}}\psi(\mathbf{x})e^{\rho\cdot\mathbf{x}}\mathrm{d}\mathbf{x},\ \mathcal{P}_{2}=k^{2}(q_{1}^{(0)}-q_{2}^{(0)})\int_{\mathcal{S}_{r_{0}}}\delta u_{2}^{-}u_{0}\mathrm{d}\mathbf{x},$
 $\mathcal{P}_{3}$ $=k^{2}(q_{1}^{(1)}-q_{2}^{(1)})u_{2}^{-}(\mathbf{0})\int_{\mathcal{S}_{r_{0}}}x_{1}u_{0}\mathrm{d}\mathbf{x}+k^{2}(q_{1}^{(2)}-q_{2}^{(2)})u_{2}^{-}(\mathbf{0})\int_{\mathcal{S}_{r_{0}}}x_{2}u_{0}\mathrm{d}\mathbf{x},$
-$\mathcal{P}_{4}$ $=k^{2}(q_{1}^{(1)}-q_{2}^{(1)})u_{2}^{-}(\mathbf{0})\int_{\mathcal{S}_{r_{0}}}x_{1}\tilde{\delta}u_{2}^{-}u_{0}\mathrm{d}\mathbf{x}+k^{2}(q_{1}^{(2)}-q_{2}^{(2)})u_{2}^{-}(\mathbf{0})\int_{\mathcal{S}_{r_{0}}}x_{2}\tilde{\delta}u_{2}^{-}u_{0}\mathrm{d}\mathbf{x},$
+$\mathcal{P}_{4}$ $=k^{2}(q_{1}^{(1)}-q_{2}^{(1)})u_{2}^{-}(\mathbf{0})\int_{\mathcal{S}_{r_{0}}}x_{1}\delta u_{2}^{-}u_{0}\mathrm{d}\mathbf{x}+k^{2}(q_{1}^{(2)}-q_{2}^{(2)})u_{2}^{-}(\mathbf{0})\int_{\mathcal{S}_{r_{0}}}x_{2}\delta u_{2}^{-}u_{0}\mathrm{d}\mathbf{x},$
 and $I_{4}$ is defined in (3.34). Then by Proposition 2.1, it is directly shown that
@@
 $\mathcal{E}_{1}$ $=k^{2}(q_{1}^{(1)}-q_{2}^{(1)})u_{2}^{-}(\mathbf{0})\int_{\mathcal{S}_{r_{0}}}x_{1}\psi(\mathbf{x})e^{\rho\cdot\mathbf{x}}\mathrm{d}\mathbf{x}+k^{2}(q_{1}^{(2)}-q_{2}^{(2)})u_{2}^{-}(\mathbf{0})\int_{\mathcal{S}_{r_{0}}}x_{2}\psi(\mathbf{x})e^{\rho\cdot\mathbf{x}}\mathrm{d}\mathbf{x},$
-$\mathcal{E}_{2}$ $=k^{2}(q_{1}^{(1)}-q_{2}^{(1)})\int_{\mathcal{S}_{r_{0}}}x_{1}\tilde{\delta}u_{2}^{-}u_{0}\mathrm{d}\mathbf{x}+k^{2}(q_{1}^{(2)}-q_{2}^{(2)})\int_{\mathcal{S}_{r_{0}}}x_{2}\tilde{\delta}u_{2}^{-}u_{0}\mathrm{d}\mathbf{x}.$
+$\mathcal{E}_{2}$ $=k^{2}(q_{1}^{(1)}-q_{2}^{(1)})\int_{\mathcal{S}_{r_{0}}}x_{1}\delta u_{2}^{-}u_{0}\mathrm{d}\mathbf{x}+k^{2}(q_{1}^{(2)}-q_{2}^{(2)})\int_{\mathcal{S}_{r_{0}}}x_{2}\delta u_{2}^{-}u_{0}\mathrm{d}\mathbf{x}.$
 Utilizing Proposition 2.1, it can be calculated estimates about $\mathcal{E}_{1}$ and $\mathcal{E}_{2}$ as follows,
 $|\mathcal{E}_{1}|$ $\lesssim\int_{\mathcal{S}_{r_{0}}}|\mathbf{x}||\psi(\mathbf{x})e^{\rho\cdot\mathbf{x}}|\mathrm{d}\mathbf{x}\lesssim\tau^{-\frac{41}{12}}+\mathcal{O}(\tau^{-1}e^{-\frac{1}{2}\varsigma r_{0}\tau}),$
-$|\mathcal{E}_{2}|$ $\lesssim\int_{\mathcal{S}_{r_{0}}}|\mathbf{x}||\tilde{\delta}u_{2}^{-}||u_{0}|\mathrm{d}\mathbf{x}\lesssim\tau^{-(3+\alpha)}+\tau^{-(\frac{53}{12}+\alpha)}+\mathcal{O}(\tau^{-1}e^{-\frac{1}{2}\varsigma r_{0}\tau}).$ (3.41)
+$|\mathcal{E}_{2}|$ $\lesssim\int_{\mathcal{S}_{r_{0}}}|\mathbf{x}||\delta u_{2}^{-}||u_{0}|\mathrm{d}\mathbf{x}\lesssim\tau^{-(2+\alpha)}+\tau^{-(\frac{41}{12}+\alpha)}+\mathcal{O}(\tau^{-1}e^{-\frac{1}{2}\varsigma r_{0}\tau}).$ (3.40)
 Then we can write the left-hand side of (3.39) as
@@
 Hence we get that $\overline{q}_{1}=0$ and $\overline{q}_{2}=0$, which implies that $q_{1}^{(1)}=q_{2}^{(1)}$ and $q_{1}^{(2)}=q_{2}^{(2)}$ by using $u_{2}^{-}(\mathbf{0})\not=0.$
-Case (II): $|\nabla u_{2}^{-}(\mathbf{0})|\not=0$ and $u(\mathbf{0})=0.$ In this scenario, we utilize the rotation transformation $\tilde{\mathbf{x}}=M(\vartheta_{m})^{\top}\mathbf{x}$, where $M(\vartheta_{m})$ is defined by (3.29).
-Consequently, in the new coordinate system, we denote the linear segment boundary of the corner $\mathcal{S}_{r_{0}}$ by $\tilde{\Gamma}_{r_{0}}^{\pm}$, and the new corner is represented as $\mathcal{S}_{r_{0}}^{{}^{\prime}}$.
-The boundary part $\Lambda_{r_{0}}$ of $\mathcal{S}_{r_{0}}^{{}^{\prime}}$ is denoted by $\Lambda_{r_{0}}^{{}^{\prime}}$.
-The polar angle of $\tilde{\Gamma}_{r_{0}}^{-}$ is 00, and the polar angle of $\tilde{\Gamma}_{r_{0}}^{+}$ is $\tilde{\vartheta}=\vartheta_{M}-\vartheta_{m}$.
-Denote
-$\tilde{q}_{i}(\tilde{\mathbf{x}})=q_{i}(M(\vartheta_{m})^{\top}\tilde{\mathbf{x}}),\ \tilde{q}_{i}=\tilde{q}_{0}^{(i)}+\tilde{q}_{1}^{(i)}\tilde{x}_{1}+\tilde{q}_{2}^{(i)}\tilde{x}_{2},\ i=1,2,3,$
-and
-$U_{j}^{\pm}(\tilde{\mathbf{x}})=u_{j}^{\pm}(M(\vartheta_{m})^{\top}\tilde{\mathbf{x}}),\ j=1,2.$
-Then we can rewrite (3.30) as follows:
-$\begin{cases}(\Delta+k^{2}\tilde{q}_{1})U_{1}^{-}=0,\hskip 91.04872pt\mbox{in}\ \mathcal{S}^{{}^{\prime}}_{r_{0}},\\ (\Delta+k^{2}\tilde{q}_{2})U_{2}^{-}=0,\hskip 91.04872pt\mbox{in}\ \mathcal{S}^{{}^{\prime}}_{r_{0}},\\ (\Delta+k^{2}\tilde{q}_{3})U^{+}_{1}=0,\hskip 91.04872pt\mbox{in}\ B_{r_{0}}\setminus\overline{\mathcal{S}^{{}^{\prime}}_{r_{0}}},\\ (\Delta+k^{2}\tilde{q}_{3})U^{+}_{2}=0,\hskip 91.04872pt\mbox{in}\ B_{r_{0}}\setminus\overline{\mathcal{S}^{{}^{\prime}}_{r_{0}}},\\ U^{+}_{1}=U^{-}_{1},\ \partial_{\nu}U^{-}_{1}=\partial_{\nu}U^{+}_{1}+\eta_{1}U^{+}_{1},\hskip 15.6491pt\mbox{on}\ \tilde{\Gamma}_{r_{0}}^{\pm},\\ U^{+}_{2}=U^{-}_{2},\ \partial_{\nu}U^{-}_{2}=\partial_{\nu}U^{+}_{2}+\eta_{2}U^{+}_{2},\hskip 15.6491pt\mbox{on}\ \tilde{\Gamma}_{r_{0}}^{\pm}.\\ \end{cases}$ (3.50)
-Denote $\tilde{U}=U_{1}^{-}-U_{2}^{-},$ then one has
-[…]
-By Lemma 2.1, we know that there exists a CGO solution $\tilde{u}_{0}$ of the form
-$\tilde{u}_{0}(\tilde{\mathbf{x}})=(1+\psi(\tilde{\mathbf{x}}))e^{\rho\cdot\tilde{\mathbf{x}}}$
-satisfying that
-$(\Delta+k^{2}\tilde{q}_{1})\tilde{u}_{0}(\tilde{\mathbf{x}})=0,\ \mbox{in}\ \mathcal{S}_{r_{0}}^{{}^{\prime}}.$
-Recall that
-$u_{2}^{-}$ $=u_{2}^{-}(\mathbf{0})+\partial_{1}u_{2}^{-}(\mathbf{0})x_{1}+\partial_{2}u_{2}^{-}(\mathbf{0})x_{2}+\delta u_{2}^{-}$
-$=u_{2}^{-}(\mathbf{0})+\mathbf{x}^{\top}\nabla_{\mathbf{x}}u_{2}^{-}(\mathbf{0})+\delta u_{2}^{-}.$
-Due to $\mathbf{x}=M(\vartheta_{m})^{\top}\tilde{\mathbf{x}}$ and $\nabla_{\mathbf{x}}u_{2}^{-}(\mathbf{0})=M(\vartheta_{m})^{\top}\nabla_{\tilde{\mathbf{x}}}U_{2}^{-}(\mathbf{0})$, then we have
-$U_{2}^{-}(\tilde{\mathbf{x}})=U_{2}^{-}(\mathbf{0})+\partial_{1}U_{2}^{-}(\mathbf{0})\tilde{x}_{1}+\partial_{2}U_{2}^{-}(\mathbf{0})\tilde{x}_{2}+\delta U_{2}^{-},|\delta U_{2}^{-}|\leq C\|U_{2}^{-}\|_{C^{1,\alpha}}.$ (3.52)
-According to Proposition 3.2, using the condition $|\nabla u_{2}^{-}(\mathbf{0})|\not=0$ and $u(\mathbf{0})=0$, we have $\nabla U_{2}^{-}(\mathbf{0})\neq 0$ and $U_{2}^{-}(\mathbf{0})=0$.
-In the following, we further divide the rest proof into the following three distinct cases.
-Case (i): $\partial_{1}U_{2}^{-}(\mathbf{0})\not=0,\ \partial_{2}U_{2}^{-}(\mathbf{0})\not=0,\ U_{2}^{-}(\mathbf{0})=0.$ Recall that $U_{2}^{-}=\partial_{1}U_{2}^{-}(\mathbf{0})\tilde{x}_{1}+\partial_{2}U_{2}^{-}(\mathbf{0})\tilde{x}_{2}+\delta U_{2}^{-}$, then we have
-$\overline{\eta}$ $\int_{\tilde{\Gamma}_{r_{0}}^{\pm}}(\partial_{1}U_{2}^{-}(\mathbf{0})\tilde{x}_{1}+\partial_{2}U_{2}^{-}(\mathbf{0})\tilde{x}_{2})e^{\rho\cdot\tilde{\mathbf{x}}}\mathrm{d}\sigma=\int_{\mathcal{S}_{r_{0}}^{{}^{\prime}}}k^{2}(q_{1}-q_{2})\tilde{u}_{0}U_{2}^{-}\mathrm{d}\tilde{\mathbf{x}}$
-$-\overline{\eta}\int_{\tilde{\Gamma}_{r_{0}}^{\pm}}(\partial_{1}U_{2}^{-}(\mathbf{0})\tilde{x}_{1}+\partial_{2}U_{2}^{-}(\mathbf{0})\tilde{x}_{2})\psi(\tilde{\mathbf{x}})e^{\rho\cdot\tilde{\mathbf{x}}}\mathrm{d}\sigma-\overline{\eta}\int_{\tilde{\Gamma}_{r_{0}}^{\pm}}\delta U_{2}^{-}\tilde{u}_{0}\mathrm{d}\sigma+\mathcal{I},$ (3.53)
-where $\overline{\eta}=\eta_{2}-\eta_{1}$. Comparing (3.53) with (3.23), then we see that the right-hand and left-hand sides of (3.53) satisfy
-$|RHS|\lesssim\tau^{-(2+\alpha)}+\tau^{-\frac{13}{6}}+\mathcal{O}(\tau^{-1}e^{-\frac{1}{2}\varsigma r_{0}\tau})+(1+\tau)(1+\tau^{-\varsigma r_{0}\tau})$
-and
-$LHS=\frac{\Gamma(2)(\eta_{2}-\eta_{1})\mathcal{Q}}{\tau^{2}((\mathbf{d}+\mathrm{i}\mathbf{d}^{\perp})\cdot\mathbf{\hat{x}}_{1})^{2}((\mathbf{d}+\mathrm{i}\mathbf{d}^{\perp})\cdot\mathbf{\hat{x}}_{2})^{2}}+\mathcal{O}(\tau^{-1}e^{-\frac{1}{2}\lambda r_{0}\tau}),$
-where
-$\mathcal{Q}$ $=(\partial_{1}U_{2}^{-}(\mathbf{0})\cos\tilde{\vartheta}+\partial_{2}U_{2}^{-}(\mathbf{0})\sin\tilde{\vartheta})+\partial_{1}U_{2}^{-}(\mathbf{0})e^{-2\mathrm{i}\tilde{\vartheta}}$
-$=(\partial_{1}U_{2}^{-}(\mathbf{0})\cos\tilde{\vartheta}+\partial_{2}U_{2}^{-}(\mathbf{0})\sin\tilde{\vartheta}+\partial_{1}U_{2}^{-}(\mathbf{0})\cos 2\tilde{\vartheta})-\mathrm{i}\partial_{1}U_{2}^{-}(\mathbf{0})\sin 2\tilde{\vartheta}.$
-Since $\partial_{1}U_{2}^{-}(\mathbf{0})\neq 0$, it follows that $\Im(\mathcal{Q})\neq 0$ when $\tilde{\vartheta}\neq\frac{\pi}{2}$.
-By employing a similar manner as in Case (b) of the proof of Theorem 3.1, it can be demonstrated that $\eta_{1}=\eta_{2}$ since $\tilde{\vartheta}=\vartheta_{M}-\vartheta_{m}\neq\frac{\pi}{2}$.
-Next, we will prove that $\tilde{q}_{1}^{(0)}=\tilde{q}_{2}^{(0)}$. Using $\eta_{1}=\eta_{2}$, the integral identity (3.53) can be rewritten as
-$\tilde{\sigma}_{0}$ $\left(\partial_{1}U_{2}^{-}(\mathbf{0})\int_{\mathcal{S}_{r_{0}}^{{}^{\prime}}}\tilde{x}_{1}e^{\rho\cdot\tilde{\mathbf{x}}}\mathrm{d}\tilde{\mathbf{x}}+\partial_{2}U_{2}^{-}(\mathbf{0})\int_{\mathcal{S}_{r_{0}}^{{}^{\prime}}}\tilde{x}_{2}e^{\rho\cdot\tilde{\mathbf{x}}}\mathrm{d}\tilde{\mathbf{x}}\right)$
-$=-\tilde{\sigma}_{0}\left(\partial_{1}U_{2}^{-}(\mathbf{0})\int_{\mathcal{S}_{r_{0}}^{{}^{\prime}}}\tilde{x}_{1}\psi(\tilde{\mathbf{x}})e^{\rho\cdot\tilde{\mathbf{x}}}\mathrm{d}\tilde{\mathbf{x}}+\partial_{2}U_{2}^{-}(\mathbf{0})\int_{\mathcal{S}_{r_{0}}^{{}^{\prime}}}\tilde{x}_{2}\psi(\tilde{\mathbf{x}})e^{\rho\cdot\tilde{\mathbf{x}}}\mathrm{d}\tilde{\mathbf{x}}\right)$
-$-\tilde{\sigma}_{0}\int_{\mathcal{S}_{r_{0}}^{{}^{\prime}}}\delta U_{2}^{-}\tilde{u}_{0}\mathrm{d}\tilde{\mathbf{x}}+k^{2}(\tilde{q}_{1}^{(1)}-\tilde{q}_{2}^{(1)})\int_{\mathcal{S}_{r_{0}}^{{}^{\prime}}}\tilde{x}_{1}U_{2}^{-}\tilde{u}_{0}\mathrm{d}\tilde{\mathbf{x}}$
-$+k^{2}(\tilde{q}_{1}^{(2)}-\tilde{q}_{2}^{(2)})\int_{\mathcal{S}_{r_{0}}^{{}^{\prime}}}\tilde{x}_{2}U_{2}^{-}\tilde{u}_{0}\mathrm{d}\tilde{\mathbf{x}}+\mathcal{I}$ (3.54)
-where $\tilde{\sigma}_{0}=k^{2}(\tilde{q}_{2}^{(0)}-\tilde{q}_{1}^{(0)})$. Initially, according to Proposition 2.1, for the right-hand side of (3.54), it follows that
-$|RHS|\lesssim\tau^{-(3+\alpha)}+\tau^{-\frac{41}{12}}+\mathcal{O}(\tau^{-1}e^{-\frac{1}{2}\varsigma r_{0}\tau})+(1+\tau)(1+\tau^{-\frac{2}{3}})e^{-\varsigma r_{0}\tau}.$
-Concerning the left-hand side of (3.54) and letting $\mathbf{d}^{\perp}$ be in the form of (2.14), it follows that
-$LHS$ $=\frac{\Gamma(3)\tilde{\sigma}_{0}}{\tau^{3}e^{3\mathrm{i}\varphi}}\left(\partial_{1}U_{2}^{-}(\mathbf{0})\int_{0}^{\tilde{\vartheta}}e^{3\mathrm{i}\vartheta}\cos\vartheta\mathrm{d}\vartheta+\partial_{2}U_{2}^{-}(\mathbf{0})\int_{0}^{\tilde{\vartheta}}e^{3\mathrm{i}\vartheta}\sin\vartheta\mathrm{d}\vartheta\right)$
-$+\mathcal{O}(\tau^{-1}e^{-\frac{1}{2}\lambda r_{0}\tau}).$
-Denote
-$\Psi:=\partial_{1}U_{2}^{-}(\mathbf{0})\int_{0}^{\tilde{\vartheta}}e^{3\mathrm{i}\vartheta}\cos\vartheta\mathrm{d}\vartheta+\partial_{2}U_{2}^{-}(\mathbf{0})\int_{0}^{\tilde{\vartheta}}e^{3\mathrm{i}\vartheta}\sin\vartheta\mathrm{d}\vartheta.$
-Multiplying both sides of (3.54) by $\tau^{3}$ and taking the limit as $\tau$ tends to infinity yields $\tilde{\sigma}_{0}\Psi=0$. Direct calculations show that
-$\Re(\Psi)$ $=-\partial_{1}U_{2}^{-}(\mathbf{0})(\sin\tilde{\vartheta}\cos 3\tilde{\vartheta}-3\sin 3\tilde{\vartheta}\cos\tilde{\vartheta})$
-$+\partial_{2}U_{2}^{-}(\mathbf{0})(\cos\tilde{\vartheta}\cos 3\tilde{\vartheta}+3\sin\tilde{\vartheta}\sin 3\tilde{\vartheta}-1),$
-$\Im(\Psi)$ $=-\partial U_{2}^{-}(\mathbf{0})(\sin\tilde{\vartheta}\sin 3\tilde{\vartheta}+3\cos\tilde{\vartheta}\cos 3\tilde{\vartheta}-3)$
-$+\partial_{2}U_{2}^{-}(\mathbf{0})(\cos\tilde{\vartheta}\sin 3\tilde{\vartheta}-3\sin\tilde{\vartheta}\cos 3\tilde{\vartheta}).$
-Next, we will show that $\Psi\not=0$. By contradiction, suppose that $\Re(\Psi)=\Im(\Psi)=0$, then we have
-$M\left(\begin{matrix}\partial_{1}U_{2}^{-}(\mathbf{0})\\ \partial_{2}U_{2}^{-}(\mathbf{0})\end{matrix}\right)=\mathbf{0},$
-where
-$M=\left(\begin{matrix}3\sin 3\tilde{\vartheta}\cos\tilde{\vartheta}-\sin\tilde{\vartheta}\cos 3\tilde{\vartheta}&\cos\tilde{\vartheta}\cos 3\tilde{\vartheta}+3\sin\tilde{\vartheta}\sin 3\tilde{\vartheta}-1\\ \sin\tilde{\vartheta}\sin 3\tilde{\vartheta}+3\cos\tilde{\vartheta}\cos 3\tilde{\vartheta}-3&\cos\tilde{\vartheta}\sin 3\tilde{\vartheta}-3\sin\tilde{\vartheta}\cos 3\tilde{\vartheta}\end{matrix}\right).$
-It can be deduced that
-$\det(M)=16\sin^{4}\tilde{\vartheta}\not=0\ \mbox{for}\ \tilde{\vartheta}\in(0,\pi),$
-where we get a contradiction. Hence, the proof of $\tilde{q}_{2}^{(0)}=\tilde{q}_{1}^{(0)}$ is complete.
-Substituting $\tilde{q}_{2}^{(0)}=\tilde{q}_{1}^{(0)}$ into (3.54), we obtain
-$\tilde{\sigma}_{1}$ $\int_{\mathcal{S}_{r_{0}}^{{}^{\prime}}}(\partial_{1}U_{2}^{-}(\mathbf{0})\tilde{x}_{1}^{2}+\partial_{2}U_{2}^{-}(\mathbf{0})\tilde{x}_{1}\tilde{x}_{2})e^{\rho\cdot\tilde{\mathbf{x}}}\mathrm{d}\tilde{\mathbf{x}}+\tilde{\sigma}_{2}\int_{\mathcal{S}_{r_{0}}^{{}^{\prime}}}(\partial_{1}U_{2}^{-}(\mathbf{0})\tilde{x}_{1}\tilde{x}_{2}+\partial_{2}U_{2}^{-}(\mathbf{0})\tilde{x}_{2}^{2})e^{\rho\cdot\tilde{\mathbf{x}}}\mathrm{d}\tilde{\mathbf{x}}$
-$\ =-\tilde{\sigma}_{1}\int_{\mathcal{S}_{r_{0}}^{{}^{\prime}}}(\partial_{1}U_{2}^{-}(\mathbf{0})\tilde{x}_{1}^{2}+\partial_{2}U_{2}^{-}(\mathbf{0})\tilde{x}_{1}\tilde{x}_{2})\psi(\tilde{\mathbf{x}})e^{\rho\cdot\tilde{\mathbf{x}}}\mathrm{d}\tilde{\mathbf{x}}-\tilde{\sigma}_{1}\int_{\mathcal{S}_{r_{0}}^{{}^{\prime}}}\tilde{x}_{1}\delta U_{2}^{-}\tilde{u}_{0}\mathrm{d}\mathbf{x}$
-$\ -\tilde{\sigma}_{2}\int_{\mathcal{S}_{r_{0}}^{{}^{\prime}}}(\partial_{1}U_{2}^{-}(\mathbf{0})\tilde{x}_{1}\tilde{x}_{2}+\partial_{2}U_{2}^{-}(\mathbf{0})\tilde{x}_{2}^{2})\psi(\tilde{\mathbf{x}})e^{\rho\cdot\tilde{\mathbf{x}}}\mathrm{d}\tilde{\mathbf{x}}-\tilde{\sigma}_{2}\int_{\mathcal{S}_{r_{0}}^{{}^{\prime}}}\tilde{x}_{2}\delta U_{2}^{-}\tilde{u}_{0}\mathrm{d}\tilde{\mathbf{x}}+\mathcal{I},$
-here $\tilde{\sigma}_{i}=k^{2}(\tilde{q}_{2}^{(i)}-\tilde{q}_{1}^{(i)}),i=1,2.$ Following a similar argument as in Case (I), we obtain a linear equation:
-$Q\left(\begin{matrix}\tilde{q}_{2}^{(1)}-\tilde{q}_{1}^{(1)}\\ \tilde{q}_{2}^{(2)}-\tilde{q}_{1}^{(2)}\end{matrix}\right)=\mathbf{0},$ (3.55)
-where
-$Q_{11}$ $=\left(\partial_{1}U_{2}^{-}(\mathbf{0})(-\frac{1}{2\mathrm{i}}+\frac{1}{3}(2\mathrm{i}\cos 2\vartheta-\sin 2\vartheta))+\frac{1}{3}\partial_{2}U_{2}^{-}(\mathbf{0})(2\mathrm{i}\sin 2\vartheta+\cos 2\vartheta)\right)e^{-4\mathrm{i}\vartheta}|_{0}^{\tilde{\vartheta}}$
-$Q_{12}$ $=\left(\frac{1}{3}\partial_{1}U_{2}^{-}(\mathbf{0})(2\mathrm{i}\sin 2\vartheta+\cos 2\vartheta)+\partial_{2}U_{2}^{-}(\mathbf{0})(-\frac{1}{2\mathrm{i}}-\frac{1}{3}(2\mathrm{i}\cos 2\vartheta-\sin 2\vartheta))\right)e^{-4\mathrm{i}\vartheta}|_{0}^{\tilde{\vartheta}}$
-$Q_{21}$ $=\left(\partial_{1}U_{2}^{-}(\mathbf{0})(\frac{1}{2\mathrm{i}}-\frac{1}{3}(\sin 2\vartheta+2\mathrm{i}\cos 2\vartheta))+\frac{1}{3}\partial_{2}U_{2}^{-}(\mathbf{0})(\cos 2\vartheta-2\mathrm{i}\sin\vartheta)\right)e^{4\mathrm{i}\vartheta}|_{0}^{\tilde{\vartheta}}$
-$Q_{22}$ $=\left(\frac{1}{3}\partial_{1}U_{2}^{-}(\mathbf{0})(\cos 2\vartheta-2\mathrm{i}\sin 2\vartheta)+\partial_{2}U_{2}^{-}(\mathbf{0})(\frac{1}{2\mathrm{i}}+\frac{1}{3}(\sin 2\vartheta+2\mathrm{i}\cos 2\vartheta))\right)e^{4\mathrm{i}\vartheta}|_{0}^{\tilde{\vartheta}}.$
-After tedious calculations, it is revealed that
-$\det(Q)$ $=\frac{\mathrm{i}}{9}(\partial_{1}U_{2}^{-}(\mathbf{0})\ \partial_{2}U_{2}^{-}(\mathbf{0}))\mathcal{A}\left(\begin{matrix}\partial_{1}U_{2}^{-}(\mathbf{0})\\ \partial_{2}U_{2}^{-}(\mathbf{0})\end{matrix}\right),$
-where
-$\mathcal{A}(:,1)=\left(\begin{matrix}-3\cos 4\tilde{\vartheta}-8\cos 2\tilde{\vartheta}-5\sin 4\tilde{\vartheta}\sin 2\tilde{\vartheta}+11\\ 6\sin 2\tilde{\vartheta}-6\sin 4\tilde{\vartheta}+3\sin 4\tilde{\vartheta}\cos 2\tilde{\vartheta}\end{matrix}\right)$
-and
-$\mathcal{A}(:,2)=\left(\begin{matrix}6\sin 2\tilde{\vartheta}-6\sin 4\tilde{\vartheta}+3\sin 4\tilde{\vartheta}\cos 2\tilde{\vartheta}\\ 3\cos 4\tilde{\vartheta}-8\cos 2\tilde{\vartheta}+\sin 4\tilde{\vartheta}\sin 2\tilde{\vartheta}+5\end{matrix}\right).$
-In the following, we will show that $\mathcal{A}$ is positive definite. In fact, we have the first order leading principle minors of $\mathcal{A}$ fulfills that
-$D_{1}(\mathcal{A})$ $=2(5\cos 2\tilde{\vartheta}+7)(\cos 2\tilde{\vartheta}-1)^{2}>0,\ \mbox{for}\ \tilde{\vartheta}\in(0,\pi).$
-When $\tilde{\vartheta}\in(\frac{\pi}{3},\frac{2\pi}{3})$, we can show that the second order leading principle minors of $\mathcal{A}$ satisfies that
-$D_{2}(\mathcal{A})$ $=8(2\cos 2\tilde{\vartheta}+1)(\cos 2\tilde{\vartheta}-1)^{5}>0.$
-Given that the conductive boundary parameter is a real constant and the refractive index is a linear polynomial with real coefficients, we can treat the PDE system (3.30) separately by considering its real and imaginary parts.
-This allows us to assume that the solutions $u_{1}$ and $u_{2}$ of (3.30) are real-valued by separating the real and imaginary parts of $u_{1}$ and $u_{2}$.
-Consequently, $(\partial_{1}u_{2}^{-}(\mathbf{0})\ \partial_{2}u_{2}^{-}(\mathbf{0}))^{\top}$ becomes a nonzero real vector.
-Leveraging the positivity of $\mathcal{A}$, we deduce that $Q$ is nonsingular, which can be used to prove that $\tilde{q}_{1}^{(1)}=\tilde{q}_{2}^{(1)}$ and $\tilde{q}_{1}^{(2)}=\tilde{q}_{2}^{(2)}$.
-By Proposition 3.1, we ascertain that $q_{1}=q_{2}$ and $\eta_{1}=\eta_{2}$.
-Case (ii): $\partial_{1}U_{2}^{-}(\mathbf{0})\not=0,\ \partial_{2}U_{2}^{-}(\mathbf{0})=U_{2}^{-}(\mathbf{0})=0.$ In view of (3.51), we observe that
-$(\eta_{2}-\eta_{1})\partial_{1}U_{2}^{-}(\mathbf{0})\int_{\tilde{\Gamma}_{r_{0}}^{\pm}}\tilde{x}_{1}e^{\rho\cdot{\tilde{\mathbf{x}}}}\mathrm{d}\sigma=\sum_{i=1}^{3}\mathcal{F}_{i}+\mathcal{I},$ (3.56)
-where
-$\mathcal{F}_{1}$ $=(\eta_{1}-\eta_{2})\partial_{1}U_{2}^{-}(\mathbf{0})\int_{\tilde{\Gamma}_{r_{0}}^{\pm}}\tilde{x}_{1}\psi(\tilde{\mathbf{x}})e^{\rho\cdot\tilde{\mathbf{x}}}\mathrm{d}\sigma,\ \mathcal{F}_{2}=(\eta_{1}-\eta_{2})\int_{\tilde{\Gamma}_{r_{0}}^{\pm}}\delta U_{2}^{-}\tilde{u}_{0}\mathrm{d}\sigma$
-$\mathcal{F}_{3}$ $=\int_{\mathcal{S}^{{}^{\prime}}_{r_{0}}}k^{2}(\tilde{q}_{1}-\tilde{q}_{2})U_{2}^{-}\tilde{u}_{0}\mathrm{d}\tilde{\mathbf{x}},\ \mathcal{I}=\int_{\Lambda_{r_{0}}^{{}^{\prime}}}\partial_{\nu}\tilde{u}_{0}\tilde{U}-\partial_{\nu}\tilde{U}\tilde{u}_{0}\mathrm{d}\sigma.$
-Similar to the proof of Case(b) in Theorem 3.1, using Proposition 2.1, it yields that the right-hand side of (3.56) satisfies
-$|RHS|\lesssim\tau^{-\frac{13}{6}}+\tau^{-(2+\alpha)}+\mathcal{O}(\tau^{-1}e^{-\frac{1}{2}\varsigma r_{0}\tau})+(1+\tau)(1+\tau^{-\frac{2}{3}})e^{-\varsigma r_{0}\tau}$ (3.57)
-and the left-hand side of (3.56) fulfills that
-$\overline{\eta}\int_{\tilde{\Gamma}_{r_{0}}^{\pm}}\tilde{x}_{1}e^{\rho\cdot\tilde{\mathbf{x}}}\mathrm{d}\sigma=\frac{\Gamma(2)\overline{\eta}}{\tau^{2}}\left(\frac{\cos\tilde{\vartheta}}{((\mathbf{d}+\mathrm{i}\mathbf{d}^{\perp})\cdot\hat{\tilde{\mathbf{x}}}_{1})^{2}}+\frac{1}{((\mathbf{d}+\mathrm{i}\mathbf{d}^{\perp})\cdot\hat{\tilde{\mathbf{x}}}_{2})^{2}}\right)+\mathcal{O}(\tau^{-1}e^{-\frac{1}{2}\lambda r_{0}\tau})$
-where $\overline{\eta}=(\eta_{2}-\eta_{1})\partial_{1}U_{2}^{-}(\mathbf{0})$ and $\hat{\tilde{\mathbf{x}}}_{1}\in\tilde{\Gamma}_{r_{0}}^{+},\hat{\tilde{\mathbf{x}}}_{2}\in\tilde{\Gamma}_{r_{0}}^{-}$.
-Let $\mathbf{d}$ and $\mathbf{d}^{\perp}$ be defined in (2.14), then one has
-$\left|\frac{\cos\tilde{\vartheta}}{((\mathbf{d}+\mathrm{i}\mathbf{d}^{\perp})\cdot\hat{\tilde{\mathbf{x}}}_{1})^{2}}+\frac{1}{((\mathbf{d}+\mathrm{i}\mathbf{d}^{\perp})\cdot\hat{\tilde{\mathbf{x}}}_{2})^{2}}\right|\not=0,\ \tilde{\vartheta}\in(0,\pi).$ (3.58)
-By virtue of (3.56)-(3.58), one can directly derive:
-$|\Gamma(2)\overline{\eta}|=0,$
-as $\tau\to\infty$. Consequently, $\eta_{1}=\eta_{2}$ follows from $\partial_{1}U_{2}^{-}(\mathbf{0})\neq 0$.
-Next, from (3.56), we have the following identity due to $\eta_{1}=\eta_{2}$:
-$k^{2}(\tilde{q}_{2}^{(0)}$ $-\tilde{q}_{1}^{(0)})\partial_{1}U_{2}^{-}(\mathbf{0})\int_{\mathcal{S}_{r_{0}}^{{}^{\prime}}}\tilde{x}_{1}e^{\rho\cdot\tilde{\mathbf{x}}}\mathrm{d}\tilde{\mathbf{x}}=k^{2}(\tilde{q}_{1}^{(0)}-\tilde{q}_{2}^{(0)})\partial_{1}U_{2}^{-}(\mathbf{0})\int_{\mathcal{S}_{r_{0}}^{{}^{\prime}}}\tilde{x}_{1}\psi(\tilde{\mathbf{x}})e^{\rho\cdot\tilde{\mathbf{x}}}\mathrm{d}\tilde{\mathbf{x}}$
-$+k^{2}(\tilde{q}_{1}^{(0)}-\tilde{q}_{2}^{(0)})\int_{\mathcal{S}_{r_{0}}^{{}^{\prime}}}\delta U_{2}^{-}\tilde{u}_{0}\mathrm{d}\tilde{\mathbf{x}}+k^{2}(\tilde{q}_{1}^{(1)}-\tilde{q}_{2}^{(1)})\int_{\mathcal{S}_{r_{0}}^{{}^{\prime}}}\tilde{x}_{1}U_{2}^{-}\tilde{u}_{0}\mathrm{d}\tilde{\mathbf{x}}$ (3.59)
-$+k^{2}(\tilde{q}_{1}^{(2)}-\tilde{q}_{2}^{(2)})\int_{\mathcal{S}_{r_{0}}^{{}^{\prime}}}\tilde{x}_{2}U_{2}^{-}\tilde{u}_{0}\mathrm{d}\tilde{\mathbf{x}}+\mathcal{I},$
-where $\mathcal{I}$ is given by (3.56). By (2.14) and direct computations, we can derive that the term on the left-hand side of (3) fulfils that
-$LHS=-\frac{\Gamma(3)((\sin\tilde{\vartheta}+3\mathrm{i}\cos\tilde{\vartheta})e^{3\mathrm{i}\tilde{\vartheta}}-3\mathrm{i})}{8\tau^{3}e^{3\mathrm{i}\varphi}}+\mathcal{O}(\tau^{-1}e^{-\frac{1}{2}\lambda r_{0}\tau}),$ (3.60)
-and the right-hand side of (3) satisfies
-$|RHS|\lesssim\tau^{-\frac{41}{12}}+\tau^{-(\alpha+3)}+(1+\tau)(1+\tau^{-\frac{2}{3}})e^{-\varsigma r_{0}\tau}+\mathcal{O}(\tau^{-1}e^{-\frac{1}{2}\varsigma r_{0}\tau})$ (3.61)
-for a sufficient large $\tau$. Additionally, since $\tilde{\vartheta}\in(0,\pi)$, we can deduce
-$(\sin\tilde{\vartheta}+3\mathrm{i}\cos\tilde{\vartheta})e^{3\mathrm{i}\tilde{\vartheta}}-3\mathrm{i}=-2\sin 2\tilde{\vartheta}(\cos 2\tilde{\vartheta}+1)+2(\cos 2\tilde{\vartheta})(\cos 2\tilde{\vartheta}-1)\mathrm{i}\not=0.$ (3.62)
-Subsequently, by combining (3) - (3.62), and multiplying both sides by $\tau^{3}$ before letting $\tau\to\infty$, we obtain
-$k^{2}|(\tilde{q}_{2}^{(0)}-\tilde{q}_{1}^{(0)})\partial_{1}U_{2}^{-}(\mathbf{0})|=0,$
-which indicates that $\tilde{q}_{1}^{(0)}=\tilde{q}_{2}^{(0)}$ since $|\partial_{1}U_{2}^{-}(\mathbf{0})|\not=0$.
-Substituting $\tilde{q}_{1}^{(0)}=\tilde{q}_{2}^{(0)}$ into (3), we have
-$k^{2}$ $(\tilde{q}_{2}^{(1)}-\tilde{q}_{1}^{(1)})\partial_{1}U_{2}^{-}(\mathbf{0})\int_{\mathcal{S}_{r_{0}}^{{}^{\prime}}}\tilde{x}_{1}^{2}e^{\rho\cdot\tilde{\mathbf{x}}}\mathrm{d}\tilde{\mathbf{x}}+k^{2}(\tilde{q}_{2}^{(2)}-\tilde{q}_{1}^{(2)})\partial_{1}U_{2}^{-}(\mathbf{0})\int_{\mathcal{S}_{r_{0}}^{{}^{\prime}}}\tilde{x}_{1}\tilde{x}_{2}e^{\rho\cdot\tilde{\mathbf{x}}}\mathrm{d}\tilde{\mathbf{x}}$
-$=k^{2}(\tilde{q}_{1}^{(1)}-\tilde{q}_{2}^{(1)})\partial_{1}U_{2}^{-}(\mathbf{0})\int_{\mathcal{S}_{r_{0}}^{{}^{\prime}}}\tilde{x}_{1}^{2}\psi(\tilde{\mathbf{x}})e^{\rho\cdot\tilde{\mathbf{x}}}\mathrm{d}\tilde{\mathbf{x}}+k^{2}(\tilde{q}_{1}^{(1)}-\tilde{q}_{2}^{(1)})\int_{\mathcal{S}_{r_{0}}^{{}^{\prime}}}\tilde{x}_{1}\delta U_{2}^{-}\tilde{u}_{0}\mathrm{d}\tilde{\mathbf{x}}$
-$+k^{2}(\tilde{q}_{1}^{(2)}-\tilde{q}_{2}^{(2)})\partial_{1}U_{2}^{-}(\mathbf{0})\int_{\mathcal{S}_{r_{0}}^{{}^{\prime}}}\tilde{x}_{1}\tilde{x}_{2}\psi(\tilde{\mathbf{x}})e^{\rho\cdot\mathbf{x}}\mathrm{d}\mathbf{x}$
-$+k^{2}(\tilde{q}_{1}^{(2)}-\tilde{q}_{2}^{(2)})\int_{\mathcal{S}_{r_{0}}^{{}^{\prime}}}\tilde{x}_{2}\delta U_{2}^{-}\tilde{u}_{0}\mathrm{d}\tilde{\mathbf{x}}+\mathcal{I}.$ (3.63)
-By Proposition 2.1, we know that the right-hand side of (3.63) satisfies
-$|RHS|\lesssim\tau^{-\frac{53}{12}}+\tau^{-(4+\alpha)}+\mathcal{O}(\tau^{-1}e^{-\frac{1}{2}\varsigma r_{0}\tau})+(1+\tau)(1+\tau^{\frac{2}{3}})e^{-\varsigma r_{0}\tau},$ (3.64)
-when $\tau$ is sufficient large. Moreover, we can deduce that the left-hand side of (3.63) is equal to
-$\sigma_{1}$ $\frac{\Gamma(4)e^{4\mathrm{i}\varphi}}{\tau^{4}}\frac{1}{4}\left(-\frac{1}{2\mathrm{i}}+\frac{1}{3}(2\mathrm{i}\cos 2\vartheta-\sin 2\vartheta)\right)e^{-4\mathrm{i}\vartheta}\bigg{|}_{0}^{\tilde{\vartheta}}$
-$\ +\sigma_{2}\frac{\Gamma(4)e^{4\mathrm{i}\varphi}}{\tau^{4}}\frac{1}{12}(2\mathrm{i}\sin 2\vartheta+\cos 2\vartheta)e^{-4\mathrm{i}\vartheta}\bigg{|}_{0}^{\tilde{\vartheta}}+\mathcal{O}(\tau^{-1}e^{-\frac{1}{2}\lambda r_{0}\tau}),$ (3.65)
-when $\mathbf{d}^{\perp}$ is defined in (2.13) and
-$\sigma_{1}$ $\frac{\Gamma(4)e^{-4\mathrm{i}\varphi}}{\tau^{4}}\frac{1}{4}\left(\frac{1}{2\mathrm{i}}-\frac{1}{3}(\sin 2\vartheta+2\mathrm{i}\cos 3\vartheta)\right)e^{4\mathrm{i}\vartheta}\bigg{|}_{0}^{\tilde{\vartheta}}$
-$\ +\sigma_{2}\frac{\Gamma(4)e^{-4\mathrm{i}\varphi}}{\tau^{4}}\frac{1}{12}(\cos 2\vartheta-2\mathrm{i}\sin 2\vartheta)e^{4\mathrm{i}\vartheta}\bigg{|}_{0}^{\tilde{\vartheta}}+\mathcal{O}(\tau^{-1}e^{-\frac{1}{2}\lambda r_{0}\tau}),$ (3.66)
-when $\mathbf{d}^{\perp}$ is defined in (2.14) , where $\sigma_{i}=k^{2}(\tilde{q}_{2}^{(i)}-\tilde{q}_{1}^{(i)})\partial_{1}U_{2}^{-}(\mathbf{0}),i=1,2.$ Combing (3.63)-(3.66) and following a similar argument for the Case (I), we derive that
-$D\left(\begin{matrix}\sigma_{1}\\ \sigma_{2}\end{matrix}\right)=\mathbf{0},$
-where
-$D=\left(\begin{matrix}\frac{1}{4}\left(-\frac{1}{2\mathrm{i}}+\frac{1}{3}(2\mathrm{i}\cos 2\vartheta-\sin 2\vartheta)\right)e^{-4\mathrm{i}\vartheta}|_{0}^{\tilde{\vartheta}}&\frac{1}{12}(2\mathrm{i}\sin 2\vartheta+\cos 2\vartheta)e^{-4\mathrm{i}\vartheta}|_{0}^{\tilde{\vartheta}}\\ \frac{1}{4}\left(\frac{1}{2\mathrm{i}}-\frac{1}{3}(\sin 2\vartheta+2\mathrm{i}\cos 2\vartheta)\right)e^{4\mathrm{i}\vartheta}|_{0}^{\tilde{\vartheta}}&\frac{1}{12}(\cos 2\vartheta-2\mathrm{i}\sin 2\vartheta)e^{4\mathrm{i}\vartheta}|_{0}^{\tilde{\vartheta}}\end{matrix}\right).$
-After careful and tedious calculations, we have
-$\det(D)=\frac{i}{72}(5\cos^{3}2\tilde{\vartheta}-3\cos^{2}\tilde{\vartheta}-9\cos 2\tilde{\vartheta}+7)\not=0,\ \mbox{for}\ \tilde{\vartheta}\in(0,\pi).$
-Therefore, it yields that $\sigma_{1}=\sigma_{2}$, which implies that $\tilde{q}_{1}^{(1)}=\tilde{q}_{2}^{(1)}$ and $\tilde{q}_{1}^{(2)}=\tilde{q}_{2}^{(2)}$.
-The proof of Case (ii) is complete.
-Case (iii): $\partial_{2}U_{2}^{-}(\mathbf{0})\not=0,\ \partial_{1}U_{2}^{-}(\mathbf{0})=\ U_{2}^{-}(\mathbf{0})=0.$ Direct calculations yield
-$(\eta_{2}-\eta_{1})\partial_{2}U_{2}^{-}$ $(\mathbf{0})\int_{\tilde{\Gamma}_{r_{0}}^{\pm}}\tilde{x}_{2}e^{\rho\cdot\tilde{\mathbf{x}}}\mathrm{d}\sigma=(\eta_{1}-\eta_{2})\partial_{2}U_{2}^{-}(\mathbf{0})\int_{\tilde{\Gamma}_{r_{0}}^{\pm}}x_{2}\psi(\tilde{\mathbf{x}})e^{\rho\cdot\tilde{\mathbf{x}}}\mathrm{d}\sigma$
-$+\int_{\mathcal{S}_{r_{0}}^{{}^{\prime}}}k^{2}(\tilde{q}_{1}-\tilde{q}_{2})(\partial_{2}U_{2}^{-}(\mathbf{0})\tilde{x}_{2}+\delta U_{2}^{-})\tilde{u}_{0}\mathrm{d}\tilde{\mathbf{x}}+\mathcal{I}.$ (3.67)
-Comparing (3.67) with (3.56), then we observe that the the right-hand side of (3.67) can be estimated by (3.57).
-Furthermore, by utilizing Proposition 2.1 again, we can deduce that the left-hand side of (3.67) satisfies
-$LHS=(\eta_{2}-\eta_{1})\frac{\Gamma(2)\partial_{2}U_{2}^{-}(\mathbf{0})}{\tau^{2}}\left\{\frac{\sin\tilde{\vartheta}}{((\mathbf{d}+\mathrm{i}\mathbf{d}^{\perp})\cdot\hat{\tilde{\mathbf{x}}})^{2}}\right\}+\mathcal{O}(\tau^{-1}e^{-\frac{1}{2}\lambda r_{0}\tau}),$ (3.68)
-where $\hat{\tilde{\mathbf{x}}}\in\tilde{\Gamma}_{r_{0}}^{+}$.
-Therefore, we can directly derive that $\eta_{1}=\eta_{2}$ since $\partial_{2}U_{2}^{-}(\mathbf{0})\neq 0$ and $\sin\tilde{\vartheta}\neq 0,\tilde{\vartheta}\in(0,\pi)$.
-Consequently, we obtain
-$k^{2}(\tilde{q}_{2}^{(0)}$ $-\tilde{q}_{1}^{(0)})\partial_{2}U_{2}^{-}(\mathbf{0})\int_{\mathcal{S}_{r_{0}}^{{}^{\prime}}}\tilde{x}_{2}e^{\rho\cdot\tilde{\mathbf{x}}}\mathrm{d}\tilde{\mathbf{x}}=k^{2}(\tilde{q}_{1}^{(0)}-\tilde{q}_{2}^{(0)})\partial_{2}U_{2}^{-}(\mathbf{0})\int_{\mathcal{S}_{r_{0}}^{{}^{\prime}}}\tilde{x}_{2}\psi(\tilde{\mathbf{x}})e^{\rho\cdot\tilde{\mathbf{x}}}\mathrm{d}\tilde{\mathbf{x}}$
-$+k^{2}(\tilde{q}_{1}^{(0)}-\tilde{q}_{2}^{(0)})\int_{\mathcal{S}_{r_{0}}^{{}^{\prime}}}\delta U_{2}^{-}\tilde{u}_{0}\mathrm{d}\tilde{\mathbf{x}}+k^{2}(\tilde{q}_{1}^{(1)}-\tilde{q}_{2}^{(1)})\int_{\mathcal{S}_{r_{0}}^{{}^{\prime}}}\tilde{x}_{1}U_{2}^{-}\tilde{u}_{0}\mathrm{d}\tilde{\mathbf{x}}$
-$+k^{2}(\tilde{q}_{1}^{(2)}-\tilde{q}_{2}^{(2)})\int_{\mathcal{S}_{r_{0}}^{{}^{\prime}}}{\tilde{x}}_{2}U_{2}^{-}\tilde{u}_{0}\mathrm{d}\tilde{\mathbf{x}}+\mathcal{I}.$ (3.69)
-Through meticulous calculation, we ascertain that the left-hand side of (3) equals
-$LHS=k^{2}(\tilde{q}_{2}^{(0)}-\tilde{q}_{1}^{(0)})\partial_{2}U_{2}^{-}(\mathbf{0})\frac{\Gamma(3)((\cos\tilde{\vartheta}-3\mathrm{i}\sin\tilde{\vartheta})e^{3\mathrm{i}\tilde{\vartheta}}-1)}{\tau^{3}e^{3\mathrm{i}\varphi}}+\mathcal{O}(\tau^{-1}e^{-\frac{1}{2}\lambda r_{0}\tau})$
-and the right-hand side of (3) fulfills that (3.61). Given that
-$(\cos\tilde{\vartheta}-3\mathrm{i}\sin\tilde{\vartheta})e^{3\mathrm{i}\tilde{\vartheta}}-1=-2\cos 2\tilde{\vartheta}(\cos 2\tilde{\vartheta}-1)+2\sin 2\tilde{\vartheta}(1-\cos 2\tilde{\vartheta})\mathrm{i}\not=0$
-for $\tilde{\vartheta}\in(0,\pi)$ and $\partial_{2}U_{2}^{-}(\mathbf{0})\not=0$, we can derive that
-$\tilde{q}_{2}^{(0)}-\tilde{q}_{1}^{(0)}=0.$
-The proof that $\tilde{q}_{2}^{(0)}=\tilde{q}_{1}^{(0)}$ is complete.
-Now we are going to prove $\tilde{q}_{1}^{(1)}=\tilde{q}_{2}^{(1)}$ and $\tilde{q}_{1}^{(2)}=\tilde{q}_{2}^{(2)}.$ It is easy to get
-$\tilde{\sigma}_{1}\int_{\mathcal{S}_{r_{0}}^{{}^{\prime}}}$ $\tilde{x}_{1}\tilde{x}_{2}e^{\rho\cdot\tilde{\mathbf{x}}}\mathrm{d}\tilde{\mathbf{x}}+\tilde{\sigma}_{2}\int_{\mathcal{S}_{r_{0}}^{{}^{\prime}}}\tilde{x}_{2}^{2}e^{\rho\cdot\tilde{\mathbf{x}}}\mathrm{d}\tilde{\mathbf{x}}=-\tilde{\sigma}_{1}\int_{\mathcal{S}_{r_{0}}^{{}^{\prime}}}\tilde{x}_{1}\tilde{x}_{2}\psi(\tilde{\mathbf{x}})e^{\rho\cdot\tilde{\mathbf{x}}}\mathrm{d}\tilde{\mathbf{x}}$
-$-\tilde{\sigma}_{2}\int_{\mathcal{S}_{r_{0}}^{{}^{\prime}}}\tilde{x}_{2}^{2}\psi(\tilde{\mathbf{x}})e^{\rho\cdot\tilde{\mathbf{x}}}\mathrm{d}\tilde{\mathbf{x}}+k^{2}(\tilde{q}_{1}^{(1)}-\tilde{q}_{2}^{(1)})\int_{\mathcal{S}_{r_{0}}^{{}^{\prime}}}\tilde{x}_{1}\delta U_{2}^{-}\tilde{u}_{0}\mathrm{d}\tilde{\mathbf{x}}$
-$+k^{2}(q_{1}^{(2)}-\tilde{q}_{2}^{(2)})\int_{\mathcal{S}_{r_{0}}^{{}^{\prime}}}\tilde{x}_{2}\delta U_{2}^{-}\tilde{u}_{0}\mathrm{d}\tilde{\mathbf{x}}+\mathcal{I},$ (3.70)
-here, $\tilde{\sigma}_{i}=k^{2}(\tilde{q}_{2}^{(i)}-\tilde{q}_{1}^{(i)})\partial_{2}U_{2}^{-}(\mathbf{0}),i=1,2$. It is clear that the right-hand side term of (3.70) satisfies (3.64).
-Moreover, the left-hand side term of (3.70) fulfills that
-$LHS$ $=\tilde{\sigma}_{1}\frac{\Gamma(4)e^{4\mathrm{i}\varphi}}{\tau^{4}}\frac{1}{12}(2\mathrm{i}\sin 2\vartheta+\cos 2\vartheta)e^{-4\mathrm{i}\vartheta}\bigg{|}_{0}^{\tilde{\vartheta}}$
-$+\tilde{\sigma}_{2}\frac{\Gamma(4)e^{4\mathrm{i}\varphi}}{\tau^{4}}\frac{1}{4}\left(-\frac{1}{2\mathrm{i}}-\frac{1}{6}(2\mathrm{i}\cos 2\vartheta-\sin 2\vartheta)\right)e^{-4\mathrm{i}\vartheta}\bigg{|}_{0}^{\tilde{\vartheta}}+\mathcal{O}(\tau^{-1}e^{-\frac{1}{2}\lambda r_{0}\tau})$
-when $\mathbf{d}^{\perp}=(-\sin\varphi,\cos\varphi)^{\top}$ and
-$LHS$ $=\tilde{\sigma}_{1}\frac{\Gamma(4)e^{-4\mathrm{i}\varphi}}{\tau^{4}}\frac{1}{12}(\cos\vartheta-\cos 2\mathrm{i}\sin\vartheta)e^{4\mathrm{i}\vartheta}\bigg{|}_{0}^{\tilde{\vartheta}}$
-$+\tilde{\sigma}_{2}\frac{\Gamma(4)e^{-4\mathrm{i}\varphi}}{\tau^{4}}\frac{1}{4}\left(\frac{1}{2\mathrm{i}}+\frac{1}{6}(2\mathrm{i}\cos 2\vartheta+\sin 2\vartheta)\right)e^{4\mathrm{i}\vartheta}\bigg{|}_{0}^{\tilde{\vartheta}}+\mathcal{O}(\tau^{-1}e^{-\frac{1}{2}\lambda r_{0}\tau})$
-when $\mathbf{d}^{\perp}=(\sin\varphi,-\cos\varphi)^{\top}$. For a sufficiently large $\tau$, it holds that
-$E\left(\begin{matrix}\tilde{\sigma}_{1}\\ \tilde{\sigma}_{2}\end{matrix}\right)=\mathbf{0},$
-where
-$E=\left(\begin{matrix}\frac{1}{12}(2\mathrm{i}\sin 2\vartheta+\cos 2\vartheta)e^{-4\mathrm{i}\vartheta}|_{0}^{\tilde{\vartheta}}&\frac{1}{4}\left(-\frac{1}{2\mathrm{i}}-\frac{1}{3}(2\mathrm{i}\cos 2\vartheta-\sin 2\vartheta)\right)e^{-4\mathrm{i}\vartheta}|_{0}^{\tilde{\vartheta}}\\ \frac{1}{12}(\cos 2\vartheta-2\mathrm{i}\sin\vartheta)e^{4\mathrm{i}\vartheta}|_{0}^{\tilde{\vartheta}}&\frac{1}{4}\left(\frac{1}{2\mathrm{i}}+\frac{1}{3}(2\mathrm{i}\cos 2\vartheta+\sin 2\vartheta)\right)e^{4\mathrm{i}\vartheta}|_{0}^{\tilde{\vartheta}}\end{matrix}\right).$
-Following meticulous calculations, it yields that
-$\det(E)=\frac{-\mathrm{i}}{144}(\cos\tilde{\vartheta}-1)^{3}\neq 0,\ \tilde{\vartheta}\in(0,\pi)$
-which indicates that $\tilde{\sigma}_{1}=\tilde{\sigma}_{2}=0$, implying $\tilde{q}_{1}^{(i)}=\tilde{q}_{2}^{(i)}$ for $i=1,2$ due to $\partial_{2}U_{2}^{-}(\mathbf{0})\neq 0$.
-Thus, the proof for Case (ii) is complete.
 Finally, we shall prove $u_{1}^{-}=u_{2}^{-}$ in $\mathcal{S}_{r_{0}}$. By the fact that $\eta_{1}=\eta_{2}$ and $q_{1}=q_{2}$ and the boundary condition, then we have
@@
 The proof is complete. ∎
-Remark 3.5.
-In the proof of Lemma 3.2, the assumption that the constant conductive boundary parameter is real and coefficients of the linear polynomial refractive index are real is only necessary for Case (i) of Case (I).
-However, if the conductive boundary parameter is a complex constant and the refractive index is a linear polynomial with complex coefficients, we speculate that the conclusion of Lemma 3.2 still holds in a generic scenario.
-In fact, in the final step of proving Case (iii), we demonstrated that the determinant of the coefficient matrix in (3.55) is quadratic, with the matrix $\mathcal{A}$ being positive.
-To address this scenario, let us consider a complex nonzero vector $(\partial_{1}U_{2}^{-}(\mathbf{0}),\ \partial_{2}U_{2}^{-}(\mathbf{0}))^{\top}$ and define
-${\bf a}=\Re\left(\left(\begin{matrix}\partial_{1}U_{2}^{-}(\mathbf{0})\\ \partial_{2}U_{2}^{-}(\mathbf{0})\end{matrix}\right)\right),\quad{\bf b}=\Im\left(\left(\begin{matrix}\partial_{1}U_{2}^{-}(\mathbf{0})\\ \partial_{2}U_{2}^{-}(\mathbf{0})\end{matrix}\right)\right).$
-Then $\det(Q)=0$ is equivalent to
-${\bf a}^{\top}\mathcal{A}{\bf a}={\bf b}^{\top}\mathcal{A}{\bf b},\quad{\bf a}^{\top}\mathcal{A}{\bf b}=0.$ (3.71)
-By defining the $\mathcal{A}$-inner product as $(\bf a,\bf b)_{\mathcal{A}}={\bf a}^{\top}\mathcal{A}{\bf b}$ and the $\mathcal{A}$-norm as $\|\bf a\|_{\mathcal{A}}=\sqrt{{\bf a}^{\top}\mathcal{A}{\bf a}}$, (3.71) can be equivalently expressed as
-$\|\bf a\|_{\mathcal{A}}=\|\bf b\|_{\mathcal{A}}$ and $({\bf a},{\bf b})_{\mathcal{A}}=0.$ (3.72)
-The entries of $\mathcal{A}$ depend solely on the opening angle of $\mathcal{S}_{r_{0}}$.
-It is important to note that $\bf a$ and $\bf b$ represent the real and imaginary parts of $(\partial_{1}U_{2}^{-}(\mathbf{0})\ \partial_{2}U_{2}^{-}(\mathbf{0}))^{\top}$, respectively, where $u_{2}^{-}(\mathbf{x})=U_{2}^{-}(M(\vartheta_{m})\mathbf{x})$ is a solution of (3.30).
-When applying Lemma 3.2 to prove Theorems 3.2 and 3.3, $u_{2}^{-}(\mathbf{x})=U_{2}^{-}(M(\vartheta_{m})\mathbf{x})$ represents the total wave in (1.7).
-From a physical perspective, (3.72) is rarely encountered, as it necessitates a peculiar algebraic relationship between the gradient of the total wave and $\mathcal{A}$.
-Consequently, $Q$ is generically nonsingular, enabling the unique determination of the linear polynomial reflective index with complex coefficients through a single far-field measurement, as demonstrated in Theorems 3.2 and 3.3.
-However, a rigorous proof of this speculation is fraught with significant difficulties and hence we defer it to our future study.
-Nevertheless, we would like to emphasize that the case with $q$ and $\eta$ being real-valued in the current article are physically unobjectionable.
 4. The proofs of Theorems 3.2 and 3.3
@@
 which is a contradiction.
-The proof is completed. ∎
+The proof is complete. ∎
 The proof of Theorem 3.3..

### v4

@@
 Let a sector $\mathcal{S}$ be defined as follows:
-$\mathcal{S}=\{\mathbf{x}\in\mathbb{R}^{2}~{}|~{}\vartheta_{m}<\arg(x_{1}+\mathrm{i}x_{2})<\vartheta_{M}\},$ (1.1)
+$\mathcal{S}=\{\mathbf{x}\in\mathbb{R}^{2}~{}|~{}\vartheta_{m}\leq\arg(x_{1}+\mathrm{i}x_{2})\leq\vartheta_{M}\},$ (1.1)
 where $-\pi\leq\vartheta_{m}<\vartheta_{M}<\pi$, $\vartheta_{M}-\vartheta_{m}\in(0,\pi)$ and $\mathrm{i}=\sqrt{-1}$. Additionally, the two boundary lines of $\mathcal{S}$ are given by
-$\Gamma^{+}=\{\mathbf{x}\in\mathbb{R}^{2}~{}|~{}\mathbf{x}=(r\cos\vartheta_{M},r\sin\vartheta_{M})^{\top},r>0\},$
-$\Gamma^{-}=\{\mathbf{x}\in\mathbb{R}^{2}~{}|~{}\mathbf{x}=(r\cos\vartheta_{m},r\sin\vartheta_{m})^{\top},r>0\}.$
+$\Gamma^{+}=\{\mathbf{x}\in\mathbb{R}^{2}~{}|~{}\mathbf{x}=(r\cos\vartheta_{M},r\sin\vartheta_{M})^{\top},r>0\},$ (1.2a)
+$\Gamma^{-}=\{\mathbf{x}\in\mathbb{R}^{2}~{}|~{}\mathbf{x}=(r\cos\vartheta_{m},r\sin\vartheta_{m})^{\top},r>0\}.$ (1.2b)
 Furthermore, define
@@
 If $\frac{\vartheta_{M}-\vartheta_{m}}{\pi}$ is a rational number, $\mathcal{S}_{r_{0}}$ is termed a rational corner; otherwise, it is an irrational corner.
-Theorem 1.1 shows a novel type of UCP result associated with a PDEs system described by (1.3). The detailed proof of Theorem 1.1 is given in Section 2.
+Theorem 1.1 shows a novel type of UCP result associated with a PDE system described by (1.4). The detailed proof of Theorem 1.1 is given in Section 2.
 Theorem 1.1.
-Let $D$ and $\Omega$ be two bounded Lipschitz domains in $\mathbb{R}^{2}$ with connected complements $\mathbb{R}^{2}\setminus\overline{D}$ and $\mathbb{R}^{2}\setminus\overline{\Omega}$ respectively, where $\Omega\subset D$.
+Let $D$ and $\Omega$ be two bounded Lipschitz domains in $\mathbb{R}^{2}$ with connected complements $\mathbb{R}^{2}\setminus\overline{D}$ and $\mathbb{R}^{2}\setminus\overline{\Omega}$ respectively, where $\overline{\Omega}\subset D$.
 Assume $\gamma_{1}\in\mathbb{R}_{+}$ and $\gamma_{2}\in L^{\infty}(D)\cap H^{2}(\Omega)$ with ${\rm supp}(\gamma_{2}-\gamma_{1})\subset\Omega$ are given.
@@
 The line segements $\Gamma^{\pm}_{r_{0}}$ are defined in (1.3).
-Consider $\mathbf{u}=(v,w)^{\top}\in H^{1}(D)\times H^{1}(D)$ solving the PDEs system:
+Consider $\mathbf{u}=(v,w)^{\top}\in H^{1}(D)\times H^{1}(D)$ solving the PDE system:
 $\begin{cases}\mathcal{P}\mathbf{u}=0\quad\mbox{in}\quad D,\\ \mathcal{B}\mathbf{u}=0\quad\mbox{on}\quad\Gamma^{\pm}_{r_{0}},\end{cases}$ (1.4)
@@
 It has important applications in the study of the inverse conductive medium scattering problem (1.8), which will be further elaborated in the subsequent subsection.
-The PDE system described by (1.3) in Theorem 1.1 originates from the invisibility phenomena associated with the time-harmonic conductive medium scattering problem (1.7), which will be further elaborated upon in subsection 1.2.
-The direct scattering problem (1.7) involves solving for the scattered wave $u^{s}$ given the incident wave $u^{i}$ and the conductive medium scatterer $(\Omega,q,\eta)$, where $\Omega$ is a bounded Lipschitz domain in $\mathbb{R}^{2}$ with a connected complement such that $\Omega\Subset B_{R}\ (R\in\mathbb{R}_{+})$ and $u^{i}$ can be either a plane wave of the form
-$u^{i}(\mathbf{x})=e^{{\rm i}k\mathbf{x}\cdot\mathbf{d}}$ (1.5)
+Remark 1.2.
+The PDE system described by (1.4) in Theorem 1.1 originates from the invisibility associated with the time-harmonic conductive medium scattering problem (1.6).
+The phenomenon of invisibility in inverse scattering problems has been extensively studied (see, e.g., [7, 8, 14, 12, 13, 23, 34, 46, 53, 56] and the references therein).
+A direct result of Theorem 1.1 is Corollary 1.2, which establishes that if a medium scatterer contains a convex irrational corner, invisibility cannot occur.
+In other words, for a scatterer to be invisible while possessing a convex corner, the corner must necessarily be rational, not irrational.
+This conclusion aligns with the specific example provided in [46, Example 1.5], where invisible anisotropic medium scatterers feature a rational corner with an angle of the form $\ell\pi/m$ for integers $m\geq 2$ and $11\leq\ell<2m-$ 1.
+Thus, Corollary 1.2 is nearly optimal, as demonstrated by the existence of invisible scatterers with rational corners.
+Moreover, the condition $\eta\neq 0$ in Theorem 1.1 can be interpreted as a nondegeneracy condition, which is necessary for establishing Theorem 1.1.
+Analogous nondegeneracy conditions have been proposed to characterize invisible medium scatterers; see [46, 12].
+The nondegeneracy condition [46, Equation (1.6)] for the invisible medium scatterer requires a certain non-vanishing property associated with the incident wave in a neighborhood of a point on the boundary of the scatterer.
+This condition is used to prove that the boundary of the invisible medium scatterer is $C^{1,\alpha}$ around the underlying point.
+On the contrary, if the invisible medium scatterer has a corner, the nondegeneracy condition is not satisfied, as demonstrated in [46, Examples 1.5 and 1.7].
+Such nondegeneracy conditions were also discussed in [12, Section 3.2].
+Therefore, the nondegeneracy condition is necessary for characterizing invisible medium scatterers in [12, 46].
+We then introduce the conductive medium scattering problem (1.6), which will be elaborated upon further in Subsection 1.2.
+Let $\Omega$ be a conductive medium scatterer suited in a homogenous isotropic background medium $\mathbb{R}^{2}$ with the medium configurations $q$ and $\eta$, which is a bounded Lipschitz domain with a connected complement $\mathbb{R}^{2}\setminus\overline{\Omega}$.
+The medium parameter $q\in L^{\infty}(\Omega)$ characterises the refractive index of the medium $\Omega$ with ${\rm supp}(q-1)\subset\Omega$ and the constant $\eta$ defined on $\partial\Omega$ signifies the conductive property of $\Omega$.
+In what follows, we write $(\Omega;q,\eta)$ to describe the conductive medium scatterer $\Omega$ with the physical configurations $q$ and $\eta$.
+For a given incident wave $u^{i}$ impinging on $\Omega$, it generates a scattered wave $u^{s}$.
+The total wave $u$ is given by $u=u^{i}+u^{s}$.
+The incident wave $u^{i}$ is a solution to the Helmholtz equation.
+The conductive medium scattering problem in $\mathbb{R}^{2}$ can be modeled as:
+[…]
+where $k\in\mathbb{R}_{+}$ is the wave number, and $\nu\in\mathbb{S}^{1}$ is the exterior unit normal vector to $\partial\Omega$.
+The third equation in (1.6) describes the conductive transmission property of the total wave $u$ across $\partial\Omega$.
+The final equation refers to the Sommerfeld radiation condition, characterizing the outgoing nature of the scattered wave $u^{s}$.
+Moreover, $u^{\pm}$ and $\partial_{\nu}^{\pm}u$ denote the limits of $u$ and $\partial_{\nu}u$ from the exterior and interior of $\Omega$, respectively.
+The direct scattering problem (1.6) involves solving for the scattered wave $u^{s}$ given the incident wave $u^{i}$ and the conductive medium scatterer $(\Omega,q,\eta)$, where $\Omega$ is a bounded Lipschitz domain in $\mathbb{R}^{2}$ with a connected complement such that $\Omega\Subset B_{R}\ (R\in\mathbb{R}_{+})$ and $u^{i}$ can be a solution to the Helmholtz equation $\Delta u^{i}+k^{2}u^{i}=0$ at least in a neighborhood of $\Omega$, including, a plane wave of the form
+$u^{i}(\mathbf{x})=e^{{\rm i}k\mathbf{x}\cdot\mathbf{d}}$
+or a Herglotz wave with kernel $g_{j}\in L^{2}(\mathbb{S}^{1})$ of the form
+$u^{i}(\mathbf{x})=\int_{\mathbb{S}^{1}}e^{\mathrm{i}k\xi\cdot\mathbf{x}}g_{j}(\xi)\mathrm{d}\sigma(\xi),\ \xi\in\mathbb{S}^{1},\ \mathbf{x}\in\mathbb{R}^{2}$
 or a point source of the form
-$u^{i}(\mathbf{x};\mathbf{z}_{0})=H_{0}^{1}(k|\mathbf{x}-\mathbf{z}_{0}|)$ (1.6)
+$u^{i}(\mathbf{x};\mathbf{z}_{0})=H_{0}^{(1)}(k|\mathbf{x}-\mathbf{z}_{0}|)$
 associated with a wave number $k\in\mathbb{R}_{+}$.
-Here, $\mathbf{d}\in\mathbb{S}^{1}$ is the incident direction of the plane wave, $H_{0}^{1}$ denotes the zeroth-order Hankel function of the first kind, and $\mathbf{z}_{0}$ signifies the location of the point source with $\mathbf{z}_{0}\in\mathbb{R}^{2}\setminus\overline{B}_{R}$ with $\Omega\Subset B_{R}$ and $R\in\mathbb{R}_{+}$.
+Here, $\mathbf{d}\in\mathbb{S}^{1}$ is the incident direction of the plane wave, $H_{0}^{(1)}$ denotes the zeroth-order Hankel function of the first kind, and $\mathbf{z}_{0}$ signifies the location of the point source with $\mathbf{z}_{0}\in\mathbb{R}^{2}\setminus\overline{B}_{R}$ with $\Omega\Subset B_{R}$ and $R\in\mathbb{R}_{+}$.
 The function $q\in L^{\infty}(\Omega)$ characterizes the refractive index of the medium scatterer $\Omega$ with ${\rm supp}(q)\subset\Omega$, and the constant conductive parameter $\eta$ defined on $\partial\Omega$ signifies the conductive property of $\Omega$.
@@
 The novel UCP derived in Theorem 1.1 has several applications in the conductive medium scattering problem (1.6).
-Specifically, in Theorem 1.2, utilizing the UCP for (1.3), it can be shown that the conductive medium scatterer in (1.7) containing a convex irrational corner cannot be invisible.
-The detailed proof of Theorem 1.2 is postponed to Section 2.
-Theorem 1.2.
-Consider the conductive medium scattering problem modeled by (1.7), associated with a given incident wave $u^{i}$ with the wave number $k$, as defined by (1.5) or (1.6).
+Specifically, in Corollary 1.2, utilizing the UCP for (1.4), it can be shown that the conductive medium scatterer in (1.6) containing a convex irrational corner cannot be invisible.
+The detailed proof of Corollary 1.2 is postponed to Section 2.
+Corollary 1.2.
+Consider the conductive medium scattering problem modeled by (1.6), associated with a given incident wave $u^{i}$ satisfying the Helmholtz equation with wave number $k$.
 Let the bounded Lipschitz domain $\Omega\subset\mathbb{R}^{2}$ with a connected complement in (1.6) describe the corresponding conductive medium scatterer $(\Omega,q,\eta)$, where ${\rm supp}(q)\subset\Omega$, $q\in L^{\infty}(\Omega)$ and $\eta$ is a nonzero constant.
@@
 Using the unique continuation property (UCP) for (1.4) as established in Theorem 1.1, we can address the unique identifiability for the inverse problem of the conductive medium scattering problem (1.6).
-Specifically, in Theorem 1.3, we present both local and global unique determination of a conductive medium scatterer based on a single measurement of the scattered wave, where the single measurement refers to that taken under a fixed incident wave.
+Specifically, in Corollary 1.3, we present both local and global unique determination of a conductive medium scatterer based on a single measurement of the scattered wave, where the single measurement refers to that taken under a fixed incident wave.
 The unique identifiability by a single measurement is challenging and has a long and colorful history in inverse scattering problems, often referred to as Schiffer problem in the literature; see [20, 19] and the references therein for more details.
-The detailed statement of Theorem 1.3 is provided in Theorem 2.1, along with the corresponding proof.
-Theorem 1.3.
-Considering the conductive medium scattering problem (1.7), if the measurements of the two scattered wave corresponding to two conductive medium scatterers under a fixed incident wave are identical, then the difference between these two scatterers cannot contain a convex irrational corner.
+The detailed statement of Corollary 1.3 is provided in Theorem 2.1, along with the corresponding proof.
+Before that, we introduce the concepts of the irrational angle, irrational polygon and convex irrational polygon.
+A similar definition for irrational polygons can be founded in [16].
+Definition 1.1.
+Suppose $\Omega\subset\mathbb{R}^{2}$ is a polygon.
+Let $\omega=\lambda\cdot\pi$, where $\lambda\in(0,2)$, be an interior angle of $\Omega$.
+The angle $\omega$ is termed irrational if $\lambda$ is an irrational number; otherwise, it is termed rational.
+Furthermore, if $\lambda\in(0,1)$ is an irrational number, the angle $\omega$ is referred to as a convex irrational angle.
+The polygon $\Omega$ is called an irrational polygon if all its interior angles are irrational; otherwise, it is considered rational.
+Additionally, the irrational polygon $\Omega$ is referred to as a convex irrational polygon if it is convex.
+Corollary 1.3.
+Considering the conductive medium scattering problem (1.6), if the measurements of the two scattered wave corresponding to two conductive medium scatterers, $\Omega_{1}$ and $\Omega_{2}$, under a fixed incident wave are identical, then the difference between these two scatterers, $\Omega_{1}\aleph\Omega:=(\text{Co}(\Omega_{1})\setminus\text{Co}({\Omega_{2}}))\cup(\text{Co}(\Omega_{2})\setminus\text{Co}({\Omega_{1}}))$, cannot contain a convex irrational corner, where $\text{Co}(\Omega_{i}),i=1,2$ is the convex hull of $\Omega_{i}$.
 Furthermore, if two convex irrational polygonal conductive medium scatterers, $\Omega_{1}$ and $\Omega_{2}$, produce the same scattered wave measurements for a fixed incident wave, then $\Omega_{1}$ and $\Omega_{2}$ must be identical.
@@
 Indeed, this kind of the conductive medium scattering problem can be derived by TM (transverse magnetic) polarisation from the corresponding Maxwell scattering system; see more details in [11, 15].
-Let $\Omega$ be a conductive medium scatterer suited in a homogenous isotropic background medium $\mathbb{R}^{2}$ with the medium configurations $q$ and $\eta$, which is a bounded Lipschitz domain with a connected complement $\mathbb{R}^{2}\setminus\overline{\Omega}$.
-The medium parameter $q\in L^{\infty}(\Omega)$ characterises the refractive index of the medium $\Omega$ with ${\rm supp}(q-1)\subset\Omega$ and the constant $\eta$ defined on $\partial\Omega$ signifies the conductive property of $\Omega$.
-In what follows, we write $(\Omega;q,\eta)$ to describe the conductive medium scatterer $\Omega$ with the physical configurations $q$ and $\eta$.
-For a given incident wave $u^{i}$ impinging on $\Omega$, it generates a scattered wave $u^{s}$.
-The total wave $u$ is given by $u=u^{i}+u^{s}$.
-The incident wave $u^{i}$ can be either a plane wave of the form (1.5) or a point source of the form (1.6).
-The conductive medium scattering problem in $\mathbb{R}^{2}$ can be modeled as:
-$\begin{cases}\Delta u^{-}+k^{2}qu^{-}=0&\text{in }\Omega,\\[5.0pt] \Delta u^{+}+k^{2}u^{+}=0&\text{in }\mathbb{R}^{2}\setminus\overline{\Omega},\\[5.0pt] u^{+}=u^{-},\quad\partial_{\nu}u^{-}=\partial_{\nu}u^{+}+\eta u^{+}&\text{on }\partial\Omega,\\[5.0pt] u^{+}=u^{i}+u^{s}&\text{in }\mathbb{R}^{2}\setminus\overline{\Omega},\\[5.0pt] \lim_{r\to\infty}r^{1/2}\left(\partial_{r}u^{s}-iku^{s}\right)=0,&\quad r=|\mathbf{x}|,\end{cases}$ (1.7)
-where $k\in\mathbb{R}_{+}$ is the wave number, and $\nu\in\mathbb{S}^{1}$ is the exterior unit normal vector to $\partial\Omega$.
-The third equation in (1.7) describes the conductive transmission property of the total wave $u$ across $\partial\Omega$.
-The final equation refers to the Sommerfeld radiation condition, characterizing the outgoing nature of the scattered wave $u^{s}$.
-Moreover, $u^{\pm}$ and $\partial_{\nu}^{\pm}u$ denote the limits of $u$ and $\partial_{\nu}u$ from the exterior and interior of $\Omega$, respectively.
 The direct problem (1.6) aims to solve the scattered wave $u^{s}$ given the conductive medium scatterer $(\Omega;q,\eta)$ and the incident wave $u^{i}$.
@@
 where $V=q-1$ and $\Gamma=\partial\Omega$.
-On the other hand, given the bounded Lipschitz domain $\Omega$, $V\in L^{\infty}(\Omega)$ and the boundary parameter $\eta\in L^{\infty}(\partial\Omega)$, consider the PDEs system (1.11).
+On the other hand, given the bounded Lipschitz domain $\Omega$, $V\in L^{\infty}(\Omega)$ and the boundary parameter $\eta\in L^{\infty}(\partial\Omega)$, consider the PDE system (1.10).
 If there exists a nontrivial pair of solutions $(w,v)$ to (1.10) associated with $k$, then $k$ is referred to as a transmission eigenvalue and $(w,v)$ is the corresponding pair of transmission eigenfunctions of (1.10).
@@
 Hence, $(\Omega;q,\eta)$ is actually a generalized conductive medium scatterer and the transmission eigenvalue problem (1.10) is generalized transmission eigenvalue problem in our study, but we shall not distinguish it.
+The boundary localization for an acoustic-elastic transmission eigenfunctions is revealed in [27].
 In the following, we shall investigate the connection between the PDE system (1.4) in Theorem 1.1 and the conductive medium scattering (1.6) when the invisibility of the conductive medium scatterer $(\Omega;q,\eta)$ occurs.
-Recall that the incident wave $u^{i}$ is given by (1.5) or (1.6).
+Recall that the incident wave $u^{i}$ satisfying the Helmholtz equation.
 Since $\Omega$ is bounded and has a connected complement, let $D:=B_{R}$ such that $\Omega\Subset B_{R}$, where $B_{R}$ is a disk centered at the origin with radius $R$.
@@
 Moreover, [52] derived UCP results for elliptic PDE systems, including strong UCP, weak UCP, and UCP for local Cauchy data.
+The UCP for the Maxwells system can be founded in [32].
 As discussed in the previous subsection, the invisibility or transparency in inverse scattering problems implies that the scatterer cannot be detected by external measurements, as the probing incident wave remains unperturbed and the corresponding scattered field is identically zero.
-However, when the scatterer exhibits geometrical singularities, such as corners or points of high curvature, as shown in [8, 12, 44, 24, 6] for time-harmonic acoustic scattering, invisibility or transparency cannot occur for any frequency associated with an incident plane wave.
+However, when the scatterer exhibits geometrical singularities, such as corners or points of high curvature, as shown in [8, 14, 24, 53, 34, 7] for time-harmonic acoustic scattering, invisibility or transparency cannot occur for any frequency associated with an incident plane wave.
 Recently, new methods based on free boundary problems have been introduced to investigate invisibility or transparency in isotropic and anisotropic acoustic medium scattering [12, 56, 13, 46].
-The quest to determine the geometrical and physical properties of scatterers through a single far-field measurement boasts a rich and extensive history in the realm of inverse scattering [2, 7, 18, 19, 34, 41].
+The quest to determine the geometrical and physical properties of scatterers through a single far-field measurement boasts a rich and extensive history in the realm of inverse scattering [2, 6, 19, 20, 43, 50, 28].
 This pursuit, often referred to as the Schiffer problem [20], aims to uniquely identify the location and shape of scatterers based on a single far-field measurement.
@@
 In the inverse obstacle scattering, Lax and Phillips [48] demonstrated that the shape of a sound-soft obstacle can be uniquely identified through infinite far-field measurements.
-However, achieving uniqueness with just a single far-field measurement has thus far required additional a-priori knowledge of the geometry, such as the size or shape of the scatterer(cf.
-[20, 17, 41]).
+However, achieving uniqueness with just a single far-field measurement has thus far required additional a-priori knowledge of the geometry, such as the size or shape of the scatterer( cf.[21, 18, 50]).
+The unique identifiability for both inverse acoustic, elastic and electromagnetic impenetrable or penetrable scatterers under a-priori geometric knowledge using a few far-field measurements or single far-field measurement can be founded in [31, 25, 26, 33, 32].
+Recently, the first stability estimating for simultaneous determination of the acoustic polygonal obstacle and the boundary impedance through a single far-field measurement is established in [30].
+Furthermore, the effective medium theory for a time-harmonic electromagnetic inhomogeneous and possibly anisotropic medium with embedded obstacles can be founded in [29].
 In the context of conductive medium scattering described by (1.6), uniqueness results for determining $\Omega$ based on the far-field patterns of all incident plane waves at a fixed frequency, namely infinitely many far-field measurements, have been obtained in [11, 10, 39].
@@
   1. a)
-    We derive a novel type of UCP result for an elliptic second order PDEs system (1.3), where the components $v$ and $w$ of the solution $\mathbf{u}$ to (1.3) satisfy a transmission condition across two intersecting line segments within the underlying domain.
+    We derive a novel type of UCP result for an elliptic second order PDE system (1.4), where the components $v$ and $w$ of the solution $\mathbf{u}$ to (1.4) satisfy a transmission condition across two intersecting line segments within the underlying domain.
 These two intersecting line segments form an irrational corner.
   2. b)
-    We reveal that when the conductive medium scatter $(\Omega,q,\eta)$ has a convex irrational corner $\mathbf{x}_{0}$ and the physical parameter $q$ satisfies $H^{2}$ regularity near $\mathbf{x}_{0}$, then $\Omega$ radiates any wave number.
+    We reveal that when the conductive medium scatter $(\Omega,q,\eta)$ has a convex irrational corner $\mathbf{x}_{0}$ and the physical parameter $q$ satisfies $H^{2}$ regularity near $\mathbf{x}_{0}$, then $\Omega$ radiates at any wave number.
 Furthermore, we demonstrate that an admissible complex irrational polygonal scatterer can be uniquely determined using only a single far-field measurement, independent of its medium content.
@@
 We will discuss this aspect in more detail in Section 2.
-To the best of our knowledge, this represents a novel type of UCP for second-order elliptic PDEs systems, playing a crucial role in investigating invisibility and uniqueness results in inverse conductive medium scattering problems.
+To the best of our knowledge, this represents a novel type of UCP for second-order elliptic PDE systems, playing a crucial role in investigating invisibility and uniqueness results in inverse conductive medium scattering problems.
 For the main contribution b), we can substitute the technical condition on the a-priori information in [23], which requires the non-vanishing of the total wave at the corner points, with an irrationality condition related to the interior angles of the conductive polygonal medium $\Omega$.
@@
 We establish this uniqueness result by using the novel UCP result described in Theorem 1.1.
-Furthermore, using the novel UCP established in Theorem 1.1, we can show that if a conductive medium scatterer $\Omega$ has a convex irrational corner on the boundary, where the reflective index $q$ has $H^{2}$ regularity near the corner, then $\Omega$ radiates any incident wave $u^{i}$ defined by (1.5) or (1.6).
+Furthermore, using the novel UCP established in Theorem 1.1, we can show that if a conductive medium scatterer $\Omega$ has a convex irrational corner on the boundary, where the reflective index $q$ has $H^{2}$ regularity near the corner, then $\Omega$ radiates any incident wave $u^{i}$ which is a solution to the Helmholtz equation with wave number $k$.
 For the main contribution c), we establish a local uniqueness result for determining the shape of the polygonal-nest or polygonal-cell conductive medium scatterer.
@@
 The remainder of the paper is organized as follows.
-In Section 2, we present the proofs of Theorems 1.1-1.3.
+In Section 2, we present the proofs of Theorems 1.1 and Corollaries 1.2-1.3.
 In Section 3, we present our main results regarding the uniqueness of recovering the conductive polygonal-nest and polygonal-cell medium scatterers, including the scatterers shape, the partition of their interior polygonal-nest and polygonal-cell structure and the material parameters.
 Section 4 is devoted to the proofs of the main theorems of Section 3.
-2. The proofs of Theorems 1.1-1.3
-In this section, we aim to provide the proofs of Theorems 1.1-1.3.
-Specifically, we present a novel UCP result for an elliptic second-order PDEs system and establish both local and global uniqueness results for the shape of an admissible complex conductive polygonal medium scatterer $\Omega$, utilizing only a single far-field measurement.
+2. The proofs of Theorem 1.1 and Corollaries 1.2-1.3
+In this section, we aim to provide the proofs of Theorems 1.1 and Corollaries 1.2-1.3.
+Specifically, we present a novel UCP result for an elliptic second-order PDE system and establish both local and global uniqueness results for the shape of an admissible complex conductive polygonal medium scatterer $\Omega$, utilizing only a single far-field measurement.
 To prove Theorem 1.1, we first introduce some auxiliary lemmas.
@@
 Lemma 2.2.
-[18, Section 3.4] Suppose that $v$ is a solution to Helmholtz equation $(\Delta+\gamma_{1})v=0$ in $\mathcal{S}_{r_{0}}$, where $\gamma_{1}\in\mathbb{R}_{+}$ and $\mathcal{S}_{r_{0}}$ is defined in (1.2).
-If $v$ is analytic in $\mathcal{S}_{r_{0}}$, then $v$ has the following spherical wave expansion in polar coordinates around the origin:
+[19, Section 3.4] Suppose that $v$ is a solution to Helmholtz equation $(\Delta+\gamma_{1})v=0$ in $B_{r_{0}}$, where $\gamma_{1}\in\mathbb{R}_{+}$.
+Then $v$ has the following spherical wave expansion in polar coordinates around the origin:
 $v=\sum_{n=0}^{\infty}(a_{n}e^{\mathrm{i}n\vartheta}+b_{n}e^{-\mathrm{i}n\vartheta})J_{n}(\sqrt{\gamma_{1}}r),\ r=|\mathbf{x}|,$ (2.6)
@@
 $\begin{cases}\Delta w+\gamma_{2}w=0\hskip 54.34486pt\mbox{in}\ \mathcal{S}_{r_{0}},\\ \Delta v+\gamma_{1}v=0\hskip 59.18178pt\mbox{in}\ \mathcal{S}_{r_{0}},\\ w=v,\ \partial_{\nu}w=\partial_{\nu}v+\eta v\hskip 11.38092pt\mbox{on}\ \Gamma_{r_{0}}^{\pm}.\end{cases}$ (2.7)
-By the interior regularity of the elliptic equation, we know that $v$ is analytic in $B_{r_{0}}$. When $v$ is analytic, it was proven in [22, Theorem 3.2] that $v(\mathbf{0})=0$. Using (2.6), one has
+By the interior regularity of the elliptic equation, we know that $v$ is analytic in $B_{r_{0}}$. When $v$ is analytic, it was proven in [24, Theorem 3.2] that $v(\mathbf{0})=0$. Using (2.6), one has
 $a_{0}+b_{0}=0,$
@@
 Namely, we shall prove that $a_{\ell}=b_{\ell}=0$ for $\ell\in\mathbb{N}$.
-Since $\gamma_{2}\in H^{2}(\mathcal{S}_{r_{0}})$, let $\widetilde{\gamma_{2}}$ be the Sobolev extension of $\gamma_{2}$ in $\mathbb{R}^{2}$.
+Since $\gamma_{2}\in H^{2}(\mathcal{S}_{r_{0}})$, let $\widetilde{\gamma_{2}}$ be the Sobolev extension of $\gamma_{2}$ in $H^{2}(\mathbb{R}^{2})$.
 According to Lemma 2.1, there exists the CGO solution $u_{0}$ given by (2.1), which satisfies
@@
 where
-$A=\left(\begin{matrix}e^{-\mathrm{i}\vartheta_{m}}+e^{-\mathrm{i}\vartheta_{M}}&e^{-3\mathrm{i}\vartheta_{m}}+e^{-3\mathrm{i}\vartheta_{M}}\\ e^{3\mathrm{i}\vartheta_{m}}+e^{-3\mathrm{i}\vartheta_{M}}&e^{\mathrm{i}\vartheta_{m}}+e^{\mathrm{i}\vartheta_{M}}\end{matrix}\right).$
+$A=\left(\begin{matrix}e^{-\mathrm{i}\vartheta_{m}}+e^{-\mathrm{i}\vartheta_{M}}&e^{-3\mathrm{i}\vartheta_{m}}+e^{-3\mathrm{i}\vartheta_{M}}\\ e^{3\mathrm{i}\vartheta_{m}}+e^{3\mathrm{i}\vartheta_{M}}&e^{\mathrm{i}\vartheta_{m}}+e^{\mathrm{i}\vartheta_{M}}\end{matrix}\right).$
 Since $\vartheta_{M}-\vartheta_{m}\not=\frac{\pi}{2}$ by noting $\mathcal{S}_{r_{0}}$ is irrational, after some calculations, it shows that
@@
 The proof is complete. ∎
-In the following we will prove Theorem 1.2.
-The proof of Theorem 1.2.
+In the following we will prove Corollary 1.2.
+The proof of Corollary 1.2.
 We prove this theorem using a contradiction argument.
@@
 Here $D:=B_{R}$ is chosen such that $\Omega\Subset D$.
-By applying Theorem 1.1, we conclude that $u^{i}\equiv 0$ in $D$, leading to a contradiction by unique continuation principle.
+By applying Theorem 1.1, we conclude that $u^{i}\equiv 0$ in $D$, which leads to a contradiction by the unique continuation principle.
 ∎
 To address the inverse problem (1.9), it is essential to define the admissible class of scatterers.
-Definition 2.1 introduces the concept of an irrational polygon, which will be used in Definition 2.2.
-Definition 2.1.
-Suppose $\Omega\subset\mathbb{R}^{2}$ is a polygon.
-Let $\omega=\lambda\cdot\pi$, where $\lambda\in(0,2)$, be an interior angle of $\Omega$.
-An angle $\omega$ is considered irrational if $\lambda$ is an irrational number; otherwise, it is considered rational.
-Furthermore, if $\lambda\in(0,1)$ is an irrational number, then the angle $\omega$ is considered convex irrational.
-The polygon $\Omega$ is termed an irrational polygon if each of its interior angles is an irrational angle; otherwise, it is considered rational.
+Recall that Definition 1.1 introduces the concept of an irrational polygon, which will be utilized in Definition 2.1.
 Definition 2.1.
 We define $\Omega$ as an admissible complex scatterer if it comprises finitely many disjointed scatterers, denoted as follows:
-$(\Omega;q,\eta)=\bigcup_{\ell=1}^{N}(\Omega_{\ell};q_{\ell},\eta_{\ell}),\ \Omega_{\ell}\cap\Omega_{m}=\emptyset\ \mbox{for}\ \ell\not=m,$
+$(\Omega;q,\eta)=\bigcup_{\ell=1}^{N}(\Omega_{\ell};q_{\ell},\eta_{\ell}),\ \overline{\Omega_{\ell}}\cap\overline{\Omega_{m}}=\emptyset\ \mbox{for}\ \ell\not=m,$
 where $N\in\mathbb{N}$, and each $\Omega_{\ell}$ is a bounded Lipschtiz domain.
@@
 Similarly, if each $\Omega_{\ell}$ is a convex irrational polygon, then $\Omega$ is considered to be an admissible complex convex irrational polygonal scatterer.
-The following theorem provides a local uniqueness result for determining the shape of a scatterer through a single far-field measurement, offering a more rigorous version of Theorem 1.3.
+The following theorem provides a local uniqueness result for determining the shape of a scatterer through a single far-field measurement, offering a more rigorous version of Corollary 1.3.
 It asserts that the difference set between two admissible complex polygonal scatterers cannot contain a convex irrational corner if the far-field measurements associated with these two scatterers, for a fixed incident wave, are identical.
@@
 Let $u_{\ell}^{s}\in H^{1}_{\rm loc}(\mathbb{R}^{2})$ represent the scattered wave associated with $u^{i}$ and an admissible complex polygonal scatterer $(\Omega_{\ell};q_{\ell},\eta_{\ell})$ be defined in Definition 2.1, where
-$(\Omega_{\ell};q_{\ell},\eta_{\ell})=\bigcup_{i=1}^{N_{\ell}}(\Omega_{i,\ell};q_{i,\ell},\eta_{i,\ell}),\quad N_{\ell}\in\mathbb{N}{+},\quad\ell=1,2.$
+$(\Omega_{\ell};q_{\ell},\eta_{\ell})=\bigcup_{i=1}^{N_{\ell}}(\Omega_{i,\ell};q_{i,\ell},\eta_{i,\ell}),\quad N_{\ell}\in\mathbb{N}_{+},\quad\ell=1,2.$
 Let $u^{\infty}_{\ell}$ $(\ell=1,2)$, defined in (1.7), be the far-field pattern corresponding to $u_{\ell}^{s}$, respectively.
@@
 $u^{\infty}_{1}(\mathbf{\hat{x}};u^{i})=u^{\infty}_{2}(\mathbf{\hat{x}};u^{i}),$ (2.18)
-for all $\mathbf{\hat{x}}\in\mathbb{S}^{1}$ and a fixed incident wave $u^{i}$, then $\Omega_{1}\Delta\Omega_{2}=(\Omega_{1}\backslash\Omega_{2})\cup(\Omega_{2}\backslash\Omega_{1})$ cannot contain a convex irrational corner.
+for all $\mathbf{\hat{x}}\in\mathbb{S}^{1}$ and a fixed incident wave $u^{i}$, then the difference $\Omega_{1}\aleph\Omega_{2}$ defined in Corollary 1.3 cannot contain a convex irrational corner.
 Moreover, if (2.18) is satisfied and $\Omega_{\ell}$ is an admissible complex convex irrational polygonal scatterer $(\ell=1,2)$, then $\Omega_{1}=\Omega_{2}$.
@@
 We prove the first conclusion of this theorem by contradiction.
-When (2.18) holds, assume that $\Omega_{1}\Delta\Omega_{2}$ contains an irrational corner.
-Underlying the fact that $\Delta$ is variant under the rigid motion and by the definition of an admissible complex irrational polygonal scatterer, we also assume that there exists an irrational corner $\mathcal{S}_{r_{0}}$ such that $\mathcal{S}_{r_{0}}\Subset\Omega_{1}\setminus\overline{\Omega_{2}}$, where $r_{0}\in\mathbb{R}_{+}$ is sufficiently small.
-Let $B_{r_{0}}$ be a disk with $r_{0}\in\mathbb{R}_{+}$ be chosen such that $B_{r_{0}}\Subset\Omega_{1}\setminus\overline{\Omega_{2}}$ and $\mathcal{S}_{r_{0}}\subset B_{r_{0}}$.
+When (2.18) holds, assume that $\Omega_{1}\aleph\Omega_{2}$ contains an irrational corner.
+Underlying the fact that the Laplacian $\Delta$ is invariant under the rigid motion and by the definition of an admissible complex irrational polygonal scatterer, we also assume that there exists an irrational corner $\mathcal{S}_{r_{0}}$ such that $\mathcal{S}_{r_{0}}\Subset\Omega_{1}\setminus\overline{\Omega_{2}}$, where $r_{0}\in\mathbb{R}_{+}$ is sufficiently small.
+Let $B_{r_{0}}$ be a disk with radius $r_{0}\in\mathbb{R}_{+}$.
 Let $u_{1}(\mathbf{x})$ and $u_{2}(\mathbf{x})$ be total wave fields associated $(\Omega_{1};q_{1},\eta_{1})$ and $(\Omega_{2};q_{2},\eta_{2})$, respectively.
@@
 Here, $u^{\infty}(\mathbf{\hat{x}};u^{i})$ is the far-field pattern given by (1.7), associated with the total wave $u$, where $u$ is the solution of (1.6) corresponding to the conductive polygonal-cell medium scatterer $\bigcup_{i=1}^{N}(\Sigma_{i};q_{i},\tilde{\eta})$.
-In order to introduce the admissible class for studying the inverse problems of the unique identification of the shape, we need to study the Hölder-regularity of the solution of the coupled PDEs system (3.7) near the corner, where $u_{1}$ and $u_{2}$ are the total wave field corresponding to the two different scatterers $\Omega_{1}$ and $\Omega_{2}$, respectively.
-Indeed, the PDEs system (3.7) is derived from the proceed of proving the unique determination of the shape of the scatterer by contradiction (see (3.12) in the proof of Theorem 3.1).
+In order to introduce the admissible class for studying the inverse problems of the unique identification of the shape, we need to study the Hölder-regularity of the solution of the coupled PDE system (3.7) near the corner, where $u_{1}$ and $u_{2}$ are the total wave field corresponding to the two different scatterers $\Omega_{1}$ and $\Omega_{2}$, respectively.
+Indeed, the PDE system (3.7) is derived from the proceed of proving the unique determination of the shape of the scatterer by contradiction (see (3.13) in the proof of Theorem 3.1).
+Recall that $\mathcal{S}_{r_{0}}$ is defined by (1.3) with two line segment boundaries $\Gamma^{\pm}_{r_{0}}$.
+We can similarly define $\mathcal{S}_{tr_{0}}$ for $t\in\mathbb{R}_{+}$.
 Lemma 3.1.
@@
 Suppose that $u_{1},u_{2}\in H^{1}(\mathcal{S}_{2r_{0}})$ satisfy
+$\begin{cases}(\Delta+k^{2}q)u_{1}=0\hskip 28.45274pt\mbox{in}\ \mathcal{S}_{2r_{0}},\\ (\Delta+k^{2})u_{2}=0\hskip 34.14322pt\mbox{in}\ B_{2r_{0}},\\ u_{1}=u_{2}\hskip 68.28644pt\mbox{on}\ \Gamma_{2r_{0}}^{\pm},\end{cases}$ (3.7)
+where $k$ is a positive constant and $q$ is in the form of (3.1).
+There exists a positive number $t_{0}^{\prime}\in\mathbb{R}_{+}$ such that $u_{1}\in C^{1,\alpha}(\overline{\mathcal{S}_{t_{0}^{\prime}r_{0}}})$, where $\alpha\in(0,1)$ depends on the opening angle of $\mathcal{S}_{2r_{0}}$.
+Proof.
+For any given $t\in\mathbb{R}_{+}$, let $\mathbf{y}_{t,+}\in\Gamma^{+}$ and $\mathbf{y}_{t,-}\in\Gamma^{-}$ be points satisfying $|\mathbf{y}_{t,+}|=|\mathbf{y}_{t,-}|=t$, where $\Gamma^{+}$ and $\Gamma^{-}$ are defined by (1.2a) and (1.2b), respectively.
+For any $t\in\mathbb{R}_{+}$, we introduce the triangle $V_{tr_{0}}$ formed by the three line segments connecting $\mathbf{0}$, $\mathbf{y}_{t,+}$, and $\mathbf{y}_{t,-}$.
+Let $\varphi\in C^{\infty}({\mathcal{S}_{2r_{0}}})$ be a smooth cutoff function satisfying that $\varphi\equiv 1$ in ${\mathcal{S}_{t_{0}r_{0}}}$ and $\varphi\equiv 0$ in $\mathcal{S}_{2r_{0}}\setminus\overline{\mathcal{S}_{t_{0}r_{0}}}$, where $t_{0}\in(0,2)$ such that $\mathcal{S}_{t_{0}r_{0}}\Subset V_{2r_{0}}$.
+Let $w=u_{1}-u_{2}$ and define $\tilde{\varphi}=w\varphi$.
+Then there holds that
+$\Delta\tilde{\varphi}=f\ \mbox{in}\ V_{2r_{0}},\quad\tilde{\varphi}=0\ \mbox{on}\ \partial{V_{2{r_{0}}}},$
+where $f=\Delta w\varphi+w\Delta\varphi+2\nabla w\cdot\nabla\varphi=(k^{2}u_{2}-k^{2}qu_{1})\varphi+w\Delta\varphi+2\nabla w\cdot\nabla\varphi.$ Since $u_{1}\in H^{1}(\mathcal{S}_{2r_{0}})$ and $u_{2}$ is analytic in $\mathcal{S}_{2r_{0}}$, it yields that $w\in H^{1}(\mathcal{S}_{2r_{0}})$, $\nabla w\in L^{2}(\mathcal{S}_{2r_{0}})$ and $f\in L^{2}(\mathcal{S}_{2r_{0}})$.
+Moreover, it is noted that the opening angle of the sector $\vartheta_{M}-\vartheta_{m}\in(0,\pi)$ and $\tilde{\varphi}\in H^{1}_{0}(V_{2r_{0}})$.
+Applying [22, Theorem 10.28] to the case of the polygonal network consists of only one subset $V_{2r_{0}}$, then we have $\tilde{\varphi}\in H^{2}(V_{2r_{0}})$.
+Consequently, we have $w\in H^{2}(\mathcal{S}_{t_{0}r_{0}})$ since $\varphi\equiv 1$ in $\mathcal{S}_{t_{0}r_{0}}$.
+Therefore, we conclude that $u_{1}\in H^{2}(\mathcal{S}_{t_{0}r_{0}})$.
+Now we focus on the triangular domain $V_{t_{0}r_{0}}$. Let $\psi\in C^{\infty}(\mathcal{S}_{t_{0}r_{0}})$ be a smooth cutoff function satisfying:
+$\psi\equiv 1\text{ in }\mathcal{S}_{t_{0}^{\prime}r_{0}}\quad\text{and}\quad\psi\equiv 0\text{ in }\mathcal{S}_{t_{0}r_{0}}\setminus\mathcal{S}_{t_{0}^{\prime}r_{0}},$
+where $t_{0}^{\prime}$ is fixed such that $0<t_{0}^{\prime}<t_{0}$ and $\mathcal{S}_{t_{0}^{\prime}r_{0}}\Subset V_{t_{0}r_{0}}$. Similarly, denote $\tilde{\psi}=w\psi$, we then have
+$\Delta\tilde{\psi}=g\ \mbox{in}\ V_{t_{0}r_{0}},\quad\tilde{\psi}=0\ \mbox{on}\ \partial V_{t_{0}r_{0}},$ (3.8)
 […]
-where $k$ is a positive constant and $q$ is in the form of (3.1).
-If $u_{2}$ is real analytic in $\mathcal{S}_{2r_{0}}$, then $u_{1}\in C^{1,\alpha}(\overline{\mathcal{S}_{r_{0}}})$, where $\alpha\in(0,1)$.
-Before proving Lemma 3.1, we first introduce a triangle $V_{2r_{0}}$ described below, which will be used in a subsequent argument.
-Let us pick up the points $\mathbf{y}_{+}$ and $\mathbf{y}_{-}$ on $\Gamma_{2r_{0}}^{+}$ and $\Gamma_{2r_{0}}^{-}$, respectively, where $|\mathbf{y}_{\pm}|\in(r_{0},2r_{0})$.
-Denote $\Lambda_{2r_{0}}^{{}^{\prime}}=\mathbf{y}_{+}-\mathbf{y}_{-}$.
-Then the triangle domain $V_{2r_{0}}$ is formed by three vectors $\mathbf{y}_{+}$, $\mathbf{y}_{-}$ and $\Lambda_{2r_{0}}^{{}^{\prime}}$.
-It is clear that $\overline{\mathcal{S}_{r_{0}}}\subset\overline{V_{2r_{0}}}\subset\overline{\mathcal{S}_{2r_{0}}}$.
-We can refer to Fig.
-2 for a geometric illustration of $V_{2r_{0}}$.
-Figure 2. A schematic illustration of $\mathcal{S}_{2r_{0}}$, $V_{2r_{0}}$ and $\mathcal{S}_{r_{0}}$.
-The proof of Lemma 3.1.
-Let $w=u_{1}-u_{2},$ then from (3.7) it can be directly calculated that
-$\begin{cases}\Delta w=f,\hskip 15.6491pt\mbox{in}\ V_{2r_{0}},\\ w=\varphi,\hskip 22.76228pt\mbox{on}\ \Lambda^{{}^{\prime}}_{2r_{0}},\\ w=0,\hskip 24.18501pt\mbox{on}\ \Gamma_{2r_{0}}^{\pm},\end{cases}$
-[…]
-$w({\bf x})=w_{\rm reg}({\bf x})+\sum_{0<\lambda_{\ell}<3-\frac{2}{p}}\zeta^{\ell}r^{\lambda_{\ell}}\sin(\lambda_{\ell}\vartheta),\ w_{\rm reg}\in W^{3,p}(V_{2r_{0}}),\ {\bf x}=r(\cos\vartheta,\sin\vartheta),$
-where $\zeta$ is a constant depending only on $f$ and $\varphi$, and $\lambda_{\ell}=\frac{\ell\pi}{\vartheta_{M}-\vartheta_{m}},\ \ell\in\mathbb{N}$.
-If $\vartheta_{M}-\vartheta_{m}\in(0,\pi/2]$, then we know that $w({\bf x})=w_{\rm reg}({\bf x})$.
-If $\vartheta_{M}-\vartheta_{m}\in(\pi/2,\pi)$, it can be verified that $\frac{2}{3-\frac{\pi}{\vartheta_{M}-\vartheta_{m}}}\in(1,2)$.
-Hence we can choose $p\in(1,\frac{2}{3-\frac{\pi}{\vartheta_{M}-\vartheta_{m}}})$.
-This indicates that $\lambda_{\ell}>3-\frac{2}{p}$ for any $\ell\in\mathbb{N}$.
-Therefore, $w=w_{\rm reg}\in W^{3,p}(V_{2r_{0}})$.
-By virtue of the Sobolev embedding theorem (cf.
-[28, Theorem 7.26]), we have that $w=w_{\rm reg}\in C^{1,\alpha}(\overline{V_{2r_{0}}})$ for $\alpha\in(0,1)$.
+$\tilde{\psi}({\bf x})=\tilde{\psi}_{\rm reg}({\bf x})+\sum_{0<\lambda_{\ell}<3-\frac{2}{p}}\zeta_{\ell}r^{\lambda_{\ell}}\sin(\lambda_{\ell}\vartheta),\ \tilde{\psi}_{\rm reg}\in W^{3,p}(V_{2r_{0}}),\ {\bf x}=r(\cos\vartheta,\sin\vartheta),$
+where $\zeta$ is a constant depending only on $f$ and $\varphi$, and $\lambda_{\ell}=\frac{\ell\pi}{\vartheta_{M}-\vartheta_{m}},\ \ell\in\mathbb{N}_{+}$.
+If $\vartheta_{M}-\vartheta_{m}\in(0,\pi/2]$, then we know that $\tilde{\psi}({\bf x})=\tilde{\psi}_{\rm reg}({\bf x})$ since $\lambda_{\ell}\geq 2\ell>3-\frac{2}{p}$.
+If $\vartheta_{M}-\vartheta_{m}\in(\pi/2,\pi)$, we then choose $p\in(1,\frac{2}{3-\frac{\pi}{\vartheta_{M}-\vartheta_{m}}})$ since it can be verified that $\frac{2}{3-\frac{\pi}{\vartheta_{M}-\vartheta_{m}}}\in(1,2)$, which also indicates that $\lambda_{\ell}>3-\frac{2}{p}$ for any $\ell\in\mathbb{N}_{+}$.
+Therefore, $\tilde{\psi}=\tilde{\psi}_{\rm reg}\in W^{3,p}(V_{t_{0}r_{0}})$ for $p\in(1,2)$.
+Recall that $\psi\equiv 1$ in $\mathcal{S}_{t_{0}^{\prime}r_{0}}$, then we deduce that $w\in W^{3,p}(\mathcal{S}_{t_{0}^{\prime}r_{0}})$ which implies that $u_{1}\in W^{3,p}(\mathcal{S}_{t_{0}^{\prime}r_{0}})$ since $u_{2}$ is analytic in $\mathcal{S}_{t_{0}^{\prime}r_{0}}$.
+It follows from the Sobolev embedding theorem (cf.
+[38, Theorem 7.26]), we have that $u_{1}\in C^{1,\alpha}(\overline{\mathcal{S}_{t_{0}^{\prime}r_{0}}})$ with $\alpha\in(0,1)$, where $\alpha$ depends on the opening angle of the sector $\mathcal{S}_{2r_{0}}$.
 The proof is complete. ∎
@@
 Definition 3.5.
-Suppose $(\Omega;q,\eta)$ represents a conductive polygonal-nest or polygonal-cell medium scatterer as defined in Definitions 3.1 or 3.2.
+Suppose $(\Omega;q,\eta)$ represents a conductive polygonal-nest or polygonal-cell medium scatterer as defined in Definitions 3.3 or 3.4.
 Let $u$ denote the total wave field of (1.6) associated with $(\Omega;q,\eta)$.
@@
       • (I)
-        $\lim_{\rho\to 0}\frac{1}{m(B(\mathbf{x}_{p},\rho))}\int_{B(\mathbf{x}_{p},\rho)}|u(\mathbf{x}_{p})|\not=0,$ (3.8)
+        $\lim_{\rho\to+0}\frac{1}{m(B(\mathbf{x}_{p},\rho))}\int_{B(\mathbf{x}_{p},\rho)}|u(\mathbf{x})|\mathrm{d}\mathbf{x}\not=0,$ (3.9)
       • (II)
         if
-        $\lim_{\rho\to 0}\frac{1}{m(B(\mathbf{x}_{p},\rho))}\int_{B(\mathbf{x}_{p},\rho)}|u(\mathbf{x}_{p})|\mathrm{d}\mathbf{x}=0,\ \mbox{then}$
-        $\lim_{\rho\to 0}\frac{1}{m(B(\mathbf{x}_{p},\rho))}\int_{B(\mathbf{x}_{p},\rho)}|\nabla u(\mathbf{x}_{p})|\mathrm{d}\mathbf{x}\neq 0,$ (3.9)
+        $\lim_{\rho\to+0}\frac{1}{m(B(\mathbf{x}_{p},\rho))}\int_{B(\mathbf{x}_{p},\rho)}|u(\mathbf{x})|\mathrm{d}\mathbf{x}=0,\ \mbox{then}$
+        $\lim_{\rho\to+0}\frac{1}{m(B(\mathbf{x}_{p},\rho))}\int_{B(\mathbf{x}_{p},\rho)}|\nabla u(\mathbf{x})|\mathrm{d}\mathbf{x}\neq 0,$ (3.10)
       where $m(B(\mathbf{x}_{p},\rho))$ is the measure of $B(\mathbf{x}_{p},\rho)$, $\mathbf{x}_{p}\in{\mathcal{V}}(\Sigma_{i})$ and ${\mathcal{V}}(\Sigma_{i})=\{\mathbf{x}_{p,i}\}_{p=1}^{\ell_{i}}(3\leq\ell_{i}\in\mathbb{Z}_{+})$ is the vertex set of the convex polygon $\Sigma_{i}$.
@@
 where $C$ is a positive number and $B_{R}$ is a ball centered at the origin with radius $R$.
-Hence, if (3.10) is satisfied and the conductive boundary parameter $\eta$ is sufficiently small, the plane incident wave $u^{i}$ dominates the total wave $u=u^{i}+u^{s}$.
-It is evident that $u^{i}$ fulfills (3.8) (or (3.9)) , which implies that the total wave $u$ generically satisfies (3.8) or (3.9).
+Hence, if (3.11) is satisfied and the conductive boundary parameter $\eta$ is sufficiently small, the incident wave $u^{i}$ dominates the total wave $u=u^{i}+u^{s}$.
+It is evident that the incident wave $u^{i}$ fulfills (3.9) (or (3.10)) , which implies that the total wave $u$ generically satisfies (3.9) or (3.10).
 On the other hand, from a physical standpoint, in conductive medium scattering, if the total wave does not satisfies (3.9) at a corner point, it implies that the scattered wave and the incident wave cancel each other at those points.
@@
   • a)
-    $\lim_{\rho\to 0}\frac{1}{m(B(\mathbf{x}_{p},\rho))}\int_{B(\mathbf{x}_{p},\rho)}|\partial_{1}u(\mathbf{x}_{p})|\mathrm{d}\mathbf{x}\neq 0$,
-    $\lim_{\rho\to 0}\frac{1}{m(B(\mathbf{x}_{p},\rho))}\int_{B(\mathbf{x}_{p},\rho)}|u(\mathbf{x}_{p})|\mathrm{d}\mathbf{x}=\lim_{\rho\to 0}\frac{1}{m(B(\mathbf{x}_{p},\rho))}\int_{B(\mathbf{x}_{p},\rho)}|\partial_{2}u(\mathbf{x}_{p})|\mathrm{d}\mathbf{x}=0$,
+    $\lim_{\rho\to+0}\frac{1}{m(B(\mathbf{x}_{p},\rho))}\int_{B(\mathbf{x}_{p},\rho)}|\partial_{1}u(\mathbf{x})|\mathrm{d}\mathbf{x}\neq 0$,
+    $\lim_{\rho\to+0}\frac{1}{m(B(\mathbf{x}_{p},\rho))}\int_{B(\mathbf{x}_{p},\rho)}|u(\mathbf{x})|\mathrm{d}\mathbf{x}=\lim_{\rho\to+0}\frac{1}{m(B(\mathbf{x}_{p},\rho))}\int_{B(\mathbf{x}_{p},\rho)}|\partial_{2}u(\mathbf{x})|\mathrm{d}\mathbf{x}=0$,
   • b)
-    $\lim_{\rho\to 0}\frac{1}{m(B(\mathbf{x}_{p},\rho))}\int_{B(\mathbf{x}_{p},\rho)}|\partial_{2}u(\mathbf{x}_{p})|\mathrm{d}\mathbf{x}\neq 0$,
-    $\lim_{\rho\to 0}\frac{1}{m(B(\mathbf{x}_{p},\rho))}\int_{B(\mathbf{x}_{p},\rho)}|u(\mathbf{x}_{p})|\mathrm{d}\mathbf{x}=\lim_{\rho\to 0}\frac{1}{m(B(\mathbf{x}_{p},\rho))}\int_{B(\mathbf{x}_{p},\rho)}|\partial_{1}u(\mathbf{x}_{p})|\mathrm{d}\mathbf{x}=0$,
+    $\lim_{\rho\to+0}\frac{1}{m(B(\mathbf{x}_{p},\rho))}\int_{B(\mathbf{x}_{p},\rho)}|\partial_{2}u(\mathbf{x})|\mathrm{d}\mathbf{x}\neq 0$,
+    $\lim_{\rho\to+0}\frac{1}{m(B(\mathbf{x}_{p},\rho))}\int_{B(\mathbf{x}_{p},\rho)}|u(\mathbf{x})|\mathrm{d}\mathbf{x}=\lim_{\rho\to+0}\frac{1}{m(B(\mathbf{x}_{p},\rho))}\int_{B(\mathbf{x}_{p},\rho)}|\partial_{1}u(\mathbf{x})|\mathrm{d}\mathbf{x}=0$,
   • c)
-    $\lim_{\rho\to 0}\frac{1}{m(B(\mathbf{x}_{p},\rho))}\int_{B(\mathbf{x}_{p},\rho)}|\partial_{1}u(\mathbf{x}_{p})|\mathrm{d}\mathbf{x}\neq 0$,
-    $\lim_{\rho\to 0}\frac{1}{m(B(\mathbf{x}_{p},\rho))}\int_{B(\mathbf{x}_{p},\rho)}|\partial_{2}u(\mathbf{x}_{p})|\mathrm{d}\mathbf{x}\neq 0$, $\lim_{\rho\to 0}\frac{1}{m(B(\mathbf{x}_{p},\rho))}\int_{B(\mathbf{x}_{p},\rho)}|u(\mathbf{x}_{p})|\mathrm{d}\mathbf{x}=0$.
+    $\lim_{\rho\to+0}\frac{1}{m(B(\mathbf{x}_{p},\rho))}\int_{B(\mathbf{x}_{p},\rho)}|\partial_{1}u(\mathbf{x})|\mathrm{d}\mathbf{x}\neq 0$,
+    $\lim_{\rho\to+0}\frac{1}{m(B(\mathbf{x}_{p},\rho))}\int_{B(\mathbf{x}_{p},\rho)}|\partial_{2}u(\mathbf{x})|\mathrm{d}\mathbf{x}\neq 0$, $\lim_{\rho\to+0}\frac{1}{m(B(\mathbf{x}_{p},\rho))}\int_{B(\mathbf{x}_{p},\rho)}|u(\mathbf{x})|\mathrm{d}\mathbf{x}=0$.
 Since (3.10) encompasses more cases compared to (3.9), it is more likely to be satisfied in practical scenarios.
@@
 Theorem 3.1.
-Let $(\Omega_{\ell};q_{\ell},\eta_{\ell}),\ell=1,2$ be two admissible conductive polygonal-nest or polygonal-cell medium scatterer satisfying (1.7), $u^{\infty}_{\ell}(\mathbf{\hat{x}};u^{i})$ be the far-field pattern associated with the scatterer $(\Omega_{\ell};q_{\ell},\eta_{\ell})$ and the incident field $u^{i}$.
+Let $(\Omega_{\ell};q_{\ell},\eta_{\ell}),\ell=1,2$ be both admissible conductive polygonal-nest or both polygonal-cell medium scatterers satisfying (1.6), $u^{\infty}_{\ell}(\mathbf{\hat{x}};u^{i})$ be the far-field pattern associated with the scatterers $(\Omega_{\ell};q_{\ell},\eta_{\ell})$ and the incident field $u^{i}$.
 If
 $u^{\infty}_{1}(\mathbf{\hat{x}};u^{i})=u^{\infty}_{2}(\mathbf{\hat{x}};u^{i})$
-is fulfilled for all $\mathbf{\hat{x}}\in\mathbb{S}^{1}$ and a fixed incident wave $u^{i}$, then $\Omega_{1}\Delta\Omega_{2}$ cannot have a convex corner.
-Moreover, if $\Omega_{1}$ and $\Omega_{2}$ are two admissible conductive polygonal-nest medium scatterers, one holds that
+is fulfilled for all $\mathbf{\hat{x}}\in\mathbb{S}^{1}$ and a fixed incident wave $u^{i}$, then the difference $\Omega_{1}\aleph\Omega_{2}$ defined in Corollary 1.3 cannot have a convex corner.
+Moreover, if $\Omega_{1}$ and $\Omega_{2}$ are both admissible conductive polygonal-nest medium scatterers, one holds that
 $\partial\Omega_{1}=\partial\Omega_{2}.$
@@
 We proceed to prove this theorem by contradiction.
-Let us assume that there exists a corner in $\Omega_{1}\Delta\Omega_{2}$.
+Let us assume that there exists a corner in $\Omega_{1}\aleph\Omega_{2}$.
 According to (3.2), we can express the admissible conductive polygonal-nest medium scatterers as $(\Omega_{\ell};q_{\ell},\eta_{\ell})$ ($\ell=1,2$) as follows
@@
 $\eta_{1}\int_{\Gamma_{r_{0}}^{\pm}}u_{1}^{-}u_{0}\mathrm{d}\mathbf{x}=k^{2}\int_{\mathcal{S}_{r_{0}}}(1-q_{1})u_{1}^{-}u_{0}\mathrm{d}\mathbf{x}+\int_{\Lambda_{r_{0}}}\partial_{\nu}u_{0}\bar{u}-\partial_{\nu}\bar{u}u_{0}\mathrm{d}\sigma.$ (3.14)
-According to Lemma 3.1, we know that $u_{1}^{-}\in C^{1,\alpha}(\overline{\mathcal{S}_{r_{0}}}),\alpha\in(0,1)$, which implies the following expansion
+According to Lemma 3.1, there exists a positive number $t_{0}^{\prime}\in\mathbb{R}_{+}$ such that $u_{1}^{-}\in C^{1,\alpha}(\overline{\mathcal{S}_{t_{0}^{\prime}r_{0}}}),\alpha\in(0,1)$, which implies the following expansion
 $u_{1}^{-}=u_{1}^{-}(\mathbf{0})+\partial_{1}u_{1}^{-}(\mathbf{0})x_{1}+\partial_{2}u_{1}^{-}(\mathbf{0})x_{2}+\delta u_{1}^{-},\ |\delta u_{1}^{-}|\leq|\mathbf{x}|^{1+\alpha}\|u_{2}^{+}\|_{C^{1,\alpha}}.$ (3.15)
+In the subsequent analysis, we need to establish an integral identity in $\mathcal{S}_{t_{0}^{\prime}r_{0}}$ that is analogous to (3.14).
+However, our analysis indicates that the parameter $t_{0}^{\prime}$ defining $\mathcal{S}_{t_{0}^{\prime}r_{0}}$ does not play a crucial role in proving this theorem.
+Therefore, we will instead use (3.14) for our further analysis.
 According to the admissibility conditions in Definition 3.5, we will divide the proof into four cases, as follows.
@@
 Remark 3.3.
-From the proof of Theorem 3.1, we can see that when the conductive boundary parameter $\eta$ is a Hölder continuous function on the boundary and $\eta$ does not vanish at the vertex of the scatterer, we can also establish the local unique result for determining the shape of an admissible conductive polygonal-nest or polygonal-cell medium scatterer by a single far-field pattern.
+From the proof of Theorem 3.1, we can see that when the conductive boundary parameter $\eta$ is a Hölder continuous function on the boundary and $\eta$ does not vanish at the vertex of the scatterer, we can also establish the local uniqueness result for determining the shape of an admissible conductive polygonal-nest or polygonal-cell medium scatterer by a single far-field pattern.
 The corresponding proof can be modified slightly from the proof of Theorem 3.1.
@@
 $q_{\ell}=\sum_{i=1}^{N}q_{i,\ell}\chi_{\Sigma_{i}},\ \eta_{\ell}=\sum_{i=1}^{N}\tilde{\eta}_{\ell}\chi_{\Sigma_{i}}.$
-The coefficients of $q_{i,\ell}$ $\in\mathbb{C}$, where $q_{i,\ell}$ follows the form specified in (3.1), while $\tilde{\eta}_{\ell}$ $\in\mathbb{C}$.
+where the coefficients of $q_{i,\ell}$ are complex valued following the form specified in (3.1) and $\tilde{\eta}_{\ell}$ $\in\mathbb{C}$.
 Let $u_{\ell}^{\infty}(\mathbf{\hat{x}};u^{i})$ be the far-field pattern associated with the incident wave $u^{i}$ corresponding to $(\Omega;q_{\ell},\eta_{\ell})$.
@@
 $q_{\ell}=\sum_{i=1}^{N_{\ell}}q_{i,\ell}\chi_{\mathcal{D}_{i,\ell}},\ \eta_{\ell}=\sum_{i=1}^{N_{\ell}}\eta_{i,\ell}\chi_{\partial\Sigma_{i,\ell}}.$
-The coefficients of $q_{i,\ell}$ $\in\mathbb{C}$ , where $q_{i,\ell}$ follows the form specified in (3.1), while $\eta_{i,\ell}$ $\in\mathbb{C}$.
+The coefficients of $q_{i,\ell}$ are complex valued, where $q_{i,\ell}$ follows the form specified in (3.1), while $\eta_{i,\ell}$ $\in\mathbb{C}$.
 Let $u_{\ell}^{\infty}(\mathbf{\hat{x}};u^{i})$ be the far-field pattern associated with the incident wave $u^{i}$ corresponding to $(\Omega_{\ell};q_{\ell},\eta_{\ell})$.
@@
 The proofs of Theorems 3.2 and 3.3 are deferred to Section 4.
-Prior to that, we require the following lemma, which establishes local uniqueness results for identifying the refractive index and conductive boundary parameter of a coupled PDE system near a corner.
+Prior to that, we require the following lemmas.
+Lemma 3.2 establishes that the solution $u^{-}\in H^{1}(\Omega)$ to the conductive scattering problem (1.6) is Hölder continuous in the vicinity of the corner.
+Furthermore, Lemma 3.3 investigates the local behavior of solutions to the coupled PDE system (3.31) near a corner.
+Furthermore, we obtain local uniqueness for identifying both the refractive index and the conductive boundary parameter of the coupled PDE system.
+Lemma 3.2.
+Consider a conductive medium scatterer $(\Omega;q,\eta)$ possessing either a polygonal-nest structure or polygonal-cell structure defined in Definitions 3.3 and 3.4 associated with the convex polygons $\Sigma_{i},i=1,\dots,N$ given in Definitions 3.1 or 3.2.
+Recall that $\mathcal{V}(\Sigma_{i})$ and $\mathcal{V}(\Omega)$ as defined in Definition 3.5 denote the vertex sets of the polygons $\Sigma_{i}$ for the polygonal-nest and of the $\Omega$ for the polygonal-cell structure, respectively.
+Let $\mathbf{x}_{p}\in\mathcal{V}(\Sigma_{i})$ or $\mathbf{x}\in\mathcal{V}(\Omega)$ be a vertex point.
+For sufficiently small radius $r_{0}$, define $\mathcal{S}_{\mathbf{x}_{p},r_{0}}=B_{\mathbf{x}_{p},r_{0}}\cap\Sigma_{i}$ for the polygonal-nest structure and $\mathcal{S}_{\mathbf{x}_{p},\rho}=B_{\mathbf{x}_{p},r_{0}}\cap\Omega$ for the polygonal-cell structure.
+If $u=u^{-}\chi_{\Omega}+u^{+}\chi_{\mathbb{R}^{2}\setminus\overline{\Omega}}\in H^{1}_{loc}(\mathbb{R}^{2})$ satisfies the conductive scattering problem (1.6), then one has $u^{-}\in C^{\alpha}(\overline{\mathcal{S}_{\mathbf{x}_{p},r_{0}}}),\ \alpha\in(0,1]$.
+Figure 2. A schematic illustration of network for polygonal-nest structure.
+Proof.
+Let $\mathcal{P}$ and $\mathcal{P}^{{}^{\prime}}$ be bounded, connected and convex polygons such that $\overline{\Omega}\subset\overline{\mathcal{P}{{}^{\prime}}}\subset\overline{\mathcal{P}}$.
+Here, $\Omega$ has either a polygonal-nest structure or polygonal-cell structure.
+First, by an argument analogous to the proof in Lemma 3.1, let $\varphi\in C^{\infty}(\mathcal{P})$ be a smooth cutoff function satisfying
+$\varphi\equiv 1\mbox{ in }\Omega,\quad\varphi\equiv 0\mbox{ in }\mathcal{P}\setminus\mathcal{P}^{{}^{\prime}}.$
+Define $\tilde{\varphi}=u\varphi$. Using the conductive scattering problem (1.6), we obtain that
+$\Delta\tilde{\varphi}=f\ \mbox{in}\ \Omega,\quad\Delta\tilde{\varphi}=g\ \mbox{in}\ \mathcal{P}\setminus{\overline{\Omega}},\quad\tilde{\varphi}=0\ \mbox{on}\ \partial\mathcal{P}.$
+where $f=(-k^{2}qu^{-})\varphi+\Delta\varphi u^{-}+2\nabla u^{-}\cdot\nabla\varphi$ and $g=(-k^{2}u^{+})\varphi+\Delta\varphi u^{+}+2\nabla u^{+}\cdot\nabla\varphi$.
+Recall that $u^{-}\in H^{1}({\Omega})$ and $u^{+}\in H^{1}(\mathbb{R}^{2}\setminus\overline{\Omega})$, it is clear that $f\in L^{2}(\Omega)$ and $g\in L^{2}(\mathcal{P}\setminus{\overline{\Omega}})$.
+Moreover, through appropriate segmentation of $\mathcal{P}$ and $\Omega$, these domains naturally constitute a polygonal network as defined in [22, Definition 8.A].
+The polygonal-cell structure case follows immediately.
+Figure 2 provides a schematic illustration of the polygonal-nest structure case.
+Moreover, it is noted that $\tilde{\varphi}$ satisfies the continuity condition at the interface within the interior of $\mathcal{P}$.
+Consequently, by [22, Theorem 10.28], we conclude that $\tilde{\varphi}\in H^{2}(\mathcal{S}_{\mathbf{x}_{p},\rho})$ due to the convexity of $\mathcal{S}_{\mathbf{x}_{p},\rho}$.
+Therefore, by virtue of $\varphi\equiv 1$ in $\Omega$, one has $u^{-}\in H^{2}(\mathcal{S}_{\mathbf{x}_{p},\rho})$.
+Thus, the Sobolev embedding theorem implies that $u^{-}\in C^{\alpha}(\overline{\mathcal{S}_{\mathbf{x}_{p},\rho}}),\alpha\in(0,1]$.
+The proof is complete. ∎
 Lemma 3.3.
 Recall that $\mathcal{S}_{r_{0}},\Gamma_{r_{0}}^{\pm}$ and $B_{r_{0}}$ are defined in (1.3).
-Suppose that $u_{1}=u_{1}^{-}\chi_{\mathcal{S}_{r_{0}}}+u_{1}^{+}\chi_{B_{r_{0}}\setminus\overline{\mathcal{S}_{r_{0}}}}\in H^{1}(B_{r_{0}})$ and $u_{2}=u^{-}_{2}\chi_{\mathcal{S}_{r_{0}}}+u^{+}_{2}\chi_{B_{r_{0}}\setminus\overline{\mathcal{S}_{r_{0}}}}\in H^{1}(B_{r_{0}})$ satisfy the following PDEs system:
+Suppose that $u_{1}=u_{1}^{-}\chi_{\mathcal{S}_{r_{0}}}+u_{1}^{+}\chi_{B_{r_{0}}\setminus\overline{\mathcal{S}_{r_{0}}}}\in H^{1}(B_{r_{0}})$ and $u_{2}=u^{-}_{2}\chi_{\mathcal{S}_{r_{0}}}+u^{+}_{2}\chi_{B_{r_{0}}\setminus\overline{\mathcal{S}_{r_{0}}}}\in H^{1}(B_{r_{0}})$ satisfy the following PDE system:
 $\begin{cases}(\Delta+k^{2}q_{1})u_{1}^{-}=0,\hskip 78.24507pt\mbox{in}\ \mathcal{S}_{r_{0}},\\ (\Delta+k^{2}q_{2})u_{2}^{-}=0,\hskip 78.24507pt\mbox{in}\ \mathcal{S}_{r_{0}},\\ (\Delta+k^{2}q_{3})u^{+}_{1}=0,\hskip 78.24507pt\mbox{in}\ B_{r_{0}}\setminus\overline{\mathcal{S}_{r_{0}}},\\ (\Delta+k^{2}q_{3})u^{+}_{2}=0,\hskip 78.24507pt\mbox{in}\ B_{r_{0}}\setminus\overline{\mathcal{S}_{r_{0}}},\\ u^{+}_{1}=u^{-}_{1},\ \partial_{\nu}u^{-}_{1}=\partial_{\nu}u^{+}_{1}+\eta_{1}u^{+}_{1},\hskip 12.80365pt\mbox{on}\ \Gamma_{r_{0}}^{\pm},\\ u^{+}_{2}=u^{-}_{2},\ \partial_{\nu}u^{-}_{2}=\partial_{\nu}u^{+}_{2}+\eta_{2}u^{+}_{2},\hskip 12.80365pt\mbox{on}\ \Gamma_{r_{0}}^{\pm},\\ \end{cases}$ (3.31)
@@
 Assume that
-$\lim_{\rho\to 0}\frac{1}{m(B(\mathbf{0},\rho))}\int_{B(\mathbf{x}_{p},\rho)}|u_{i}(\mathbf{0})|\mathrm{d}\mathbf{x}\neq 0,\ i=1\ \mbox{or}\ 2.$ (3.31)
+$\lim_{\rho\to 0}\frac{1}{m(B(\mathbf{x}_{p},\rho))}\int_{B(\mathbf{x}_{p},\rho)}|u_{i}(\mathbf{x})|\mathrm{d}\mathbf{x}\neq 0,\ i=1\ \mbox{or}\ 2.$ (3.32)
 If
@@
 Since $u^{+}_{1}=u_{2}^{+}$ in $B_{r_{0}}\setminus\overline{\mathcal{S}_{r_{0}}}$, we have that $u^{+}_{1}=u^{+}_{2}\ \mbox{on}\ \Gamma_{r_{0}}^{\pm}$ by the trace theorem, which indicates that $u_{1}^{-}=u_{1}^{+}=u_{2}^{+}=u_{2}^{-}\mbox{on}\ \Gamma_{r_{0}}^{\pm}$ according to the transmission boundary condition.
-Without loss of generality, we assume that $u_{2}^{-}$ fulfills (3.31).
+Without loss of generality, we assume that $u_{2}$ fulfills (3.32).
 Denote $\tilde{u}:=u^{-}_{1}-u_{2}^{-}$, then it can be shown directly that
@@
 $\int_{\mathcal{S}_{r_{0}}}k^{2}(q_{2}-q_{1})u_{2}^{-}u_{0}\mathrm{d}\mathbf{x}=(\eta_{1}-\eta_{2})\int_{\Gamma_{r_{0}}^{\pm}}u_{2}^{-}u_{0}\mathrm{d}\sigma+\int_{\Lambda_{r_{0}}}u_{0}\partial_{\nu}\tilde{u}-\tilde{u}\partial_{\nu}u_{0}\mathrm{d}\sigma.$ (3.33)
-First, we demonstrate that the solution $u_{i}^{-}\in H_{1}(\mathcal{S}_{r_{0}}),i=1,2$ of the system (3.30) belong to $C^{\alpha}(\overline{\mathcal{S}_{r_{0}}})$.
-From [39, 43], we know that $u_{i}^{-}$ admits a decomposition into a regular part and a singular part in the following form:
-$u_{i}^{-}=u_{i,reg}^{-}+u_{i,sing}^{-}.$
-The regular part $u_{i,reg}^{-}\in H^{2}(\mathcal{S}_{r_{0}})$ and hence by the Sobolev embedding theorem, it follows that $u_{i,reg}^{-}\in C^{\beta}(\mathcal{S}_{r_{0}}),\beta\in(0,1]$.
-Moreover, $u_{{i,sing}}^{-}\in C^{\alpha}(\overline{\mathcal{S}_{r_{0}}})$, where $\alpha\in(0,1)$ is the Hölder exponent depending on the opening angle of $\mathcal{S}_{r_{0}}$.
-Therefore, $u_{i}^{-}\in C^{\alpha}(\overline{\mathcal{S}_{r_{0}}}).$ Thus, $u_{2}^{-}$ satisfies the following expansion:
+By virtue of Lemma 3.2, we know that $u_{i}^{-}\in C^{\alpha}(\overline{\mathcal{S}_{r_{0}}}),\alpha\in(0,1],i=1,2.$ Thus, $u_{2}^{-}$ satisfies the following expansion:
 $u_{2}^{-}$ $=u_{2}^{-}(\mathbf{0})+\delta u_{2}^{-},\ |\delta u_{2}^{-}|\leq|\mathbf{x}|^{\alpha}\|u_{2}^{-}\|_{C^{\alpha}}.$ (3.34)
@@
 Figure 3. Schematic illustration of the polygonal-cell structure.
-Recall that $u_{1}$ and $u_{2}$ be the total wave fields associated with the admissible conductive polygonal-cell medium scatterers $(\Omega;q_{\ell},\eta_{\ell}),\ell=1,2$.
+Recall that $u_{1}$ and $u_{2}$ are the total wave fields associated with the admissible conductive polygonal-cell medium scatterers $(\Omega;q_{\ell},\eta_{\ell}),\ell=1,2$.
 Since $u_{1}^{\infty}(\mathbf{\hat{x}};u^{i})=u_{2}^{\infty}(\mathbf{\hat{x}};u^{i})$ for all $\mathbf{\hat{x}}\in\mathbb{S}^{1},$ by Rellich lemma and the unique continuation principle, we have
@@
 Diao is supported by by National Natural Science Foundation of China (No.
-12371422) and the Fundamental Research Funds for the Central Universities, JLU (No.
-93Z172023Z01).
+12371422) and the Fundamental Research Funds for the Central Universities, JLU.
 The work of X.
@@
 The work of H.
-Liu is supported by the NSFC/RGC Joint Research Scheme, N_CityU101/21; ANR/RGC Joint Research Scheme, A_CityU203/19; and the Hong Kong RGC General Research Funds (projects 11311122, 12301420 and 11300821).
+Liu was supported by the Hong Kong RGC General Research Funds (projects 11311122, 11300821 and 11304224), NSF/RGC Joint Research Fund (project N_CityU101/21) and the ANR/RGC Joint Research Fund (project A_CityU203/19).